### v3

# Corks, exotic 4-manifolds and genus functions

Abstract.
We prove that every 4-dimensional oriented handlebody without 3- and 4-handles can be modified to admit infinitely many exotic smooth structures, and further that their genus functions are pairwise equivalent. We furthermore show that, for any 4-manifold admitting an embedding into a symplectic 4-manifold with weakly convex boundary, its genus function is algebraically realized as those of infinitely many pairwise exotic 4-manifolds. We also prove that, for any two 4-manifolds with non-degenerate intersection forms, algebraic inequivalences of genus functions are stable under connected sums and boundary sums with a certain type of 4-manifolds having arbitrarily large second Betti numbers. Furthermore, we introduce a notion of genus function type for diffeomorphism invariants and show that similar results hold for any such invariant. As an application of our exotic 4-manifolds, we also prove that for any (possibly non-orientable) 4-manifold, every submanifold of codimension less than two satisfying a mild condition can be modified to admit infinitely many exotically knotted copies.
Key words and phrases:
4-manifolds; genus functions; corks; knotted submanifolds; cobordisms
2020 Mathematics Subject Classification:
Primary 57R55, Secondary 57K41, 57R65, 57K43
1. Introduction
The genus function of a 4-manifold is a powerful invariant of smooth structures and has been extensively studied dating back to the work of Kervaire and Milnor [44] in 1961. In general, it is quite hard to determine values of genus functions, and they are completely determined only for few 4-manifolds (see [48], [49], [20]). Nevertheless, genus functions can detect (infinitely many) exotic smooth structures for numerous 4-manifolds (e.g. [49], [11], [67], [35], [12], [68], [70]) and have various applications to low dimensional topology (e.g. [6], [7], [50], [69], [70], [71], [42], [1], [65], [19]).
Lawson conjectured in his 1997 survey [49] that, for simply connected compact 4-manifolds, genus functions determine smooth structures. Indeed, for rational surfaces and ruled surfaces, genus functions uniquely determine their smooth structures among symplectic 4-manifolds (see Proposition 2.6). It later turned out that his conjecture is false ([8], [36]), but the counterexamples are not interesting, because their second homology groups are trivial, and hence the underlying topological structures uniquely determine genus functions. Thus, the conjecture remains intriguing, especially for 4-manifolds with large second Betti numbers. It would also be natural to ask whether a 4-manifold can simultaneously admit exotic smooth structures with the same genus functions and those with distinct genus functions.
Let us recall that the genus function of a connected oriented smooth 4-manifold $X$ is defined to be the map $H_{2}(X)\to\mathbb{Z}$ which sends a second homology class to the minimal genus of a smoothly embedded surface representing the class. Genus functions of two 4-manifolds will be called equivalent if there exists a homeomorphism between them whose induced isomorphism preserves the genus functions. They will be called algebraically equivalent if there exists an isomorphism between their second homology groups that preserves the genus functions and the intersection forms.
In this paper, we prove that every 4-manifold with boundary satisfying mild conditions can be modified to admit infinitely many exotic smooth structures. We then show that exotic 4-manifolds with equivalent genus functions exist in abundance. Furthermore, we study behavior of genus functions and properties of related invariants. We also produce exotically knotted submanifolds.
Convention 1.1.
Throughout this paper, all manifolds, submanifolds, boundaries, handlebodies and surfaces are assumed to be compact, connected, oriented and smooth unless otherwise stated. Also, two 4-manifolds are said to be exotic to each other if one is orientedly homeomorphic to but not unorientedly diffeomorphic to the other.
1.1. Exotic 4-manifolds and genus functions
To state our results shortly, we introduce the following terminology.
Definition 1.2.
For two 4-manifolds $X,Y$ having non-empty (possibly disconnected) boundary, we will say that $Y$ topologically resembles $X$ if $Y$ is homotopy equivalent to $X$, the intersection form of $Y$ is isomorphic to that of $X$, and $\partial Y$ is homology cobordant to $\partial X$.
Let us recall that a $2$-handlebody is a handlebody consisting of finitely many handles with indices not greater than two. They form a very large class of 4-manifolds with boundary, realizing arbitrary large integers as their second Betti numbers $b_{2}$. (Beware that in some papers, 2-handlebodies mean handlebodies consisting of only 0- and 2-handles, which form a strictly smaller class than ours.) We prove that every 4-dimensional 2-handlebody can be modified to admit infinitely many exotic smooth structures with pairwise equivalent genus functions, by utilizing corks.
Theorem 1.3.
For any $4$-dimensional $2$-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise equivalent.
Even if the condition on genus functions is removed, this result is still new and shows the existence of infinitely many exotic smooth structures for a very large class of 4-manifolds, strongly improving the following earlier results: every 2-handlebody with non-trivial second homology group can be modified to admit arbitrarily many exotic smooth structures (Akbulut and the author [11]); there exist contractible 4-manifolds admitting infinitely many exotic smooth structures (Gompf [36]).
The corollary below is straightforward from the above theorem.
Corollary 1.4.
$(1)$ For any integral symmetric bilinear form $Q$, there exist infinitely many pairwise exotic simply connected $4$-manifolds with pairwise equivalent genus functions such that their intersection forms are isomorphic to $Q$.
$(2)$ For any finitely presented group $G$, there exist infinitely many pairwise exotic $4$-manifolds with pairwise equivalent genus functions such that their fundamental groups are isomorphic to $G$.
$(3)$ For any closed $3$-manifold $M$, there exist infinitely many pairwise exotic simply connected $4$-manifolds with pairwise equivalent genus functions such that their boundaries are homology cobordant to $M$.
We furthermore prove that various genus functions can not determine smooth structures.
Theorem 1.5.
Suppose that a $4$-manifold $X$ admits an embedding into a symplectic $4$-manifold with weakly convex boundary. Then, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
The assumption of this theorem holds for a large class of smooth 4-manifolds. For example, blow ups of 4-manifolds admitting Stein structures and their codimension-00 submanifolds satisfy the assumption. Furthermore, these 4-manifolds provide varieties of genus functions even within fixed homeomorphism types ([11], [68], [70]). We here state only simple examples.
Corollary 1.6.
$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((((with boundary)))) such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
$(2)$ If $X$ is the total space of a $D^{2}$-bundle over a closed surface, then there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
The claim (1) contrasts with the aforementioned fact that the genus function of $\mathbb{C}P^{2}$ determines its smooth structure among closed symplectic 4-manifolds.
We next show that a fixed 4-manifold can simultaneously admit exotic smooth structures with pairwise equivalent genus functions and those with pairwise inequivalent genus functions. In fact, we prove the much stronger theorem below.
Theorem 1.7.
There exist infinitely many pairwise exotic simply connected $4$-manifolds with pairwise algebraically inequivalent genus functions such that each of their genus functions is equivalent to those of another infinitely many pairwise exotic $4$-manifolds.
We can construct similar infinite families for varieties of homeomorphism types. For example, any finitely presented group is realized as the fundamental group of infinite families of such exotic 4-manifolds. Here we give similar examples for much more homeomorphism types, by relaxing the condition “infinitely many” to “arbitrarily many”.
Theorem 1.8.
For any $4$-dimensional $2$-handlebody $X$ with non-trivial second homology group, there exist arbitrarily many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise algebraically inequivalent for any orientations and that each of their genus functions is equivalent to those of infinitely many pairwise exotic $4$-manifolds.
Exotic 4-manifolds in Theorems 1.3, 1.5 and 1.8 have more properties inherited from a given 4-manifold $X$, see Theorems 9.4, 9.5 and 9.8 for details.
It remains interesting to ask whether exotic closed 4-manifolds with equivalent genus functions exist. For closed 4-manifolds, we instead show that algebraic equivalent classes do not determine homeomorphism types even within 4-manifolds having isomorphic cohomology rings and that there are many such 4-manifolds, realizing arbitrarily large $b_{2}$. See Proposition 8.13 for details. During the preparation of this paper, Stipsicz and Szabó [61] proved that infinitely many pairwise non-diffeomorphic closed 4-manifolds with the cohomology ring of $S^{2}\times S^{2}$ obtained in [29] have pairwise algebraically equivalent genus functions. It is not known whether their 4-manifolds are pairwise homeomorphic. They showed algebraic equivalence by explicitly determining all the values of the genus functions, but our proof is completely different from their approach.
After the first version of this paper appeared on arXiv on January 2025, Nakamura submitted the paper [53] to arXiv which gives a different example of Corollary 1.4.(1) in the case where the intersection form is represented by the $1\times 1$ matrix 00.
1.2. Stabilities of algebraic inequivalences of genus functions
To show that our exotic 4-manifolds have equivalent genus functions, we prove that attachment of a cobordism of certain type preserves genus functions (See Sections 6 and 7). As an application, we discuss behavior of genus functions under connected sums and boundary sums with a certain type of 4-manifolds.
For an $S^{2}$-link $L$ in $S^{4}$, let $S^{4}(L)$ denote the closed 4-manifold obtained from $S^{4}$ by performing the surgery on $L$, that is, by removing the interior of the tubular neighborhood $S^{2}\times D^{2}$ for each component of $L$ and gluing $S^{1}\times D^{3}$ along the resulting boundary. A simple example of $S^{4}(L)$ is $n(S^{1}\times S^{3})$. Each $S^{4}(L)$ has varieties of codimension-00 submanifolds having arbitrary large $b_{2}$ (see Section 8). For example, $\Sigma_{g}\times D^{2}$ $(g\geq 0)$ with 2-handles attached along a strongly slice link with 0-framings are such examples, where $\Sigma_{g}$ denotes the closed surface of genus $g$.
In the rest of this subsection, for each $i=1,2$, let $X_{i}$ be a ((((possibly closed)))) 4-manifold, $L_{i}$ be an $S^{2}$-link in $S^{4}$, and $Z_{i}$ be a connected summand or a codimension-00 submanifold of $S^{4}(L_{i})$.
We will prove that the values of the genus functions of 4-manifolds are preserved under connected sums and boundary sums with $Z_{i}$’s, see Theorem 8.4. Applying this result, we show that algebraic inequivalences of genus functions are stable under these operations.
Theorem 1.9.
Suppose $H_{2}(Z_{1})=H_{2}(Z_{2})=0$. Then the following hold.
$(1)$ The genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
$(2)$ In the case where each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty, the genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
This result contrasts with the fact that the Seiberg–Witten invariants of closed 4-manifolds vanish after taking a connected sum with $S^{1}\times S^{3}$ and hence cannot detect them after the connected sum.
Even when $Z_{i}$’s have non-trivial $b_{2}$, a similar stability still holds under additional assumptions.
Theorem 1.10.
Suppose that the intersection form of each $X_{i}$ is non-degenerate and that each $H_{2}(Z_{i})$ is torsion-free. Then the following hold.
$(1)$ If the genus functions of $X_{1}$ and $X_{2}$ are algebraically inequivalent, then those of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically inequivalent.
$(2)$ Assume each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty. If the genus functions of $X_{1}$ and $X_{2}$ are algebraically inequivalent, then those of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically inequivalent.
We also give similar results without the above torsion-free assumption, see Theorems 8.10 and 8.11. These stabilities for algebraic inequivalences show that for any family of exotic 4-manifolds with algebraically inequivalent genus functions, the family remain (orientedly) exotic under connected sums and boundary sums with homeomorphic 4-manifolds satisfying the assumption. We can apply these stabilities to various families of exotic 4-manifolds having algebraically inequivalent genus functions (e.g. [11], [67], [68], [70]). Furthermore, we give similar stabilities for attachments of certain cobordisms, see Sections 6 and 7.
As an application of our arguments, we give exotic 4-manifolds which are stably exotic under boundary sums with arbitrary Stein manifolds having non-degenerate intersection forms.
Theorem 1.11.
For any positive integer $n$, there exist pairwise exotic simply connected $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ admitting Stein structures such that, for any pairwise orientedly homeomorphic $4$-manifolds $Y_{1},Y_{2},\dots,Y_{n}$ having non-degenerate intersection forms and admitting Stein structures, the boundary sums $X_{1}\natural Y_{1}$, $X_{2}\natural Y_{2}$, $\dots$, $X_{n}\natural Y_{n}$ remain pairwise exotic.
Note that the summands $Y_{i}$’s are not necessarily pairwise diffeomorphic. As seen from the proof, we can realize various homeomorphism types as those of $X_{i}$’s. Although the above theorem assumes that each $Y_{i}$ admits a Stein structure, the above theorem holds under the weaker condition that each $Y_{i}$ admits an embedding into a symplectic 4-manifold with weakly convex boundary. This will be discussed in the paper [72] announced in [71]. Moreover, in [72], we will produce various families of exotic 4-manifolds which are stably exotic under connected sums with arbitrary closed definite 4-manifolds.
1.3. Diffeomorphism invariants of genus function type
In a different direction, we introduce a notion of genus function type for diffeomorphism invariants of 4-manifolds by extracting a simple property of genus functions, see Definition 11.1. We then give their examples and show that any such invariant shares many properties of genus functions.
Specifically, we give examples of this type which can detect families of exotic 4-manifolds, by utilizing the Seiberg–Witten invariants of closed 4-manifolds. The lasagna $s$-invariant recently introduced by Ren and Willis [58] is also an example of this type according to their result. See Example 11.6 for these examples.
Furthermore, we show that an invariant of genus function type yields lower bounds for values of genus functions, which can be sharp depending on the invariant, see Proposition 11.8. We also prove stabilities of algebraic inequivalences similar to Theorems 1.9 and 1.10, see Theorems 11.12 and 11.13. These stabilities imply that if (algebraic equivalent classes of) an invariant of genus function type can detect a family of exotic 4-manifolds, then the invariant can detect infinitely many families of exotic 4-manifolds, see Corollary 11.14. By contrast, we also prove that there exist various families of exotic 4-manifolds which cannot be detected by any invariant of genus function type, see Theorem 11.15. We will give further related results in Section 11.
1.4. Exotically knotted submanifolds
Applying our exotic 4-manifolds, we also discuss exotically knotted submanifolds. It is natural to ask whether a given submanifold of a 4-manifold admits exotically knotted copies and/or exotically embedded copies, and many examples of submanifolds admitting such copies have been constructed, see, for example, [26], [27], [11], [14], [59], [40], [52], [45] and the references therein. Here a pair of submanifolds of a manifold will be called exotically knotted if they are diffeomorphic, topologically ambient isotopic, but not smoothly ambient isotopic. Also, a pair of submanifolds will be called exotically embedded if they are diffeomorphic, the ambient manifold admits a self-homeomorphism that maps one to the other, but the ambient manifold does not admit such a self-diffeomorphism. We furthermore require natural conditions for orientations and boundaries as well (see Definition 10.1), but our results still hold under the above simpler conditions.
Many examples of exotically knotted submanifolds are also exotically embedded, but interestingly, exotically knotted submanifolds are not necessarily exotically embedded (e.g. [59], [45]). In this paper, we focus on submanifolds which are both exotically knotted and exotically embedded. Such examples indicate strong differences between smooth and topological categories. We remark that our results in this subsection are new as exotically knotted submanifolds and also as exotically embedded submanifolds.
For a codimension-00 submanifold of a 4-manifold, the closure of the complement will be called the exterior of the submanifold. Utilizing our exotic 4-manifolds, we show that every codimension-00 submanifold satisfying a mild condition can be modified to admit infinitely many exotically knotted and exotically embedded copies.
Theorem 1.12.
Let $X$ be a codimension-00 submanifold of a $4$-manifold $Z$. These manifolds are possibly non-orientable and may have ((((possibly disconnected)))) boundaries. Suppose the exterior of $X$ is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-00 submanifolds topologically resembling $X$.
Exotically knotted submanifolds in this theorem have more properties inherited from a given submanifold $X$, see Theorem 10.2 for details. This result improves an earlier result of Akbulut and the author [11], which modified $X$ into arbitrarily many exotically embedded submanifolds additionally assuming that the exterior is a 2-handlebody with non-trivial second homology group. Note that their examples were not shown to be exotically knotted in general.
Since every 4-manifold has numerous 2-handlebodies as their submanifolds, the corollary below is straightforward.
Corollary 1.13.
Every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded ((((possibly non-orientable)))) submanifolds of codimension-00.
We also discuss exotically knotted codimension-1 submanifolds by applying Theorem 1.12. Prior results in the literature are realizations of many 3-manifolds as exotically knotted codimension-1 submanifolds of some orientable 4-manifolds (see e.g. [40] and the references therein). By contrast, we show that for a given (possibly non-orientable) 4-manifold, every codimension-$1$ submanifold satisfying a mild condition can be modified to admit infinitely many exotically knotted (and exotically embedded) copies, using homology cobordisms embedded in the given 4-manifold.
Theorem 1.14.
Let $Z$ be a ((((possibly non-orientable)))) $4$-manifold, and let $M$ be a closed $3$-manifold embedded in the interior of $Z$ as a separating submanifold. Suppose the exterior of the tubular neighborhood of $M$ has a connected component which is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$.
As corollaries, we show that there exist infinitely many closed $3$-manifolds such that every ((((possibly non-orientable)))) $4$-manifold realizes each of them as infinitely many exotically knotted submanifolds.
Corollary 1.15.
For any closed $3$-manifold embedded in $S^{4}$, every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to the $3$-manifold.
Corollary 1.16.
There exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold of them, every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to the $3$-manifold.
Note that many closed 3-manifolds do not have the above property, since there are many closed 3-manifolds not admitting any embedding into $S^{4}$ (see e.g. [41] and the references therein).
For 1-manifolds embedded in a 4-manifold, the existence of a topological ambient isotopy guarantees that of a smooth ambient isotopy, so there exist no exotically knotted 1-manifolds. It would be natural to ask whether the above corollary holds for other codimensions.
Question 1.17.
For $n=2,4$, does there exist an $n$-manifold such that every 4-manifold admits a pair of exotically knotted (or exotically embedded) submanifolds diffeomorphic to the $n$-manifold?
Note that manifolds are compact due to our convention, and thus any $n=4$ example is assumed to have non-empty boundary. This question is open even if ambient 4-manifolds are restricted to be simply connected and closed. A partial answer to the $n=4$ case is Corollary 1.13, but, to the best of the author’s knowledge, the corresponding result is not known in the $n=2$ case.
As is well-known, every closed 3-manifold admits an embeddding into a simply connected closed 4-manifold (see e.g. [2, Theorem 2.1]). Hence, for a given closed 3-manifold, it would be natural to ask what is the minimum second Betti number of a simply connected closed 4-manifold for which the 3-manifold is realized as a pair of exotically knotted (or exotically embedded) submanifolds. We show that every non-negative even integer is realized as such a minimum second Betti number by using Theorem 1.14 and a result of Kawauchi [43].
Corollary 1.18.
For each non-negative even integer $n$, there exists a closed $3$-manifold $M$ and a simply connected closed $4$-manifold $Z$ with $b_{2}=n$ such that $Z$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$ and that $M$ admits no embedding into any simply connected closed $4$-manifold with $b_{2}<n$.
We remark that the above $Z$ can be chosen to be $\frac{n}{2}(S^{2}\times S^{2})$ for each $n$. It is likely that our construction gives infinitely many distinct homology cobordism classes of 3-manifolds for each $n$, but we do not pursue this point.
Acknowledgements
Parts of this work were presented at conferences MSJ Autumn Meeting 2023, Four Dimensional Topology, and Differential Topology ’24. The author would like to thank the organizers and participants for useful discussions. The author was partially supported by JSPS KAKENHI Grant Numbers 19K03491, 23K03090 and 19H01788.
2. Genus functions
In this section, we introduce a few notions of equivalences for genus functions. We also observe that genus functions can determine smooth structures among symplectic 4-manifolds.
We here fix notations of this paper. Every homology group is integral coefficient. For a continuous map $f:X\to Y$, the induced homomorphism $H_{i}(X)\to H_{i}(Y)$ is denoted by $f_{*}$. For a manifold $Z$, the manifold equipped with the reverse orientation is denoted by $\overline{Z}$.
Throughout this section, let $X,Y$ be 4-manifolds which may have possibly disconnected boundaries. We recall the definition of the genus function.
Definition 2.1.
For a second homology class $\alpha$ of $X$, the integer $g_{X}(\alpha)$ is defined to be the minimal genus of a closed surface smoothly embedded in $X$ that represents $\alpha$. The map $g_{X}:H_{2}(X)\to\mathbb{Z}$ is called the genus function of $X$.
This function is well-defined, because every second homology class of $X$ is represented by a closed surface (see [38, Exercise 4.5.12(b)]). We will say that a continuous map $f:X\to Y$ preserves the genus functions if the induced homomorphism $f_{*}:H_{2}(X)\to H_{2}(Y)$ preserves the genus functions, that is, $f_{*}$ satisfies $g_{X}(\alpha)=g_{Y}(f_{*}(\alpha))$ for any $\alpha\in H_{2}(X)$. We here introduce a notion of equivalence for genus functions of 4-manifolds.
Definition 2.2.
The genus functions of $X$ and $Y$ will be called equivalent if there exists a homeomorphism $f:X\to Y$ that preserves the genus functions. If $f$ furthermore preserves the orientations of $X$ and $Y$, then the genus functions will be called orientedly equivalent.
It is easy to see that if two 4-manifolds are diffeomorphic (resp. orientedly diffeomorphic), then their genus functions are equivalent (resp. orientedly equivalent). Hence, the genus function is an invariant of smooth structures.
We also introduce a notion of algebraic equivalence for genus functions of 4-manifolds. We say that a homomorphism $\varphi:H_{2}(X)\to H_{2}(Y)$ preserves the intersection forms of $X$ and $Y$ if $\varphi(\alpha)\cdot\varphi(\beta)=\alpha\cdot\beta$ for any $\alpha,\beta\in H_{2}(X)$, where $\alpha\cdot\beta$ denotes the intersection number of $\alpha,\beta$.
Definition 2.3.
(1) The genus functions of $X$ and $Y$ will be called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the genus functions and the intersection forms.
(2) The genus functions of $X$ and $Y$ will be called algebraically equivalent for some orientations if $g_{X}$ is algebraically equivalent to $g_{Y}$ or to $g_{\overline{Y}}$.
(3) For subgroups $F_{X}\subset H_{2}(X)$ and $F_{Y}\subset H_{2}(Y)$, the restrictions of the genus functions of $X$ and $Y$ to $F_{X}$ and $F_{Y}$ will be called algebraically equivalent if there exists an isomorphism $F_{X}\to F_{Y}$ that preserves the genus functions and the intersection forms.
It is clear that if the genus functions of two 4-manifolds are algebraically inequivalent (resp. algebraically inequivalent for any orientations), then they are orientedly inequivalent (resp. (unorientedly) inequivalent) and hence orientedly non-diffeomorphic. We remark that the definition of equivalence in the paper [61] of Stipsicz and Szabó is the same as our algebraic equivalence, not our (ordinary) equivalence.
As we will see, it is useful to consider the genus functions modulo torsion part.
Definition 2.4.
For a class $\widehat{\alpha}$ of $H_{2}(X)/\mathrm{Tor}$, we define an integer $g^{*}_{X}(\widehat{\alpha})$ by
$g^{*}_{X}(\widehat{\alpha})=\min\{g_{X}(\alpha)\mid\textnormal{$\alpha\in H_{2}(X)$ represents $\widehat{\alpha}$.}\}.$
The map $g^{*}_{X}:H_{2}(X)/\mathrm{Tor}\to\mathbb{Z}$ will be called the torsion-free genus function of $X$.
Clearly, $g^{*}_{X}$ is completely determined by $g_{X}$. We note that $g^{*}_{X}$ is often easier to deal with, since adjunction inequalities give lower bounds for values of $g^{*}_{X}$ rather than $g_{X}$.
We define their equivalences and algebraic equivalences in the same way as before.
Definition 2.5.
(1) The torsion-free genus functions of $X$ and $Y$ will be called equivalent if there exists a homeomorphism $f:X\to Y$ such that the isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ induced by $f$ preserves $g^{*}_{X}$ and $g^{*}_{Y}$. If $f$ furthermore preserves orientations of $X$ and $Y$, then the torsion-free genus functions will be called orientedly equivalent.
(2) The torsion-free genus functions of $X$ and $Y$ will be called algebraically equivalent if there exists an isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ that preserves $g^{*}_{X},g^{*}_{Y}$ and the intersection forms.
(3) The torsion-free genus functions of $X$ and $Y$ will be called algebraically equivalent for some orientations if $g^{*}_{X}$ is algebraically equivalent to $g^{*}_{Y}$ or to $g^{*}_{\overline{Y}}$.
(4) For subgroups $F_{X}\subset H_{2}(X)/\mathrm{Tor}$ and $F_{Y}\subset H_{2}(Y)/\mathrm{Tor}$, the restrictions of the torsion-free genus functions of $X$ and $Y$ to $F_{X}$ and $F_{Y}$ will be called algebraically equivalent if there exists an isomorphism $F_{X}\to F_{Y}$ that preserves the torsion-free genus functions and the intersection forms.
Let us recall that a rational surface is orientedly diffeomorphic to $\mathbb{C}P^{2}\#n\overline{\mathbb{C}P^{2}}$, and that a ruled surface is diffeomorphic to the total space of an $S^{2}$-bundle over a closed surface or to its blow-ups. We observe that genus functions uniquely determine smooth structures of rational surfaces and ruled surfaces among symplectic 4-manifolds. This fact easily follows from [50, Proposition 4.3].
Proposition 2.6 (cf.  [50]).
If the genus function of a given closed symplectic $4$-manifold is equivalent to that of a rational surface or a ruled surface, then the given $4$-manifold is diffeomorphic to the surface.
Proof.
Let $N_{g}$ denote the total space of a non-trivial $S^{2}$-bundle over the closed surface $\Sigma_{g}$ of genus $g\geq 0$, and let $R$ be a rational surface or a ruled surface. Then $R$ is orientedly diffeomorphic to one of the $4$-manifolds $\mathbb{C}P^{2}\#n\overline{\mathbb{C}P^{2}}$, $(S^{2}\times\Sigma_{g})\#n\overline{\mathbb{C}P^{2}}$ and $N_{g}\#n\overline{\mathbb{C}P^{2}}$ for some non-negative integers $n,g$. We easily see that $R$ contains a closed surface representing a non-torsion second homology class whose self-intersection number is larger than $22g^{\prime}-$ 2, where $g^{\prime}$ denotes the genus of the surface. In the case where $b_{2}^{+}(\overline{R})\geq 1$, the orientation reversal $\overline{R}$ also contains such a surface as seen from its diffeomorphism type.
Now suppose that the genus function of a closed symplectic 4-manifold $X$ is equivalent to $R$. Then we have $b_{2}^{+}(X)\geq 1$, and it thus follows from the above paragraph that $X$ contains a closed surface of genus $g^{\prime}$ representing a non-torsion second homology class whose self-intersection number is larger than $22g^{\prime}-$ 2. Hence, by [50, Proposition 4.3], $X$ is orientedly diffeomorphic to a rational surface or to a ruled surface. We easily see that two 4-manifolds in these classes of 4-manifolds are diffeomorphic if and only if they are homeomorphic, by comparing their homology groups and intersection forms. Since $R$ is also diffeomorphic to one of these 4-manifolds and is homeomorphic to $X$, it follows that $X$ is diffeomorphic to $R$. ∎
The above proposition holds for torsion-free genus functions as well, because the second homology groups of rational surfaces and ruled surfaces are torsion-free.
3. Corks, homology cobordisms and exotic 4-manifolds
In this section, we review (generalized) corks, homology cobordisms and constructions of exotic 4-manifolds that utilize corks and invertible homology cobordisms. We also give a minor improvement of the method of Akbulut and Ruberman [8] that modifies a generalized cork into exotic 4-manifolds.
3.1. Generalized corks
In this paper, we define generalized corks as follows (cf. [9], [63]).
Definition 3.1.
(1) Let $C$ be a (not necessarily contractible) 4-manifold and let $f:\partial C\to\partial C$ be a self-diffeomorphism of the boundary $\partial C$ that preserves the orientation. The pair $(C,f)$ will be called a generalized cork if $f$ extends to a self-homeomorphism of $C$, but does not extend to any self-diffeomorphism of $C$. The order of a generalized cork $(C,f)$ is defined to be the minimum positive integer $k$ for which the power $f^{k}$ extends to a self-diffeomorphism of $C$. If $f^{k}$ does not extend to any self-diffeomorphism of $C$ for any positive integer $k$, then $(C,f)$ will be called an infinite order generalized cork. A generalized cork $(C,f)$ is called a cork if $C$ is contractible.
(2) Let $(C,f)$ be a generalized cork, and assume $C$ is embedded in a 4-manifold $X$. We remove $C$ from $X$ and glue it back along the resulting boundary via the gluing map $f$. This operation will be called a generalized cork twist along $(C,f)$.
(3) Two generalized corks $(C,f)$ and $(C,g)$ will be called equivalent if the self-diffeomorphism $g^{-1}\circ f$ of $\partial C$ extends to a self-diffeomorphism of $C$.
Clearly, a generalized cork twist always preserves the homeomorphism type of an ambient 4-manifold. We note that in the case where $C$ is contractible, every self-diffeomorphism of $\partial C$ extends to a self-homeomorphism of $C$ ([32], [16]).
3.2. Homology cobordisms
Next, we recall the definitions of a cobordism and related terminologies and fix conventions. Throughout this paper, we denote the interval $[0,1]$ by $I$. Let $M,N$ be closed (possibly disconnected) 3-manifolds, and let $X$ be a 4-manifold having $M$ as an oriented boundary component. For convenience, we allow $N$ to be empty.
Definition 3.2.
(1) A (possibly disconnected) 4-manifold $P$ is called a cobordism from $M$ to $N$ if $\partial P$ is orientedly diffeomorphic to the disjoint union $\overline{M}\sqcup N$, and every connected component of $P$ has a non-empty boundary component that belongs to $\overline{M}$. In this case, $M$ and $N$ will be respectively called a negative boundary and a positive boundary of $P$. We often identify $\partial P$ with the disjoint union $\overline{M}\sqcup N$ by fixing a diffeomorphism. For an orientation preserving self-diffeomorphism $f$ of $M$, let $X\cup_{f}P$ denote the 4-manifold constructed from $X$ by attaching $P$ along $M$, where we glue each $x\in M\subset\partial X$ with $f(x)\in M\subset\overline{\partial P}$. When we do not specify the gluing map $f$, we often denote a glued 4-manifold by $X\cup_{M}P$. Unless otherwise stated, we always attach a cobordism to a 4-manifold along a negative boundary.
(2) For a cobordism $P$ from $M$ to $N$, the cobordism $\overline{P}$ from ${N}$ to ${M}$ is called the upside down cobordism of $P$.
(3) For a cobordism $P$ from $M$ to $N$, if the homomorphisms $H_{*}({M})\to H_{*}(P)$ and $H_{*}(N)\to H_{*}(P)$ induced by the inclusions are isomorphisms, then $P$ is called a homology cobordism. If there exists a homology cobordism from $M$ to $N$, then $M$ and $N$ are called homology cobordant.
(4) A cobordism $P$ from $M$ to $N$ is called invertible if there exists a cobordism $Q$ from $N$ to $M$ such that a glued 4-manifold $P\cup_{N}Q$ is orientedly diffeomorphic to the product $I\times M$. The cobordism $Q$ will be called an inverse cobordism of $P$.
3.3. Modifying 2-handlebodies into exotic Stein manifolds
In this subsection, we review the method of Akbulut and the author [11] that modifies a given 4-dimensional 2-handlebody with non-trivial second homology group into arbitrarily many pairwise exotic 4-manifolds admitting Stein Structures.
We say that a 4-dimensional $2$-handlebody is a Legendrian handlebody if the attaching circles of 2-handles form a Legendrian link in the standard tight contact structure on the boundary $k(S^{1}\times S^{2})$ of the sub 1-handlebody $\natural_{k}S^{1}\times D^{3}$. We note that every 4-dimensional 2-handlebody becomes a Legendrian handlebody by isotoping the attaching circles of 2-handles. We say that a 4-dimensional 2-handlebody is a Stein handlebody if it is Legendrian handlebody such that the framing of the attaching circle of each 2-handle is one less than the Thurston–Bennequin number. By results of Eliashberg [21] and Gompf [34], a 4-manifold admits a Stein structure if and only if it is orientedly diffeomorphic to a Stein handlebody.
Let $X$ be a 4-dimensional 2-handlebody, and let $K$ be the attaching circle of a 2-handle of $X$. We consider a handlebody diagram of $X$. Akbulut and the author [11] introduced two local diagrammatic operations shown in Figure 1. For a positive integer $p$, the left side operation is called a $W^{-}(p)$-modification to $K$, and the right side operation is called a $W^{+}(p)$-modification to $K$. We note that these operations do not change the framing coefficient $m$ of $K$. When we do not specify the coefficient $p$, these operations are called a $W^{-}$-modification and a $W^{+}$-modification, respectively.
Figure 1. $W^{-}(p)$- and $W^{+}(p)$-modifications
Each operation creates an embedded copy of the Akbulut cork $W$ ([3]) in the resulting handlebody, and these two operations are related to each other by the cork twist along the embedded $W$ as shown in the figure. So, the 4-manifold obtained by the $W^{+}(p)$-modification to $K$ is orientedly homeomorphic to the 4-manifold obtained by the $W^{-}(p)$-modification to $K$.
Clearly, a $W^{-}$-modification attaches an algebraically canceling pair of 1- and 2-handles to $X$. Thus, this operation is an attachment of an invertible homology cobordism to $X$, and its inverse cobordism is its upside down cobordism. Furthermore, the homotopy extension property (see [39, Proposition 0.1.8]) implies that this invertible cobordism has a deformation retraction onto $\partial X$. Consequently, the result of any $W^{-}$-modification and hence also any $W^{+}$-modification topologically resembles $X$. Furthermore, the result of any $W^{-}$-modification admits a deformation retraction onto $X$. We note that the 4-manifold obtained by a $W^{+}$-modification becomes diffeomorphic to the original 4-manifold by attaching the upside down cobordism of the invertible cobordism given by the corresponding $W^{-}$-modification ([11, Proposition 4.5]).
By drawing Legendrian pictures, we obtain Legendrian versions of $W^{-}$- and $W^{+}$-modifications for Legendrian handlebodies, and these have the following two effects ([11, Proposition 4.7]). (1) Each of these two operations modifies a Stein handlebody into another Stein handlebody. (2) If we apply a $W^{+}$-modification to a Legendrian knot $K$ of a Legendrian handlebody, then the Thurston–Bennequin number of the result of $K$ increases by $p$, and the Thurston–Bennequin number for the attaching circle of the 2-handle of $W$ is $+2$.
Thus, by applying $W^{+}$-modifications to a 2-handlebody $X$ and adding zig-zags to each Legendrian knot (i.e., stabilizations), we obtain a Stein handlebody, implying the following.
Proposition 3.3 ([11]).
For any $4$-dimensional $2$-handlebody $X$, there exists a $4$-dimensional $2$-handlebody $X^{\prime}$ having the following properties.
  • $X^{\prime}$ admits a Stein structure.
  • $X^{\prime}$ topologically resembles $X$ and is homeomorphic to a $4$-manifold that admits a deformation retraction onto $X$.
  • $X^{\prime}$ becomes diffeomorphic to $X$ after attaching the upside down cobordism of an invertible homology cobordism along $\partial X^{\prime}$.
Applying $W^{-}$-modifications several times and replacing some of them with $W^{+}$-operations by cork twists, we obtain arbitrarily many pairwise homeomorphic 2-hanldebodies. In the case where $H_{2}(X)$ is a non-trivial group, by carefully choosing coefficients of $W^{+}$- and $W^{-}$-modifications and applying adjunction inequalities, Akbulut and the author showed that the resulting 2-handlebodies have pairwise algebraically inequivalent genus functions for any orientations ([11, Lemma 5.13 and Proposition 6.2]), giving the following result.
Theorem 3.4 ([11]).
For any $4$-dimensional $2$-handlebody $X$ with a non-trivial second homology group and any positive integer $n$, there exist pairwise exotic $4$-manifolds $X_{0},X_{1},\dots,X_{n}$ having the following properties.
  • Each $X_{i}$ topologically resembles $X$ and admits an embedding into $X$.
  • $X_{0}$ admits a deformation retraction onto $X$ and is obtained from $X$ by attaching an invertible homology cobordism.
  • $X_{i}$ admits a Stein structure for any $i\neq 0$.
  • The genus functions of $X_{0},X_{1},\dots,X_{n}$ are pairwise algebraically inequivalent for any orientations.
3.4. Modifying generalized corks into exotic 4-manifolds
In this subsection, we review the method of Akbulut and Ruberman [8] that modifies a given generalized cork into exotic 4-manifolds. We also give minor improvements of their results.
Throughout this subsection, we fix an arbitrary closed 3-manifold $M$. Note that $M$ is connected due to our convention. Akbulut and Ruberman [8] gave a particularly useful invertible homology cobordism.
Theorem 3.5 ([8]).
There exist a closed $3$-manifod $N$ and an invertible homology cobordism $\mathbb{P}$ from $M$ to $N$ having the following properties.
  • Every self-diffeomorphism of $N$ extends to a self-diffeomorphism of $\mathbb{P}$ whose restriction to $M$ is isotopic to the identity map.
  • The homomorphism $\pi_{1}(M)\to\pi_{1}(\mathbb{P})$ induced by the inclusion is an isomorphism.
Utilizing this cobordism, Akbulut and Ruberman [8] constructed an exotic “pair” of 4-manifolds from a given generalized cork by attaching the above cobordism and then twisting the embedded generalized cork. In order to produce infinitely many pairwise exotic 4-manifolds, they constructed another invertible homology cobordism having properties different from the above cobordism. They constructed infinitely many 4-manifolds from a given infinite order cork by attaching the latter cobordism and then twisting the embedded cork, and showed that at least infinitely many of them are pairwise exotic using the pigeonhole principle. However, for a non-simply connected generalized cork of infinite order, it is unclear whether exotic 4-manifolds given by the latter construction have the same fundamental groups as that of the generalized cork, since the latter cobordism has a weaker property on the fundamental groups.
Here we observe that, in the former construction, all 4-manifolds constructed by twisting a given generalized cork are pairwise exotic, by modifying their proof. We prove this claim in a more general form.
Theorem 3.6.
Let $f_{1},f_{2}$ be self-diffeomorphisms of $M$ that preserve the orientations, and let $X_{1},X_{2}$ be $4$-manifolds with $\partial X_{1}=\partial X_{2}=M$. If $f_{2}^{-1}\circ f_{1}$ does not extend to any diffeomorphism $X_{1}\to X_{2}$, then the following hold.
  • The $4$-manifolds $X_{1}\cup_{f_{1}}\mathbb{P}$ and $X_{2}\cup_{f_{2}}\mathbb{P}$ are not diffeomorphic to each other.
  • Each $X_{i}\cup_{f_{i}}\mathbb{P}$ admits a deformation retraction onto $X_{i}$ and topologically resembles $X_{i}$.
Remark 3.7.
Akbulut and Ruberman constructed a cobordism $\mathbb{P}$ in Theorem 3.5 using an arbitrary doubly slice knot in $S^{3}$ having certain properties, and they showed that the specific knot $J$ in $S^{3}$ given by [8, Figure 2] can be used for the construction. In the above theorem, we use the cobordism $\mathbb{P}$ corresponding to the knot $J$ in order to guarantee the homotopy equivalence. The rest of the properties hold for any $\mathbb{P}$ in Theorem 3.5 as seen from our proof.
Proof.
Let $\mathbb{Q}$ be an inverse cobordism of $\mathbb{P}$. Then the 4-manifold $\mathbb{P}\cup_{q}\mathbb{Q}$ is diffeomorphic to the product $I\times M$ for some self-diffeomorphism $q$ of $N$. For convenience, we set $\partial\mathbb{P}=\overline{M_{\mathbb{P}}}\sqcup N_{\mathbb{P}}$ and $\partial\mathbb{Q}=\overline{N_{\mathbb{Q}}}\sqcup M_{\mathbb{Q}}$, where $M_{\mathbb{P}},M_{\mathbb{Q}}$ are identified with $M$, and $N_{\mathbb{P}},N_{\mathbb{Q}}$ are identified with $N$. We may assume that the restriction $M_{\mathbb{P}}\to M(=0\times M)$ of the diffeomorphism $\mathbb{P}\cup_{q}\mathbb{Q}\to I\times M$ is the identity map if necessary by composing a self-diffeomorphism of $I\times M$. Let $\varphi:M_{\mathbb{Q}}\to M(=1\times M)$ denote the restriction of this diffeomorphism $\mathbb{P}\cup_{q}\mathbb{Q}\to I\times M$.
We set $X^{\prime}_{i}=X\cup_{f_{i}}\mathbb{P}$ for each $i$. Then each $X^{\prime}_{i}$ deformation retracts to $X$ by Lemma 3.8 which will be given later. Since $\mathbb{P}$ is a homology cobordism, we see that $X^{\prime}_{i}$ topologically resembles $X_{i}$.
We will show that $X^{\prime}_{1}$ is not diffeomorphic to $X^{\prime}_{2}$. Suppose to the contrary that there exists a diffeomorphism $X^{\prime}_{1}\to X^{\prime}_{2}$. Since every self-diffeomorphism of $N_{\mathbb{P}}$ extends to a self-diffeomorphism of $\mathbb{P}$ whose restriction to $M_{\mathbb{P}}$ is isotopic to the identity, we may assume that the diffeomorphism $X^{\prime}_{1}\to X^{\prime}_{2}$ is the identity on $N_{\mathbb{P}}$ if necessary by composing a self-diffeomorphism of $X^{\prime}_{2}$. We thus have a diffeomorphism $\Phi:X^{\prime}_{1}\cup_{q}\mathbb{Q}\to X_{2}^{\prime}\cup_{q}\mathbb{Q}$ whose restriction to $M_{\mathbb{Q}}$ is the identity map.
For each $i$, we also use a diffeomorphism $X^{\prime}_{i}\cup_{q}\mathbb{Q}\to X_{i}\cup_{\mathrm{id}_{M}}(I\times M)$ whose restriction to the boundary is $f_{i}^{-1}\circ\varphi:M_{\mathbb{Q}}\to M$. This diffeomorphism is obtained as follows. We have a diffeomorphism $X^{\prime}_{i}\cup_{q}\mathbb{Q}\to X_{i}\cup_{f_{i}}(I\times M)$ whose restriction to the boundary is the diffeomorphism $\varphi$. Since $f_{i}^{-1}$ extends to a self-diffeomorphism of $I\times M$, we have a diffeomorphism $X_{i}\cup_{f_{i}}(I\times M)\to X_{i}\cup_{\mathrm{id}_{M}}(I\times M)$ whose restriction to the boundary $1\times M$ is $f_{i}^{-1}$. By composing these two diffeomorphisms, we obtain a desired diffeomorphism $X^{\prime}_{i}\cup_{q}\mathbb{Q}\to X_{i}\cup_{\mathrm{id}_{M}}(I\times M)$.
Composing the last diffeomorphisms and $\Phi$, we obtain a diffeomorphism $X_{1}\cup_{\mathrm{id}_{M}}(I\times M)\to X_{2}\cup_{\mathrm{id}_{M}}(I\times M)$ whose restriction to the boundary is $f_{2}^{-1}\circ f_{1}$. Since each $X_{i}\cup_{\mathrm{id}_{M}}(I\times M)$ is diffeomorphic to $X_{i}$ fixing the boundary $M$ pointwise, it follows that $f_{2}^{-1}\circ f_{1}$ extends to a diffeomorphism $X_{1}\to X_{2}$, giving a contradiction. Therefore, $X^{\prime}_{1}$ is not diffeomorphic to $X^{\prime}_{2}$. ∎
The homotopy equivalence of $X_{i}\cup_{f_{i}}\mathbb{P}$ and $X_{i}$ is claimed in [8, Theorem A], but the proof is stated only in the case where $X_{i}$ is contractible. For completeness, we give a proof of the general case for $\mathbb{P}$ used in Theorem 3.6 (see Remark 3.7) by showing the lemma below.
Lemma 3.8 (cf. [8]).
The cobordism $\mathbb{P}$ used in Theorem 3.6 admits a deformation retraction onto the boundary component $M$.
Proof.
In order to construct $\mathbb{P}$, Akbulut and Ruberman [8] used a certain link $L$ in $M$ and a certain concordance $C(\subset I\times S^{3})$ from the unknot $U$ to the knot $J$ in [8, Figure 2]. Let $\nu(L)$ and $\nu(C)$ respectively denote the tubular neighborhoods of $L$ and $C$, and let $n$ denote the number of the components of $L$.
They constructed $\mathbb{P}$ from $I\times(M-\textnormal{int}\,\nu(L))$ by attaching $n$ copies of $(I\times S^{3})-\textnormal{int}\,\nu(C)$ along $I\times\partial\nu(L)$ so that the resulting negative boundary remains $M$. Note that there is a diffeomorphism from $[-1,0]\times M\cup_{0\times(M-\textnormal{int}\,\nu(L))}I\times(M-\textnormal{int}\,\nu(L))$ to $I\times M$ that maps $(0\times\nu(L))\cup(I\times\partial\nu(L))$ to $1\times\nu(L)$. Thus, $\mathbb{P}$ is obtained from $I\times M$ by attaching $n$ copies of the closure of $(I\times S^{3})-\nu(C)$ using a diffeomorphism from each component of $1\times\nu(L)$ to $\nu(m)$, where $m$ denotes the meridian of the unknot $U$ in $0\times S^{3}$, and $\nu(m)$ denotes a tubular neighborhood of $m$ in $\partial(I\times S^{3}-\textnormal{int}\,\nu(C))$.
Akbulut and Ruberman gave a handlebody diagram of the exterior of a slice disk for $J$ in $D^{4}$ ([8, Figure 4.(a)]), and it is easy to see that this 4-manifold is diffeomorphic to the closure of $(I\times S^{3})-\nu(C)$. Since the attaching circle of the unique 2-handle of this handlebody is homotopic to a meridian of the dotted circle, the closure of $(I\times S^{3})-\nu(C)$ deformation retracts to the tubular neighborhood $\nu(m)$. Therefore, $\mathbb{P}$ admits a deformation retraction onto $0\times M$. ∎
Remark 3.9.
(1) The inverse cobordism of $\mathbb{P}$ used in Theorem 3.6 (see Remark 3.7) is its upside down cobordism $\overline{\mathbb{P}}$, though we do not use this fact in this section. This claim is not stated in [8], but it is implicit and follows from their claim that the knot $J$ in [8, Figure 2] is superslice, that is, the double of a slice disk for $J$ is an unknotted 2-sphere in $S^{4}$ (see the proof of [8, Proposition 2.6]). The supersliceness can be alternatively proved using [51, Proposition 4.3], and it is clear that the slice disk obtained in [8] gives a concodance in $I\times S^{3}$ from the unknot to $J$ such that the double is isotopic to the product concordance fixing $\{0,1\}\times S^{3}$ pointwise. Therefore, it follows from their construction of $\mathbb{P}$ that the double $\mathbb{P}\cup_{\mathrm{id}_{M}}\overline{\mathbb{P}}$ admits a diffeomorphism to the product $I\times M$ whose restriction to the boundary $\overline{M}\sqcup M$ is the identity map.
(2) The cobordism $\mathbb{P}$ is obtained from $I\times M$ by attaching algebraically canceling pairs of 1- and 2-handles, and Lemma 3.8 follows from this fact as well. This fact is implicit in [8, Example 4.1] and can be proved using Akbulut’s cylinder method [4, Section 3.2] as follows. The proof of Lemma 3.8 implies that the cobordism $\mathbb{P}$ is obtained from the disjoint union of $I\times M$ and copies of the exterior of the slice disk for $J$ by attaching copies of the cylinder $I\times S^{1}\times D^{2}$ along $\partial I\times S^{1}\times D^{2}$. Since this cylinder is a union of a 1-handle and a 2-handle, the cobordism $\mathbb{P}$ is obtained from the boundary sum of $I\times M$ with the copies of the exterior by attaching 2-handles. Due to the handlebody diagram of the exterior ([8, Figure 4.(a)]), we can prove the above fact by showing that the attaching circle of the 2-handle in each cylinder cancels a 1-handle of the exterior. We remark that this argument gives a method for drawing handlebody diagrams of $\mathbb{P}$ and each glued 4-manifold $X_{i}\cup_{f_{i}}\mathbb{P}$, once the link $L$ in $M$ is specified.
The corollary below is straightforward from Theorem 3.6.
Corollary 3.10.
Let $\{(C,f)\}_{f\in A}$ be a family of pairwise inequivalent generalized corks with $\partial C=M$. Then there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M}\}}$ of pairwise exotic $4$-manifolds topologically resembling $C$ such that each $C_{f}$ admits a deformation retraction onto $C$.
We note that for any generalized cork $(C,f)$ of infinite order, $\{(C,f^{k})\}_{k\in\mathbb{Z}-\{0\}}$ satisfies the assumption of this corollary. Also, the assumption holds for families given by finite order generalized corks.
3.5. Infinite order corks
In this subsection, we briefly review Gompf’s infinite order corks ([36]). For integers $r,s,m>0$, let $C(r,s;m)$ be the 4-manifold in Figure 2. (In this notation, we changed the sign of $m$ from the original one in [36] so that we may assume $m>0$.) It is easy to see that $C(r,s;m)$ is a contractible 4-manifold of Mazur type, that is, it admits a handle decomposition consisting of a 0-handle and an algebraically canceling pair of 1- and 2-handles.
Figure 2. $C(r,s;m)$
Gompf showed that the pair $(C(r,s;m),\tau)$ is an infinite order cork, where the self-diffeomorphism $\tau$ on the boundary is the torus twist on a certain embedded torus. Roughly speaking, a torus twist is (Dehn twist on an $(\textnormal{Dehn twist on an annulus})\times\mathrm{id}_{S^{1}}$, and hence $\tau$ is the identity map outside the tubular neighborhood of the torus. For the precise definition of a torus twist, see [37, Definition 2.1]. We recall an important property of Gompf’s corks. Let $E$ be the 4-manifold given in Figure 3. We denote the obvious torus $T^{2}\times 0\subset T^{2}\times D^{2}$ in the figure by $T$. Also, for an integer $k$, we denote the twist knot in Figure 4 by $T_{k}$. Let $E_{k}$ be the 4-manifold obtained from $E$ by applying the Fintushel–Stern knot surgery ([31]) along the torus $T$ using the knot $T_{k}$.
Figure 3. $E$
Gompf [36] proved the following theorem to show that each $(C(r,s;m),\tau)$ is an infinite order cork. For details, see [37, the proof of Theorem 3.5].
Theorem 3.11 ([36], [37]).
Let $r,s,m$ be positive integers, and set $N=r+s+m-3$. Then there exists an embedding of $C(r,s;m)$ into $E\#N\overline{\mathbb{C}P^{2}}$ such that, for each integer $k$, the knot surgered $4$-manifold $E_{k}\#N\overline{\mathbb{C}P^{2}}$ is diffeomorphic to the $4$-manifold obtained from $E\#N\overline{\mathbb{C}P^{2}}$ by removing the embedded $C(r,s;m)$ and gluing it back via $\tau^{k}$. Furthermore, this diffeomorphism is isotopic to the identity map on the boundary $\partial E$ of these two $4$-manifolds.
We note that, in the last claim of this theorem, we identified the boundaries of these two 4-manifolds with $\partial E$, since the knot surgery and the cork twist are applied in the interior of $E\#N\overline{\mathbb{C}P^{2}}$.
Figure 4. The twist knot $T_{k}$
4. Producing infinite order generalized corks
In this section, we show that infinite order generalized corks exist for a large class of homeomorphism types, which is of independent interest.
The first example of an infinite order generalized cork was given by Tange [63, Theorem 1.6], and his example is simply connected and satisfies $b_{2}=2$. Gompf [36] later gave the first examples of infinite order contractible corks. We show that every 2-handlebody can be modified into an infinite order generalized cork, showing that infinite order generalized corks exist in abundance.
Theorem 4.1.
For any $4$-dimensional $2$-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ topologically resembles $X$. Furthermore, in the case where $H_{2}(X)$ is a non-trivial group, there exist arbitrarily many generalized corks $(X_{i},f_{i})$’s of infinite order such that $X_{i}$’s are pairwise exotic and topologically resemble $X$.
We will prove this theorem together with further properties in Section 9, see Theorem 9.1.
Let us recall that a symplectic 4-manifold $(Z,\omega)$ is said to have weakly convex boundary if $\partial Z$ admits a positive contact structure $\xi$ satisfying $\omega|_{\xi}>0$, that is, $\omega(v,w)>0$ for any oriented basis $v,w$ for $\xi$ (see [24]). Here we prove the following proposition to prove the above theorem and main results of this paper.
Proposition 4.2.
Suppose that a $4$-manifold $X$ admits an embedding into a symplectic $4$-manifold with weakly convex boundary. Then, there exists an infinite order generalized cork $(X^{\prime},f)$ such that $X^{\prime}$ is obtained from $X$ by attaching an invertible homology cobordism and admits a deformation retraction onto $X$. Furthermore, $f$ extends to a self-homeomorphism of $X^{\prime}$ that induces the identity map on $H_{2}(X^{\prime})$.
Let $E$ be the 4-manifold given by Figure 3. We show the lemma below.
Lemma 4.3.
$E$ admits a Stein structure.
Proof.
By canceling the 1-handles and isotoping the resulting 2-handles, we obtain the Stein handlebody diagram in Figure 5, where the coefficients are framings relative to the Seifert framings. Therefore, $E$ admits a Stein structure. ∎
Figure 5. Stein handlebody diagram of $E$
Proof of Proposition 4.2.
We fix positive integers $r,s,m$ and abbreviate the contractible 4-manifold $C(r,s;m)$ in Subsection 3.5 by $C$, and we set $N=r+s+m-3$. Since $C$ is of Mazur type, its double $C\cup\overline{C}$ is diffeomorphic to $S^{4}$.
Let $H$ be a homology cobordism obtained from $I\times\partial X$ by taking a boundary sum with $C$, where the sum is taken outside the tubular neighborhood of the torus on which Gompf’s torus twist $\tau$ is applied. We easily see that the double $H\cup\overline{H}$ is diffeomorphic to $(I\times\partial X)\#(C\cup\overline{C})$ and hence to $I\times\partial X$. Thus, $H$ is an invertible homology cobordism from $\partial X$ to $\partial X\#\partial C$.
Let $X^{\prime}$ be the 4-manifold obtained from $X$ by attaching the invertible homology cobordism $H$. We easily see that $X^{\prime}$ deformation retracts to $X$ and is diffeomorphic to $X\natural C$. Let $f$ be the self-diffeomorphism of $\partial X^{\prime}$ that extends the torus twist $\tau$ of $\partial C$ as the identity on the rest of $\partial X^{\prime}$. Since $\tau$ extends to a self-homeomorphism of $C$, the self-diffeomorphism $f$ extends to a self-homeomorphism of $X^{\prime}$ that is identity on $X$. We thus see that this self-homeomorphism induces the identity map on $H_{2}(X^{\prime})$.
We will show that $(X^{\prime},f)$ is an infinite order generalized cork. By the assumption, $X$ admits an embedding into a symplectic 4-manifold $Y$ with weakly convex boundary. Let $E^{\prime}$ be the 4-manifold obtained from $E$ by attaching a 2-handle along a meridian of the 0-framed knot with sufficiently large negative framing. Then $Y\natural E^{\prime}$ is obtained from $Y$ by attaching 2-handles along a Legendrian link with the contact $(-1)$-framings, and thus $Y\natural E^{\prime}$ and its blow-ups $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ are symplectic 4-manifolds with weakly convex boundary ([25, Theorem 2.5], [66], see also [24, Theorem 5.8]). We here construct an embedding of $X^{\prime}$ into $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ by taking a boundary sum of $X\subset Y$ and $C\subset E^{\prime}$ using a thickened path from $\partial X$ to $\partial C$.
By a result of Eliashberg [22] and Etnyre [23], we can embed $Y\natural E^{\prime}$ into a closed symplectic 4-manifold with $b_{2}^{+}>1$. We note that the torus $T$ in $E$ represents a second homology class of $Z$ having infinite order, since $T$ algebraically intersects the 2-sphere coming from the above 2-handle of $E^{\prime}$. Let $Z_{k}$ denote the 4-manifold obtained from $Z$ by applying the knot surgery along $T$ using the twist knot $T_{k}$. It follows from the blow-up formula ([30]) and the knot surgery formula ([31], [28], see also [62]) that $Z_{k}\#N\overline{\mathbb{C}P^{2}}$ yields pairwise non-diffeomorphic 4-manifolds by varying $k$. On the other hand, Theorem 3.11 and the definition of $f$ show that each $Z_{k}\#N\overline{\mathbb{C}P^{2}}$ is diffeomorphic to the 4-manifold obtained from $Z\#N\overline{\mathbb{C}P^{2}}$ by twisting $X^{\prime}$ using the diffeomorphism $f^{k}$. Therefore, $f^{k}$ $(k\neq 0)$ does not extend to any self-diffeomorphism of $X^{\prime}$, and hence $(X^{\prime},f)$ is an infinite order generalized cork. This completes the proof. ∎
Remark 4.4.
For Proposition 4.2, we can relax the assumption that $X$ admits an embedding into symplectic 4-manifold with weakly convex boundary. Indeed, as seen from the proof, if $X\natural E$ admits an embedding into a closed 4-manifold $Z$ with $b_{2}^{+}>1$ and non-vanishing Seiberg–Witten invariant such that the torus $T$ in $E$ represents an infinite order second homology class of $Z$, then the conclusion still holds.
5. Isomorphisms between symmetric bilinear forms
In this section, under certain conditions, we construct isomorphisms between symmetric bilinear forms that preserve certain maps modeled on genus functions and related invariants.
Throughout this section, we use the following setting. For each $i=1,2$, let $A_{i},B_{i}$ be finitely generated $\mathbb{Z}$-modules, and set $A^{\prime}_{i}=A_{i}\oplus B_{i}$. Let $p_{A_{i}}:A^{\prime}_{i}\to A_{i}$ and $p_{B_{i}}:A^{\prime}_{i}\to B_{i}$ denote the projections. Take an integral symmetric bilinear form $Q_{A_{i}}:A_{i}\times A_{i}\to\mathbb{Z}$, and define symmetric bilinear forms $Q_{B_{i}}:B_{i}\times B_{i}\to\mathbb{Z}$ and $Q_{i}:A^{\prime}_{i}\times A^{\prime}_{i}\to\mathbb{Z}$ by $Q_{B_{i}}(b,b^{\prime})=0$ and $Q_{i}(a+b,a^{\prime}+b^{\prime})=Q_{A_{i}}(a,a^{\prime})$, where $a,a^{\prime}\in A_{i}$ and $b,b^{\prime}\in B_{i}$. We assume that the symmetric bilinear form $Q_{\widetilde{A}_{i}}:(A_{i}/\mathrm{Tor})\times(A_{i}/\mathrm{Tor})\to\mathbb{Z}$ induced by each $Q_{A_{i}}$ is non-degenerate. These maps are models of intersection forms of 4-manifolds.
We show the following lemma.
Lemma 5.1.
Suppose the modules $A_{1}$ and $A_{2}$ or the modules $B_{1}$ and $B_{2}$ are torsion-free. If there exists an isomorphism $\Phi:A^{\prime}_{1}\to A^{\prime}_{2}$ that preserves $Q_{1}$ and $Q_{2}$, then $p_{A_{2}}\circ\Phi|_{A_{1}}:A_{1}\to A_{2}$ is an isomorphism that preserves $Q_{A_{1}}$ and $Q_{A_{2}}$, and $p_{B_{2}}\circ\Phi|_{B_{1}}:B_{1}\to B_{2}$ is an isomorphism that preserves $Q_{B_{1}}$ and $Q_{B_{2}}$.
Proof.
Due to the definition of $Q_{i}$, we can check that $\mathrm{rank}\,A_{i}$ is equal to the maximal rank of a submodule $C_{i}$ of $A^{\prime}_{i}$ such that the symmetric bilinear form $(C_{i}/\mathrm{Tor})\times(C_{i}/\mathrm{Tor})\to\mathbb{Z}$ induced by $Q_{i}$ is non-degenerate. This fact implies $\mathrm{rank}\,A_{1}=\mathrm{rank}\,A_{2}$ and also $\mathrm{rank}\,B_{1}=\mathrm{rank}\,B_{2}$, since the bilinear forms on $A^{\prime}_{1}/\mathrm{Tor}$ and $A^{\prime}_{2}/\mathrm{Tor}$ are isomorphic. Due to the assumption that $A_{1},A_{2}$ or $B_{1},B_{2}$ are torsion-free, it follows that $A_{1},B_{1}$ are respectively isomorphic to $A_{2},B_{2}$.
Let each $T_{i}$ be the torsion submodule of $A_{i}$. Clearly, $T_{1}=T_{2}=0$ in the case where $A_{1},A_{2}$ are torsion-free. Let each $F_{i}$ be a free submodule of $A_{i}$ such that $A_{i}=F_{i}\oplus T_{i}$. We note that $F_{1},T_{1}$ are isomorphic to $F_{2},T_{2}$, respectively. Let $M_{i}$ be a matrix representing the bilinear form $F_{i}\times F_{i}\to\mathbb{Z}$ induced by $Q_{A_{i}}$. Then each $\det M_{i}$ is independent of the choices of $F_{i}$ and its basis, and $\det M_{i}$ is non-zero due to the non-degeneracy of $Q_{A_{i}}$ on $A_{i}/\mathrm{Tor}$.
Let each $p_{F_{i}}:A^{\prime}_{i}=F_{i}\oplus T_{i}\oplus B_{i}\to F_{i}$ be the projection. We here show that $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$ is an isomorphism. Let $u_{1},\dots,u_{k}$ be a basis of $F_{1}$, and define $u_{1}^{\prime},\dots,u_{k}^{\prime}\in F_{2}$ by $u_{j}^{\prime}=p_{F_{2}}(\Phi(u_{j}))$. We define the matrix $M_{1}$ in the last paragraph by $M_{1}=(Q_{A_{1}}(u_{i},u_{j}))$. Since the projection $p_{F_{2}}$ preserves $Q_{2},Q_{A_{2}}$, the homomorphism $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$ preserves $Q_{A_{1}},Q_{A_{2}}$, showing that the matrix $(Q_{2}(u_{i}^{\prime},u_{j}^{\prime}))$ is equal to $M_{1}$. Due to the condition $\det M_{1}\neq 0$, we easily see that $u_{1}^{\prime},\dots,u_{k}^{\prime}$ is linearly independent (cf. [38, Corollary 1.2.13.]). We thus have a basis $v_{1},\dots,v_{k}$ of $F_{2}$ and positive integers $a_{1},\dots,a_{k}$ such that $a_{1}v_{1},\dots,a_{k}v_{k}$ generates the submodule generated by $u_{1}^{\prime},\dots,u_{k}^{\prime}$. It follows that $\det M_{1}=m\cdot\det M_{2}$ for some positive integer $m$. By reversing the roles of $A_{1}$ and $A_{2}$ and applying the same argument, we see that $\det M_{1}=\det M_{2}$. Consequently, $u_{1}^{\prime},\dots,u_{k}^{\prime}$ is a basis of $F_{2}$, and thus $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$ is an isomorphism.
We now show that $p_{A_{2}}\circ\Phi|_{A_{1}}:A_{1}\to A_{2}$ is an isomorphism that preserves $Q_{A_{1}},Q_{A_{2}}$. The projection $p_{A_{2}}$ clearly preserves $Q_{2},Q_{A_{2}}$, and hence $p_{A_{2}}\circ\Phi|_{A_{1}}$ preserves $Q_{A_{1}},Q_{A_{2}}$. Due to the assumption that $A_{1},A_{2}$ or $B_{1},B_{2}$ are torsion-free, the isomorphism $\Phi$ bijectively maps $T_{1}$ to $T_{2}$, implying that $p_{A_{2}}\circ\Phi|_{T_{1}}:T_{1}\to T_{2}$ is an isomorphism. We define a submodule $F_{2}^{\prime}$ of $A_{2}$ by $F_{2}^{\prime}=p_{A_{2}}(\Phi(F_{1}))$. Then the composition of $p_{A_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}^{\prime}$ and the projection $A_{2}=F_{2}\oplus T_{2}\to F_{2}$ is equal to the isomorphism $p_{F_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}$. This fact implies that the surjection $p_{A_{2}}\circ\Phi|_{F_{1}}:F_{1}\to F_{2}^{\prime}$ is an isomorphism, and thus $p_{F_{2}}|_{F_{2}^{\prime}}:F_{2}^{\prime}\to F_{2}$ is also an isomorphism. It follows $A_{2}=F_{2}^{\prime}\oplus T_{2}$, implying $p_{A_{2}}\circ\Phi|_{A_{1}}(A_{1})=A_{2}$. Therefore, $p_{A_{2}}\circ\Phi|_{A_{1}}:A_{1}\to A_{2}$ is an isomorphism.
Finally, we show that $p_{B_{2}}\circ\Phi|_{B_{1}}:B_{1}\to B_{2}$ is an isomorphism that preserves $Q_{B_{1}},Q_{B_{2}}$. Due to the assumption that $A_{1},A_{2}$, or $B_{1},B_{2}$, are torsion-free, $\Phi$ bijectively maps the torsion submodule of $B_{1}$ to that of $B_{2}$. We can thus easily check $p_{B_{2}}\circ\Phi|_{B_{1}}(B_{1})$ is an isomorphism that preserves $Q_{B_{1}},Q_{B_{2}}$, by using the non-degeneracy assumption of each $Q_{A_{i}}$ on $A_{i}/\mathrm{Tor}$. ∎
In the rest of this section, let $R$ be an ordered set, and let $G_{i}:A^{\prime}_{i}\to R$ be an arbitrary map for each $i=1,2$. These maps are models of genus functions and related invariants of 4-manifolds. We will construct isomorphisms that preserve the restrictions of $Q_{1},Q_{2}$ and $G_{1},G_{2}$ by using Lemma 5.1.
Proposition 5.2.
Suppose there exists an isomorphism $\Phi:A^{\prime}_{1}\to A^{\prime}_{2}$ that preserves $Q_{1},Q_{2}$ and $G_{1},G_{2}$. If $A_{1}$ and $A_{2}$ are torsion-free, then there exists an isomorphism between $B_{1}$ and $B_{2}$ that preserves $Q_{B_{1}},Q_{B_{2}}$ and $G_{1}|_{B_{1}},G_{2}|_{B_{2}}$.
Proof.
By Lemma 5.1, the map $p_{B_{2}}\circ\Phi|_{B_{1}}:B_{1}\to B_{2}$ is an isomorphism that preserves $Q_{B_{1}}$, $Q_{B_{2}}$. Since $A_{1},A_{2}$ are torsion-free, and $Q_{A_{1}},Q_{A_{2}}$ are non-degenerate, we see $\Phi(B_{1})=B_{2}$. The isomorphism $p_{B_{2}}\circ\Phi|_{B_{1}}$ thus preserves $G_{1}|_{B_{1}},G_{2}|_{B_{2}}$. ∎
Proposition 5.3.
Assume that each $G_{i}$ satisfies $G_{i}(a)\leq G_{i}(a+b)$ for any $a\in A_{i}$ and $b\in B_{i}$, and further that there exists an isomorphism $\Phi:A^{\prime}_{1}\to A^{\prime}_{2}$ that preserves $Q_{1},Q_{2}$ and $G_{1},G_{2}$. If the modules $A_{1}$ and $A_{2}$ or the modules $B_{1}$ and $B_{2}$ are torsion-free, then there exists an isomorphism between $A_{1}$ and $A_{2}$ that preserves $Q_{A_{1}},Q_{A_{2}}$ and $G_{1}|_{A_{1}},G_{2}|_{A_{2}}$.
Proof.
By Lemma 5.1, we see that $p_{A_{2}}\circ\Phi|_{A_{1}}:A_{1}\to A_{2}$ is an isomorphism that preserves $Q_{A_{1}}$, $Q_{A_{2}}$. Let $a$ be an element of $A_{1}$ and set $\widehat{a}=p_{A_{2}}\circ\Phi(a)$. We will prove $G_{1}(a)=G_{2}(\widehat{a})$. We see that $\Phi(a)=\widehat{a}+\widehat{a}_{B}$ for some $\widehat{a}_{B}\in B_{2}$. We consider two cases.
(1) The case where $A_{1},A_{2}$ are torsion-free. In this case, we see $\widehat{a}_{B}=\Phi(b)$ for some $b\in B_{1}$ due to the non-degeneracy of $Q_{\widetilde{A}_{1}}=Q_{A_{1}}$. By the assumptions on $G_{1},G_{2},\Phi$, we see
$G_{1}(a)=G_{2}(\widehat{a}+\widehat{a}_{B})\geq G_{2}(\widehat{a})=G_{1}(a-b)\geq G_{1}(a),$
giving $G_{1}(a)=G_{2}(\widehat{a})$.
(2) The case where $B_{1},B_{2}$ are torsion-free. We see $\widehat{a}_{B}=\Phi(b+t)$ for some element $b\in B_{1}$ and some torsion element $t\in A_{1}$ due to the non-degeneracy of $Q_{\widetilde{A}_{1}}$. We denote the torsion element $\Phi(t)\in A_{2}$ by $\widehat{t}$. For each integer $n$, we have $\Phi(a+nt)=\widehat{a}+n\widehat{t}+\widehat{a}_{B}$ and $\Phi(a+(n-1)t-b)=\widehat{a}+n\widehat{t}$. By the assumptions on $G_{1},G_{2},\Phi$, we thus see
$G_{1}(a+nt)=G_{2}(\widehat{a}+n\widehat{t}+\widehat{a}_{B})$ $\geq G_{2}(\widehat{a}+n\widehat{t})$
$=G_{1}(a+(n-1)t-b)\geq G_{1}(a+(n-1)t),$
and hence we have
$G_{1}(a)\geq G_{2}(\widehat{a})\geq G_{1}(a-t)\geq G_{1}(a-2t)\cdots\geq G_{1}(a-nt)\geq\cdots.$
Since $t$ is a torsion, this implies $G_{1}(a)=G_{2}(\widehat{a})$.
Therefore, the isomorphism $p_{A_{2}}\circ\Phi|_{A_{1}}:A_{1}\to A_{2}$ preserves $Q_{A_{1}}$, $Q_{A_{2}}$ and $G_{A_{1}},G_{A_{2}}$. ∎
We also discuss maps which are models of the torsion-free versions of the genus functions and related invariants. For each $i=1,2$, we set $\widetilde{A}_{i}=A_{i}/\mathrm{Tor}$, $\widetilde{B}_{i}=B_{i}/\mathrm{Tor}$ and $\widetilde{A}^{\prime}_{i}=\widetilde{A}_{i}\oplus\widetilde{B}_{i}$. Let $\widetilde{G}_{i}^{A}:\widetilde{A}_{i}\to R$, $\widetilde{G}_{i}^{B}:\widetilde{B}_{i}\to R$ and $\widetilde{G}_{i}:\widetilde{A}^{\prime}_{i}\to R$ be the maps defined by
$\widetilde{G}_{i}^{A}(\alpha)$ $=\min\{G_{i}(a)\mid a\in A_{i},\;[a]=\alpha\},$
$\widetilde{G}_{i}^{B}(\beta)$ $=\min\{G_{i}(b)\mid b\in B_{i},\;[b]=\beta\},$
$\widetilde{G}_{i}(\gamma)$ $=\min\{G_{i}(r)\mid r\in A_{i}^{\prime},\;[r]=\gamma\},$
where $[a]$, $[b]$ and $[r]$ denote their equivalent classes. We prove the following proposition.
Proposition 5.4.
Assume that each $G_{i}$ satisfies $G_{i}(a)\leq G_{i}(a+b)$ for any $a\in A_{i}$ and $b\in B_{i}$. If there exists an isomorphism $\Psi:\widetilde{A}^{\prime}_{1}\to\widetilde{A}^{\prime}_{2}$ that preserves $\widetilde{Q}_{1},\widetilde{Q}_{2}$ and $\widetilde{G}_{1},\widetilde{G}_{2}$, then there exists an isomorphism between $\widetilde{A}_{1}$ and $\widetilde{A}_{2}$ that preserves $Q_{\widetilde{A}_{1}},Q_{\widetilde{A}_{2}}$ and $\widetilde{G}_{1}^{A},\widetilde{G}_{2}^{A}$.
Proof.
Let $p_{\widetilde{A}_{i}}:\widetilde{A}^{\prime}_{i}\to\widetilde{A}_{i}$ and $p_{\widetilde{B}_{i}}:\widetilde{A}^{\prime}_{i}\to\widetilde{B}_{i}$ be the projections. By Lemma 5.1, the composition $p_{\widetilde{A}_{2}}\circ\Psi|_{\widetilde{A}_{1}}:\widetilde{A}_{1}\to\widetilde{A}_{2}$ is an isomorphism that preserves $Q_{\widetilde{A}_{1}}$ and $Q_{\widetilde{A}_{2}}$. For each $i=1,2$, we can easily check that every $\alpha\in\widetilde{A}_{i}$ and $\beta\in\widetilde{B}_{i}$ satisfy the inequality $\widetilde{G}_{i}(\alpha+\beta)\geq\widetilde{G}_{i}(\alpha)$ due to the assumption on $G_{i}$. This inequality implies $\widetilde{G}_{i}|_{\widetilde{A}_{i}}=\widetilde{G}^{A}_{i}$. Therefore, Proposition 5.3 shows that the above isomorphism preserves $\widetilde{G}_{1}^{A}$ and $\widetilde{G}_{2}^{A}$. ∎
6. Quasi-invertible cobordisms
In this section, we introduce the notion of quasi-invertible cobordisms to study genus functions. We also provide their examples and study their algebraic properties. Throughout this section, we allow manifolds, submanifolds, and their boundaries to be disconnected. For such manifolds, the uniqueness of connected sums and boundary sums do not necessarily hold, but our results hold for any choices of sums unless otherwise stated. So, we will use ##\## and $\natural$ for these operations. Let $M,N$ be closed 3-manifolds.
6.1. Definitions and examples
Let us recall that a 4-dimensional 1-handlebody of genus $k$ $(k\geq 0)$ is the $k$-fold boundary sum $\natural_{k}(S^{1}\times D^{3})$. For convenience, we use the following terminologies.
Definition 6.1.
Suppose a 4-dimensional 1-handlebody of genus $k$ is embedded in the interior of a 4-manifold. An operation of removing the 1-handlebody from the 4-manifold and gluing $\natural_{k}S^{2}\times D^{2}$ back along the resulting boundary component will be called a surgery along a $4$-dimensional $1$-handlebody. The reverse operation will be called a surgery along an embedded copy of $\natural_{k}(S^{2}\times D^{2})$.
Note that the way to glue $\natural_{k}S^{2}\times D^{2}$ is not unique in general, but we use a convenient way in each context. By contrast, the reverse operation is uniquely determined by an embedded $\natural_{k}(S^{2}\times D^{2})$ ([47]).
Here we introduce notions of a quasi-product cobordism, a quasi-invertible cobordism and an $H_{2}$-surjective cobordism.
Definition 6.2.
(1) A cobordism $R$ from $M$ to $N$ will be called quasi-product if it is orientedly diffeomorphic to the product $I\times M$ possibly after performing surgeries along finitely many embedded $4$-dimensional $1$-handlebodies. If furthermore the homomorphism $H_{2}(M)\to H_{2}(R)$ induced by the inclusion is an isomorphism, then $R$ will be called strongly quasi-product.
(2) A cobordism $P$ from $M$ to $N$ will be called quasi-invertible (resp. strongly quasi-invertible) if it admits a cobordism $Q$ from $N$ to $M$ such that a glued 4-manifold $P\cup_{N}Q$ is a quasi-product cobordism (resp. strongly quasi-product cobordism) from the boundary component $M$ of $P$ to that of $Q$. The cobordism $Q$ will be called a quasi-inverse cobordism (resp. strongly quasi-inverse cobordism) of $P$.
(3) A cobordism $P$ from $M$ to $N$ will be called $H_{2}$-surjective if the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is surjective.
As we will later show, there are numerous examples of quasi-invertible cobordisms. To produce examples using intensively studied objects knotted 2-spheres, let us recall a few terminologies. A surgery on a circle embedded in a 4-manifold is an operation of removing the tubular neighborhood $S^{1}\times D^{3}$ of the circle and gluing $S^{2}\times D^{2}$ back along the resulting boundary component. In general, there are two ways to glue $S^{2}\times D^{2}$ along the boundary component ([33]), but we use a convenient way in each context. Clearly, any surgery on a circle is realized by a surgery along an embedded $4$-dimensional $1$-handlebody of genus one. The surgery on a $2$-sphere embedded in a 4-manifold with trivial normal bundle is an operation of removing the tubular neighborhood $S^{2}\times D^{2}$ of the 2-sphere from the 4-manifold and gluing $S^{1}\times D^{3}$ back along the resulting boundary. We note that the resulting 4-manifold is uniquely determined by the embedded 2-sphere ([47]). An $S^{2}$-link in a 4-manifold is a submanifold of the interior of the 4-manifold whose each connected component is a 2-sphere.
We use the following notation.
Definition 6.3.
For an $S^{2}$-link $L$ in $S^{4}$, let $S^{4}(L)$ be the closed 4-manifold obtained from $S^{4}$ by performing the surgery on $L$, that is, by performing the surgery on each component of $L$. Note that the normal bundle of any 2-sphere embedded in $S^{4}$ is trivial.
It is easy to see that $n(S^{1}\times S^{3})$ is an example of $S^{4}(L)$ for each $n\geq 0$. We here study quasi-product cobordisms. Clearly, $I\times M$ is strongly quasi-product. Furthermore, we have simple examples below.
Lemma 6.4.
$(1)$ A $4$-manifold is a quasi-product cobordism from $M$ to $M$ if and only if the $4$-manifold is diffeomorphic to a $4$-manifold obtained from $I\times M$ by performing surgeries along finitely many embedded copies of the boundary sums of $S^{2}\times D^{2}$’s.
$(2)$ For any $S^{2}$-link $L$ in $S^{4}$, the connected sum $(I\times M)\#S^{4}(L)$ is a strongly quasi-invertible cobordism from $M$ to $M$.
Proof.
The claim (1) is straightforward. We have $H_{2}(S^{4}(L))=0$ by Lemma 8.3, so the claim (2) follows from (1). ∎
Remark 6.5.
Quasi-product cobordisms are not necessarily strongly quasi-product. For example, let $P$ be the 4-manifold obtained from $I\times S^{1}\times S^{2}$ by performing the surgery on the 2-sphere $a\times b\times S^{2}$ ($a\in\mathrm{int}\,I$, $b\in S^{1}$). Then $P$ is a quasi-product cobordism from $S^{1}\times S^{2}$ to $S^{1}\times S^{2}$, but the homomorphism $H_{2}(0\times S^{1}\times S^{2})\to H_{2}(R)$ induced by the inclusion is the zero-map, and thus $P$ is not strongly quasi-product.
Now we produce examples of quasi-invertible cobordisms. Clearly, invertible cobordisms are strongly quasi-invertible, and quasi-product cobordisms are quasi-invertible. To produce non-trivial examples, we recall the notion of a strongly slice link.
Definition 6.6.
A (1-dimensional) link $\ell$ in $N$ will be called strongly slice if $\ell$ in $1\times N$ bounds a disjoint union of smoothly and properly embedded disks in $I\times N$. We note that each component of a strongly slice link is null-homologous, and thus its 0-framing is defined as the Seifert framing.
We can generate various examples of quasi-invertible cobordisms from the above trivial examples by the lemma below.
Lemma 6.7.
Let $L$ be an $S^{2}$-link in $S^{4}$, and let $V$ be a codimension-00 submanifold of $S^{4}(L)$. Then, for any quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism $P$ from $M$ to $N$, the following hold.
$(1)$ Every codimension-00 submanifold of $P$ containing the boundary component $M$ of $P$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the other boundary component.
$(2)$ $P\#V$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the disjoint union $N\sqcup\partial V$.
$(3)$ If $\partial V$ is non-empty, then $P\natural V$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to $N\#\partial V$, where the boundary sum is taken for the boundary component $N$ of $P$.
$(4)$ Every $4$-manifold obtained by attaching $1$-handles to $P$ along a connected component of $N$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the new boundary component.
$(5)$ Every $4$-manifold obtained by attaching an algebraically canceling pair of $1$- and $2$-handles to $P$ along $N$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the new boundary component.
$(6)$ Every $4$-manifold obtained by attaching $2$-handles to $P$ along a strongly slice link $\ell$ in $N$ with 00-framings is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the 00-surgery on $\ell$ in $N$.
Furthermore, the quasi-invertible cobordisms given by these claims are invertible in the case where $P,L$ are respectively an invertible cobordism and an empty set.
Proof.
The claim (1) is straightforward, and the claims (2) and (3) are straightforward from Lemma 6.4 and the claim (1). For (4), by further attaching a canceling 2-handle to each 1-handle, we obtain a 4-manifold diffeomorphic to $P$. The claim (4) thus follows from (1). For (5), by further attaching the upside down cobordism of the cobordism consisting of the handle pair, we obtain a 4-manifold diffeomorphic to $P$. The claim (5) thus follows from (1). The claim (6) is easily checked as follows. The link $\ell$ in $0\times N$ bounds a disjoint union of smoothly embedded disks in $I\times N$, and the tubular neighborhood of these disks are the 2-handles attached to $P$. Thus the 4-manifold in (6) can be embedded into $P\cup_{\mathrm{id}_{N}}(I\times N)$, and thus the claim (6) follows from (1). ∎
See also Section 8 for examples of codimension-00 submanifolds of $S^{4}(L)$.
6.2. Algebraic properties
We study algebraic properties of quasi-invertible cobordisms. We begin with quasi-product cobordisms.
Lemma 6.8.
Let $R$ be a quasi-product cobordism from $M$ to $N$. Then $R$ is $H_{2}$-surjective, and the homomorphism $H_{1}(M)\to H_{1}(R)$ induced by the inclusion is injective. Consequently, $R$ is strongly quasi-product if and only if the homomorphism $H_{2}(M)\to H_{2}(R)$ induced by the inclusion is injective.
Proof.
Let $\widehat{R}$ be the 4-manifold obtained from $R$ by performing surgeries along finitely many embedded 1-handlebodies $\natural_{k_{i}}S^{1}\times D^{3}$’s so that $\widehat{R}$ is diffeomorphic to $I\times M$. We remove the interior of these 1-handlebodies from $R$ and denote the resulting 4-manifold by $R_{c}$. Let ${j}:M\hookrightarrow{R}$, $j_{c}:M\hookrightarrow R_{c}$, $h:R_{c}\hookrightarrow R$, $\widehat{h}:R_{c}\hookrightarrow\widehat{R}$ and $\widehat{j}:M\hookrightarrow\widehat{R}$ be the inclusions.
By the excision theorem and the Poincaré duality, we see that $H_{3}(\widehat{R},R_{c})\cong\oplus H_{3}(\natural_{k_{i}}(S^{2}\times D^{2}),k_{i}(S^{1}\times S^{2}))\cong 0$. The homology exact sequence for the pair $(\widehat{R},R_{c})$ thus shows that $\widehat{h}_{*}:H_{2}(R_{c})\to H_{2}(\widehat{R})$ is injective. Since the inclusion $\widehat{j}=\widehat{h}\circ j_{c}:M\hookrightarrow\widehat{R}$ is a homotopy equivalence, the injection $\widehat{h}_{*}$ and also ${j_{c}}_{*}:H_{2}(M)\to H_{2}(R_{c})$ are isomorphisms.
By the excision theorem and the Poincaré duality, we see that $H_{2}(R,R_{c})\cong\oplus H_{2}(\natural_{k_{i}}(S^{1}\times D^{3}),k_{i}(S^{1}\times S^{2}))\cong 0$. The homology exact sequence for the pair $({R},R_{c})$ thus shows that ${h}_{*}:H_{2}(R_{c})\to H_{2}(R)$ is surjective and that ${h}_{*}:H_{1}(R_{c})\to H_{1}(R)$ is injective.
Now we can easily prove the claims. Since ${h}_{*}:H_{2}(R_{c})\to H_{2}(R)$ is surjective, and ${j_{c}}_{*}:H_{2}(M)\to H_{2}(R_{c})$ is bijective, the composition $j_{*}=h_{*}\circ{j_{c}}_{*}:H_{2}(M)\to H_{2}(R)$ is also surjective. Consequently, $R$ is $H_{2}$-surjective. Since the inclusion $\widehat{j}=\widehat{h}\circ j_{c}:M\hookrightarrow\widehat{R}$ is a homotopy equivalence, we see that ${j_{c}}_{*}:H_{1}(M)\hookrightarrow H_{1}(R_{c})$ is injective, and thus the composition $j_{*}=h_{*}\circ{j_{c}}_{*}:H_{1}(M)\to H_{1}(R)$ is also injective. ∎
In a special case, we can easily check the strongness.
Corollary 6.9.
If $M$ is a disjoint union of rational homology $3$-spheres, then any quasi-product ((((resp. quasi-invertible)))) cobordism from $M$ to $N$ is strongly quasi-product ((((resp. strongly quasi-invertible)))).
Proof.
Since the rank of $H_{1}(M)$ is zero, we see $H_{2}(M)=0$ by the Poincaré duality and the universal coefficient theorem. The claim thus follows from Lemma 6.8. ∎
We next discuss quasi-invertible cobordisms. In the rest of this subsection, let $P$ be a quasi-invertible cobordism from $M$ to $N$, and let $Q$ be a quasi-inverse cobordism of $P$. We denote a quasi-product cobordism $P\cup_{N}Q$ by $R$. In the case where $P$ is strongly quasi-invertible, we assume $R$ is strongly quasi-product. Furthermore, let $A^{M}_{P}$ and $B_{P}$ respectively denote the image of the homomorphism $H_{2}(M)\to H_{2}(P)$ and the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$, where these are the homomorphisms induced by the inclusions. Then the following lemma holds.
Lemma 6.10.
$(1)$ The homomorphism $H_{1}(M)\to H_{1}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=A^{M}_{P}+B_{P}$ holds.
$(2)$ If $P$ is strongly quasi-invertible, then the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=A^{M}_{P}\oplus B_{P}$ holds.
Proof.
Since the homomorphism $H_{*}(M)\to H_{*}(R)$ is the composition of homomorphisms $H_{*}(M)\to H_{*}(P)$ and $H_{*}(P)\to H_{*}(R)$, the claims follow from Lemma 6.8. ∎
Finally, we discuss algebraic properties of an attachment of a quasi-invertible cobordisms. Let $X$ be a 4-manifold having a boundary component orientedly diffeomorphic to $M$, and let $X^{\prime}$ denote a 4-manifold $X\cup_{M}P$. We attach $Q$ to $X^{\prime}$ so that the submanifold $R=P\cup_{N}Q$ is a quasi-product cobordism. Let $X^{\prime\prime}$ denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$. Let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions. Then the following properties hold.
Lemma 6.11.
$(1)$ The homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is surjective.
$(2)$ $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(B_{P})$.
$(3)$ The restriction of the intersection form of $X^{\prime}$ to ${\iota_{P}}_{*}(B_{P})$ is the zero-map.
$(4)$ $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is surjective if and only if $P$ is $H_{2}$-surjective.
Proof.
$(1)$ By the homology exact sequence for the pair $(R,M)$ and Lemma 6.8, we see $H_{2}(R,M)=0$. The excision theorem shows $H_{2}(X^{\prime\prime},X)\cong H_{2}(R,M)=0$. The homology exact sequence for the pair $(X^{\prime\prime},X)$ thus shows that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is surjective.
$(2)$ The claim (1) and the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$ implies $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+B_{X^{\prime}}$, where $B_{X^{\prime}}$ denotes the kernel of the homomorphism $H_{2}(X^{\prime})\to H_{2}(X^{\prime\prime})$ induced by the inclusion. We note ${\iota_{P}}_{*}(B_{P})\subset B_{X^{\prime}}$. Using the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}P$ and Lemma 6.10, we can check $\iota_{*}(H_{2}(X))+B_{X^{\prime}}=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(B_{P})$, showing the claim.
$(3)$ By Lemma 6.8, the intersection form of $R$ is the zero-map, and thus the claim follows.
$(4)$ By Lemma 6.10 and the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}P$, the homomorphism $\iota_{*}:H_{1}(X)\to H_{1}(X^{\prime})$ is injective. Thus, due to the homology exact sequence for the pair $(X^{\prime},X)$, the homomorphism $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is surjective if and only if $H_{2}(X^{\prime},X)=0$. Similarly, by Lemma 6.10, $P$ is $H_{2}$-surjective if and only if $H_{2}(P,M)=0$. The claim thus follows from the fact $H_{2}(X^{\prime},X)\cong H_{2}(P,M)$. ∎
The corollary below is straightforward.
Corollary 6.12.
Suppose $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is injective. Then, $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is an isomorphism if and only if $P$ is $H_{2}$-surjective.
We also have the following decomposition.
Lemma 6.13.
If the intersection form of $X$ is non-degenerate ((((on $H_{2}(X)/\mathrm{Tor}$)))), then $H_{2}(X^{\prime})/\mathrm{Tor}=\iota_{*}(H_{2}(X))/\mathrm{Tor}\,\oplus\,{\iota_{P}}_{*}(B_{P})/\mathrm{Tor}$.
Proof.
Since every non-torsion class of $H_{2}(X)$ algebraically intersects with some class of $H_{2}(X)$ under the assumption, the claim easily follows from Lemma 6.11. ∎
In the case where the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective, nice properties hold.
Proposition 6.14.
Suppose the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is injective. Then, $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is injective, and the decomposition $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))\oplus{\iota_{P}}_{*}(B_{P})$ holds.
Proof.
The injectivity of $\iota_{*}$ is clear from the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$. Since the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective, we easily see $\iota_{*}(H_{2}(X))\cap{\iota_{P}}_{*}(B_{P})=0$. Lemma 6.11 thus shows $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))\oplus{\iota_{P}}_{*}(B_{P})$. ∎
We give sufficient conditions that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective.
Proposition 6.15.
Suppose at least one of the following conditions (i)(iii) holds.
  1. (i)
    $P$ is strongly quasi-invertible.
  2. (ii)
    The homomorphism $H_{2}(M)\to H_{2}(X)$ induced by the inclusion is the zero-map.
  3. (iii)
    The intersection form of $X$ is non-degenerate, and $H_{2}(X)$ has no torsion.
Then the homomorphisms $H_{2}(X)\to H_{2}(X^{\prime})$ and $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusions are injective.
Proof.
In the case where (i) or (ii) holds, by the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}R$ and Lemma 6.8, we see that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective, and thus the homomorphism $H_{2}(X)\to H_{2}(X^{\prime})$ is also injective. In the case where (iii) holds, we easily see that the condition (ii) also holds, and hence the claim holds. ∎
The property below is useful for studying genus functions.
Lemma 6.16.
Suppose at least one of the following conditions (i)(iii) holds.
  1. (i)
    $P$ is strongly quasi-product.
  2. (ii)
    $P$ is invertible.
  3. (iii)
    $P$ is strongly quasi-invertible, and $N$ is a disjoint union of rational homology $3$-spheres.
Then, $H_{2}(P)$ and ${\iota_{P}}_{*}(B_{P})$ are torsion-free.
Proof.
By the Poincaré duality and the universal coefficient theorem, we see that $H_{2}(N)$ and $H_{2}(M)$ are torsion-free. Since $P$ is strongly quasi-inverible, $H_{2}(R)$ is also torsion-free.
We consider the Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$, and we will show that the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ in the sequence is injective. In the case where (i) holds, we may assume that $Q=I\times N$, and hence this homomorphism is injective. In the case where (iii) holds, we see $H_{2}(N)=0$ as shown in the proof of Corollary 6.9, and hence the homomorphism is injective. So, we prove the injectivity in the case where (ii) holds. In this case, we may assume $R$ is diffeomorphic to $I\times M$, and hence the inclusion $M\hookrightarrow R$ is a homotopy equivalence. Due to the sequence $M\subset P\subset R$, this fact implies that the homomorphism $H_{3}(P)\to H_{3}(R)$ induced by the inclusion is surjective. The Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$ thus shows that the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ is injective.
Now we are ready to prove the claim. Since $H_{2}(N)$ and $H_{2}(R)$ are torsion-free, and the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ is injective, the Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$ implies that $H_{2}(P)$ is torsion-free, and hence $B_{p}$ is also torsion-free. By Lemma 6.10, we have $H_{2}(P)=A^{M}_{P}\oplus B_{P}$. The Mayer–Vietoris exact sequence for $X^{\prime}=X\cup_{M}P$ thus shows that the restriction of ${\iota_{P}}_{*}:H_{2}(P)\to H_{2}(X^{\prime})$ to $B_{P}$ is an isomorphism. Therefore, ${\iota_{P}}_{*}(B_{P})$ is torsion-free. ∎
7. Genus functions and quasi-invertible cobordisms
In this section, we study behavior of genus functions of 4-manifolds with quasi-invertible cobordisms attached. We also show stabilities of algebraic inequivalences of genus functions under attachments of quasi-invertible cobordisms. Throughout this section, we allow manifolds, submanifolds, and their boundaries to be disconnected unless otherwise stated, but we still assume surfaces are connected.
7.1. Behavior of genus functions
We first study behavior of genus functions under attachments of quasi-invertible cobordisms. Throughout this section, we use the following setting. Let $M,N$ be closed 3-manifolds, and let $P$ be a quasi-invertible cobordism from $M$ to $N$. Let $X$ be a connected 4-manifold having a boundary component orientedly diffeomorphic to $M$, and we denote a glued 4-manifold $X\cup_{M}P$ by $X^{\prime}$. Let $\iota:X\hookrightarrow X^{\prime}$ be the inclusion.
Under these assumptions, we prove the following theorem.
Theorem 7.1.
The inclusion $\iota:X\hookrightarrow X^{\prime}$ preserves the genus functions.
To prove this theorem, we show the following lemma, which reveals an important property of genus functions.
Lemma 7.2.
$(1)$ If a $4$-manifold $Z^{\prime}$ is obtained from a connected $4$-manifold $Z$ by attaching a $3$- or a $4$-handle, then the inclusion $Z\hookrightarrow Z^{\prime}$ preserves the genus functions.
$(2)$ If a $4$-manifold $Z^{\prime}$ is obtained from a connected $4$-manifold $Z$ by attaching $\natural_{n}(S^{1}\times D^{3})$ along a boundary component diffeomorphic to $n(S^{1}\times S^{2})$ for some $n\geq 0$, then the inclusion $Z\hookrightarrow Z^{\prime}$ preserves the genus functions.
Proof.
(1) We first consider the case where the attached handle is a 3-handle $D^{3}\times D^{1}$. Let $j:Z\hookrightarrow Z^{\prime}$ denote the inclusion. We note that the kernel of $j_{*}:H_{2}(Z)\to H_{2}(Z^{\prime})$ is the image of the homomorphism $H_{2}(S^{2}\times D^{1})\to H_{2}(Z)$ induced by the inclusion as seen from the Mayer–Vietoris exact sequence, where $S^{2}\times D^{1}$ denotes the attaching region of the 3-hanlde.
Let $\alpha$ be a class of $H_{2}(Z)$ and define the class $\alpha^{\prime}$ of $H_{2}(Z^{\prime})$ by $\alpha^{\prime}={j}_{*}(\alpha)$. We set $g=g_{Z}(\alpha)$ and $g^{\prime}=g_{Z^{\prime}}(\alpha^{\prime})$. Then $\alpha$ (resp. $\alpha^{\prime}$) is represented by a closed surface $\Sigma$ (resp. $\Sigma^{\prime}$) of genus $g$ (resp. $g^{\prime}$) embedded in the interior of $Z$ (resp. $Z^{\prime}$). Clearly, $\alpha^{\prime}$ is also represented by the surface $\Sigma$, showing $g\geq g^{\prime}$.
We next show $g\leq g^{\prime}$. Since the cocore of the 3-handle is $0\times D^{1}$, we may assume that $\Sigma^{\prime}$ is disjoint from a tubular neighborhood of the cocore by taking a general position. Therefore, we can isotope $\Sigma^{\prime}$ into the interior of $Z$. We thus see that $\alpha^{\prime}=j_{*}([\Sigma^{\prime}])$ and hence $j_{*}(\alpha)=j_{*}([\Sigma^{\prime}])$. It follows that $\alpha=[\Sigma^{\prime}]+\beta$ in $H_{2}(Z)$ for some $\beta\in\mathrm{Ker}\,j_{*}$. Here note that any class of $\mathrm{Ker}\,j_{*}$ is represented by a disjoint union of 2-spheres in the attaching region of the 3-handle. By pushing these 2-spheres into the interior of $Z$ and taking a connected sum of $\Sigma^{\prime}$ with these 2-spheres, we can construct a closed surface of genus $g^{\prime}$ embedded in $Z$ which represents $\alpha$. Hence, we see $g\leq g^{\prime}$. Therefore, we have $g=g^{\prime}$, showing that $j$ preserves the genus functions.
In the case where the attached handle is a 4-handle, the claim follows by a much simpler argument.
(2) Since $Z^{\prime}$ is obtained from $Z$ by attaching 3-and 4-handles, the claim follows from (1). ∎
Now we prove Theorem 7.1.
Proof of Theorem 7.1.
We attach a quasi-inverse cobordism $Q$ of $P$ to $X^{\prime}$ so that the submanifold $R=P\cup_{N}Q$ is a quasi-product cobordism. Let $X^{\prime\prime}$ denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$. The cobordism $R$ becomes diffeomorphic to $I\times M$ after performing surgeries along finitely many embedded 4-dimensional 1-handlebodies. Thus, the 4-manifold $\widehat{X}^{\prime\prime}$ obtained from $X^{\prime\prime}$ by these surgeries is diffeomorphic to $X\cup_{\mathrm{id}_{M}}(I\times M)$. We denote by $\widehat{j}:X\hookrightarrow\widehat{X}^{\prime\prime}$ the inclusion. By shrinking $I\times M$ together with the collar neighborhood of $\partial X$, we can easily construct a diffeomorphism $\varphi:\widehat{X}^{\prime\prime}\to X$ such that $\varphi\circ\widehat{j}:X\to X$ is homotopic to $\mathrm{id}_{X}$. Since the diffeomorphisms $\varphi$ and $\mathrm{id}_{X}$ preserves the genus functions, it follows that $\widehat{j}:X\hookrightarrow\widehat{X}^{\prime\prime}$ also preserves the genus functions.
We remove the interior of the aforementioned 1-handlebodies in $R$ from $X^{\prime\prime}$ and denote the resulting 4-manifold by $X^{\prime\prime}_{c}$. Let $j:X\hookrightarrow X^{\prime\prime}$, $j^{\prime}:X^{\prime}\hookrightarrow X^{\prime\prime}$, $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$, $h:X^{\prime\prime}_{c}\hookrightarrow X^{\prime\prime}$ and $\widehat{h}:X^{\prime\prime}_{c}\hookrightarrow\widehat{X}^{\prime\prime}$ be the inclusions. Since $j_{c}$ and $\widehat{h}$ are inclusions, we see that the inequalities $g_{X}(\alpha)\geq g_{X^{\prime\prime}_{c}}({j_{c}}_{*}(\alpha))$ and $g_{X^{\prime\prime}_{c}}(\beta)\geq g_{\widehat{X}^{\prime\prime}}(\widehat{h}_{*}(\beta))$ hold for any $\alpha\in H_{2}(X)$ and $\beta\in H_{2}(X^{\prime\prime}_{c})$. As noted before, $\widehat{j}=\widehat{h}\circ j_{c}:X\hookrightarrow\widehat{X}^{\prime\prime}$ preserves the genus functions. The above inequalities thus imply that $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$ also preserves the genus functions. By Lemma 7.2, the inclusion $h:X^{\prime\prime}_{c}\hookrightarrow X^{\prime\prime}$ preserves the genus functions, and thus $j=h\circ j_{c}:X\hookrightarrow X^{\prime\prime}$ also preserves the genus functions. This fact implies that $\iota:X\hookrightarrow X^{\prime}$ preserves the genus functions due to the relation $j=j^{\prime}\circ\iota$. ∎
Remark 7.3.
(1) By Theorem 7.1 and Lemma 6.7, the genus functions are preserved under attachment of a 1-handle whose attaching region belongs to a connected boundary component and also under attachment of 2-handles along a strongly slice link with 0-framings.
(2) To prove a result similar to Theorem 7.1 for other invariants, we note that the following properties (i)–(iii) of genus functions are used in the above proof. (i) Every diffeomorphism preserves the genus functions. (ii) For any orientation preserving embedding, the homomorphism induced by the embedding does not strictly increase the values of their genus functions. (iii) Lemma 7.2.(1). Even when we do not use Lemma 7.2.(1), we can still prove Theorem 7.1 in the case where $P$ is an invertible cobordism as seen from the above proof.
The corollary below is straightforward.
Corollary 7.4.
If ${\iota}_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is an isomorphism, then the genus function of ${X^{\prime}}$ is algebraically equivalent to that of ${X}$.
The corollary below provides sufficient conditions that $g_{X^{\prime}}$ is algebraically equivalent to $g_{X}$.
Corollary 7.5.
Suppose that $P$ is $H_{2}$-surjective and that at least one of the following conditions (i)(iii) holds.
  1. (i)
    $P$ is strongly quasi-invertible.
  2. (ii)
    The homomorphism $H_{2}(M)\to H_{2}(X)$ induced by the inclusion is the zero-map.
  3. (iii)
    The intersection form of $X$ is non-degenerate, and $H_{2}(X)$ has no torsion.
Then the genus function of ${X^{\prime}}$ is algebraically equivalent to that of ${X}$.
Proof.
This is straightforward from Corollary 7.4, Lemma 6.11, and Proposition 6.15. ∎
In the case where $P$ is $H_{2}$-surjective, the genus function $g_{X^{\prime}}$ is completely determined by $g_{X}$ and $\iota_{*}$ due to Theorem 7.1 and Lemma 6.11. In the case where $P$ is not $H_{2}$-surjective, we give estimates on the values of $g_{X^{\prime}}$ for all classes. As in Subsection 6.2, let $B_{P}$ be the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$ induced by the inclusion, where $R$ is the quasi-product cobordism defined in the proof of Theorem 7.1. Also, let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions. Due to Lemma 6.11, we have $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(B_{P})$. We prove the following estimates for $g_{X^{\prime}}$.
Proposition 7.6.
The following inequality holds for any $\alpha\in H_{2}(X)$ and any $\beta\in B_{P}$.
$g_{X}(\alpha)\leq g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\leq g_{X}(\alpha)+g_{P}(\beta).$
Proof.
The right inequality is obvious, so we prove the left inequality. Let $X^{\prime\prime}$ be as in the proof of Theorem 7.1, and let $j:X\hookrightarrow X^{\prime\prime}$ and $j^{\prime}:X^{\prime}\hookrightarrow X^{\prime\prime}$ be the inclusions. By the definition of $B_{P}$, we see $j^{\prime}_{*}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))=j_{*}(\alpha)$, showing $g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\geq g_{X^{\prime\prime}}(j_{*}(\alpha))$. Since $j_{*}$ preserves the genus functions due to Theorem 7.1, we obtain $g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\geq g_{X}(\alpha)$. ∎
There are many examples of quasi-invertible cobordisms $P$ such that $g_{P}$ is the zero-map, and for such examples, $g_{X^{\prime}}$ is completely determined even when $P$ is not $H_{2}$-surjective due to Proposition 7.6.
Similarly to ordinary genus functions, torsion-free genus functions are preserved by the inclusions as easily seen from Proposition 7.6 and Lemma 6.11.
Corollary 7.7.
The inclusion $\iota:X\hookrightarrow X^{\prime}$ preserves the torsion-free genus functions.
7.2. Stabilities for algebraic inequivalences of genus functions
We next discuss stabilities for algebraic inequivalences of genus functions under attachments of quasi-invertible cobordisms.
In the rest of this section, we use the following setting. For $i=1,2$, let $M_{i},N_{i}$ be closed 3-manifolds, let $P_{i}$ be a quasi-invertible cobordism from $M_{i}$ to $N_{i}$, and let $Q_{i}$ be its quasi-inverse cobordism. Also, let $X_{i}$ be a connected 4-manifolds having a boundary component diffeomorphic to $M_{i}$, and we denote a glued 4-manifold $X_{i}\cup_{M_{i}}P_{i}$ by $X^{\prime}_{i}$. We attach $Q_{i}$ to $X_{i}^{\prime}$ so that the submanifold $R_{i}=P_{i}\cup_{N_{i}}Q_{i}$ is a quasi-product cobordism. Let $X^{\prime\prime}$ denote the resulting 4-manifold $X_{i}\cup_{M_{i}}P_{i}\cup_{N_{i}}Q_{i}$. Let $\iota:X_{i}\hookrightarrow X^{\prime}_{i}$ and $j:X_{i}\hookrightarrow X^{\prime\prime}_{i}$ be the inclusions. Furthermore, we define the subgroup ${\iota_{P_{i}}}_{*}(B_{P_{i}})$ of $H_{2}(X_{i}^{\prime})$ as in Subsection 6.2.
The theorem below shows a stability for algebraic inequivalence of genus functions.
Theorem 7.8.
$(1)$ Suppose each ${\iota_{i}}_{*}:H_{2}(X_{i})\to H_{2}(X^{\prime}_{i})$ is an isomorphism. Then the genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
$(2)$ Suppose each ${\iota_{i}}_{*}:H_{2}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime}_{i})/\mathrm{Tor}$ is an isomorphism. Then the torsion-free genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
Proof.
The claim (1) is straightforward from Corollary 7.4. For (2), the torsion-free genus functions of each $X_{i}^{\prime}$ is algebraically equivalent to that of $X_{i}$ due to Corollary 7.7. The claim (2) thus follows. ∎
For general quasi-invertible cobordisms, the stability of algebraic inequivalences still holds under additional assumptions. See Propositions 6.14 and 6.15 and Lemma 6.16 for sufficient conditions that the assumption below holds.
Theorem 7.9.
Suppose the intersection form of each $X_{i}$ is non-degenerate. Suppose further that the groups $H_{2}(X_{1})$ and $H_{2}(X_{2})$ or the groups ${\iota_{P_{1}}}_{*}(B_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(B_{P_{2}})$ are torsion-free and that each ${j_{i}}_{*}:H_{2}(X_{i})\to H_{2}(X^{\prime\prime}_{i})$ is injective. If the genus functions of ${X^{\prime}_{1}}$ and $X^{\prime}_{2}$ are algebraically equivalent, then those of ${X_{1}}$ and $X_{2}$ are algebraically equivalent. If furthermore $H_{2}(X_{1})$ and $H_{2}(X_{2})$ are torsion-free, then the restrictions of the genus functions of ${X^{\prime}_{1}}$ and $X^{\prime}_{2}$ respectively to ${\iota_{P_{1}}}_{*}(B_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(B_{P_{2}})$ are also algebraically equivalent.
Proof.
Due to Propositions 5.3 and 5.2, the claim follows from Lemma 6.11, Propositions 6.14 and 7.6 and, Theorem 7.1. ∎
For torsion-free genus functions, a similar stability holds under a much simpler assumption.
Theorem 7.10.
Suppose the intersection form of each $X_{i}$ is non-degenerate. If the torsion-free genus functions of ${X^{\prime}_{1}}$ and $X^{\prime}_{2}$ are algebraically equivalent, then those of ${X_{1}}$ and $X_{2}$ are algebraically equivalent.
Proof.
Since the intersection form of each $X_{i}$ is non-degenerate, we see that $\iota_{*}:H_{"}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime}_{i})/\mathrm{Tor}$ is injective. Due to Proposition 5.4, the claim follows from Lemmas 6.11 and 6.13, Proposition 7.6 and Corollary 7.7. ∎
8. Genus functions, connected sums and boundary sums
In this section, we study behavior of genus functions under connected sums and boundary sums with a certain type of 4-manifolds, by applying results in the last section. Throughout this section, we allow manifolds, submanifolds, and their boundaries to be disconnected unless otherwise stated. Our results for boundary sums hold for any choice of a boundary component unless otherwise stated, and hence we do not specify the boundary component.
8.1. Submanifolds of $S^{4}(L)$
Before discussing genus functions, we show that various 4-manifolds are realized as codimension-00 submanifolds of the surgered 4-manifold $S^{4}(L)$ (see Definition 6.3) for some $S^{2}$-link $L$ in $S^{4}$.
Lemma 8.1.
Let $L,L^{\prime}$ be $S^{2}$-links in $S^{4}$, and let $V,V^{\prime}$ be connected codimension-00 submanifolds of $S^{4}(L),S^{4}(L^{\prime})$ with non-empty boundaries, respectively. Then, the following hold. Note that $S^{4}(L)=S^{4}$ in the case where $L$ is an empty set.
$(1)$ The boundary sum $V\natural V^{\prime}$ admits an embedding into $S^{4}(L\sqcup L^{\prime})$, where $L\sqcup L^{\prime}$ denotes the disjoint union of $L$ and $L^{\prime}$ in $S^{4}\#S^{4}\cong S^{4}$.
$(2)$ Attach 1-handles to $V$. Then the resulting $4$-manifold admits an embedding into $S^{4}(L)$.
$(3)$ Attach an algebraically canceling pair of $1$- and $2$-handles to $V$. Then the resulting $4$-manifold admits an embedding into $S^{4}(L)$.
$(4)$ Attach $2$-handles to $V$ along a strongly slice link in $\partial V$ with 00-framings. Then the resulting $4$-manifold admits an embedding into $S^{4}(L)$.
$(5)$ Let $\ell$ be a link in $\partial(\natural_{n}S^{2}\times D^{2})$ that bounds a disjoint union of smoothly and properly embedded disks in $\natural_{n}S^{2}\times D^{2}$. Attach $2$-handles to $\natural_{n}S^{1}\times D^{3}$ along a link $\ell$ with the framings induced from the disks, where $\ell$ is regarded as a link in $\partial(\natural_{n}S^{1}\times D^{3})$ using an arbitrary diffeomorphism $\partial(\natural_{n}S^{2}\times D^{2})\to\partial(\natural_{n}S^{1}\times D^{3})$. Then the resulting $4$-manifold admits an embedding into $S^{4}$.
Proof.
The claim (1) is straightforward, and the claims (2)–(4) are proved similarly to the proof of Lemma 6.7, so we prove (5). The tubular neighborhood of the disks bounding $\ell$ form 2-handles attached $\natural_{n}S^{1}\times D^{3}$. Since $\natural_{n}S^{2}\times D^{2}$ becomes diffeomorphic to $S^{4}$ by attaching $\natural_{n}S^{1}\times D^{3}$ for any gluing map ([47]), the claim (5) follows. ∎
Using the above lemma, we can generate numerous examples. Here we provide simple examples.
Example 8.2.
Let $\Sigma_{g}$ denote a closed surface of genus $g$. Then $\Sigma_{g}\times D^{2}$ $(g\geq 0)$, $\natural_{n}S^{1}\times D^{3}$ $(n\geq 0)$, Mazur-type contractible 4-manifolds, their boundary sums, and their connected sums admit embeddings into $S^{4}$ and hence realized as codimension-00 submanifolds of $S^{4}$. Furthermore, these 4-manifolds with 2-handles attached along strongly slice links with 0-framings are realized as codimension-00 submanifolds of $S^{4}$.
We here give sufficient conditions that the second homology group of a submanifold of $S^{4}(L)$ is torsion-free.
Lemma 8.3.
$(1)$ $H_{2}(S^{4}(L))=0$.
$(2)$ For any codimension-00 submanifold of $S^{4}$, its second homology group is torsion-free.
$(3)$ If the boundary of a codimension-00 submanifold $Z$ of $S^{4}(L)$ is a disjoint union of rational homology $3$-spheres, then $H_{2}(Z)$ is torsion-free.
Proof.
We first prove the claim (1). Let $E(L)$ be the exterior of $L$ in $S^{4}$. By the homology exact sequence for the pair $(S^{4},E(L))$ and the excision theorem, we see that $H_{2}(E(L))=0$. We can thus easily check the claim using the homology exact sequence for the pair $(S^{4}(L),E(L))$ and the excision theorem.
We next prove the claims (2) and (3). In the case where $Z=S^{4}$, the claim (2) is trivial. So, we assume $\partial Z$ is non-empty. Since $Z-\mathrm{int}\,D^{4}$ is an invertible cobordism from $S^{3}$ to $\partial Z$, and $H_{2}(Z-\mathrm{int}\,D^{4})$ is isomorphic to $H_{2}(Z)$, the claims (2) and (3) follow from Lemma 6.16. ∎
8.2. Behavior of genus functions
Now we discuss genus functions. Throughout this subsection, let $X$ be a ((((possibly closed)))) connected 4-manifold, let $L$ be an $S^{2}$-link in $S^{4}$, and let $Z$ be a connected summand or a codimension-00 connected submanifold of the surgered 4-manifold $S^{4}(L)$. We show that genus functions are preserved under connected sums and boundary sums with $Z$.
Theorem 8.4.
The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the genus functions of $X$ and $X\#Z$. Furthermore, in the case where $\partial X$ and $\partial Z$ are non-empty, the inclusion $X\hookrightarrow X\natural Z$ also preserves the genus functions.
For the first statement, we identify $H_{2}(X\#Z)$ with $H_{2}(X)\oplus H_{2}(Z)$ by using the isomorphisms $H_{2}(X-\mathrm{int}\,D^{4})\to H_{2}(X)$ and $H_{2}(Z-\mathrm{int}\,D^{4})\to H_{2}(Z)$ induced by the inclusions.
Proof of Theorem 8.4.
Since $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is diffeomorphic to $Z\#(I\times S^{3})$, Lemma 6.7 shows that $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is a (strongly) quasi-invertible cobordism from $S^{3}$ to the disjoint union $S^{3}\sqcup\partial Z$. The 4-manifold $X\#(Z-\mathrm{int}\,D^{4})$ is constructed from $X-\mathrm{int}\,D^{4}$ by attaching this cobordism, so the inclusion $X-\mathrm{int}\,D^{4}\hookrightarrow X\#(Z-\mathrm{int}\,D^{4})$ preserves the genus functions due to Theorem 7.1. We note that the inclusions $X-\mathrm{int}\,D^{4}\hookrightarrow X$ and $Z-\mathrm{int}\,D^{4}\hookrightarrow Z$ preserve the genus functions by Lemma 7.2. Therefore, the inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)$ preserves the genus functions.
Now we assume $\partial X$ and $\partial Z$ are non-empty. By Lemma 6.7, any boundary sum $(I\times\partial X)\natural Z$ is a (strongly) quasi-invertible cobordism, where the sum is taken along $1\times\partial X$. Since $X\natural Z$ is constructed from $X$ by attaching this cobordism, the inclusion $X\hookrightarrow X\natural Z$ preserves the genus functions due to Theorem 7.1. ∎
In the case where $H_{2}(Z)=0$, the above inclusions in Theorem 8.4 are isomorphisms, and hence $g_{X\#Z}$ and $g_{X}$ are completely determined by $g_{X}$. In the case where $H_{2}(Z)\neq 0$, we give the following estimates.
Proposition 8.5.
Every $\alpha\in H_{2}(X)$ and $\beta\in H_{2}(Z)$ satisfy
$g_{X}(\alpha)\leq g_{X\#Z}(\alpha+\beta)\leq g_{X}(\alpha)+g_{Z}(\beta).$
Furthermore, in the case where $\partial X$ and $\partial Z$ are non-empty, they also satisfy
$g_{X}(\alpha)\leq g_{X\natural Z}(\alpha+\beta)\leq g_{X}(\alpha)+g_{Z}(\beta).$
Proof.
Due to the constructions of $X\#Z$ and $X\natural Z$ in the proof of Theorem 8.4, the claim follows from Proposition 7.6. ∎
Due to the above proposition, torsion-free genus functions are preserved under connected sums and boundary sums with $Z$.
Corollary 8.6.
The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the torsion-free genus functions of $X$ and $X\#Z$. Furthermore, in the case where both $\partial X$ and $\partial Z$ are non-empty, the inclusion $X\hookrightarrow X\natural Z$ also preserves the torsion-free genus functions.
Remark 8.7.
The inclusions $H_{2}(Z)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ and $Z\hookrightarrow X\natural Z$ do not necessarily preserve the genus functions. For example, in the case where $X$ is $S^{2}\times S^{2}$ or $S^{2}\times S^{2}-\mathrm{int}\,D^{4}$, we have many counterexamples of $Z$ as easily seen from results of [10], [11]. See also Remark 8.12.
8.3. Stabilities for algebraic inequivalences of genus functions
Now we discuss stabilities for algebraic inequivalences of genus functions. In the rest of this section, for each $i=1,2$, let $X_{i}$ be a ((((possibly closed)))) connected 4-manifold, let $L_{i}$ be an $S^{2}$-link in $S^{4}$, and let $Z_{i}$ be a connected summand or a codimension-00 connected submanifold of $S^{4}(L_{i})$.
We show that algebraic inequivalences of genus functions are stable under connected sums and boundary sums in the case where $H_{2}(Z_{i})=0$. The theorem below is a restatement of Theorem 1.9.
Theorem 8.8.
Suppose each $H_{2}(Z_{i})=0$. Then the following hold.
$(1)$ The genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
$(2)$ Assume each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty. Then, the genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
Proof.
Under the assumptions, the genus functions of each $X_{i}\#Z_{i}$ and $X_{i}\natural Z_{i}$ are algebraically equivalent to those of $X_{i}$ due to Theorem 8.4. The claims thus immediately follow. ∎
The corresponding results hold for torsion-free genus functions as well.
Theorem 8.9.
Suppose each $b_{2}(Z_{i})=0$. Then the following hold.
$(1)$ The torsion-free genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
$(2)$ Assume each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty. Then, the torsion-free genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
Proof.
Due to Corollary 8.6, the claims follow from the same argument as in the proof of Theorem 8.8. ∎
In the case where $b_{2}(Z_{i})\neq 0$, similar stabilities still hold under additional assumptions. The theorem below is a refinement of Theorem 1.10. See Lemma 8.3 for sufficient conditions that each $H_{2}(Z_{i})$ is torsion-free.
Theorem 8.10.
Assume that the intersection form of each $X_{i}$ is non-degenerate and that $H_{2}(X_{i})$’s or $H_{2}(Z_{i})$’s are torsion-free. Then the following hold.
$(1)$ Suppose that the genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent. Then, those of $X_{1}$ and $X_{2}$ are algebraically equivalent. If furthermore $H_{2}(X_{i})$’s are torsion-free, then those of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ restricted to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
$(2)$ Suppose that each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty and that the genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent. Then, those of $X_{1}$ and $X_{2}$ are algebraically equivalent. If furthermore $H_{2}(X_{i})$’s are torsion-free, then those of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ restricted to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
Proof.
Since $H_{2}(S^{4}(L_{i}))=0$ for each $i$, the intersection form of each $Z_{i}$ is the zero-map. Due to Propositions 5.3 and 5.2, the claims thus follow from Theorem 8.4 and Proposition 8.5. ∎
For torsion-free genus functions, the stability holds without the torsion-free assumption.
Theorem 8.11.
Assume the intersection form of each $X_{i}$ is non-degenerate. Then the following hold.
$(1)$ If the torsion-free genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
$(2)$ If each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty, and the torsion-free genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
Proof.
Since the intersection form of each $Z_{i}$ is the zero-map, the claims follow from Propositions 5.4 and 8.5 and Corollary 8.6. ∎
Remark 8.12.
Theorems 8.10 and 8.11 do not hold in the case where the intersection forms of $X_{1},X_{2}$ are degenerate by the following trivial reason. Let $X_{1},X_{2}$ be an exotic pair of 4-manifolds with algebraically inequivalent genus functions, and assume they admits embeddings into $S^{4}$. Such examples were produced by Akbulut and the author ([10], [11]), see also Theorem 3.3. Set $Z_{1}=X_{2}$ and $Z_{2}=X_{1}$. In this case $X_{1},X_{2},Z_{1},Z_{2}$ satisfy the assumption of the above theorem except the non-degeneracy assumption, but clearly the conclusions do not hold.
As an application of Theorem 8.8, we show that algebraic equivalent classes of genus functions do not determine homeomorphism types even within 4-manifolds having isomorphic cohomology rings and arbitrarily large $b_{2}$.
Proposition 8.13.
For any closed connected $4$-manifold $X$, there exist infinitely many pairwise non-homeomorphic closed $4$-manifolds with pairwise algebraically equivalent genus functions such that their cohomology rings are isomorphic to those of $X\#S^{1}\times S^{3}$.
Proof.
Let $K_{1},K_{2},\dots$ be infinitely many $S^{2}$-knots in $S^{4}$ such that their complements $S^{4}-K_{1},S^{4}-K_{2},\dots$ have pairwise non-isomorphic fundamental groups. As is well-known, such families of $S^{2}$-knots exist in abundance and easily produced, for example, as follows. Let $k$ be a knot in $S^{3}$, and let $K$ be the $S^{2}$-knot in $S^{4}$ derived from $k$ by Artin’s spinning construction ([13]). Then $\pi_{1}(S^{4}-K)$ is isomorphic to $\pi_{1}(S^{3}-k)$ (see e.g. [18, Section 2.1]), and thus we obtain a desired infinite family of $S^{2}$-knots by varying a knot $k$. By the Seifert van-Kampen theorem, we see that each $\pi_{1}(S^{4}(K_{i}))$ is isomorphic to $\pi_{1}(S^{4}-K_{i})$. It thus easily follows from Lemma 8.3 that the homology groups of $S^{4}(K_{i})$ are isomorphic to those of $S^{1}\times S^{3}$. Furthermore, we can easily show that the cohomology ring of $S^{4}(K_{i})$ is isomorphic to that of $S^{1}\times S^{3}$ (cf. [17, Section 11]).
Now let $X$ be a closed 4-manifold, and let $Y_{i}$ be the connected sum $X\#S^{4}(K_{i})$. Note that $\pi_{1}(Y_{i})$ is isomorphic to the free product $\pi_{1}(X)*\pi_{1}(S^{4}(K_{i}))$. It thus follows from the Grushko decomposition theorem (see e.g. [60]) that $\pi_{1}(Y_{i})$’s are pairwise non-isomorphic, showing that $Y_{i}$’s are pairwise non-homeomorphic. On the other hand, it is clear that their homology groups, cohomology rings and intersection forms are isomorphic to those of $X\#S^{1}\times S^{3}$. Furthermore, by Theorem 8.4, their genus functions are algebraically equivalent to that of $X$ and hence to that of $X\#S^{1}\times S^{3}$. The claim thus follows. ∎
9. Exotic 4-manifolds and genus functions
In this section, we produce various families of exotic 4-manifolds with respect to genus functions. Note that we follow Convention 1.1.
We first show that any $4$-dimensional $2$-handlebody can be modified into infinite order generalized corks. The theorem below is a refinement of Theorem 4.1.
Theorem 9.1.
For any $4$-dimensional $2$-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ topologically resembles $X$, admits an embedding into $X$, and is homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$. Furthermore, in the case where $H_{2}(X)$ is a non-trivial group, there exist arbitrarily many generalized corks $(X_{i},f_{i})$’s of infinite order such that $X_{i}$’s have the properties of $X^{\prime}$ and are pairwise exotic, and that the genus functions of $X_{i}$’s are pairwise algebraically inequivalent for any orientations.
Proof.
Let $X$ be a 4-dimensional 2-handlebody. By Proposition 3.3, we have a 4-manifold $Y$ admitting a Stein structure such that $Y$ topologically resembles $X$, admits an embedding into $X$, and is homeomorphic to a 4-manifold admitting a deformation retraction onto $X$. Therefore, by Proposition 4.2, we obtain an infinite order generalized cork $(X^{\prime},f)$ having the required properties.
Now we assume $H_{2}(X)$ is non-trivial. Then, by Theorem 3.4, we obtain arbitrarily many pairwise exotic 4-manifolds $Y_{i}$’s such that each $Y_{i}$ have the above properties of $Y$ and that the genus functions of $Y_{i}$’s are pairwise algebraically inequivalent for any orientations. Therefore, by the construction in the proof of Proposition 4.2, we obtain arbitrarily many generalized corks $(X_{i},f_{i})$’s of infinite order such that each $X_{i}$ have the above properties of $X^{\prime}$ and that $X_{i}$’s are pairwise homeomorphic. Since each $X_{i}$ is obtained by attaching an invertible homology cobordism to $Y_{i}$, the genus function of each $X_{i}$ is algebraically equivalent to that of $Y_{i}$ due to Corollary 7.5, and thus $X_{i}$’s are pairwise exotic. This completes the proof. ∎
Remark 9.2.
As seen from the construction, the self-homeomorphism $f$ of $\partial X^{\prime}$ extends to a self-homeomorphism of $X^{\prime}$ that induces the identity on $H_{2}(X^{\prime})$, and the same claim holds for the self-homeomorphism $f_{i}$ of each $\partial X_{i}$.
We here give a method for producing exotic 4-manifolds with equivalent genus functions from generalized corks.
Proposition 9.3.
Let $\{(C,f)\}_{f\in A}$ be a family of pairwise inequivalent generalized corks with $\partial C=M$, and assume that each $f\in A$ extends to a self-homeomorphism of $C$ that preserves the genus function. Then there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M}\}}$ of pairwise exotic $4$-manifolds topologically resembling $C$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $C$. Furthermore, they admit embeddings into $C$ and deformation retractions onto $C$.
Proof.
By Corollary 3.10, we have a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M}\}}$ of pairwise exotic 4-manifolds having all the desired properties except those on genus functions. We note that $C_{\mathrm{id}_{M}}$ is obtained from $C$ by attaching an invertible cobordism $\mathbb{P}$, and each $C_{f}$ is obtained from $C_{\mathrm{id}_{M}}$ by removing $C$ and gluing it back using $f$. By Corollary 7.5, the genus function of each $C_{f}$ is algebraically equivalent to that of $C$, and this equivalence is induced by the inclusion $C\hookrightarrow C_{f}$. Since each $f\in A$ preserves the orientation and extends to a self-homeomorphism of $C$ that preserves the genus function, this self-homeomorphism of $C$ extends to an orientation-preserving homeomorphism $C_{\mathrm{id}_{M}}\to C_{f}$ that preserves the genus functions. ∎
Now we will prove results on exotic 4-manifolds given in Section 1. The theorem below is a refinement of Theorem 1.3.
Theorem 9.4.
For any $4$-dimensional $2$-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise orientedly equivalent. Furthermore, they admit embeddings into $X$ and are homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
Proof.
This is straightforward from Theorem 9.1, Remark 9.2, and Proposition 9.3. ∎
The theorem below is a refinement of Theorem 1.5.
Theorem 9.5.
Suppose that a $4$-manifold $X$ admits an embedding into a symplectic $4$-manifold with weakly convex boundary. Then, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$. Furthermore, they admit embeddings into $X$ and are homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
Proof.
This is straightforward from Propositions 4.2 and 9.3. ∎
The corollary below is a slight refinement of Corollary 1.6.
Corollary 9.6.
$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((((with boundary)))) such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
$(2)$ If $X$ is the total space of a $D^{2}$-bundle over a closed surface, then there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$.
Proof.
(1) Since $D^{4}$ is a Stein manifold, its blow-ups $n\overline{\mathbb{C}P^{2}}-\textnormal{int}D^{4}$ is a symplectic $4$-manifold with weakly convex boundary. Also, the genus function of $n\overline{\mathbb{C}P^{2}}-\textnormal{int}D^{4}$ is equivalent to that of $n\overline{\mathbb{C}P^{2}}$ due to Lemma 7.2. The claim (1) thus follows from Theorem 9.5.
(2) Let $X$ denote the total space of $D^{2}$-bundle over a closed surface of genus $g$ with Euler number $n$. Then $X$ admits a Stein structure for $n\leq 2g-2$ (see [38, Exercises 11.2.5.(a)]). Hence the claim (2) follows from Theorem 9.5 in the case where $n\leq 2g-2$. In the case where $n\geq 2g-1$ and $g\geq 1$, the claim (2) follows from the case where $n\leq 2g-2$ by reversing the orientation. The same argument also shows the claim (2) in the case where $n\geq 2$ and $g=0$. The remaining cases are $(g,n)=(0,0),(0,\pm 1)$. In the case where $(g,n)=(0,\pm 1)$, the claim (2) follows from (1). In the case where $(g,n)=(0,0)$, the 4-manifold $X$ is diffeomorphic to $S^{2}\times D^{2}$ and hence admits an embedding into $D^{4}$. The claim (2) thus follows from Theorem 9.5 in this case. This completes the proof. ∎
The theorem below is a slight refinement of Theorem 1.7.
Theorem 9.7.
There exist infinitely many pairwise exotic simply connected $4$-manifolds such that their genus functions are pairwise algebraically inequivalent for any orientations and that each of their genus functions is orientedly equivalent to those of another infinitely many pairwise exotic $4$-manifolds.
Proof.
Let $\{X_{i}\}_{i\in\mathbb{Z}}$ be an infinite family of pairwise exotic 4-manifolds admitting Stein structures such that the genus functions of the members are algebraically inequivalent for any orientations. Such examples were obtained by Akbulut and the author in [12] and the author in [68]. By Proposition 4.2, we obtain an infinite family $\{(Y_{i},f_{i})\}_{i\in\mathbb{Z}}$ of infinite order generalized corks such that each $Y_{i}$ is obtained from $X_{i}$ by attaching the same invertible homology cobordism, and that $Y_{i}$’s are pairwise homeomorphic. By Corollary 7.5, the genus function of each $Y_{i}$ is algebraically equivalent to that of $X_{i}$. Due to the proof of Proposition 9.3, by attaching another fixed invertible homology cobordism to each $Y_{i}$ and twisting $Y_{i}$ in the resulting 4-manifold $X^{\prime}_{i,0}$ using each $f_{i}^{j}$, we obtain an infinite family $\{X^{\prime}_{i,j}\}_{(i,j)\in\mathbb{Z}\times\mathbb{Z}}$ of pairwise exotic simply connected 4-manifolds. By Proposition 9.3, for each fixed $i$, the genus functions of $X^{\prime}_{i,j}$’s are pairwise equivalent, and furthermore for each fixed $j$, the genus functions of $X^{\prime}_{i,j}$’s are pairwise algebraically inequivalent for any orientations. ∎
By the proof of the above theorem, we can construct similar infinite families for varieties of homeomorphism types by using exotic Stein manifolds of the author [68] as a building block. For example, any finitely presented group is realized as the fundamental group of infinitely many exotic 4-manifolds satisfying the properties in the above theorem.
Here we construct similar examples for much more homeomorphism types, by relaxing the condition “infinitely many” to “arbitrarily many”. The theorem below is a refinement of Theorem 1.8.
Theorem 9.8.
For any $4$-dimensional $2$-handlebody $X$ with non-trivial second homology group, there exist arbitrarily many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise algebraically inequivalent for any orientations and that each of their genus functions is orientedly equivalent to those of infinitely many pairwise exotic $4$-manifolds. Furthermore, all of them admit embeddings into $X$ and are homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
Proof.
Using infinite order generalized corks of Theorem 9.1 as a building block, we can prove the claim similarly to the proof of Theorem 9.7. ∎
As an application of our results on genus functions, we give exotic 4-manifolds which are stably exotic under boundary sums with arbitrary Stein manifolds having non-degenerate intersection forms. The theorem below is a restatement of Theorem 1.11.
Theorem 9.9.
For any positive integer $n$, there exist pairwise exotic simply connected $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ admitting Stein structures such that, for any pairwise orientedly homeomorphic $4$-manifolds $Y_{1},Y_{2},\dots,Y_{n}$ having non-degenerate intersection forms and admitting Stein structures, the boundary sums $X_{1}\natural Y_{1}$, $X_{2}\natural Y_{2}$, $\dots$, $X_{n}\natural Y_{n}$ are pairwise exotic.
Proof.
Let $X$ be a 4-dimensional 2-handlebody with non-trivial second homology group such that its intersection form is represented by a zero-matrix, and let $Y_{1},Y_{2},\dots,Y_{n}$ be 4-manifolds having non-degenerate intersection forms and admitting Stein structures. By Theorem 3.4, we have pairwise exotic $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ admitting Stein structures such that they topologically resemble $X$ and that their genus functions are pairwise algebraically inequivalent for any orientations. Since each $X_{i}$ and $Y_{i}$ admit Stein structures, we can apply the argument on estimating genus functions in the proof of Theorem 3.4, and thus we see that the restrictions of the genus functions of $X_{i}\natural Y_{i}$’s to $H_{2}(X_{i})$’s remain pairwise algebraically inequivalent. Note that the second homology group of any 2-handlebody is torsion-free. It thus follows from Proposition 5.2 that the genus functions of $X_{i}\natural Y_{i}$’s are pairwise inequivalent. ∎
As seen from the above proof, we can realize various homeomorphism types as those of $X_{i}$’s.
10. Exotically knotted submanifolds
As an application of our exotic 4-manifolds in the last section, we modify a submanifold of a 4-manifold with codimension less than two into infinitely many pairwise exotically knotted submanifolds.
For a codimension-00 submanifold $X$ of a 4-manifold $Z$, the closure of $Z-X$ will be called the exterior of $X$ and denoted by $E(X)$. To state our results shortly, we use the following terminologies.
Definition 10.1.
Let $X_{1},X_{2}$ be (possibly non-orientable) submanifolds of a (possibly non-orientable) 4-manifold $Z$.
$(1)$ The pair $X_{1},X_{2}$ will be called exotically knotted if they have the following properties.
  • (i)
    $X_{1}$ is diffeomorphic to $X_{2}$. In the case where $Z$ is orientable, and their codimensions are zero, the diffeomorphism is assumed to preserve the orientations induced from $Z$.
  • (ii)
    $X_{1}$ is topologically ambient isotopic to $X_{2}$. In the case where $\partial Z$ is non-empty, the topological isotopy is assumed to fix $\partial Z$ pointwise.
  • (iii)
    $X_{1}$ is not smoothly isotopic to $X_{2}$. (We do not assume a smooth isotopy fixes $\partial Z$ pointwise.)
$(2)$ The pair $X_{1},X_{2}$ will be called exotically embedded if they have the following properties.
  • (i)
    The condition (i) of (1) holds.
  • (ii)
    There exists a self-homeomorphism of $Z$ that maps $X_{1}$ to $X_{2}$. In the case where $Z$ is orientable, the self-homeomorphism is assumed to be orientation-preserving.
  • (iii)
    There exists no self-diffeomorphism of $Z$ that maps $X_{1}$ to $X_{2}$.
We produce exotically knotted submanifolds of codimension-0. The theorem below is a refinement of Theorem 1.12.
Theorem 10.2.
Let $X$ be a codimension-00 submanifold of a $4$-manifold $Z$. These manifolds are possibly non-orientable and may have ((((possibly disconnected)))) boundaries. Suppose the exterior of $X$ is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-00 submanifolds topologically resembling $X$ such that each of these $4$-manifolds contains $X$ as a submanifold and deformation retracts to $X$.
To prove this theorem, we give a method for producing exotically knotted submanifolds by using generalized corks obtained from contractible corks. We use the following setting only for the lemma and the proposition below. Let $\{(C_{0},f)\}_{f\in A}$ be a family of contractible corks with connected boundary having a fixed embedded 3-ball $B_{0}$ in $\partial C_{0}$ such that every $f\in A$ is identity on a neighborhood of the 3-ball $B_{0}$. Let $Y$ be a 4-manifold with $\partial Y=M$. Take the boundary sum $(I\times M)\natural C_{0}$ for the boundary component $1\times M$ using the 3-ball $B_{0}$ of $\partial C_{0}$. This manifold $P$ is a homology cobordism from $M$ to $M\#\partial C_{0}$. We set $M^{\prime}=M\#\partial C_{0}$. Abusing the notation, let $f$ denote the self-diffeomorphism of $M^{\prime}$ that extends the self-diffeomorphism $f$ of $\partial C_{0}$ as the identity on the rest of $M^{\prime}$. We denote the glued 4-manifold $Y\cup_{\mathrm{id}_{M}}P$ by $C$, and assume $\{(C,f)\}_{f\in A}$ is a family of pairwise inequivalent generalized corks. We furthermore assume that, for each $f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}$, the twisted double $C_{0}\cup_{f}\overline{C_{0}}$ of $C_{0}$ is diffeomorphic to $S^{4}$.
We identify the double of $I\times M$ with $[0,2]\times M$. Then, for each $f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}$, the twisted double $P\cup_{f}\overline{P}$ can be regarded as the connected sum of $[0,2]\times M$ and the twisted double $C_{0}\cup_{f}\overline{C_{0}}$. Since $f$ is identity on a neighborhood of the 3-ball $B_{0}$ in $\partial C_{0}$, it follows that the double $P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ has an embedded 4-ball $D_{P}$ such that every twisted double $P\cup_{f}\overline{P}$ is obtained from $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{P}$ by attaching a “twisted partial double” of $C_{0}$, which is obtained from $C_{0}$ by attaching $\overline{C_{0}}$ using the restriction of $f$ to the exterior of $B_{0}$. We here observe the following lemma.
Lemma 10.3.
$(1)$ $P$ admits a deformation retraction onto the boundary component $M$.
$(2)$ For each $f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}$, there exists a diffeomorphism $P\cup_{f}\overline{P}\to I\times M$ such that it fixes the boundary $M\sqcup\overline{M}$ pointwise and that its restriction to the submanifold $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{P}$ is independent of $f$.
Proof.
The claim $(1)$ is straightforward from the construction of $P$. Since each twisted double $C_{0}\cup_{f}\overline{C_{0}}$ is diffeomorphic to $S^{4}$, the disk theorem shows that the twisted partial double of $C_{0}$ is diffeomorphic to $D^{4}$. The claim (2) thus follows from the decomposition of the twisted double of $P$ in the paragraph preceding this lemma. ∎
Proposition 10.4.
Let $X$ be a codimension-00 submanifold of a $4$-manifold $Z$. These manifolds are possibly non-orientable and may have ((((possibly disconnected)))) boundaries. If the exterior $E(X)$ is diffeomorphic to $Y$, then $Z$ admits a family $\{X_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}}$ of pairwise exotically knotted codimension-00 submanifolds topologically resembling $X$ such that each $X_{f}$ contains $X$ as a submanifold and admits a deformation retraction onto $X$.
Proof.
We identify the boundary $\partial E(X)$ with $M$ using a fixed diffeomorphism $Y\to E(X)$. Note that $Y$ is oriented due to our convention. Let $\mathbb{P}$ be the invertible homology cobordism from $M^{\prime}$ to some closed (connected) 3-manifold $N$ used in Theorem 3.6. We denote the 4-manifold $C\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}$ by $S_{\mathrm{id}_{M^{\prime}}}$. We attach the upside down cobordism $\overline{\mathbb{P}}$ to $S_{\mathrm{id}_{M^{\prime}}}$ along $N$ using $\mathrm{id}_{N}$, and then attach the upside down cobordism $\overline{P}$ of $P$ along $M^{\prime}$ using $\mathrm{id}_{M^{\prime}}$. We denote the resulting 4-manifold $Y\cup_{\mathrm{id}_{M}}P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ by $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$. Since the double $\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}$ admits a diffeomorphism to $I\times M^{\prime}$ that fixes the boundary $\overline{M^{\prime}}\sqcup M^{\prime}$ pointwise (see Remark 3.9), the submanifold $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ admits a diffeomorphism to the double $P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ that fixes the boundary $\overline{M}\sqcup M$ pointwise. Let $\widehat{D}_{P}$ denote the preimage of $D_{P}$ under this diffeomorphism. By Lemma 10.3, the 4-manifold $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ is diffeomorphic to $Y$ and hence to the exterior $E(X)$, and this diffeomorphism $\varphi_{\mathrm{id}_{M^{\prime}}}:\widehat{Y}_{\mathrm{id}_{M^{\prime}}}\to E(X)$ fixes the boundary $M$ pointwise. We set $D_{P}^{0}=\varphi_{\mathrm{id}_{M^{\prime}}}(\widehat{D}_{P})$.
For each $f\in A$, we apply generalized cork twists to $S_{\mathrm{id}_{M^{\prime}}}$ and $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ along $(C,f)$, and denote the resulting 4-manifolds $C\cup_{f}\mathbb{P}$ and $C\cup_{f}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ by $S_{f}$ and $\widehat{Y}_{f}$, respectively. By the diffeomorphism $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}\to P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ in the last paragraph, each cobordism $P\cup_{f}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ is obtained from the exterior of $\widehat{D}_{P}$ in $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ by attaching a homotopy 4-ball diffeomorphic to the partial twisted double of $C_{0}$. Thus, by Lemma 10.3, we have a diffeomorphism $P\cup_{f}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}\to P\cup_{{\mathrm{id}_{M^{\prime}}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ that fixes the exterior of $\widehat{D}_{P}$ and hence the boundary $\overline{M}\sqcup M$ pointwise. Extending this diffeomorphism trivially, we obtain a diffeomorphism $\psi_{f}:\widehat{Y}_{f}\to\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that is identity on $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{P}$. Let $\varphi_{f}:\widehat{Y}_{f}\to E(X)$ denote the diffeomorphism defined by $\varphi_{f}=\varphi_{\mathrm{id}_{M^{\prime}}}\circ\psi_{f}$.
For each $f\in A\sqcup\{{\mathrm{id}_{M^{\prime}}}\}$, we denote the submanifolds $\varphi_{f}(S_{f})$, $\varphi_{f}(\overline{\mathbb{P}})$ and $\varphi_{f}(\overline{P})$ of $E(X)$ respectively by $S_{f}^{0}$, ¯$\overline{\mathbb{P}}^{0}_{f}$ and $\overline{P}^{0}_{f}$. Let $X_{f}$ be the submanifold ¯$\overline{\mathbb{P}}^{0}_{f}\cup\overline{P}^{0}_{f}\cup X$ of $Z$. Then, $X_{f}$’s are pairwise diffeomorphic since each $\varphi_{f}$ fixes the boundary $M$ pointwise. Note that they are pairwise orientedly diffeomorphic in the case where $Z$ is orientable. Clearly, the submanifold $\overline{\mathbb{P}^{0}_{f}}\cup\overline{P}^{0}_{f}$ of $Z$ is an invertible homology cobordism from $\overline{M}$ to $\overline{N}$, and each $X_{f}$ is obtained from $X$ by attaching this cobordism along $\overline{M}$. We thus see that each $X_{f}$ deformation retracts to $X$ and topologically resembles $X$ by Lemmas 3.8 and 10.3.
We will show that the pairs $X_{f}$’s are pairwise exotically knotted and exotically embedded. We note that $Z=S_{f}^{0}\cup X_{f}$ for each $f\in A\sqcup\{{\mathrm{id}_{M^{\prime}}}\}$. By Theorem 3.6, the exteriors of $X_{f}$’s are pairwise exotic, and thus the pairs $(Z,X_{f})$’s are pairwise non-diffeomorphic. Consequently, $X_{f}$’s are pairwise smoothly non-isotopic.
In the rest, we will prove that $X_{f}$’s are pairwise topologically ambient isotopic by showing that each $X_{f}$ is topologically ambient isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$. Due to the assumption, each $f^{-1}\in A$ extends to a self-homeomorphism of $C$ that is identity on the exterior of the contractible 4-manifold $C_{0}$. Extending this homeomorphism trivially, we get a homeomorphism $\widehat{Y}_{f}\to\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that maps $S_{f}$ to $S_{\mathrm{id}_{M^{\prime}}}$. Note that this homeomorphism is identity on the exterior $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{P}$. Composing this homeomorphism with the diffeomorphism $\psi_{f}$, we obtain a self-homeomorphism of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that fixes the $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{P}$ and hence the boundary $M$ pointwise. Therefore, by composing this self-homeomorphism of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ with the diffeomorphisms $\varphi_{\mathrm{id}_{M^{\prime}}}$ and $\varphi_{\mathrm{id}_{M^{\prime}}}^{-1}$, we obtain a self-homeomorphism of $E(X)$ that maps $S_{f}^{0}$ to $S_{{\mathrm{id}_{M^{\prime}}}}^{0}$ and fixes $E(X)-\textnormal{int}\,D_{P}^{0}$ and the boundary $M$ pointwise. Extending this self-homeomorphism trivially, we obtain a self-homeomorphism ${\Phi}_{f}$ of $Z$ that maps $X_{f}$ to $X_{{\mathrm{id}_{M^{\prime}}}}$ and fixes the exterior of $D_{P}^{0}$ pointwise. Note that ${\Phi}_{f}$ fixes $\partial Z$ pointwise in the case where $\partial Z$ is non-empty.
The restriction of ${\Phi}_{f}$ to $D_{P}^{0}$ fixes the boundary pointwise. Since $D_{P}^{0}$ is diffeomorphic to $D^{4}$, it thus follows from a result of Orson and Powell [55, Corollary D] that the restriction to $D_{P}^{0}$ is topologically isotopic to the identity map on $D_{P}^{0}$ fixing the boundary pointwise. Thus, ${\Phi}_{f}$ is topologically isotopic to the identity map on $Z$. Note that this isotopy fixes $\partial Z$ pointwise in the case where $\partial Z$ is non-empty. Therefore, each submanifold $X_{f}$ of $Z$ is topologically isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$. This completes the proof. ∎
We can now easily prove Theorem 10.2 using our generalized corks.
Proof of Theorem 10.2.
Let $Y$ be a 4-manifold diffeomorphic to the exterior $E(X)$. We first consider the case where $Y$ is diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Let $(C_{0},\tau)$ be an infinite order contractible cork of Gompf (see Subsection 3.5). By results of Akbulut [5], Tange [64] and Naylor and Schwartz [54], the twisted double $C_{0}\cup_{\tau^{i}}\overline{C_{0}}$ is diffeomorphic to $S^{4}$ for each integer $i$. Therefore, we can apply Proposition 10.4 using a generalized cork obtained from $Y$ by Proposition 4.2. The claim thus follows in this case.
In the case where $Y$ is a 2-handlebody, we first modify $Y$ into a Stein handlebody $Y^{\prime}$ by using Proposition 3.3. Note that $Y$ is obtained from $Y^{\prime}$ by attaching the upside down cobordism of an invertible cobordism. Furthermore, $\partial Y$ is a deformation retract of this invertible homology cobordism. We attach this invertible cobordism to $X$ and denote the resulting 4-manifold by $X^{\prime}$. We can regard $X^{\prime}$ as a submanifold of $Z$ containing $X$ as a submanifold. Clearly, $X$ is a deformation retract of $X^{\prime}$. Therefore, we can prove the claim by applying the argument of the first case to the submanifold $X^{\prime}$. ∎
Remark 10.5.
As shown in the proof, each $X_{i}$ in Theorem 10.2 is obtained from $X$ by attaching an invertible homology cobordism, and thus the genus function of $X_{i}$ is algebraically equivalent to that of $X$.
We next discuss codimension-1 submanifolds. We use the following terminology.
Definition 10.6.
Let $M_{1},M_{2}$ be codimension-1 submanifolds of a (possibly non-orientable) 4-manifold $Z$. The manifold $Z$ is possibly non-orientable and may have ((((possibly disconnected)))) boundary. The submanifold $M_{1}$ will be called homology cobordant to $M_{2}$ through $Z$ if $Z$ admits a codimension-00 submanifold $H$ with $\partial H=\overline{M_{1}}\sqcup M_{2}$ such that $H$ is a homology cobordism from $M_{1}$ to $M_{2}$.
As an application of Theorem 10.2, we modify separating codimension-$1$ submanifolds into infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds. The theorem below is a refinement of Theorem 1.14.
Theorem 10.7.
Let $Z$ be a ((((possibly non-orientable)))) $4$-manifold which may have ((((possibly disconnected)))) boundary, and let $M$ be a closed $3$-manifold embedded in the interior of $Z$ as a separating submanifold. Suppose the exterior of the tubular neighborhood of $M$ has a connected component which is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$ through $Z$.
Proof.
Due to the assumption, we can apply Theorem 10.2 to a connected component $X$ of the exterior of the tubular neighborhood for $M$. Let $X_{i}$’s be the resulting infinitely many pairwise exotically knotted and exotically embedded codimension-0 submanifolds of $Z$, and let each $M_{i}$ be the boundary $\partial E(X_{i})$ of the exterior $E(X_{i})$. By the proof of Theorem 10.2, each $M_{i}$ is homology cobordant to $M$ through $Z$. Since $X_{i}$’s are pairwise topologically ambient isotopic, their boundary components $M_{i}$’s are also pairwise topologically ambient isotopic. By the proof of Theorem 10.2, the exteriors $E(X_{i})$’s are pairwise non-diffeomorphic, and $X_{i}$’s are pairwise diffeomorphic. Note that the exterior of the tubular neighborhood of each $M_{i}$ is the disconnected 4-manifold $E(X_{i})\sqcup X_{i}$. If a pair $(Z,M_{i})$ is diffeomorphic to some $(Z,M_{j})$, then $E(X_{i})\sqcup X_{i}$ is diffeomorphic to $E(X_{j})\sqcup X_{j}$ due to the uniqueness of the tubular neighborhood. However, in the case where $E(X_{i})$ is not homeomorphic to $X_{i}$, it follows that all $(E(X_{i})\sqcup X_{i})$’s are pairwise non-diffeomorphic, and thus the pairs $(Z,M_{i})$’s are pairwise non-diffeomorphic. Therefore, $M_{i}$’s are pairwise exotically embedded and exotically knotted. In the case where $E(X_{i})$ is homeomorphic to $X_{i}$, at least infinitely many of $(E(X_{i})\sqcup X_{i})$’s are pairwise non-diffeomorphic by the pigeonhole principle. Therefore, at least infinitely many of $M_{i}$’s are pairwise exotically embedded and exotically knotted. ∎
The corollaries below are restatements of Corollaries 1.15 and 1.16.
Corollary 10.8.
Suppose a closed $3$-manifold admits an embedding into $S^{4}$. Then every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds homology cobordant to the $3$-manifold.
Proof.
Let $M$ be a closed 3-manifold embedded in $S^{4}$. It follows from the Mayer–Vietoris exact sequence that $M$ is separating. Let $Y$ be a connected component of the exterior of the tubular neighborhood for $M$. Since $M$ is separating, the boundary $\partial Y$ is diffeomorphic to $M$. Also, $Y$ admits an embedding into $D^{4}$ and hence into any symplectic 4-manifold with weakly convex boundary. Now let $Z$ be an arbitrary 4-manifold. Then $Y$ admits an embedding into the interior of $Z$. Let $M$ denote the boundary of $Y$. Then we can apply Theorem 10.7 to this $M$, and thus the claim follows. ∎
Corollary 10.9.
There exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold of them, every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to the $3$-manifold.
Proof.
There are infinitely many closed 3-manifolds admitting embeddings into $S^{4}$ such that their homology cobordism classes are pairwise distinct. For example, $n(S^{1}\times S^{2})$ admits an embedding into $S^{4}$ for each $n\geq 0$. The claim thus follows from Corollary 10.8. ∎
The corollary below is a refinement of Corollary 1.18. Note that every closed 3-manifold embedded in a simply connected closed 4-manifold is separating as easily seen from the Mayer–Vietoris exact sequence.
Corollary 10.10.
For each non-negative even integer $n$, there exists a closed $3$-manifold $M$ and a simply connected closed $4$-manifold $Z$ with $b_{2}=n$ such that $Z$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$ and that $M$ admits no embedding into any closed $4$-manifold with $b_{2}<n$ as a separating submanifold.
Proof.
Let $n$ be a non-negative even integer, and set $r=2/n$. In the $n=0$ case, the claim immediately follows from Corollary 10.9, so we prove the $n>0$ case. Let $K$ be an invertible knot in $S^{3}$ with $|\sigma(K)|\geq 2n-1$, where $\sigma(K)$ denotes the signature of $K$. Let $Y_{1}$ be the 4-manifold obtained from $D^{4}$ by attaching a 2-handle along $K$ with 00-framing, and let $M_{r}$ be the boundary of the $r$-fold boundary sum $Y_{r}$ of $Y_{1}$. Note that $M_{r}$ is the $r$-fold connected sum of $M_{1}$. This 3-manifold $M_{r}$ admits an embedding into the closed 4-manifold $r(S^{2}\times S^{2})$ by taking the double of $Y_{r}$. By Theorem 10.7, we obtain a closed 3-manifold $M^{\prime}_{r}$ homology cobordant to $M_{r}$ such that $r(S^{2}\times S^{2})$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M^{\prime}_{r}$. Note that $r(S^{2}\times S^{2})$ is simply connected and satisfies $b_{2}=n$.
We will show that $M^{\prime}_{r}$ does not admit any embedding into a closed 4-manifold with $b_{2}<n$ as a separating submanifold. Since $M^{\prime}_{r}$ is homology cobordant to $M_{r}$, we have a homology cobordism from $M^{\prime}_{r}$ to $M^{\prime}_{r}$ that contains $M_{r}$ in the interior as a separating submanifold. Now suppose to the contrary that $M^{\prime}_{r}$ is diffeomorphic to a separating submanifold of a closed 4-manifold with $b_{2}<n$. We remove the tubular neighborhood of $M^{\prime}_{r}$ from the closed 4-manifold and glue the homology cobordism from $M^{\prime}_{r}$ to $M^{\prime}_{r}$ along the resulting boundary. As seen from the Mayer–Vietoris exact sequence, the resulting closed 4-manifold satisfies $b_{2}<n$. Clearly, $M_{r}$ is embedded in the resulting closed 4-manifold as a separating submanifold. This contradicts a result of Kawauchi that $M_{r}$ admits no such an embedding (see [43, the proof of Theorem 2.5]). Therefore, the desired claim follows. ∎
11. Diffeomorphism invariants of genus function type
In this section, we introduce a notion of genus function type for diffeomorphism invariants of 4-manifolds and give examples of such invariants which can detect exotic 4-manifolds. Furthermore, we show that an invariant of genus function type yields lower bounds for values of genus functions and shares many properties of genus functions. Throughout this section, we allow boundaries of $4$-manifolds to be disconnected.
11.1. Definitions and examples
To state the notion, we consider the category of 4-manifolds whose morphisms are orientation preserving diffeomorphisms. Note that manifolds are oriented due to our convention. For a fixed ordered set $R$, we also consider the category of $R$-valued maps whose morphisms are maps between the source sets of $R$-valued maps.
Definition 11.1.
Let $R$ be an ordered set. A functor $G$ from the category of 4-manifolds to the category of $R$-valued maps will be called an oriented diffeomorphism invariant of genus function type if the following conditions hold.
  • (i)
    $G$ assigns to each $4$-manifold $X$ a map $G_{X}:H_{2}(X)\to R$, which will be called the $G$ invariant of $X$.
  • (ii)
    Let $X,Y$ be arbitrary 4-manifolds. For any orientation preserving embedding $\varphi:X\to Y$ and any class $\alpha\in H_{2}(X)$, the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ satisfies $G_{X}(\alpha)\geq G_{Y}(\varphi_{*}(\alpha))$.
If furthermore $G$ satisfies the condition (ii) for any embedding $\varphi$ that does not necessarily preserves the orientations, then $G$ will be called a ((((unoriented)))) diffeomorphism invariant of genus function type.
Remark 11.2.
It is straightforward from the condition (ii) that any oriented diffeomorphism $\varphi:X\to Y$ preserves the values of $G$ invariants, that is, the equality $G_{X}=G_{Y}\circ\varphi_{*}$ holds. The above $G$ is an oriented diffeomorphism invariant in this sense. Similarly, if $G$ is an unoriented diffeomorphism invariant, then any unoriented diffeomorphism preserves the values of $G$ invariants.
Clearly, genus functions of 4-manifolds form an unoriented diffeomorphism invariant of genus function type. As in the case of genus functions, we use the following terminologies.
Definition 11.3.
Let $G$ be a (possibly oriented) diffeomorphism invariant of genus function type, and let $X,Y$ be $4$-manifolds.
(1) We will say that a continuous map $\varphi:X\to Y$ preserves $G$ invariants if the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ preserves the values of the $G$ invariants.
(2) The $G$ invariants of $X$ and $Y$ will be called equivalent if there exists a homeomorphism $\varphi:X\to Y$ that preserves the $G$ invariants. If $\varphi$ furthermore preserves the orientations, then the $G$ invariants of $X$ and $Y$ will be called orientedly equivalent.
(3) The $G$ invariants of $X$ and $Y$ will be called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the intersection forms and $G$ invariants.
Due to Remark 11.2, $G$ invariants of orientedly diffeomorphic 4-manifolds are orientedly equivalent. Also, if $G$ invariants of two 4-manifolds are orientedly equivalent, then they are algebraically equivalent. Furthermore, if $G$ is an unoriented diffeomorphism invariant of genus function type, then $G$ invariants of unorientedly diffeomorphic 4-manifolds are equivalent.
We also introduce the following terminologies, extracting properties of genus functions.
Definition 11.4.
Let $G$ be an oriented diffeomorphism invariant of genus function type.
(1) Let $R$ be an additive semigroup and suppose that the $G$ invariant of each 4-manifold is an $R$-valued map. We will say that $G$ is subadditive with respect to gluing if $G$ satisfies the following property for any 4-manifold $X$ with non-empty (possibly disconnected) boundary $M$ and for any cobordism $P$ from $M$ to a (possibly disconnected) closed 3-manifold. Set $X^{\prime}=X\cup_{M}P$, and let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions. Then the inequality $G_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\leq G_{X}(\alpha)+G_{P}(\beta)$ holds for any $\alpha\in H_{2}(X)$ and any $\beta\in H_{2}(P)$.
(2) We will say that $G$ is preserved under attachment of a $3$-handle ((((resp. a $4$-handle)))) if the following property holds. For any 4-manifold $X^{\prime}$ obtained from a 4-manifold $X$ by attaching a 3-handle (resp. a 4-handle), the inclusion $X\hookrightarrow X^{\prime}$ preserves the $G$ invariants.
Genus functions clearly have the above subadditivity and are preserved by 3- and 4-handles due to Lemma 7.2. We remark that all the properties of genus functions used for the proofs of the results in Sections 7 and 8 are only those in Definitions 11.1 and 11.4. We also define “torsion-free $G$ invariants” and related terminologies in the same way as the torsion-free genus functions.
The genus functions have a property that $g_{X}(0)=g_{D^{4}}(0)$ for any 4-manifold $X$. This property gives the following simple obstructions for having the properties on 3- and 4-handles. As we will see in Example 11.6, these obstructions are effective.
Lemma 11.5.
$(1)$ If $G$ is preserved under attachment of a $4$-handle, then $G_{D^{4}}(0)=G_{S^{4}}(0)$.
$(2)$ If $G$ is preserved under attachment of a $3$-handle, then every $\alpha\in H_{2}(S^{2}\times D^{2})$ satisfies $G_{S^{2}\times D^{2}}(\alpha)=G_{D^{4}}(0)=G_{S^{1}\times S^{3}-\textnormal{int}D^{4}}(0)$.
Proof.
The claim (1) and the equality $G_{S^{2}\times D^{2}}(\alpha)=G_{D^{4}}(0)$ are straightforward. Since attachment of a 1-handle preserves $G$ invariants due to Remark 11.11, the equality $G_{D^{4}}(0)=G_{S^{1}\times S^{3}-\textnormal{int}D^{4}}(0)$ follows. ∎
Other than genus functions, we give the following examples of genus function type which can detect exotic 4-manifolds.
Example 11.6.
(1) For a 4-manifold $X$, we define the map $G_{X}:H_{2}(X)\to\mathbb{Z}\cup\{-\infty\}$ as follows. Suppose $X$ admits an orientation preserving embedding $\varphi:X\to Z$ into a closed 4-manifold $Z$ with $b_{2}^{+}>1$ having a non-vanishing Seiberg–Witten invariant. We furthermore assume $Z$ is of simple type in the sense of [56]. Let $K\in H^{2}(Z)$ be a Seiberg-Witten basic class of $Z$. Then, for $\alpha\in H_{2}(X)$, we define $G_{X}(\alpha)$ as the maximal value of the integer $|\langle K,\varphi_{*}(\alpha)\rangle|$ for all $Z$, $\varphi$ and $K$. Note that $G_{X}(\alpha)<\infty$ for any $\alpha$ due to the adjunction inequality ([46], [56]). In the case where $X$ admits no embedding into such a closed 4-manifold $Z$, we set $G_{X}(\alpha)=-\infty$ for any $\alpha\in H_{2}(X)$. It is easy to see that $G_{X}$’s form an oriented diffeomorphism invariant $G$ of genus function type. This invariant can detect various families of orientedly exotic 4-manifolds, as implicitly shown in [11], [12], [67], [68] and [70]. We can define several variations of the above invariant $G$ by using the Seiberg–Witten invariant and the Bauer–Furuta invariant. These invariants will be further discussed in [72].
The above $G$ clearly has the subadditivity with respect to gluing, but it is preserved under an attachment of neither 3- nor 4-handles. The claim for a 4-handle follows from the equalities $G_{D^{4}}(0)=0$ and $G_{S^{4}}(0)=-\infty$ due to Lemma 11.5. Similarly, the claim for a 3-handle follows from the equalities $G_{D^{4}}(0)=0$ and $G_{S^{1}\times S^{3}-\textnormal{int}D^{4}}(0)=-\infty$ due to Lemma 11.5. Note that for any closed 4-manifold $Z$ with $b_{2}^{+}>1$, the Seiberg–Witten invariant of $Z\#(S^{1}\times S^{3})$ vanishes, since the parities of $b_{2}^{+}-b_{1}$ of these two 4-manifolds are distinct.
(2) Recently, Ren and Willis [58] introduced an interesting new invariant of 4-manifolds called the lasagna $s$-invariant by using skein lasagna modules. The lasagna $s$-invariant (with empty boundary link) of a 4-manifold $X$ is a map $s_{X}:H_{2}(X)\to\mathbb{Z}\cup\{-\infty\}$. According to their result [58, Proposition 1.11], this invariant is an oriented diffeomorphism invariant of genus function type. Furthermore, they showed that this invariant can detect many orientedly exotic pairs of 4-manifolds.
(3) Let $G$ be an oriented diffeomorphism invariant of genus function type. We can modify $G$ into an unoriented invariant, for example, as follows. For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, we set $G^{\prime}_{X}(\alpha)=\max\{G_{X}(\alpha),G_{\overline{X}}(\alpha)\}$. The resulting $G^{\prime}$ is an unoriented diffeomorphism invariant of genus function type. For example, in the case of the invariant defined in (1), the unoriented version can distinguish various families of unorientedly exotic 4-manifolds as implicitly shown in [11] and [68].
11.2. Properties
Now we will show that invariants of genus function type share many properties of genus functions. Throughout this subsection, let $G$ be an oriented diffeomorphism invariant of genus function type, and assume that the invariant $G_{X}$ of each 4-manifold $X$ is an $R$-valued map.
We first show that $G$ yields a function $A_{G}:R\times\mathbb{Z}\to\mathbb{Z}\cup\{\infty\}$ that gives lower bounds for values of genus functions. For integers $g,n$ with $g\geq 0$, let $S(g,n)$ be the total space of the $D^{2}$-bundle over the closed surface of genus $g$ with Euler number $n$, and let $\gamma(g,n)$ denote a generator of $H_{2}(S(g,n))(\cong\mathbb{Z})$. For $(r,n)\in R\times\mathbb{Z}$, we define an element $A_{G}(r,n)$ of ${}\mathbb{Z}\cup\{\infty\}$ as follows. In the case where $G_{S(g,n)}(\gamma(g,n))<r<G_{S(g+1,n)}(\gamma(g+1,n))$ for some integer $g\geq 0$, we set $A_{G}(r,n)=g+1$. In the case where $r=G_{S(g,n)}(\gamma(g,n))$ for some integer $g\geq 0$, we set
$A_{G}(r,n)=\min\{g\in\mathbb{Z}_{\geq 0}\mid r=G_{S(g,n)}(\gamma(g,n))\}.$
In the case where $r<G_{S(g,n)}(\gamma(g,n))$ (resp. $r>G_{S(g,n)}(\gamma(g,n))$) for any integer $g\geq 0$, we set $A_{G}(r,n)=0$ (resp. $A_{G}(r,n)=\infty$). The resulting function $A_{G}:R\times\mathbb{Z}\to\mathbb{Z}\cup\{\infty\}$ is well-defined due to the following lemma.
Lemma 11.7.
$(1)$ $G_{S(g,n)}(-\gamma(g,n))=G_{S(g,n)}(\gamma(g,n))$ for any integers $g,n$ with $g\geq 0$.
$(2)$ For any integers $g_{1},g_{2},n$ with $g_{2}\geq g_{1}\geq 0$, the inequality $G_{S(g_{2},n)}(\gamma(g_{2},n))\geq G_{S(g_{1},n)}(\gamma(g_{1},n))$ holds.
Proof.
As easily seen from a handlebody diagram, the 4-manifold $S(g,n)$ admits an orientation-preserving self-diffeomorphism whose induced isomorphism maps $\gamma(g,n)$ to $-\gamma(g,n)$. The claim (1) thus follows. Due to the assumption $g_{2}\geq g_{1}$, we easily see that $\gamma(g_{1},n)$ is represented by a closed surface of genus $g_{2}$ with self-intersection number $n$, which is obtained from the base surface of $S(g_{1},n)$ by taking a connected sum with a locally embedded surface. We thus have an embedding of $S(g_{2},n)$ into $S(g_{1},n)$ whose induced homomorphism maps $\gamma(g_{2},n)$ to $\gamma(g_{1},n)$ or to $-\gamma(g_{1},n)$. The claim (2) thus follows from the definition of genus function type and the claim (1). ∎
We prove that the function $A_{G}$ gives lower bounds for values of genus functions.
Proposition 11.8.
Every second homology class $\alpha$ of a $4$-manifold $X$ satisfies $g_{X}(\alpha)\geq A_{G}(G_{X}(\alpha),\alpha\cdot\alpha)$.
In the above claim, $\alpha\cdot\alpha$ denotes the self-intersection number of $\alpha$. We remark that this inequality can be sharp, depending on the invariant $G$. For example, for the invariant in Example 11.6.(1) defined by using the Seiberg–Witten invariant, there are many examples of 4-manifolds and their homology classes for which the above bound gives sharp estimates, realizing arbitrary large values of genus functions. This claim can be easily checked using the adjunction inequality for the Seiberg–Witten invariant.
Proof of Proposition 11.8.
We set $g_{0}=g_{X}(\alpha)$. Then $\alpha$ is represented by a closed surface of genus $g_{0}$ with the self-intersection number $n$. We thus have an embedding of $S(g_{0},n)$ into $X$ whose induced homomorphism maps $\gamma(g_{0},n)$ to $\alpha$, showing that $G_{S(g_{0},n)}(\gamma(g_{0},n))\geq G_{X}(\alpha)$. This inequality and the definition of $A_{G}(G_{X}(\alpha),\alpha\cdot\alpha)$ shows $g_{0}\geq A_{G}(G_{X}(\alpha),\alpha\cdot\alpha)$. The claim thus follows. ∎
We note that for each fixed $n$, the function $A_{G}(\;,n):R\to\mathbb{Z}$ is non-decreasing. We observe that this function is the best possible one among similar functions derived from $G$ in the following sense.
Proposition 11.9.
Let $n$ be an integer. Suppose a non-increasing function $A:R\to\mathbb{Z}\cup\{\infty\}$ satisfies $g_{X}(\alpha)\geq A(G_{X}(\alpha))$ for any $4$-manifold $X$ and any $\alpha\in H_{2}(X)$ with $\alpha\cdot\alpha=n$. Then the inequality $A_{G}(G_{X}(\alpha),n)\geq A(G_{X}(\alpha))$ holds for any $4$-manifold $X$ and any $\alpha\in H_{2}(X)$ with $\alpha\cdot\alpha=n$.
Proof.
Suppose to the contrary that the inequality $A_{G}(G_{X}(\alpha),n)<A(G_{X}(\alpha))$ holds for some 4-manifold $X$ and some $\alpha\in H_{2}(X)$ with $\alpha\cdot\alpha=n$. We set $g_{A}=A(G_{X}(\alpha))$. Then, we have $g_{A}>A_{G}(G_{X}(\alpha),n)$, which implies $G_{S(g_{A}-1,n)}(\gamma(g_{A}-1,n))\geq G_{X}(\alpha)$. By the assumptions on $A$, this inequality implies $g_{A}-1\geq A(G_{S(g_{A}-1,n)}(\gamma(g_{A}-1,n)))\geq g_{A}$, giving a contradiction. Therefore, the claim follows. ∎
We next show that $G$ is preserved under attachment of a (quasi-)invertible cobordism.
Theorem 11.10.
Let $M,N$ be ((((possibly disconnected)))) closed $3$-manifolds, let $P$ be a cobordism from $M$ to $N$, and let $X$ be a $4$-manifold with $\partial X=M$. Then the following hold.
$(1)$ If $P$ is invertible, then the inclusion $X\hookrightarrow X\cup_{M}P$ preserves the $G$ invariants.
$(2)$ Suppose that $G$ is preserved by attachments of $3$- and $4$-handles. If $P$ is quasi-invertible, then the inclusion $X\hookrightarrow X\cup_{M}P$ preserves the $G$ invariants.
Proof.
We can prove these claims similarly to Theorem 7.1. See also Remark 7.3. ∎
Remark 11.11.
(1) By Theorem 11.10.(1) and Lemma 6.7, we see that $G$ is preserved under attachment of a 1-handle whose attaching region belongs to a connected boundary component and also under attachment of 2-handles along a strongly slice link with 0-framings.
(2) The conclusion of Theorem 11.10.(2) still holds even if we replace the assumption on 3- and 4-handles with the weaker assumption that $G$ is preserved under attachment of $\natural_{n}(S^{1}\times D^{3})$ along $n(S^{1}\times S^{2})$ for any $n\geq 0$.
We show that the stabilities of algebraic inequivalences also hold for $G$ invariants. More precisely, we prove the following theorem for attachments of (quasi-)invertible cobordisms.
Theorem 11.12.
$(1)$ Theorems 7.1, 7.8, 7.9 and 7.10 and Corollaries 7.4, 7.5 and 7.7 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $P$ is invertible.
$(2)$ Suppose that $G$ is preserved under attachments of $3$- and $4$-handles. Then, Theorems 7.1, 7.8, 7.9 and 7.10 and Corollaries 7.4, 7.5 and 7.7 hold even when “genus functions” in each statement is replaced with “$G$ invariants”.
Proof.
Due to Theorem 11.10.(2), we see that the left inequality in Proposition 7.6 holds for $G$ invariants by the identical argument. Therefore, we can prove the claims of (2) by the identical arguments. Due to Theorem 11.10.(1), we can prove the claims of (1) in the same way. ∎
For boundary sums and connected sums, we prove the following theorem.
Theorem 11.13.
$(1)$ For boundary sums, Theorems 8.4, 8.8, 8.9, 8.10 and 8.11 and Corollary 8.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets.
$(2)$ Suppose that $G$ is preserved under attachment of a $4$-handle. Then, Theorems 8.4, 8.8, 8.9, 8.10 and 8.11 and Corollary 8.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets. If furthermore $G$ is preserved by attachment of a $3$-handle, then these claims and Proposition 8.13 for $G$ invariants hold without any condition on $L$, $L_{1}$ and $L_{2}$.
Proof.
Due to Theorem 11.10, we see that the second claim of Theorem 8.4 and the left inequality in the second claim of Proposition 8.5 hold for $G$ invariants by the identical arguments. Therefore, we can prove the claim (1) by the identical arguments. In the case where $G$ is preserved under attachment of a $4$-handle, we can prove the claim (2) in the same way. We remark that the left inequality in the first claim of Proposition 8.5 also holds for $G$ invariants in this case. ∎
The above theorems imply that if algebraic equivalent classes of $G$ invariants can detect a family of exotic 4-manifolds, then they can detect infinitely many families of exotic 4-manifolds.
Corollary 11.14.
Suppose $G$ admits a family $\{X_{i}\}_{i\in A}$ of pairwise orientedly homeomorphic $4$-manifolds whose $G$ invariants are pairwise algebraically inequivalent. Then the following hold.
$(1)$ If the members $X_{i}$’s have non-empty boundary, then there exists a doubly indexed family $\{X_{i,j}\}_{(i,j)\in A\times\mathbb{Z}}$ of $4$-manifolds having the following properties.
  • For each fixed $j\in\mathbb{Z}$, the members of $\{X_{i,j}\}_{i\in A}$ are pairwise orientedly exotic and have pairwise algebraically inequivalent $G$ invariants.
  • For each fixed $i\in A$, the members of $\{X_{i,j}\}_{j\in\mathbb{Z}}$ are pairwise not homeomorphic.
$(2)$ If $G$ is preserved by a $4$-handle, then the conclusion of the above $(1)$ holds even in the case where $X_{i}$’s are closed $4$-manifolds.
Proof.
We first consider the case (1). By Theorems 11.12 and 11.13, we can construct many examples of desired families, for example, by taking boundary sums of $X_{i}$’s with $S^{1}\times D^{3}$’s. Their homeomorphism types are detected by their first homology groups. In the (2) case, we can similarly construct many examples of desired families, for example, by taking connected sums of $X_{i}$’s with $S^{1}\times D^{3}$’s. We remark that, if $G$ is furthermore preserved by a 3-handle, then we can give desired families of closed 4-manifolds as well, by taking connected sums with $S^{4}(L)$’s. ∎
In contrast to the above corollary, we also show that there exist various infinite families of exotic 4-manifolds which can not be detected by their $G$ invariants. More precisely, we show the following theorem.
Theorem 11.15.
Proposition 9.3 and Theorems 9.4 and 9.5 hold even when “genus functions” in each statement is replaced with “$G$ invariants”.
Proof.
Due to Theorem 11.10, the proofs are identical with the original ones. Note that the boundary of a 4-manifold is assumed to be connected as in the original claims. ∎
Remark 11.16.
$(1)$ The topological genus function of a 4-manifold is defined similarly to the (smooth) genus function by using locally flat embedded surfaces instead of smoothly embedded surfaces, and this function also has long been studied, see [57] and the references therein, for example. Topological genus functions clearly form an (unoriented) diffeomorphism invariant of genus function type. Due to the transversality theorem for locally flat proper submanifolds (see [15, Theorem 1.5] and [57]), the proof of Lemma 7.2 shows that topological genus functions are preserved by 3- and 4-handles. Since two homeomorphic 4-manifolds have equivalent topological genus functions, they cannot be used for detecting exotic 4-manifolds. However, results in this section for topological genus functions might be useful for determining their values and also for detecting homeomorphism types by using stabilities of algebraic inequivalences.
$(2)$ Due to our convention, an invariant $G$ of genus function type is defined only for compact oriented 4-manifolds, but $G$ can be defined for non-compact 4-manifolds as well at least for genus functions. For such an invariant $G$, it easily follows from the definition of genus function type that the $G$ invariants of a 4-manifold and its interior are algebraically equivalent. Thus, for any two 4-manifolds having algebraically inequivalent $G$ invariants, their interiors are not orientedly diffeomorphic.

### v5

@@
 Abstract.
-We prove that every 4-dimensional oriented handlebody without 3- and 4-handles can be modified to admit infinitely many exotic smooth structures, and further that their genus functions are pairwise equivalent.
-We furthermore show that, for any 4-manifold admitting an embedding into a symplectic 4-manifold with weakly convex boundary, its genus function is algebraically realized as those of infinitely many pairwise exotic 4-manifolds.
-We also prove that, for any two 4-manifolds with non-degenerate intersection forms, algebraic inequivalences of genus functions are stable under connected sums and boundary sums with a certain type of 4-manifolds having arbitrarily large second Betti numbers.
-Furthermore, we introduce a notion of genus function type for diffeomorphism invariants and show that similar results hold for any such invariant.
-As an application of our exotic 4-manifolds, we also prove that for any (possibly non-orientable) 4-manifold, every submanifold of codimension less than two satisfying a mild condition can be modified to admit infinitely many exotically knotted copies.
+We prove that every 4-dimensional oriented handlebody without 3- and 4-handles can be modified to admit infinitely many exotic smooth structures, and moreover prove that their genus functions are pairwise equivalent.
+We furthermore show that for any 4-manifold admitting an embedding into a symplectic 4-manifold with weakly convex boundary, its genus function is algebraically realized as those of infinitely many pairwise exotic 4-manifolds.
+In addition, we prove that algebraic inequivalences of genus functions are stable under connected sums and boundary sums with a certain type of 4-manifolds having arbitrarily large second Betti numbers.
+Besides, we introduce a notion of genus function type for diffeomorphism invariants, and show that any such invariant shares properties similar to all the preceding results and yields lower bounds for the values of genus functions.
+As an application of our exotic 4-manifolds, we also prove that for any (possibly non-orientable) 4-manifold, every submanifold of codimension at most one satisfying a mild condition can be modified to admit infinitely many exotically knotted copies.
 Key words and phrases:
-4-manifolds; genus functions; corks; knotted submanifolds; cobordisms
+4-manifolds; genus functions; corks; knotted submanifolds
 2020 Mathematics Subject Classification:
-Primary 57R55, Secondary 57K41, 57R65, 57K43
+Primary 57R55, Secondary 57K41, 57R65, 57K45
 1. Introduction
@@
 Lawson conjectured in his 1997 survey [51] that, for simply connected compact 4-manifolds, genus functions determine smooth structures.
-Indeed, for rational surfaces and ruled surfaces, genus functions uniquely determine their smooth structures among symplectic 4-manifolds (see Proposition 2.6).
-It later turned out that his conjecture is false ([8], [36]), but the counterexamples are not interesting, because their second homology groups are trivial, and hence the underlying topological structures uniquely determine genus functions.
-Thus, the conjecture remains intriguing, especially for 4-manifolds with large second Betti numbers.
-It would also be natural to ask whether a 4-manifold can simultaneously admit exotic smooth structures with the same genus functions and those with distinct genus functions.
-Let us recall that the genus function of a connected oriented smooth 4-manifold $X$ is defined to be the map $H_{2}(X)\to\mathbb{Z}$ which sends a second homology class to the minimal genus of a smoothly embedded surface representing the class.
+Indeed, for rational surfaces and ruled surfaces, genus functions uniquely determine their smooth structures among symplectic 4-manifolds (Proposition 2.7).
+It later turned out that his conjecture is false ([8], [37]), but the counterexamples are not particularly interesting, because their second homology groups are trivial, and hence the underlying topological structures uniquely determine genus functions.
+Thus, the conjecture remains intriguing especially for 4-manifolds with large second Betti numbers, and naturally raises the problem whether a 4-manifold can simultaneously admit exotic structures with the same genus function and those with distinct genus functions.
+Let us recall that the genus function of a connected oriented smooth 4-manifold $X$ is defined to be the map $g_{X}:H_{2}(X)\to\mathbb{Z}$ which sends a class $\alpha$ to the minimum genus of a smoothly embedded surface representing $\alpha$.
 Genus functions of two 4-manifolds will be called equivalent if there exists a homeomorphism between them whose induced isomorphism preserves the genus functions.
-They will be called algebraically equivalent if there exists an isomorphism between their second homology groups that preserves the genus functions and the intersection forms.
+They will be called algebraically equivalent if there exists an isomorphism between their second homology groups that preserves their genus functions and intersection forms.
 In this paper, we prove that every 4-manifold with boundary satisfying mild conditions can be modified to admit infinitely many exotic smooth structures.
-We then show that exotic 4-manifolds with equivalent genus functions exist in abundance.
-Furthermore, we study behavior of genus functions and properties of related invariants.
-We also produce exotically knotted submanifolds.
+Applying this result, we answer the aforementioned problems on genus functions and also modify submanifolds of 4-manifolds into exotically knotted ones.
+Furthermore, we reveal several interesting properties of genus functions.
+Besides, we introduce a notion of genus function type for diffeomorphism invariants of 4-manifolds, and show that any such invariant shares many properties of genus functions.
 Convention 1.1.
-Throughout this paper, all manifolds, submanifolds, boundaries, handlebodies and surfaces are assumed to be compact, connected, oriented and smooth unless otherwise stated.
-Also, two 4-manifolds are said to be exotic to each other if one is orientedly homeomorphic to but not unorientedly diffeomorphic to the other.
+Throughout this paper, unless otherwise stated, all manifolds, submanifolds, boundaries, handlebodies, and surfaces are assumed to be compact, connected, oriented, and smooth, and all embeddings of manifolds are assumed to be smooth but not necessarily proper.
+A pair of 4-manifolds is called exotic (resp.
+orientedly exotic) if they are orientedly homeomorphic but not unorientedly diffeomorphic (resp.
+not orientedly diffeomorphic).
+Here, “orientedly” and “unorientedly” respectively mean “orientation-preserving” and “not necessarily orientation-preserving”.
+We require exotic 4-manifolds to be orientedly homeomorphic, since this assumption is necessary to ensure that their connected sums and boundary sums with a common 4-manifold always yield homeomorphic 4-manifolds.
+We note that boundary 3-manifolds of two exotic 4-manifolds are always orientedly diffeomorphic.
 1.1. Exotic 4-manifolds and genus functions
-To state our results shortly, we introduce the following terminology.
+For convenience, we introduce a notion of topological similarity for 4-manifolds, which we call HIHC-equivalence.
 Definition 1.2.
-For two 4-manifolds $X,Y$ having non-empty (possibly disconnected) boundary, we will say that $Y$ topologically resembles $X$ if $Y$ is homotopy equivalent to $X$, the intersection form of $Y$ is isomorphic to that of $X$, and $\partial Y$ is homology cobordant to $\partial X$.
-Let us recall that a $2$-handlebody is a handlebody consisting of finitely many handles with indices not greater than two.
-They form a very large class of 4-manifolds with boundary, realizing arbitrary large integers as their second Betti numbers $b_{2}$.
-(Beware that in some papers, 2-handlebodies mean handlebodies consisting of only 0- and 2-handles, which form a strictly smaller class than ours.) We prove that every 4-dimensional 2-handlebody can be modified to admit infinitely many exotic smooth structures with pairwise equivalent genus functions, by utilizing corks.
+Two 4-manifolds with (possibly disconnected) boundary are called HIHC-equivalent if they are homotopy equivalent, their intersection forms are isomorphic, and their boundaries are homology cobordant.
+We note that two orientedly homeomorphic 4-manifolds are necessarily HIHC-equivalent, and the converse also holds for simply connected closed (smooth) 4-manifolds due to Freedman’s celebrated theorem.
+Let us recall that a $2$-handlebody is a handlebody consisting of finitely many handles whose indices are at most two.
+$2$-handlebodies form a very large class of 4-manifolds with boundary, realizing arbitrarily large integers as their second Betti numbers $b_{2}$.
+(Note that in some papers, 2-handlebodies mean handlebodies consisting only of 0- and 2-handles, which form a much smaller class.)
+We prove that every 4-dimensional 2-handlebody can be modified to admit infinitely many exotic smooth structures by using corks, and moreover show that these exotic structures have pairwise equivalent genus functions.
 Theorem 1.3.
-For any $4$-dimensional $2$-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise equivalent.
-Even if the condition on genus functions is removed, this result is still new and shows the existence of infinitely many exotic smooth structures for a very large class of 4-manifolds, strongly improving the following earlier results: every 2-handlebody with non-trivial second homology group can be modified to admit arbitrarily many exotic smooth structures (Akbulut and the author [11]); there exist contractible 4-manifolds admitting infinitely many exotic smooth structures (Gompf [36]).
+For any $4$-dimensional 22-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds which are HIHC-equivalent to $X$.
+Moreover, they can be chosen so that their genus functions are pairwise equivalent.
+Even without the conclusion on genus functions, this result is new and shows the existence of infinitely many exotic smooth structures for a very large class of 4-manifolds, substantially strengthening the following earlier results: every 2-handlebody with non-trivial second homology group can be modified to admit arbitrarily many exotic structures (Akbulut and the author [11]); there exist contractible 4-manifolds admitting infinitely many exotic structures (Gompf [37]).
 The corollary below is straightforward from the above theorem.
@@
 $(3)$ For any closed $3$-manifold $M$, there exist infinitely many pairwise exotic simply connected $4$-manifolds with pairwise equivalent genus functions such that their boundaries are homology cobordant to $M$.
-We furthermore prove that various genus functions can not determine smooth structures.
+We furthermore show that a given 4-manifold embedded in a symplectic $4$-manifold with weakly convex boundary can be modified to admit infinitely many exotic smooth structures.
+We moreover prove that the genus function of the given 4-manifold is algebraically realized as their genus functions, showing that there are various genus functions which cannot determine smooth structures.
 Theorem 1.5.
-Suppose that a $4$-manifold $X$ admits an embedding into a symplectic $4$-manifold with weakly convex boundary.
-Then, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
+For any $4$-manifold $X$ admitting an embedding into a symplectic $4$-manifold with weakly convex boundary, there exist infinitely many pairwise exotic $4$-manifolds which are HIHC-equivalent to $X$.
+Moreover, they can be chosen so that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
 The assumption of this theorem holds for a large class of smooth 4-manifolds.
-For example, blow ups of 4-manifolds admitting Stein structures and their codimension-00 submanifolds satisfy the assumption.
+For example, blow ups of 4-manifolds admitting Stein structures and their codimension-0 submanifolds satisfy the assumption.
 Furthermore, these 4-manifolds provide varieties of genus functions even within fixed homeomorphism types ([11], [70], [72]).
@@
 Corollary 1.6.
-$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((((with boundary)))) such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
-$(2)$ If $X$ is the total space of a $D^{2}$-bundle over a closed surface, then there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
+$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((with boundary)) such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
+$(2)$ For any $D^{2}$-bundle over a closed surface, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to the total space $X$ of the bundle such that their genus functions are pairwise equivalent and are algebraically equivalent to that of $X$.
 The claim (1) contrasts with the aforementioned fact that the genus function of $\mathbb{C}P^{2}$ determines its smooth structure among closed symplectic 4-manifolds.
@@
 Theorem 1.7.
-There exist infinitely many pairwise exotic simply connected $4$-manifolds with pairwise algebraically inequivalent genus functions such that each of their genus functions is equivalent to those of another infinitely many pairwise exotic $4$-manifolds.
+There exist infinitely many pairwise exotic simply connected $4$-manifolds with pairwise algebraically inequivalent genus functions such that, for each $4$-manifold $X$ among them, there exist infinitely many pairwise exotic $4$-manifolds homeomorphic to $X$ whose genus functions are all equivalent to that of $X$.
 We can construct similar infinite families for varieties of homeomorphism types.
@@
 Theorem 1.8.
-For any $4$-dimensional $2$-handlebody $X$ with non-trivial second homology group, there exist arbitrarily many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise algebraically inequivalent for any orientations and that each of their genus functions is equivalent to those of infinitely many pairwise exotic $4$-manifolds.
-Exotic 4-manifolds in Theorems 1.3, 1.5 and 1.8 have more properties inherited from a given 4-manifold $X$, see Theorems 9.4, 9.5 and 9.8 for details.
-It remains interesting to ask whether exotic closed 4-manifolds with equivalent genus functions exist.
+For any $4$-dimensional 22-handlebody $X$ with non-trivial second homology group, there exist arbitrarily many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ and having pairwise algebraically inequivalent genus functions such that, for each $4$-manifold $X^{\prime}$ among them, there exist infinitely many pairwise exotic $4$-manifolds homeomorphic to $X^{\prime}$ whose genus functions are all equivalent to that of $X^{\prime}$.
+We remark that exotic 4-manifolds in Theorems 1.3, 1.5 and 1.8 have more properties inherited from a given 4-manifold $X$, see Theorems 9.4, 9.5 and 9.8.
+It remains interesting to see whether exotic closed 4-manifolds with equivalent genus functions exist.
 For closed 4-manifolds, we instead show that algebraic equivalent classes do not determine homeomorphism types even within 4-manifolds having isomorphic cohomology rings and that there are many such 4-manifolds, realizing arbitrarily large $b_{2}$.
-See Proposition 8.13 for details.
+See Theorem 8.12 for details.
 During the preparation of this paper, Stipsicz and Szabó [63] proved that infinitely many pairwise non-diffeomorphic closed 4-manifolds with the cohomology ring of $S^{2}\times S^{2}$ obtained in [32] have pairwise algebraically equivalent genus functions.
@@
 They showed algebraic equivalence by explicitly determining all the values of the genus functions, but our proof is completely different from their approach.
-After the first version of this paper appeared on arXiv on January 2025, Nakamura submitted the paper [53] to arXiv which gives a different example of Corollary 1.4.(1) in the case where the intersection form is represented by the $1\times 1$ matrix 00.
+We remark that, after the first version of this paper appeared on arXiv on January 2025, Nakamura submitted the paper [56] to arXiv which gives an example of Corollary 1.4.(1) in the case where the intersection form is represented by the $1\times 1$ matrix 0, by using a different method.
 1.2. Stabilities of algebraic inequivalences of genus functions
-To show that our exotic 4-manifolds have equivalent genus functions, we prove that attachment of a cobordism of certain type preserves genus functions (See Sections 6 and 7).
+To show that our exotic 4-manifolds have equivalent genus functions, we prove that attachments of a class of cobordisms called quasi-invertible cobordisms always preserve the genus functions, which is of independent interest (see Sections 6 and 7).
 As an application, we discuss behavior of genus functions under connected sums and boundary sums with a certain type of 4-manifolds.
@@
 A simple example of $S^{4}(L)$ is $n(S^{1}\times S^{3})$.
-Each $S^{4}(L)$ has varieties of codimension-00 submanifolds having arbitrary large $b_{2}$ (see Section 8).
-For example, $\Sigma_{g}\times D^{2}$ $(g\geq 0)$ with 2-handles attached along a strongly slice link with 0-framings are such examples, where $\Sigma_{g}$ denotes the closed surface of genus $g$.
-In the rest of this subsection, for each $i=1,2$, let $X_{i}$ be a ((((possibly closed)))) 4-manifold, $L_{i}$ be an $S^{2}$-link in $S^{4}$, and $Z_{i}$ be a connected summand or a codimension-00 submanifold of $S^{4}(L_{i})$.
-We will prove that the values of the genus functions of 4-manifolds are preserved under connected sums and boundary sums with $Z_{i}$’s, see Theorem 8.4.
+As we will show in Subsection 8.1, each $S^{4}(L)$ has rich examples of codimension-0 submanifolds, realizing arbitrarily large $b_{2}$.
+For instance, $\Sigma_{g}\times D^{2}$ $(g\geq 0)$ with 2-handles attached along a strongly slice link (see Definition 6.6) with 0-framings are such examples, where $\Sigma_{g}$ denotes the closed surface of genus $g$.
+In the rest of this subsection, for each $i=1,2$, let $X_{i}$ be a ((possibly closed)) 4-manifold, $L_{i}$ be an $S^{2}$-link in $S^{4}$, and $Z_{i}$ be a connected summand or a codimension-0 submanifold of $S^{4}(L_{i})$.
+We will prove that the values of the genus functions of 4-manifolds are preserved under connected sums and boundary sums with $Z_{1}$ and $Z_{2}$ (Theorem 8.4).
 Applying this result, we show that algebraic inequivalences of genus functions are stable under these operations.
+In the case where $H_{2}(Z_{i})=0$, we prove the following theorem.
 Theorem 1.9.
@@
 $(2)$ In the case where each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty, the genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
-This result contrasts with the fact that the Seiberg–Witten invariants of closed 4-manifolds vanish after taking a connected sum with $S^{1}\times S^{3}$ and hence cannot detect them after the connected sum.
-Even when $Z_{i}$’s have non-trivial $b_{2}$, a similar stability still holds under additional assumptions.
+This result contrasts with the fact that the Seiberg–Witten invariants of closed symplectic 4-manifolds vanish after taking a connected sum with $S^{1}\times S^{3}$ and hence cannot detect them after the connected sum.
+In the general case, we prove similar stabilities under additional assumptions.
 Theorem 1.10.
@@
 If the genus functions of $X_{1}$ and $X_{2}$ are algebraically inequivalent, then those of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically inequivalent.
-We also give similar results without the above torsion-free assumption, see Theorems 8.10 and 8.11.
-These stabilities for algebraic inequivalences show that for any family of exotic 4-manifolds with algebraically inequivalent genus functions, the family remain (orientedly) exotic under connected sums and boundary sums with homeomorphic 4-manifolds satisfying the assumption.
-We can apply these stabilities to various families of exotic 4-manifolds having algebraically inequivalent genus functions (e.g.
-[11], [67], [68], [70]).
-Furthermore, we give similar stabilities for attachments of certain cobordisms, see Sections 6 and 7.
+Furthermore, we give similar results and refinements without assuming the torsion-free conditions (Theorems 8.9 and 8.10).
+However, the above theorem does not hold if the non-degeneracy assumption is removed (Remark 8.11).
+These stabilities for algebraic inequivalences show that for any family of exotic 4-manifolds with algebraically inequivalent genus functions, the family remain (orientedly) exotic under connected sums and boundary sums with homeomorphic 4-manifolds satisfying the assumptions.
+One can apply these stabilities to various families of exotic 4-manifolds (e.g.
+[11], [69], [70], [72]) and produce further examples of exotic families by connected sums and boundary sums.
+Furthermore, we give similar stabilities for attachments of quasi-invertible cobordisms, see Subsection 7.2.
+We remark that the proofs of these stabilities use no gauge theory.
 As an application of our arguments, we give exotic 4-manifolds which are stably exotic under boundary sums with arbitrary Stein manifolds having non-degenerate intersection forms.
@@
 For any positive integer $n$, there exist pairwise exotic simply connected $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ admitting Stein structures such that, for any pairwise orientedly homeomorphic $4$-manifolds $Y_{1},Y_{2},\dots,Y_{n}$ having non-degenerate intersection forms and admitting Stein structures, the boundary sums $X_{1}\natural Y_{1}$, $X_{2}\natural Y_{2}$, $\dots$, $X_{n}\natural Y_{n}$ remain pairwise exotic.
-Note that the summands $Y_{i}$’s are not necessarily pairwise diffeomorphic.
-As seen from the proof, we can realize various homeomorphism types as those of $X_{i}$’s.
+In fact, we realize various homeomorphism types including not simply connected ones as those of $X_{1},\dots,X_{n}$, see Theorem 9.9.
+Note that the summands $Y_{1},\dots,Y_{n}$ can be taken as pairwise exotic 4-manifolds due to a result of [11] (see also Theorem 3.7).
 Although the above theorem assumes that each $Y_{i}$ admits a Stein structure, the above theorem holds under the weaker condition that each $Y_{i}$ admits an embedding into a symplectic 4-manifold with weakly convex boundary.
@@
 See Example 11.5 for these examples.
-Furthermore, we show that an invariant of genus function type yields lower bounds for values of genus functions, which can be sharp depending on the invariant, see Proposition 11.8.
+Furthermore, we show that every invariant of genus function type yields lower bounds for values of genus functions, which can be sharp depending on the invariant, see Proposition 11.7.
 We also prove stabilities of algebraic inequivalences similar to Theorems 1.9 and 1.10, see Theorems 11.12 and 11.13.
-These stabilities imply that if (algebraic equivalent classes of) an invariant of genus function type can detect a family of exotic 4-manifolds, then the invariant can detect infinitely many families of exotic 4-manifolds, see Corollary 11.14.
-By contrast, we also prove that there exist various families of exotic 4-manifolds which cannot be detected by any invariant of genus function type, see Theorem 11.15.
+These stabilities imply that if (algebraic equivalent classes of) an invariant of genus function type can detect a family of orientedly exotic 4-manifolds, then the invariant can detect infinitely many families of orientedly exotic 4-manifolds, see Corollary 11.14.
+By contrast, we also prove that there exist various families of exotic 4-manifolds which can not be detected by any invariant of genus function type, see Theorems 11.15 and 11.16.
 We will give further related results in Section 11.
 1.4. Exotically knotted submanifolds
-Applying our exotic 4-manifolds, we also discuss exotically knotted submanifolds.
-It is natural to ask whether a given submanifold of a 4-manifold admits exotically knotted copies and/or exotically embedded copies, and many examples of submanifolds admitting such copies have been constructed, see, for example, [26], [27], [11], [14], [59], [40], [52], [45] and the references therein.
+Applying our exotic 4-manifolds, we also discuss exotically knotted and exotically embedded submanifolds.
 Here a pair of submanifolds of a manifold will be called exotically knotted if they are diffeomorphic, topologically ambient isotopic, but not smoothly ambient isotopic.
@@
 We remark that our results in this subsection are new as exotically knotted submanifolds and also as exotically embedded submanifolds.
-For a codimension-00 submanifold of a 4-manifold, the closure of the complement will be called the exterior of the submanifold.
-Utilizing our exotic 4-manifolds, we show that every codimension-00 submanifold satisfying a mild condition can be modified to admit infinitely many exotically knotted and exotically embedded copies.
+It is a major problem in 4-dimensional topology to see whether a given submanifold of a 4-manifold admits exotically knotted and/or exotically embedded copies.
+Many examples of submanifolds admitting such copies have been constructed (see, for example, [27], [30], [11], [15], [61], [41], [55], [46] and the references therein), but this problem remains open in general.
+Utilizing our exotic 4-manifolds, here we show that every codimension-0 submanifold satisfying a mild condition on the exterior (i.e., the closure of the complement) can be modified to admit infinitely many exotically knotted and exotically embedded copies.
 Theorem 1.12.
-Let $X$ be a codimension-00 submanifold of a $4$-manifold $Z$.
-These manifolds are possibly non-orientable and may have ((((possibly disconnected)))) boundaries.
-Suppose the exterior of $X$ is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary.
-Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-00 submanifolds topologically resembling $X$.
-Exotically knotted submanifolds in this theorem have more properties inherited from a given submanifold $X$, see Theorem 10.2 for details.
-This result improves an earlier result of Akbulut and the author [11], which modified $X$ into arbitrarily many exotically embedded submanifolds additionally assuming that the exterior is a 2-handlebody with non-trivial second homology group.
-Note that their examples were not shown to be exotically knotted in general.
+Let $X$ be a $4$-manifold embedded in a $4$-manifold $Z$.
+These manifolds may be non-orientable and may have ((possibly disconnected)) boundaries.
+Suppose the exterior of $X$ either is a ((orientable)) 22-handlebody or admits an embedding into a symplectic $4$-manifold with weakly convex boundary, and assume the boundary of the exterior is connected.
+Then, $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-0 submanifolds which are HIHC-equivalent to $X$.
+Furthermore, we show that each of these exotically knotted submanifolds contains the given submanifold $X$ as a submanifold and admits a deformation retraction onto $X$, indicating that the modification to $X$ is small, see Theorem 10.2 and also Remark 10.5.
+Note that the modification is necessary, since the 4-ball admits neither exotically knotted nor exotically embedded copies due to the disk theorem.
+The above theorem improves an earlier result of Akbulut and the author [11], which modified $X$ into arbitrarily many exotically embedded submanifolds additionally assuming that the exterior is a 2-handlebody with non-trivial second homology group.
+Their examples were not shown to be exotically knotted in general.
 Since every 4-manifold has numerous 2-handlebodies as their submanifolds, the corollary below is straightforward.
 Corollary 1.13.
-Every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded ((((possibly non-orientable)))) submanifolds of codimension-00.
-We also discuss exotically knotted codimension-1 submanifolds by applying Theorem 1.12.
+Every ((possibly non-orientable)) 44-manifold admits infinitely many ((possibly non-orientable)) codimension-0 submanifolds which are pairwise exotically knotted and exotically embedded.
+We also discuss codimension-1 submanifolds by applying Theorem 1.12.
 Prior results in the literature are realizations of many 3-manifolds as exotically knotted codimension-1 submanifolds of some orientable 4-manifolds (see e.g.
 [41] and the references therein).
-By contrast, we show that for a given (possibly non-orientable) 4-manifold, every codimension-$1$ submanifold satisfying a mild condition can be modified to admit infinitely many exotically knotted (and exotically embedded) copies, using homology cobordisms embedded in the given 4-manifold.
+By contrast, we show that for a given (possibly non-orientable) 4-manifold, every codimension-$1$ separating submanifold satisfying a mild condition can be modified to admit infinitely many exotically knotted (and exotically embedded) copies, using homology cobordisms embedded in the given 4-manifold.
 Theorem 1.14.
-Let $Z$ be a ((((possibly non-orientable)))) $4$-manifold, and let $M$ be a closed $3$-manifold embedded in the interior of $Z$ as a separating submanifold.
-Suppose the exterior of the tubular neighborhood of $M$ has a connected component which is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary.
-Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$.
-As corollaries, we show that there exist infinitely many closed $3$-manifolds such that every ((((possibly non-orientable)))) $4$-manifold realizes each of them as infinitely many exotically knotted submanifolds.
+Let $M$ be a closed $3$-manifold embedded in the interior of a ((possibly non-orientable)) 44-manifold $Z$ which may have ((possibly disconnected)) boundary.
+Suppose the exterior of the tubular neighborhood of $M$ has a connected component with connected boundary which either is a $4$-dimensional ((orientable)) 22-handlebody or admits an embedding into a symplectic $4$-manifold with weakly convex boundary.
+Then, $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$.
+As corollaries, we show that there exist infinitely many closed $3$-manifolds such that every ((possibly non-orientable)) 44-manifold realizes each of them as infinitely many exotically knotted submanifolds.
 Corollary 1.15.
-For any closed $3$-manifold embedded in $S^{4}$, every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to the $3$-manifold.
+For any closed $3$-manifold $M$ embedded in $S^{4}$, every ((possibly non-orientable)) 44-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$.
 Corollary 1.16.
-There exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold of them, every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to the $3$-manifold.
-Note that many closed 3-manifolds do not have the above property, since there are many closed 3-manifolds not admitting any embedding into $S^{4}$ (see e.g. [41] and the references therein).
-For 1-manifolds embedded in a 4-manifold, the existence of a topological ambient isotopy guarantees that of a smooth ambient isotopy, so there exist no exotically knotted 1-manifolds.
-It would be natural to ask whether the above corollary holds for other codimensions.
-Question 1.17.
+There exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold $M$ among them, every ((possibly non-orientable)) 44-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$.
+Note that many closed 3-manifolds do not have the above property. Indeed, there are many closed 3-manifolds not admitting any embedding into $S^{4}$ (see e.g. [42] and the references therein).
+For a given closed 3-manifold $M$, we consider the question of what is the minimum second Betti number of a simply connected closed 4-manifold which contains a pair of exotically knotted (or exotically embedded) submanifolds diffeomorphic to $M$ (cf.
+[2, Definition 2.2], [41, Question 7.7]).
+This question naturally arises from the well-known fact that every closed 3-manifold admits an embedding into a simply connected closed 4-manifold (see e.g.
+[2, Theorem 2.1]).
+Here we show that every non-negative even integer is realized as such a minimum second Betti number by using Theorem 1.14.
+Corollary 1.17.
+For each even integer $n\geq 0$, there exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold $M$ among them, a simply connected closed $4$-manifold with $b_{2}=n$ contains infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$, but no simply connected closed $4$-manifold with $b_{2}<n$ contains a submanifold diffeomorphic to $M$.
+In fact, we show that the above $4$-manifold with $b_{2}=n$ can be taken as $\frac{n}{2}(S^{2}\times S^{2})$, independently of the choice of the 3-manifold $M$.
+We close this subsection by posing a question.
+As is well-known, every 4-manifold admits no pair of exotically knotted codimension-3 submanifolds.
+By contrast, Corollary 1.16 gives closed 3-manifolds such that every 4-manifold admits pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to them.
+Thus, the following question naturally arises for the remaining codimensions.
+Question 1.18.
 For $n=2,4$, does there exist an $n$-manifold such that every 4-manifold admits a pair of exotically knotted (or exotically embedded) submanifolds diffeomorphic to the $n$-manifold?
-Note that manifolds are compact due to our convention, and thus any $n=4$ example is assumed to have non-empty boundary.
 This question is open even if ambient 4-manifolds are restricted to be simply connected and closed.
-A partial answer to the $n=4$ case is Corollary 1.13, but, to the best of the author’s knowledge, the corresponding result is not known in the $n=2$ case.
-As is well-known, every closed 3-manifold admits an embeddding into a simply connected closed 4-manifold (see e.g.
-[2, Theorem 2.1]).
-Hence, for a given closed 3-manifold, it would be natural to ask what is the minimum second Betti number of a simply connected closed 4-manifold for which the 3-manifold is realized as a pair of exotically knotted (or exotically embedded) submanifolds.
-We show that every non-negative even integer is realized as such a minimum second Betti number by using Theorem 1.14 and a result of Kawauchi [43].
-Corollary 1.18.
-For each non-negative even integer $n$, there exists a closed $3$-manifold $M$ and a simply connected closed $4$-manifold $Z$ with $b_{2}=n$ such that $Z$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$ and that $M$ admits no embedding into any simply connected closed $4$-manifold with $b_{2}<n$.
-We remark that the above $Z$ can be chosen to be $\frac{n}{2}(S^{2}\times S^{2})$ for each $n$.
-It is likely that our construction gives infinitely many distinct homology cobordism classes of 3-manifolds for each $n$, but we do not pursue this point.
+Corollary 1.13 gives a partial answer to the $n=4$ case, but to the best of the author’s knowledge, it is not known whether the corresponding result holds in the $n=2$ case.
 Acknowledgements
@@
 2. Genus functions
-In this section, we introduce a few notions of equivalences for genus functions. We also observe that genus functions can determine smooth structures among symplectic 4-manifolds.
+In this section, we introduce notions of equivalences for genus functions, as well as a notion of the torsion-free genus function.
+We then prove that the values of the genus function do depend on the torsion part of the second homology group, and hence the genus function is not necessarily determined by the torsion-free genus function.
+Finally, we show that genus functions can determine smooth structures among symplectic 4-manifolds.
 We here fix notations of this paper.
@@
 Definition 2.1.
-For a second homology class $\alpha$ of $X$, the integer $g_{X}(\alpha)$ is defined to be the minimal genus of a closed surface smoothly embedded in $X$ that represents $\alpha$.
+For a second homology class $\alpha$ of $X$, the integer $g_{X}(\alpha)$ is defined to be the minimum genus of a closed surface smoothly embedded in $X$ that represents $\alpha$.
 The map $g_{X}:H_{2}(X)\to\mathbb{Z}$ is called the genus function of $X$.
-This function is well-defined, because every second homology class of $X$ is represented by a closed surface (see [38, Exercise 4.5.12(b)]).
+This function is well-defined, because every second homology class of $X$ is represented by a closed surface (see [39, Exercise 4.5.12(b)]).
+We occasionally treat disconnected 4-manifolds as well, and for such 4-manifolds, we allow a representative surface to be disconnected and define the genus of such a surface as the sum of the genera of its connected components.
+Then, we define the genus function of a disconnected 4-manifold in the same way as above.
+Note that, for a (connected) 4-manifold, these two definitions give the same function as easily seen.
 We will say that a continuous map $f:X\to Y$ preserves the genus functions if the induced homomorphism $f_{*}:H_{2}(X)\to H_{2}(Y)$ preserves the genus functions, that is, $f_{*}$ satisfies $g_{X}(\alpha)=g_{Y}(f_{*}(\alpha))$ for any $\alpha\in H_{2}(X)$.
@@
 Definition 2.2.
-The genus functions of $X$ and $Y$ will be called equivalent if there exists a homeomorphism $f:X\to Y$ that preserves the genus functions.
-If $f$ furthermore preserves the orientations of $X$ and $Y$, then the genus functions will be called orientedly equivalent.
+The genus functions of $X$ and $Y$ are called equivalent if there exists a homeomorphism $f:X\to Y$ that preserves the genus functions.
+If $f$ furthermore preserves the orientations of $X$ and $Y$, then the genus functions are called orientedly equivalent.
 It is easy to see that if two 4-manifolds are diffeomorphic (resp.
@@
 orientedly equivalent).
-Hence, the genus function is an invariant of smooth structures.
+Hence, the genus function is a diffeomorphism invariant of 4-manifolds.
 We also introduce a notion of algebraic equivalence for genus functions of 4-manifolds.
@@
 Definition 2.3.
-(1) The genus functions of $X$ and $Y$ will be called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the genus functions and the intersection forms.
-(2) The genus functions of $X$ and $Y$ will be called algebraically equivalent for some orientations if $g_{X}$ is algebraically equivalent to $g_{Y}$ or to $g_{\overline{Y}}$.
-(3) For subgroups $F_{X}\subset H_{2}(X)$ and $F_{Y}\subset H_{2}(Y)$, the restrictions of the genus functions of $X$ and $Y$ to $F_{X}$ and $F_{Y}$ will be called algebraically equivalent if there exists an isomorphism $F_{X}\to F_{Y}$ that preserves the genus functions and the intersection forms.
-It is clear that if the genus functions of two 4-manifolds are algebraically inequivalent (resp.
-algebraically inequivalent for any orientations), then they are orientedly inequivalent (resp.
-(unorientedly) inequivalent) and hence orientedly non-diffeomorphic.
+(1) The genus functions of $X$ and $Y$ are called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the genus functions and the intersection forms.
+(2) The genus functions of $X$ and $Y$ are called algebraically equivalent for some orientations if $g_{X}$ is algebraically equivalent to $g_{Y}$ or to $g_{\overline{Y}}$.
+(3) For subgroups $F_{X}\subset H_{2}(X)$ and $F_{Y}\subset H_{2}(Y)$, the restrictions of the genus functions of $X$ and $Y$ respectively to $F_{X}$ and $F_{Y}$ are called algebraically equivalent if there exists an isomorphism $F_{X}\to F_{Y}$ that preserves the restrictions of the genus functions and the intersection forms.
+It is clear that if the genus functions of two 4-manifolds are algebraically inequivalent, then they are orientedly inequivalent.
+Hence, if they are algebraically inequivalent for any orientations, then they are inequivalent.
 We remark that the definition of equivalence in the paper [63] of Stipsicz and Szabó is the same as our algebraic equivalence, not our (ordinary) equivalence.
-As we will see, it is useful to consider the genus functions modulo torsion part.
+As we will show in Sections 7 and 8, it is useful to consider the genus function modulo torsion part. We thus introduce the following function.
 Definition 2.4.
-For a class $\widehat{\alpha}$ of $H_{2}(X)/\mathrm{Tor}$, we define an integer $g^{*}_{X}(\widehat{\alpha})$ by
-$g^{*}_{X}(\widehat{\alpha})=\min\{g_{X}(\alpha)\mid\textnormal{$\alpha\in H_{2}(X)$ represents $\widehat{\alpha}$.}\}.$
-The map $g^{*}_{X}:H_{2}(X)/\mathrm{Tor}\to\mathbb{Z}$ will be called the torsion-free genus function of $X$.
-Clearly, $g^{*}_{X}$ is completely determined by $g_{X}$.
-We note that $g^{*}_{X}$ is often easier to deal with, since adjunction inequalities give lower bounds for values of $g^{*}_{X}$ rather than $g_{X}$.
-We define their equivalences and algebraic equivalences in the same way as before.
+For a class $A$ of $H_{2}(X)/\mathrm{Tor}$, we define an integer $g^{*}_{X}(A)$ by
+$g^{*}_{X}(A)=\min\{g_{X}(\alpha)\mid\textnormal{$\alpha\in H_{2}(X)$ represents $A$.}\}.$
+The map $g^{*}_{X}:H_{2}(X)/\mathrm{Tor}\to\mathbb{Z}$ is called the torsion-free genus function of $X$.
+Clearly, the torsion-free genus function is completely determined by the genus function.
+The torsion-free genus function is often easier to deal with, since the adjunction inequality for the Seiberg–Witten invariant gives a lower bound for $g_{X}(\alpha)$ up to torsion.
+By contrast, here we show that the values of the genus functions do depend on torsion parts even for closed 4-manifolds having non-vanishing Seiberg–Witten invariants, which is of independent interest.
+Hence, the genus functions are not necessarily determined by the torsion-free genus functions.
+Theorem 2.5.
+There exists a closed $4$-manifold $X$ with $b_{2}^{+}>1$ having non-vanishing Seiberg-Witten invariant that admits a class $\alpha$ and a torsion class $\tau$ of $H_{2}(X)$ satisfying $g_{X}(\alpha)\neq g_{X}(\alpha+\tau)$.
+Furthermore, there exist such closed $4$-manifolds having arbitrarily large $b_{2}^{+}$ and $b_{2}^{-}$.
+Proof.
+Let $Y$ be an arbitrary closed spin 4-manifold, and let $R$ be the closed spin 4-manifold that is the total space of an $S^{2}$-bundle over $\mathbb{R}P^{2}$ (see [39, Exercises 5.7.7.(a)]).
+We note that $R$ is a rational homology 4-sphere satisfying $H_{2}(R)\cong\mathbb{Z}/2\mathbb{Z}$.
+We consider the closed 4-manifold $X=Y\#2\overline{\mathbb{C}P^{2}}\#R$.
+In the case where $Y$ satisfies $b_{2}^{+}>1$ and has a non-vanishing Seiberg–Witten invariant, the 4-manifold $X$ also satisfies these conditions due to the blow-up formula [47, Proposition 2].
+Since each elliptic surface $E(2n)$ ($n\in\mathbb{N}$) is an example of such a 4-manifold $Y$, the 4-manifold $X$ realizes arbitrarily large $b_{2}^{+}$ and $b_{2}^{-}$.
+Let $e_{1},e_{2}$ be a standard orthogonal basis of $H_{2}(2\overline{\mathbb{C}P^{2}})$.
+Then, the mod 2 reduction of the class $\alpha=3e_{1}+e_{2}$ of $H_{2}(X)$ is the Poincaré dual of $w_{2}(X)\in H^{2}(X;\mathbb{Z}/2\mathbb{Z})$.
+Due to Rochlin’s signature theorem, the signature $\sigma(X)$ of $X$ satisfies $\alpha\cdot\alpha-\sigma(X)\equiv-8\not\equiv 0\pmod{16}$.
+We thus obtain $g_{X}(\alpha)>0$ by [39, Exercises 5.7.7.(b)]).
+(In fact, $g_{X}(\alpha)=1$ as seen from the rest of this proof.) Let $V$ be the 4-manifold obtained from the 4-ball by attaching a 2-handle along the (10)(-10)-framed figure-eight knot.
+This manifold V{V} admits an embedding into $2\overline{\mathbb{C}P^{2}}$ that sends a generator of $H_{2}(V)$ to the class $\alpha$ due to a result of Aceto et.al.
+[1].
+Furthermore, by a result of Levine [53, Theorem 1.1 and Remark 1.2], the figure-eight knot bounds a disk in the rational homology 4-ball $R-\mathrm{int}\,D^{4}$.
+The above inequality thus implies $g_{X}(\alpha+\tau)=g_{2\overline{\mathbb{C}P^{2}}\#R}(\alpha+\tau)=0<g_{X}(\alpha)$, where $\tau$ denotes the torsion class of $H_{2}(R)$.
+Therefore, the desired claim follows.
+∎
+We define equivalences for torsion-free genus functions in the same way as before. We state two of them for accuracy.
 Definition 2.6.
-(1) The torsion-free genus functions of $X$ and $Y$ will be called equivalent if there exists a homeomorphism $f:X\to Y$ such that the isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ induced by $f$ preserves $g^{*}_{X}$ and $g^{*}_{Y}$.
-If $f$ furthermore preserves orientations of $X$ and $Y$, then the torsion-free genus functions will be called orientedly equivalent.
-(2) The torsion-free genus functions of $X$ and $Y$ will be called algebraically equivalent if there exists an isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ that preserves $g^{*}_{X},g^{*}_{Y}$ and the intersection forms.
-(3) The torsion-free genus functions of $X$ and $Y$ will be called algebraically equivalent for some orientations if $g^{*}_{X}$ is algebraically equivalent to $g^{*}_{Y}$ or to $g^{*}_{\overline{Y}}$.
-(4) For subgroups $F_{X}\subset H_{2}(X)/\mathrm{Tor}$ and $F_{Y}\subset H_{2}(Y)/\mathrm{Tor}$, the restrictions of the torsion-free genus functions of $X$ and $Y$ to $F_{X}$ and $F_{Y}$ will be called algebraically equivalent if there exists an isomorphism $F_{X}\to F_{Y}$ that preserves the torsion-free genus functions and the intersection forms.
-Let us recall that a rational surface is orientedly diffeomorphic to $\mathbb{C}P^{2}\#n\overline{\mathbb{C}P^{2}}$, and that a ruled surface is diffeomorphic to the total space of an $S^{2}$-bundle over a closed surface or to its blow-ups.
+(1) The torsion-free genus functions of $X$ and $Y$ are called equivalent if there exists a homeomorphism $f:X\to Y$ such that the isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ induced by $f$ preserves the torsion-free genus functions.
+(2) The torsion-free genus functions of $X$ and $Y$ are called algebraically equivalent if there exists an isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ that preserves the torsion-free genus functions and the intersection forms.
+Let us recall that a rational surface is orientedly diffeomorphic to $\mathbb{C}P^{2}\#n\overline{\mathbb{C}P^{2}}$, and that a ruled surface is orientedly diffeomorphic to the total space of an $S^{2}$-bundle over a closed surface or to its blow-ups.
 We observe that genus functions uniquely determine smooth structures of rational surfaces and ruled surfaces among symplectic 4-manifolds.
-This fact easily follows from [50, Proposition 4.3].
+This fact easily follows from [52, Proposition 4.3].
 Proposition 2.7 (cf. [52]).
@@
 Then $R$ is orientedly diffeomorphic to one of the $4$-manifolds $\mathbb{C}P^{2}\#n\overline{\mathbb{C}P^{2}}$, $(S^{2}\times\Sigma_{g})\#n\overline{\mathbb{C}P^{2}}$ and $N_{g}\#n\overline{\mathbb{C}P^{2}}$ for some non-negative integers $n,g$.
-We easily see that $R$ contains a closed surface representing a non-torsion second homology class whose self-intersection number is larger than $22g^{\prime}-$ 2, where $g^{\prime}$ denotes the genus of the surface.
+Note that $R$ satisfies $b_{2}^{+}=1$.
+We easily see that $R$ contains a closed surface representing a non-torsion second homology class whose self-intersection number is larger than $2g^{\prime}-2$, where $g^{\prime}$ denotes the genus of the surface.
 In the case where $b_{2}^{+}(\overline{R})\geq 1$, the orientation reversal $\overline{R}$ also contains such a surface as seen from its diffeomorphism type.
-Now suppose that the genus function of a closed symplectic 4-manifold $X$ is equivalent to $R$.
-Then we have $b_{2}^{+}(X)\geq 1$, and it thus follows from the above paragraph that $X$ contains a closed surface of genus $g^{\prime}$ representing a non-torsion second homology class whose self-intersection number is larger than $22g^{\prime}-$ 2.
-Hence, by [50, Proposition 4.3], $X$ is orientedly diffeomorphic to a rational surface or to a ruled surface.
-We easily see that two 4-manifolds in these classes of 4-manifolds are diffeomorphic if and only if they are homeomorphic, by comparing their homology groups and intersection forms.
+Now suppose that the genus function of a closed symplectic 4-manifold $X$ is equivalent to that of $R$.
+Then we have $b_{2}^{+}(X)\geq 1$, and it thus follows from the above paragraph that $X$ contains a closed surface of genus $g^{\prime}$ representing a non-torsion second homology class whose self-intersection number is larger than $2g^{\prime}-2$.
+Hence, by [52, Proposition 4.3], $X$ is orientedly diffeomorphic to a rational surface or to a ruled surface.
+We see that two 4-manifolds in these classes of 4-manifolds are diffeomorphic if and only if they are homeomorphic, by comparing their homology groups and intersection forms.
 Since $R$ is also diffeomorphic to one of these 4-manifolds and is homeomorphic to $X$, it follows that $X$ is diffeomorphic to $R$.
@@
 The above proposition holds for torsion-free genus functions as well, because the second homology groups of rational surfaces and ruled surfaces are torsion-free.
-3. Corks, homology cobordisms and exotic 4-manifolds
-In this section, we review (generalized) corks, homology cobordisms and constructions of exotic 4-manifolds that utilize corks and invertible homology cobordisms.
-We also give a minor improvement of the method of Akbulut and Ruberman [8] that modifies a generalized cork into exotic 4-manifolds.
+3. Corks, invertible cobordisms and exotic 4-manifolds
+In this section, we review (generalized) corks, invertible cobordisms and constructions of exotic 4-manifolds that utilize corks and invertible homology cobordisms.
+Furthermore, we improve the method of Akbulut and Ruberman [8] that modifies generalized corks into exotic 4-manifolds.
 3.1. Generalized corks
@@
 (1) Let $C$ be a (not necessarily contractible) 4-manifold and let $f:\partial C\to\partial C$ be a self-diffeomorphism of the boundary $\partial C$ that preserves the orientation.
-The pair $(C,f)$ will be called a generalized cork if $f$ extends to a self-homeomorphism of $C$, but does not extend to any self-diffeomorphism of $C$.
+The pair $(C,f)$ is called a generalized cork if $f$ extends to a self-homeomorphism of $C$, but does not extend to any self-diffeomorphism of $C$.
 The order of a generalized cork $(C,f)$ is defined to be the minimum positive integer $k$ for which the power $f^{k}$ extends to a self-diffeomorphism of $C$.
-If $f^{k}$ does not extend to any self-diffeomorphism of $C$ for any positive integer $k$, then $(C,f)$ will be called an infinite order generalized cork.
+If $f^{k}$ does not extend to any self-diffeomorphism of $C$ for any positive integer $k$, then $(C,f)$ is called an infinite order generalized cork.
 A generalized cork $(C,f)$ is called a cork if $C$ is contractible.
@@
 We remove $C$ from $X$ and glue it back along the resulting boundary via the gluing map $f$.
-This operation will be called a generalized cork twist along $(C,f)$.
-(3) Two generalized corks $(C,f)$ and $(C,g)$ will be called equivalent if the self-diffeomorphism $g^{-1}\circ f$ of $\partial C$ extends to a self-diffeomorphism of $C$.
+This operation is called a generalized cork twist along $(C,f)$.
+(3) Two generalized corks $(C,f)$ and $(C,g)$ are called equivalent if the self-diffeomorphism $g^{-1}\circ f$ of $\partial C$ extends to a self-diffeomorphism of $C$.
 Clearly, a generalized cork twist always preserves the homeomorphism type of an ambient 4-manifold.
 We note that in the case where $C$ is contractible, every self-diffeomorphism of $\partial C$ extends to a self-homeomorphism of $C$ ([33], [17]).
-3.2. Homology cobordisms
-Next, we recall the definitions of a cobordism and related terminologies and fix conventions.
+3.2. Invertible cobordisms
+We recall the definitions of an invertible cobordism and related terminologies and fix conventions.
 Throughout this paper, we denote the interval $[0,1]$ by $I$.
@@
 Definition 3.2.
-(1) A (possibly disconnected) 4-manifold $P$ is called a cobordism from $M$ to $N$ if $\partial P$ is orientedly diffeomorphic to the disjoint union $\overline{M}\sqcup N$, and every connected component of $P$ has a non-empty boundary component that belongs to $\overline{M}$.
-In this case, $M$ and $N$ will be respectively called a negative boundary and a positive boundary of $P$.
+(1) A possibly disconnected 4-manifold $P$ is called a ((44-dimensional)) cobordism from $M$ to $N$ if $\partial P$ is orientedly diffeomorphic to the disjoint union $\overline{M}\sqcup N$, and every connected component of $P$ has a non-empty boundary component that belongs to $\overline{M}$.
+In this case, $M$ and $N$ are respectively called a negative boundary and a positive boundary of $P$.
 We often identify $\partial P$ with the disjoint union $\overline{M}\sqcup N$ by fixing a diffeomorphism.
-For an orientation preserving self-diffeomorphism $f$ of $M$, let $X\cup_{f}P$ denote the 4-manifold constructed from $X$ by attaching $P$ along $M$, where we glue each $x\in M\subset\partial X$ with $f(x)\in M\subset\overline{\partial P}$.
+For an orientation-preserving self-diffeomorphism $f$ of $M$, let $X\cup_{f}P$ denote the 4-manifold constructed from $X$ by attaching $P$ along $M$, where we glue each $x\in M\subset\partial X$ with $f(x)\in M\subset\overline{\partial P}$.
 When we do not specify the gluing map $f$, we often denote a glued 4-manifold by $X\cup_{M}P$.
-Unless otherwise stated, we always attach a cobordism to a 4-manifold along a negative boundary.
-(2) For a cobordism $P$ from $M$ to $N$, the cobordism $\overline{P}$ from ${N}$ to ${M}$ is called the upside down cobordism of $P$.
+Unless otherwise stated, we always attach a cobordism to a 4-manifold along a negative boundary of the cobordism.
+(2) For a cobordism $P$ from $M$ to $N$, the cobordism $\overline{P}$ from N{N} to M{M} is called the upside down cobordism of $P$.
 (3) For a cobordism $P$ from $M$ to $N$, if the homomorphisms $H_{*}({M})\to H_{*}(P)$ and $H_{*}(N)\to H_{*}(P)$ induced by the inclusions are isomorphisms, then $P$ is called a homology cobordism.
@@
 (4) A cobordism $P$ from $M$ to $N$ is called invertible if there exists a cobordism $Q$ from $N$ to $M$ such that a glued 4-manifold $P\cup_{N}Q$ is orientedly diffeomorphic to the product $I\times M$.
-The cobordism $Q$ will be called an inverse cobordism of $P$.
+The cobordism $Q$ is called an inverse cobordism of $P$.
+We use the following terminology.
+Definition 3.3.
+For a positive integer $n$, attach $n$ 1-handles $h_{1}^{1},h_{2}^{1},\dots,h_{n}^{1}$ to a 4-manifold $X$, and then attach $n$ 2-handles to the resulting 4-manifold $X^{\prime}$.
+The set of these 1- and 2-handles is called homotopically canceling pairs of $1$- and $2$-handles if the link in $\partial X^{\prime}$ consisting of the attaching circles of 2-handles is homotopic to a link consisting of knots $K_{1},K_{2},\dots,K_{n}$ in $\partial X^{\prime}$, such that each $K_{i}$ intersects the belt sphere of the 1-handle $h_{i}^{1}$ geometrically exactly once.
+Here we give a simple construction of an invertible cobordism as well as its properties. We will later give much more flexible constructions, see Lemma 6.7.
+Lemma 3.4.
+Suppose a cobordism $P$ from $M$ to $N$ is obtained from $I\times M$ by attaching homotopically canceling pairs of $1$- and $2$-handles, none of which are attached to $0\times M$.
+Then $P$ is invertible, and the upside down cobordism $\overline{P}$ is its inverse cobordism.
+Moreover, $P\cup_{\mathrm{id}_{N}}\overline{P}$ is diffeomorphic to $I\times M$ fixing the boundary $\overline{M}\sqcup M$ pointwise, where $\overline{M}\sqcup M$ is identified with $\{0,1\}\times M$ via the natural identification.
+Proof.
+We consider the 5-manifold $P\times I$.
+This manifold is obtained from $(I\times M)\times I$ by attaching homotopically canceling pairs of $1$- and $2$-handles in $1\times M\times I$.
+Since a homotopy of attaching circles of these 2-handles can be approximated by a smooth isotopy, we can eliminate these pairs.
+We thus have a diffeomorphism between $P\times I$ and $(I\times M)\times I$ that fixes $(0\times M)\times I$ pointwise.
+By restricting this diffeomorphism to the exteriors of $(0\times M)\times I$ in the boundaries of these two 5-manifolds, we see that $P\cup_{\mathrm{id}_{N}}\overline{P}$ is diffeomorhic to $I\times M$ fixing the boundary $\overline{M}\sqcup M$ pointwise.
+∎
+Remark 3.5.
+Under the assumption of this lemma, $P$ is a homology cobordism.
+Furthermore, the homotopy extension property (see [40, Proposition 0.18]) implies that $P$ admits a deformation retraction onto the boundary component $M$.
 3.3. Modifying 2-handlebodies into exotic Stein manifolds
@@
 We consider a handlebody diagram of $X$.
-Akbulut and the author [11] introduced two local diagrammatic operations shown in Figure 1.
+Akbulut and the author [11] introduced two local diagrammatic operations for a subarc of $K$ shown in Figure 1.
 For a positive integer $p$, the left side operation is called a $W^{-}(p)$-modification to $K$, and the right side operation is called a $W^{+}(p)$-modification to $K$.
@@
 So, the 4-manifold obtained by the $W^{+}(p)$-modification to $K$ is orientedly homeomorphic to the 4-manifold obtained by the $W^{-}(p)$-modification to $K$.
-Clearly, a $W^{-}$-modification attaches an algebraically canceling pair of 1- and 2-handles to $X$.
-Thus, this operation is an attachment of an invertible homology cobordism to $X$, and its inverse cobordism is its upside down cobordism.
-Furthermore, the homotopy extension property (see [39, Proposition 0.1.8]) implies that this invertible cobordism has a deformation retraction onto $\partial X$.
-Consequently, the result of any $W^{-}$-modification and hence also any $W^{+}$-modification topologically resembles $X$.
-Furthermore, the result of any $W^{-}$-modification admits a deformation retraction onto $X$.
-We note that the 4-manifold obtained by a $W^{+}$-modification becomes diffeomorphic to the original 4-manifold by attaching the upside down cobordism of the invertible cobordism given by the corresponding $W^{-}$-modification ([11, Proposition 4.5]).
+Every $W^{-}$-modification attaches a homotopically canceling pair of 1- and 2-handles to $X$.
+Thus, this operation is an attachment of an invertible homology cobordism to $X$, and its upside down cobordism is its inverse cobordism (see Lemma 3.4).
+Furthermore, this invertible cobordism has a deformation retraction onto $\partial X$ (see Remark 3.5), and hence the result of any $W^{-}$-modification admits a deformation retraction onto $X$.
+Consequently, the result of any $W^{-}$-modification and hence that of any $W^{+}$-modification is HIHC-equivalent to $X$.
+In addition, the 4-manifold obtained by any $W^{+}$-modification becomes diffeomorphic to the original 4-manifold $X$ by attaching the upside down cobordism of the invertible cobordism given by the corresponding $W^{-}$-modification ([11, Proposition 4.5]).
 By drawing Legendrian pictures, we obtain Legendrian versions of $W^{-}$- and $W^{+}$-modifications for Legendrian handlebodies, and these have the following two effects ([11, Proposition 4.7]).
 (1) Each of these two operations modifies a Stein handlebody into another Stein handlebody.
-(2) If we apply a $W^{+}$-modification to a Legendrian knot $K$ of a Legendrian handlebody, then the Thurston–Bennequin number of the result of $K$ increases by $p$, and the Thurston–Bennequin number for the attaching circle of the 2-handle of $W$ is $+2$.
-Thus, by applying $W^{+}$-modifications to a 2-handlebody $X$ and adding zig-zags to each Legendrian knot (i.e., stabilizations), we obtain a Stein handlebody, implying the following.
+(2) If we apply a $W^{+}$-modification to a Legendrian knot $K$ of a Legendrian handlebody, then the Thurston–Bennequin number of the result of $K$ increases by $p$, and the Thurston–Bennequin number for the attaching circle of the 2-handle of $W$ is +2+2.
+Thus, by applying $W^{+}$-modifications to a Legendrian handlebody and adding zig-zags to each Legendrian knot (i.e., stabilizations), we obtain a Stein handlebody, implying the following.
 Proposition 3.6 ([11]).
-For any $4$-dimensional $2$-handlebody $X$, there exists a $4$-dimensional $2$-handlebody $X^{\prime}$ having the following properties.
+For any $4$-dimensional 22-handlebody $X$, there exists a $4$-dimensional 22-handlebody $X^{\prime}$ having the following properties.
   • $X^{\prime}$ admits a Stein structure.
-  • $X^{\prime}$ topologically resembles $X$ and is homeomorphic to a $4$-manifold that admits a deformation retraction onto $X$.
+  • $X^{\prime}$ is HIHC-equivalent to $X$ and is homeomorphic to a $4$-manifold that deformation retracts onto an embedded copy of $X$.
   • $X^{\prime}$ becomes diffeomorphic to $X$ after attaching the upside down cobordism of an invertible homology cobordism along $\partial X^{\prime}$.
-Applying $W^{-}$-modifications several times and replacing some of them with $W^{+}$-operations by cork twists, we obtain arbitrarily many pairwise homeomorphic 2-hanldebodies.
-In the case where $H_{2}(X)$ is a non-trivial group, by carefully choosing coefficients of $W^{+}$- and $W^{-}$-modifications and applying adjunction inequalities, Akbulut and the author showed that the resulting 2-handlebodies have pairwise algebraically inequivalent genus functions for any orientations ([11, Lemma 5.13 and Proposition 6.2]), giving the following result.
+Applying $W^{-}$-modifications several times and replacing some of them with $W^{+}$-operations by cork twists, we obtain arbitrarily many pairwise homeomorphic 2-handlebodies.
+In the case where $H_{2}(X)\neq 0$, by carefully choosing coefficients of $W^{+}$- and $W^{-}$-modifications and applying adjunction inequalities, Akbulut and the author showed that the resulting 2-handlebodies have pairwise algebraically inequivalent genus functions for any orientations ([11, Lemma 5.13 and Proposition 6.2]), giving the following result.
 Theorem 3.7 ([11]).
-For any $4$-dimensional $2$-handlebody $X$ with a non-trivial second homology group and any positive integer $n$, there exist pairwise exotic $4$-manifolds $X_{0},X_{1},\dots,X_{n}$ having the following properties.
-  • Each $X_{i}$ topologically resembles $X$ and admits an embedding into $X$.
-  • $X_{0}$ admits a deformation retraction onto $X$ and is obtained from $X$ by attaching an invertible homology cobordism.
+For any $4$-dimensional 22-handlebody $X$ with $H_{2}(X)\neq 0$ and any positive integer $n$, there exist pairwise exotic $4$-manifolds $X_{0},X_{1},\dots,X_{n}$ having the following properties.
+  • Each $X_{i}$ is HIHC-equivalent to $X$ and admits an embedding into $X$.
+  • $X_{0}$ is obtained from $X$ by attaching an invertible homology cobordism, and admits a deformation retraction onto $X$.
   • $X_{i}$ admits a Stein structure for any $i\neq 0$.
   • The genus functions of $X_{0},X_{1},\dots,X_{n}$ are pairwise algebraically inequivalent for any orientations.
+The proof of this theorem yields a concrete handlebody diagram of each $X_{i}$ from that of $X$ and hence produces various explicit examples of exotic 4-manifolds.
 3.4. Modifying generalized corks into exotic 4-manifolds
-In this subsection, we review the method of Akbulut and Ruberman [8] that modifies a given generalized cork into exotic 4-manifolds. We also give minor improvements of their results.
+In this subsection, we review the method of Akbulut and Ruberman [8] that modifies generalized corks into exotic 4-manifolds. We also improve their results.
 Throughout this subsection, we fix an arbitrary closed 3-manifold $M$.
@@
 Theorem 3.8 ([8]).
-There exist a closed $3$-manifod $N$ and an invertible homology cobordism $\mathbb{P}$ from $M$ to $N$ having the following properties.
+There exist a closed $3$-manifod NN and an invertible homology cobordism $\mathbb{P}$ from $M$ to $N$ that satisfy the following.
   • Every self-diffeomorphism of $N$ extends to a self-diffeomorphism of $\mathbb{P}$ whose restriction to $M$ is isotopic to the identity map.
@@
 Theorem 3.9.
-Let $f_{1},f_{2}$ be self-diffeomorphisms of $M$ that preserve the orientations, and let $X_{1},X_{2}$ be $4$-manifolds with $\partial X_{1}=\partial X_{2}=M$.
+Let $f_{1},f_{2}$ be self-diffeomorphisms of $M$, and let $X_{1},X_{2}$ be $4$-manifolds with $\partial X_{1}=\partial X_{2}=M$.
 If $f_{2}^{-1}\circ f_{1}$ does not extend to any diffeomorphism $X_{1}\to X_{2}$, then the following hold.
   • The $4$-manifolds $X_{1}\cup_{f_{1}}\mathbb{P}$ and $X_{2}\cup_{f_{2}}\mathbb{P}$ are not diffeomorphic to each other.
-  • Each $X_{i}\cup_{f_{i}}\mathbb{P}$ admits a deformation retraction onto $X_{i}$ and topologically resembles $X_{i}$.
+  • Each $X_{i}\cup_{f_{i}}\mathbb{P}$ is HIHC-equivalent to $X_{i}$ and admits a deformation retraction onto $X_{i}$.
 Remark 3.10.
-Akbulut and Ruberman constructed a cobordism $\mathbb{P}$ in Theorem 3.5 using an arbitrary doubly slice knot in $S^{3}$ having certain properties, and they showed that the specific knot $J$ in $S^{3}$ given by [8, Figure 2] can be used for the construction.
-In the above theorem, we use the cobordism $\mathbb{P}$ corresponding to the knot $J$ in order to guarantee the homotopy equivalence.
+Akbulut and Ruberman constructed a cobordism $\mathbb{P}$ in Theorem 3.8 using an arbitrary doubly slice knot in $S^{3}$ having certain properties, and they showed that the Kinoshita–Terasaka knot $J$ (see [8, Figure 2]) can be used for the construction.
+In Theorem 3.9, the cobordism $\mathbb{P}$ denotes the one corresponding to the knot $J$ in order to guarantee the homotopy equivalence.
 The rest of the properties hold for any $\mathbb{P}$ in Theorem 3.8 as seen from our proof.
-Proof.
+Proof of Theorem 3.9.
 Let $\mathbb{Q}$ be an inverse cobordism of $\mathbb{P}$.
@@
 We set $X^{\prime}_{i}=X\cup_{f_{i}}\mathbb{P}$ for each $i$.
-Then each $X^{\prime}_{i}$ deformation retracts to $X$ by Lemma 3.8 which will be given later.
-Since $\mathbb{P}$ is a homology cobordism, we see that $X^{\prime}_{i}$ topologically resembles $X_{i}$.
+Then each $X^{\prime}_{i}$ deformation retracts to $X$ by the forthcoming Lemma 3.11.
+Since $\mathbb{P}$ is a homology cobordism, it follows that $X^{\prime}_{i}$ is HIHC-equivalent to $X_{i}$.
 We will show that $X^{\prime}_{1}$ is not diffeomorphic to $X^{\prime}_{2}$.
@@
 By composing these two diffeomorphisms, we obtain a desired diffeomorphism $X^{\prime}_{i}\cup_{q}\mathbb{Q}\to X_{i}\cup_{\mathrm{id}_{M}}(I\times M)$.
-Composing the last diffeomorphisms and $\Phi$, we obtain a diffeomorphism $X_{1}\cup_{\mathrm{id}_{M}}(I\times M)\to X_{2}\cup_{\mathrm{id}_{M}}(I\times M)$ whose restriction to the boundary is $f_{2}^{-1}\circ f_{1}$.
+Composing the last diffeomorphisms for $i=1,2$ and $\Phi$, we obtain a diffeomorphism $X_{1}\cup_{\mathrm{id}_{M}}(I\times M)\to X_{2}\cup_{\mathrm{id}_{M}}(I\times M)$ whose restriction to the boundary is $f_{2}^{-1}\circ f_{1}$.
 Since each $X_{i}\cup_{\mathrm{id}_{M}}(I\times M)$ is diffeomorphic to $X_{i}$ fixing the boundary $M$ pointwise, it follows that $f_{2}^{-1}\circ f_{1}$ extends to a diffeomorphism $X_{1}\to X_{2}$, giving a contradiction.
@@
 ∎
-The homotopy equivalence of $X_{i}\cup_{f_{i}}\mathbb{P}$ and $X_{i}$ is claimed in [8, Theorem A], but the proof is stated only in the case where $X_{i}$ is contractible.
-For completeness, we give a proof of the general case for $\mathbb{P}$ used in Theorem 3.6 (see Remark 3.7) by showing the lemma below.
-Lemma 3.8 (cf. [8]).
+The homotopy equivalence of $X_{i}\cup_{f_{i}}\mathbb{P}$ and $X_{i}$ is claimed in [8, Theorem A], but the proof is given only in the case where $X_{i}$ is contractible, and the argument does not seem to extend to the general case.
+For completeness, we prove the general case for $\mathbb{P}$ used in Theorem 3.9 (see Remark 3.10) by showing the lemma below.
+Lemma 3.11.
 The cobordism $\mathbb{P}$ used in Theorem 3.9 admits a deformation retraction onto the boundary component $M$.
 Proof.
-In order to construct $\mathbb{P}$, Akbulut and Ruberman [8] used a certain link $L$ in $M$ and a certain concordance $C(\subset I\times S^{3})$ from the unknot $U$ to the knot $J$ in [8, Figure 2].
-Let $\nu(L)$ and $\nu(C)$ respectively denote the tubular neighborhoods of $L$ and $C$, and let $n$ denote the number of the components of $L$.
-They constructed $\mathbb{P}$ from $I\times(M-\textnormal{int}\,\nu(L))$ by attaching $n$ copies of $(I\times S^{3})-\textnormal{int}\,\nu(C)$ along $I\times\partial\nu(L)$ so that the resulting negative boundary remains $M$.
-Note that there is a diffeomorphism from $[-1,0]\times M\cup_{0\times(M-\textnormal{int}\,\nu(L))}I\times(M-\textnormal{int}\,\nu(L))$ to $I\times M$ that maps $(0\times\nu(L))\cup(I\times\partial\nu(L))$ to $1\times\nu(L)$.
-Thus, $\mathbb{P}$ is obtained from $I\times M$ by attaching $n$ copies of the closure of $(I\times S^{3})-\nu(C)$ using a diffeomorphism from each component of $1\times\nu(L)$ to $\nu(m)$, where $m$ denotes the meridian of the unknot $U$ in $0\times S^{3}$, and $\nu(m)$ denotes a tubular neighborhood of $m$ in $\partial(I\times S^{3}-\textnormal{int}\,\nu(C))$.
-Akbulut and Ruberman gave a handlebody diagram of the exterior of a slice disk for $J$ in $D^{4}$ ([8, Figure 4.(a)]), and it is easy to see that this 4-manifold is diffeomorphic to the closure of $(I\times S^{3})-\nu(C)$.
-Since the attaching circle of the unique 2-handle of this handlebody is homotopic to a meridian of the dotted circle, the closure of $(I\times S^{3})-\nu(C)$ deformation retracts to the tubular neighborhood $\nu(m)$.
-Therefore, $\mathbb{P}$ admits a deformation retraction onto $0\times M$.
+In order to construct $\mathbb{P}$, Akbulut and Ruberman [8] used a certain link $L$ in $M$ and a certain concordance $C(\subset I\times S^{3})$ from the unknot $U$ to the Kinoshita–Terasaka knot $J$.
+Let $\nu(L)$ and $\nu(C)$ respectively denote closed tubular neighborhoods of $L$ and $C$, and let $\mathring{\nu}(L)$ and $\mathring{\nu}(C)$ denote the open tubular neighborhoods.
+Let $n$ be the number of the components of $L$.
+They constructed $\mathbb{P}$ from $I\times(M-\mathring{\nu}(L))$ by attaching $n$ copies of $(I\times S^{3})-\mathring{\nu}(C)$ along $I\times\partial\nu(L)$ so that the resulting negative boundary remains $M$.
+Note that there is a diffeomorphism from $1,0]\times M\cup_{0\times(M-\mathring{\nu}(L))}I\times(M-\mathring{\nu}(L))$ to $I\times M$ that maps $(0\times\nu(L))\cup(I\times\partial\nu(L))$ to $1\times\nu(L)$.
+Thus, $\mathbb{P}$ is obtained from $I\times M$ by attaching $n$ copies of $(I\times S^{3})-\mathring{\nu}(C)$ using a diffeomorphism from each component of $1\times\nu(L)$ to $\nu(m)$, where $m$ denotes the meridian of the unknot $U$ in $0\times S^{3}$, and $\nu(m)$ denotes a closed tubular neighborhood of $m$ in $\partial(I\times S^{3}-\mathring{\nu}(C))$.
+Akbulut and Ruberman gave a handlebody diagram of the exterior of a slice disk for $J$ in $D^{4}$ ([8, Figure 4.(a)]), and it is easy to see that this 4-manifold is diffeomorphic to $(I\times S^{3})-\mathring{\nu}(C)$.
+Note that $m$ is isotopic to the meridian of a dotted circle in the diagram.
+Since the unique 2-handle of this handlebody and the 1-handle corresponding to the other dotted circle form a homotopically canceling pair, Remark 3.5 implies that $(I\times S^{3})-\mathring{\nu}(C)$ deformation retracts to the tubular neighborhood $\nu(m)$.
+Therefore, $\mathbb{P}$ admits a deformation retraction onto the boundary component $0\times M$.
 ∎
 Remark 3.12.
-(1) The inverse cobordism of $\mathbb{P}$ used in Theorem 3.6 (see Remark 3.7) is its upside down cobordism $\overline{\mathbb{P}}$, though we do not use this fact in this section.
-This claim is not stated in [8], but it is implicit and follows from their claim that the knot $J$ in [8, Figure 2] is superslice, that is, the double of a slice disk for $J$ is an unknotted 2-sphere in $S^{4}$ (see the proof of [8, Proposition 2.6]).
-The supersliceness can be alternatively proved using [51, Proposition 4.3], and it is clear that the slice disk obtained in [8] gives a concodance in $I\times S^{3}$ from the unknot to $J$ such that the double is isotopic to the product concordance fixing $\{0,1\}\times S^{3}$ pointwise.
+(1) An inverse cobordism of the cobordism $\mathbb{P}$ used in Theorem 3.9 (see Remark 3.10) is its upside down cobordism $\overline{\mathbb{P}}$, though we do not use this fact in this section.
+This claim is not stated in [8], but it is implicit and follows from their claim that the Kinoshita–Terasaka knot $J$ is superslice, that is, the double of a slice disk for $J$ is an unknotted 2-sphere in $S^{4}$ (see the proof of [8, Proposition 2.6]).
+The supersliceness can be alternatively proved using [54, Proposition 4.3], and it is clear from the latter proof that the slice disk obtained in [8] gives a concodance in $I\times S^{3}$ from the unknot to $J$ such that the double is isotopic to the product concordance fixing $\{0,1\}\times S^{3}$ pointwise.
 Therefore, it follows from their construction of $\mathbb{P}$ that the double $\mathbb{P}\cup_{\mathrm{id}_{M}}\overline{\mathbb{P}}$ admits a diffeomorphism to the product $I\times M$ whose restriction to the boundary $\overline{M}\sqcup M$ is the identity map.
-(2) The cobordism $\mathbb{P}$ is obtained from $I\times M$ by attaching algebraically canceling pairs of 1- and 2-handles, and Lemma 3.8 follows from this fact as well.
-This fact is implicit in [8, Example 4.1] and can be proved using Akbulut’s cylinder method [4, Section 3.2] as follows.
+(2) The cobordism $\mathbb{P}$ is obtained from $I\times M$ by attaching homotopically canceling pairs of 1- and 2-handles, and Lemma 3.11 and the claims in the above (1) follow from this fact and Lemma 3.4 as well.
+This fact is implicit from the handlebody diagram of an example of a contractible 4-manifold in [8, Example 4.1] and can be proved using Akbulut’s cylinder method [4, Section 3.2] as follows.
 The proof of Lemma 3.11 implies that the cobordism $\mathbb{P}$ is obtained from the disjoint union of $I\times M$ and copies of the exterior of the slice disk for $J$ by attaching copies of the cylinder $I\times S^{1}\times D^{2}$ along $\partial I\times S^{1}\times D^{2}$.
 Since this cylinder is a union of a 1-handle and a 2-handle, the cobordism $\mathbb{P}$ is obtained from the boundary sum of $I\times M$ with the copies of the exterior by attaching 2-handles.
-Due to the handlebody diagram of the exterior ([8, Figure 4.(a)]), we can prove the above fact by showing that the attaching circle of the 2-handle in each cylinder cancels a 1-handle of the exterior.
+Due to the handlebody diagram of the exterior ([8, Figure 4.(a)]), we can prove the aforementioned fact by canceling a 1-handle of each copy of the exterior using the 2-handle of the cylinder.
 We remark that this argument gives a method for drawing handlebody diagrams of $\mathbb{P}$ and each glued 4-manifold $X_{i}\cup_{f_{i}}\mathbb{P}$, once the link $L$ in $M$ is specified.
-The corollary below is straightforward from Theorem 3.6.
+We obtain the following corollary, improving a result of Akbulut and Ruberman.
 Corollary 3.13.
-Let $\{(C,f)\}_{f\in A}$ be a family of pairwise inequivalent generalized corks with $\partial C=M$.
-Then there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M}\}}$ of pairwise exotic $4$-manifolds topologically resembling $C$ such that each $C_{f}$ admits a deformation retraction onto $C$.
+For any family $\{(C,f)\}_{f\in A}$ of pairwise inequivalent generalized corks, there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C}\}}$ of pairwise exotic $4$-manifolds which are HIHC-equivalent to $C$.
+Proof.
+We set $M=\partial C$, and let $\mathbb{P}$ be the cobordism used in Theorem 3.9.
+For each $f\in A\sqcup\{\mathrm{id}_{\partial C}\}$, let $C_{f}$ denote the 4-manifold $C\cup_{f}\mathbb{P}$.
+Then, it follows from Theorem 3.9 that $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C}\}}$ is a desired family.
+∎
 We note that for any generalized cork $(C,f)$ of infinite order, $\{(C,f^{k})\}_{k\in\mathbb{Z}-\{0\}}$ satisfies the assumption of this corollary.
-Also, the assumption holds for families given by finite order generalized corks.
+The assumption holds for families given by finite order generalized corks as well.
 3.5. Infinite order corks
@@
 For integers $r,s,m>0$, let $C(r,s;m)$ be the 4-manifold in Figure 2.
-(In this notation, we changed the sign of $m$ from the original one in [36] so that we may assume $m>0$.) It is easy to see that $C(r,s;m)$ is a contractible 4-manifold of Mazur type, that is, it admits a handle decomposition consisting of a 0-handle and an algebraically canceling pair of 1- and 2-handles.
+(In this notation, we changed the sign of $m$ from the original one in [37] so that we may assume $m>0$.) It is easy to see that $C(r,s;m)$ is a contractible 4-manifold of Mazur type, that is, it admits a handle decomposition consisting of a 0-handle and an algebraically canceling pair of 1- and 2-handles (i.e., a 1-handle and a 2-handle whose attaching circle intersects the belt sphere of the 1-handle algebraically once).
+Note that such a pair attached to the 0-handle is necessarily a homotopically canceling pair.
+The boundary $\partial C(r,s;m)$ gives infinitely many pairwise non-homeomorphic homology 3-spheres by varying $r,s,m$ ([37, p.2477]).
 Figure 2. $C(r,s;m)$
-Gompf showed that the pair $(C(r,s;m),\tau)$ is an infinite order cork, where the self-diffeomorphism $\tau$ on the boundary is the torus twist on a certain embedded torus.
-Roughly speaking, a torus twist is (Dehn twist on an $(\textnormal{Dehn twist on an annulus})\times\mathrm{id}_{S^{1}}$, and hence $\tau$ is the identity map outside the tubular neighborhood of the torus.
-For the precise definition of a torus twist, see [37, Definition 2.1].
+Gompf proved that each pair $(C(r,s;m),\tau)$ is an infinite order cork, where the self-diffeomorphism $\tau$ on the boundary is the torus twist on a certain embedded torus.
+Roughly speaking, a torus twist is $(\textnormal{Dehn twist on an annulus})\times\mathrm{id}_{S^{1}}$, and hence $\tau$ is the identity map outside the tubular neighborhood of the torus.
+For the precise definition of a torus twist, see [38, Definition 2.1].
 We recall an important property of Gompf’s corks.
@@
 Let $E_{k}$ be the 4-manifold obtained from $E$ by applying the Fintushel–Stern knot surgery ([31]) along the torus $T$ using the knot $T_{k}$.
+Gompf [37] proved the following theorem to show that each $(C(r,s;m),\tau)$ is an infinite order cork.
+For details, see [38, the proof of Theorem 3.5].
 Figure 3. $E$
-Gompf [36] proved the following theorem to show that each $(C(r,s;m),\tau)$ is an infinite order cork. For details, see [37, the proof of Theorem 3.5].
+Figure 4. Twist knot $T_{k}$
 Theorem 3.14 ([37], [38]).
-Let $r,s,m$ be positive integers, and set $N=r+s+m-3$.
-Then there exists an embedding of $C(r,s;m)$ into $E\#N\overline{\mathbb{C}P^{2}}$ such that, for each integer $k$, the knot surgered $4$-manifold $E_{k}\#N\overline{\mathbb{C}P^{2}}$ is diffeomorphic to the $4$-manifold obtained from $E\#N\overline{\mathbb{C}P^{2}}$ by removing the embedded $C(r,s;m)$ and gluing it back via $\tau^{k}$.
+Let $r,s,m$ be positive integers, and set N=r+s+m3N=r+s+m-3.
+Then there exists an embedding of $C(r,s;m)$ into $E\#N\overline{\mathbb{C}P^{2}}$ such that for each integer $k$, the knot surgered $4$-manifold $E_{k}\#N\overline{\mathbb{C}P^{2}}$ is diffeomorphic to the $4$-manifold obtained from $E\#N\overline{\mathbb{C}P^{2}}$ by removing the embedded $C(r,s;m)$ and gluing it back via $\tau^{k}$.
 Furthermore, this diffeomorphism is isotopic to the identity map on the boundary $\partial E$ of these two $4$-manifolds.
 We note that, in the last claim of this theorem, we identified the boundaries of these two 4-manifolds with $\partial E$, since the knot surgery and the cork twist are applied in the interior of $E\#N\overline{\mathbb{C}P^{2}}$.
-Figure 4. The twist knot $T_{k}$
 4. Producing infinite order generalized corks
 In this section, we show that infinite order generalized corks exist for a large class of homeomorphism types, which is of independent interest.
-The first example of an infinite order generalized cork was given by Tange [63, Theorem 1.6], and his example is simply connected and satisfies $b_{2}=2$.
+The first example of an infinite order generalized cork was given by Tange [65, Theorem 1.6], and his example is simply connected and satisfies $b_{2}=2$.
 Gompf [37] later gave the first examples of infinite order contractible corks.
@@
 Theorem 4.1.
-For any $4$-dimensional $2$-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ topologically resembles $X$.
-Furthermore, in the case where $H_{2}(X)$ is a non-trivial group, there exist arbitrarily many generalized corks $(X_{i},f_{i})$’s of infinite order such that $X_{i}$’s are pairwise exotic and topologically resemble $X$.
+For any $4$-dimensional 22-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ is HIHC-equivalent to $X$.
+Furthermore, in the case where $H_{2}(X)\neq 0$, there exist arbitrarily many generalized corks of infinite order such that the $4$-manifolds constituting these corks are pairwise exotic and are HIHC-equivalent to $X$.
 We will prove this theorem together with further properties in Section 9, see Theorem 9.1.
@@
 Proof of Proposition 4.2.
-We fix positive integers $r,s,m$ and abbreviate the contractible 4-manifold $C(r,s;m)$ in Subsection 3.5 by $C$, and we set $N=r+s+m-3$.
+We fix positive integers $r,s,m$ and abbreviate the contractible 4-manifold $C(r,s;m)$ in Subsection 3.5 by $C$, and we set N=r+s+m3N=r+s+m-3.
 Since $C$ is of Mazur type, its double $C\cup\overline{C}$ is diffeomorphic to $S^{4}$.
 Let $H$ be a homology cobordism obtained from $I\times\partial X$ by taking a boundary sum with $C$, where the sum is taken outside the tubular neighborhood of the torus on which Gompf’s torus twist $\tau$ is applied.
-We easily see that the double $H\cup\overline{H}$ is diffeomorphic to $(I\times\partial X)\#(C\cup\overline{C})$ and hence to $I\times\partial X$.
+We see that the double $H\cup\overline{H}$ is diffeomorphic to $(I\times\partial X)\#(C\cup\overline{C})$ and hence to $I\times\partial X$.
 Thus, $H$ is an invertible homology cobordism from $\partial X$ to $\partial X\#\partial C$.
 Let $X^{\prime}$ be the 4-manifold obtained from $X$ by attaching the invertible homology cobordism $H$.
-We easily see that $X^{\prime}$ deformation retracts to $X$ and is diffeomorphic to $X\natural C$.
+Then, $X^{\prime}$ deformation retracts to $X$ and is diffeomorphic to $X\natural C$.
 Let $f$ be the self-diffeomorphism of $\partial X^{\prime}$ that extends the torus twist $\tau$ of $\partial C$ as the identity on the rest of $\partial X^{\prime}$.
@@
 Let $E^{\prime}$ be the 4-manifold obtained from $E$ by attaching a 2-handle along a meridian of the 0-framed knot with sufficiently large negative framing.
-Then $Y\natural E^{\prime}$ is obtained from $Y$ by attaching 2-handles along a Legendrian link with the contact $(-1)$-framings, and thus $Y\natural E^{\prime}$ and its blow-ups $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ are symplectic 4-manifolds with weakly convex boundary ([25, Theorem 2.5], [66], see also [24, Theorem 5.8]).
-We here construct an embedding of $X^{\prime}$ into $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ by taking a boundary sum of $X\subset Y$ and $C\subset E^{\prime}$ using a thickened path from $\partial X$ to $\partial C$.
+Then $Y\natural E^{\prime}$ is obtained from $Y$ by attaching 2-handles along a Legendrian link with the contact (1)(-1)-framings, and thus $Y\natural E^{\prime}$ and its blow-ups $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ are symplectic 4-manifolds with weakly convex boundary ([26, Theorem 2.5], [68], see also [25, Theorem 5.8]).
+We embed $X^{\prime}$ into $Y\natural E^{\prime}\#N\overline{\mathbb{C}P^{2}}$ by taking a boundary sum of $X\subset Y$ and $C\subset E^{\prime}$ using a thickened path from $\partial X$ to $\partial C$.
 By a result of Eliashberg [23] and Etnyre [24], we can embed $Y\natural E^{\prime}$ into a closed symplectic 4-manifold with $b_{2}^{+}>1$.
@@
 Remark 4.4.
-For Proposition 4.2, we can relax the assumption that $X$ admits an embedding into symplectic 4-manifold with weakly convex boundary.
+For each $X$, the above proof gives infinitely many pairwise non-homeomorphic 4-manifolds having the properties of $X^{\prime}$ in Proposition 4.2 by varying $r,s,m$ in the proof.
+Also, we can relax the assumption of Proposition 4.2 that $X$ admits an embedding into a symplectic 4-manifold with weakly convex boundary.
 Indeed, as seen from the proof, if $X\natural E$ admits an embedding into a closed 4-manifold $Z$ with $b_{2}^{+}>1$ and non-vanishing Seiberg–Witten invariant such that the torus $T$ in $E$ represents an infinite order second homology class of $Z$, then the conclusion still holds.
@@
 Take an integral symmetric bilinear form $Q_{A_{i}}:A_{i}\times A_{i}\to\mathbb{Z}$, and define symmetric bilinear forms $Q_{B_{i}}:B_{i}\times B_{i}\to\mathbb{Z}$ and $Q_{i}:A^{\prime}_{i}\times A^{\prime}_{i}\to\mathbb{Z}$ by $Q_{B_{i}}(b,b^{\prime})=0$ and $Q_{i}(a+b,a^{\prime}+b^{\prime})=Q_{A_{i}}(a,a^{\prime})$, where $a,a^{\prime}\in A_{i}$ and $b,b^{\prime}\in B_{i}$.
-We assume that the symmetric bilinear form $Q_{\widetilde{A}_{i}}:(A_{i}/\mathrm{Tor})\times(A_{i}/\mathrm{Tor})\to\mathbb{Z}$ induced by each $Q_{A_{i}}$ is non-degenerate.
+We set $\widetilde{A}_{i}=A_{i}/\mathrm{Tor}$ and assume that the symmetric bilinear form $Q_{\widetilde{A}_{i}}:\widetilde{A}_{i}\times\widetilde{A}_{i}\to\mathbb{Z}$ induced by each $Q_{A_{i}}$ is non-degenerate.
 These maps are models of intersection forms of 4-manifolds.
@@
 Proof.
-Due to the definition of $Q_{i}$, we can check that $\mathrm{rank}\,A_{i}$ is equal to the maximal rank of a submodule $C_{i}$ of $A^{\prime}_{i}$ such that the symmetric bilinear form $(C_{i}/\mathrm{Tor})\times(C_{i}/\mathrm{Tor})\to\mathbb{Z}$ induced by $Q_{i}$ is non-degenerate.
+Due to the definition of $Q_{i}$, we can check that $\mathrm{rank}\,A_{i}$ is equal to the maximum rank of a submodule $C_{i}$ of $A^{\prime}_{i}$ such that the symmetric bilinear form $(C_{i}/\mathrm{Tor})\times(C_{i}/\mathrm{Tor})\to\mathbb{Z}$ induced by $Q_{i}$ is non-degenerate.
 This fact implies $\mathrm{rank}\,A_{1}=\mathrm{rank}\,A_{2}$ and also $\mathrm{rank}\,B_{1}=\mathrm{rank}\,B_{2}$, since the bilinear forms on $A^{\prime}_{1}/\mathrm{Tor}$ and $A^{\prime}_{2}/\mathrm{Tor}$ are isomorphic.
@@
 Let each $T_{i}$ be the torsion submodule of $A_{i}$.
-Clearly, $T_{1}=T_{2}=0$ in the case where $A_{1},A_{2}$ are torsion-free.
+Note that $T_{1}=T_{2}=0$ in the case where $A_{1},A_{2}$ are torsion-free.
 Let each $F_{i}$ be a free submodule of $A_{i}$ such that $A_{i}=F_{i}\oplus T_{i}$.
@@
 Let $M_{i}$ be a matrix representing the bilinear form $F_{i}\times F_{i}\to\mathbb{Z}$ induced by $Q_{A_{i}}$.
-Then each $\det M_{i}$ is independent of the choices of $F_{i}$ and its basis, and $\det M_{i}$ is non-zero due to the non-degeneracy of $Q_{A_{i}}$ on $A_{i}/\mathrm{Tor}$.
+Then each $\det M_{i}$ is independent of the choices of $F_{i}$ and its basis, and $\det M_{i}$ is non-zero due to the non-degeneracy of $Q_{\widetilde{A}_{i}}$.
 Let each $p_{F_{i}}:A^{\prime}_{i}=F_{i}\oplus T_{i}\oplus B_{i}\to F_{i}$ be the projection.
@@
 Due to the condition $\det M_{1}\neq 0$, we easily see that $u_{1}^{\prime},\dots,u_{k}^{\prime}$ is linearly independent (cf.
-[38, Corollary 1.2.13.]).
+[39, Corollary 1.2.13.]).
 We thus have a basis $v_{1},\dots,v_{k}$ of $F_{2}$ and positive integers $a_{1},\dots,a_{k}$ such that $a_{1}v_{1},\dots,a_{k}v_{k}$ generates the submodule generated by $u_{1}^{\prime},\dots,u_{k}^{\prime}$.
@@
 Due to the assumption that $A_{1},A_{2}$, or $B_{1},B_{2}$, are torsion-free, $\Phi$ bijectively maps the torsion submodule of $B_{1}$ to that of $B_{2}$.
-We can thus easily check $p_{B_{2}}\circ\Phi|_{B_{1}}(B_{1})$ is an isomorphism that preserves $Q_{B_{1}},Q_{B_{2}}$, by using the non-degeneracy assumption of each $Q_{A_{i}}$ on $A_{i}/\mathrm{Tor}$.
+We can thus easily check $p_{B_{2}}\circ\Phi|_{B_{1}}(B_{1})$ is an isomorphism that preserves $Q_{B_{1}},Q_{B_{2}}$, by using the non-degeneracy assumption of each $Q_{\widetilde{A}_{i}}$.
 ∎
 In the rest of this section, let $R$ be an ordered set, and let $G_{i}:A^{\prime}_{i}\to R$ be an arbitrary map for each $i=1,2$.
-These maps are models of genus functions and related invariants of 4-manifolds.
+These maps are not necessarily homomorphisms and are models of genus functions and related invariants.
 We will construct isomorphisms that preserve the restrictions of $Q_{1},Q_{2}$ and $G_{1},G_{2}$ by using Lemma 5.1.
@@
 Proof.
-By Lemma 5.1, we see that $p_{A_{2}}\circ\Phi|_{A_{1}}:A_{1}\to A_{2}$ is an isomorphism that preserves $Q_{A_{1}}$, $Q_{A_{2}}$.
-Let $a$ be an element of $A_{1}$ and set $\widehat{a}=p_{A_{2}}\circ\Phi(a)$.
-We will prove $G_{1}(a)=G_{2}(\widehat{a})$.
-We see that $\Phi(a)=\widehat{a}+\widehat{a}_{B}$ for some $\widehat{a}_{B}\in B_{2}$.
-We consider two cases.
+[…]
 (1) The case where $A_{1},A_{2}$ are torsion-free.
@@
 $G_{1}(a)=G_{2}(\widehat{a}+\widehat{a}_{B})\geq G_{2}(\widehat{a})=G_{1}(a-b)\geq G_{1}(a),$
-giving $G_{1}(a)=G_{2}(\widehat{a})$.
+[…]
 (2) The case where $B_{1},B_{2}$ are torsion-free.
@@
 By the assumptions on $G_{1},G_{2},\Phi$, we thus see
-$G_{1}(a+nt)=G_{2}(\widehat{a}+n\widehat{t}+\widehat{a}_{B})$ $\geq G_{2}(\widehat{a}+n\widehat{t})$
+[…]
 $=G_{1}(a+(n-1)t-b)\geq G_{1}(a+(n-1)t),$
@@
 $G_{1}(a)\geq G_{2}(\widehat{a})\geq G_{1}(a-t)\geq G_{1}(a-2t)\cdots\geq G_{1}(a-nt)\geq\cdots.$
-Since $t$ is a torsion, this implies $G_{1}(a)=G_{2}(\widehat{a})$.
-Therefore, the isomorphism $p_{A_{2}}\circ\Phi|_{A_{1}}:A_{1}\to A_{2}$ preserves $Q_{A_{1}}$, $Q_{A_{2}}$ and $G_{A_{1}},G_{A_{2}}$. ∎
-We also discuss maps which are models of the torsion-free versions of the genus functions and related invariants.
-For each $i=1,2$, we set $\widetilde{A}_{i}=A_{i}/\mathrm{Tor}$, $\widetilde{B}_{i}=B_{i}/\mathrm{Tor}$ and $\widetilde{A}^{\prime}_{i}=\widetilde{A}_{i}\oplus\widetilde{B}_{i}$.
-Let $\widetilde{G}_{i}^{A}:\widetilde{A}_{i}\to R$, $\widetilde{G}_{i}^{B}:\widetilde{B}_{i}\to R$ and $\widetilde{G}_{i}:\widetilde{A}^{\prime}_{i}\to R$ be the maps defined by
-$\widetilde{G}_{i}^{A}(\alpha)$ $=\min\{G_{i}(a)\mid a\in A_{i},\;[a]=\alpha\},$
-$\widetilde{G}_{i}^{B}(\beta)$ $=\min\{G_{i}(b)\mid b\in B_{i},\;[b]=\beta\},$
-$\widetilde{G}_{i}(\gamma)$ $=\min\{G_{i}(r)\mid r\in A_{i}^{\prime},\;[r]=\gamma\},$
-where $[a]$, $[b]$ and $[r]$ denote their equivalent classes. We prove the following proposition.
-Proposition 5.4.
-Assume that each $G_{i}$ satisfies $G_{i}(a)\leq G_{i}(a+b)$ for any $a\in A_{i}$ and $b\in B_{i}$.
-If there exists an isomorphism $\Psi:\widetilde{A}^{\prime}_{1}\to\widetilde{A}^{\prime}_{2}$ that preserves $\widetilde{Q}_{1},\widetilde{Q}_{2}$ and $\widetilde{G}_{1},\widetilde{G}_{2}$, then there exists an isomorphism between $\widetilde{A}_{1}$ and $\widetilde{A}_{2}$ that preserves $Q_{\widetilde{A}_{1}},Q_{\widetilde{A}_{2}}$ and $\widetilde{G}_{1}^{A},\widetilde{G}_{2}^{A}$.
-Proof.
-Let $p_{\widetilde{A}_{i}}:\widetilde{A}^{\prime}_{i}\to\widetilde{A}_{i}$ and $p_{\widetilde{B}_{i}}:\widetilde{A}^{\prime}_{i}\to\widetilde{B}_{i}$ be the projections.
-By Lemma 5.1, the composition $p_{\widetilde{A}_{2}}\circ\Psi|_{\widetilde{A}_{1}}:\widetilde{A}_{1}\to\widetilde{A}_{2}$ is an isomorphism that preserves $Q_{\widetilde{A}_{1}}$ and $Q_{\widetilde{A}_{2}}$.
-For each $i=1,2$, we can easily check that every $\alpha\in\widetilde{A}_{i}$ and $\beta\in\widetilde{B}_{i}$ satisfy the inequality $\widetilde{G}_{i}(\alpha+\beta)\geq\widetilde{G}_{i}(\alpha)$ due to the assumption on $G_{i}$.
-This inequality implies $\widetilde{G}_{i}|_{\widetilde{A}_{i}}=\widetilde{G}^{A}_{i}$.
-Therefore, Proposition 5.3 shows that the above isomorphism preserves $\widetilde{G}_{1}^{A}$ and $\widetilde{G}_{2}^{A}$.
-∎
+[…]
 6. Quasi-invertible cobordisms
@@
 For such manifolds, the uniqueness of connected sums and boundary sums do not necessarily hold, but our results hold for any choices of sums unless otherwise stated.
-So, we will use ##\## and $\natural$ for these operations.
+So, we will use #\# and $\natural$ for these operations.
 Let $M,N$ be closed 3-manifolds.
@@
 Suppose a 4-dimensional 1-handlebody of genus $k$ is embedded in the interior of a 4-manifold.
-An operation of removing the 1-handlebody from the 4-manifold and gluing $\natural_{k}S^{2}\times D^{2}$ back along the resulting boundary component will be called a surgery along a $4$-dimensional $1$-handlebody.
-The reverse operation will be called a surgery along an embedded copy of $\natural_{k}(S^{2}\times D^{2})$.
+An operation of removing the 1-handlebody from the 4-manifold and gluing $\natural_{k}S^{2}\times D^{2}$ back along the resulting boundary component is called a surgery along a $4$-dimensional 11-handlebody.
+The reverse operation is called a surgery along an embedded copy of $\natural_{k}(S^{2}\times D^{2})$.
 Note that the way to glue $\natural_{k}S^{2}\times D^{2}$ is not unique in general, but we use a convenient way in each context.
@@
 Definition 6.2.
-(1) A cobordism $R$ from $M$ to $N$ will be called quasi-product if it is orientedly diffeomorphic to the product $I\times M$ possibly after performing surgeries along finitely many embedded $4$-dimensional $1$-handlebodies.
-If furthermore the homomorphism $H_{2}(M)\to H_{2}(R)$ induced by the inclusion is an isomorphism, then $R$ will be called strongly quasi-product.
-(2) A cobordism $P$ from $M$ to $N$ will be called quasi-invertible (resp.
+(1) A cobordism $R$ from $M$ to $N$ is called quasi-product if it is orientedly diffeomorphic to the product $I\times M$ possibly after performing surgeries along finitely many embedded $4$-dimensional 11-handlebodies.
+If furthermore the homomorphism $H_{2}(M)\to H_{2}(R)$ induced by the inclusion is an isomorphism, then $R$ is called strongly quasi-product.
+(2) A cobordism $P$ from $M$ to $N$ is called quasi-invertible (resp.
 strongly quasi-invertible) if it admits a cobordism $Q$ from $N$ to $M$ such that a glued 4-manifold $P\cup_{N}Q$ is a quasi-product cobordism (resp.
 strongly quasi-product cobordism) from the boundary component $M$ of $P$ to that of $Q$.
-The cobordism $Q$ will be called a quasi-inverse cobordism (resp.
+The cobordism $Q$ is called a quasi-inverse cobordism (resp.
 strongly quasi-inverse cobordism) of $P$.
-(3) A cobordism $P$ from $M$ to $N$ will be called $H_{2}$-surjective if the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is surjective.
-As we will later show, there are numerous examples of quasi-invertible cobordisms.
-To produce examples using intensively studied objects knotted 2-spheres, let us recall a few terminologies.
+(3) A cobordism $P$ from $M$ to $N$ is called $H_{2}$-surjective if the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is surjective.
+As we will later show, there are rich examples of quasi-invertible cobordisms.
+To produce examples, let us recall a few terminologies.
 A surgery on a circle embedded in a 4-manifold is an operation of removing the tubular neighborhood $S^{1}\times D^{3}$ of the circle and gluing $S^{2}\times D^{2}$ back along the resulting boundary component.
 In general, there are two ways to glue $S^{2}\times D^{2}$ along the boundary component ([34]), but we use a convenient way in each context.
-Clearly, any surgery on a circle is realized by a surgery along an embedded $4$-dimensional $1$-handlebody of genus one.
+Clearly, any surgery on a circle is realized by a surgery along an embedded $4$-dimensional 11-handlebody of genus one.
 The surgery on a $2$-sphere embedded in a 4-manifold with trivial normal bundle is an operation of removing the tubular neighborhood $S^{2}\times D^{2}$ of the 2-sphere from the 4-manifold and gluing $S^{1}\times D^{3}$ back along the resulting boundary.
@@
 Lemma 6.4.
-$(1)$ A $4$-manifold is a quasi-product cobordism from $M$ to $M$ if and only if the $4$-manifold is diffeomorphic to a $4$-manifold obtained from $I\times M$ by performing surgeries along finitely many embedded copies of the boundary sums of $S^{2}\times D^{2}$’s.
+$(1)$ A $4$-manifold is a quasi-product cobordism from $M$ to $M$ if and only if the $4$-manifold is diffeomorphic to a $4$-manifold obtained from $I\times M$ by performing surgeries along finitely many submanifolds, each of which is diffeomorphic to the boundary sum of copies of $S^{2}\times D^{2}$.
 $(2)$ For any $S^{2}$-link $L$ in $S^{4}$, the connected sum $(I\times M)\#S^{4}(L)$ is a strongly quasi-invertible cobordism from $M$ to $M$.
@@
 Clearly, invertible cobordisms are strongly quasi-invertible, and quasi-product cobordisms are quasi-invertible.
-To produce non-trivial examples, we recall the notion of a strongly slice link.
+To produce varieties of examples, we recall the notion of a strongly slice link.
 Definition 6.6.
-A (1-dimensional) link $\ell$ in $N$ will be called strongly slice if $\ell$ in $1\times N$ bounds a disjoint union of smoothly and properly embedded disks in $I\times N$.
+A (1-dimensional) link $\ell$ in $N$ is called strongly slice if $\ell$ in $1\times N$ bounds a disjoint union of smoothly and properly embedded disks in $I\times N$.
 We note that each component of a strongly slice link is null-homologous, and thus its 0-framing is defined as the Seifert framing.
-We can generate various examples of quasi-invertible cobordisms from the above trivial examples by the lemma below.
+In Subsection 8.1, we will produce various examples of codimension-0 submanifolds of $S^{4}(L)$.
+Using those examples, Lemma 6.4 and the lemma below, we can generate various examples of (strongly) quasi-invertible cobordisms from product cobordisms.
 Lemma 6.7.
-Let $L$ be an $S^{2}$-link in $S^{4}$, and let $V$ be a codimension-00 submanifold of $S^{4}(L)$.
-Then, for any quasi-invertible ((((resp.
-strongly quasi-invertible)))) cobordism $P$ from $M$ to $N$, the following hold.
-$(1)$ Every codimension-00 submanifold of $P$ containing the boundary component $M$ of $P$ is a quasi-invertible ((((resp.
-strongly quasi-invertible)))) cobordism from $M$ to the other boundary component.
-$(2)$ $P\#V$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the disjoint union $N\sqcup\partial V$.
-$(3)$ If $\partial V$ is non-empty, then $P\natural V$ is a quasi-invertible ((((resp.
-strongly quasi-invertible)))) cobordism from $M$ to $N\#\partial V$, where the boundary sum is taken for the boundary component $N$ of $P$.
-$(4)$ Every $4$-manifold obtained by attaching $1$-handles to $P$ along a connected component of $N$ is a quasi-invertible ((((resp.
-strongly quasi-invertible)))) cobordism from $M$ to the new boundary component.
-$(5)$ Every $4$-manifold obtained by attaching an algebraically canceling pair of $1$- and $2$-handles to $P$ along $N$ is a quasi-invertible ((((resp.
-strongly quasi-invertible)))) cobordism from $M$ to the new boundary component.
-$(6)$ Every $4$-manifold obtained by attaching $2$-handles to $P$ along a strongly slice link $\ell$ in $N$ with 00-framings is a quasi-invertible ((((resp.
-strongly quasi-invertible)))) cobordism from $M$ to the 00-surgery on $\ell$ in $N$.
+Let $L$ be an $S^{2}$-link in $S^{4}$, and let $V$ be a codimension-0 submanifold of $S^{4}(L)$.
+Then, for any quasi-invertible ((resp.
+strongly quasi-invertible)) cobordism $P$ from $M$ to $N$, the following hold.
+$(1)$ Every codimension-0 submanifold of $P$ containing the boundary component $M$ of $P$ is a quasi-invertible ((resp. strongly quasi-invertible)) cobordism from $M$ to the other boundary component.
+$(2)$ P#VP\#V is a quasi-invertible ((resp. strongly quasi-invertible)) cobordism from $M$ to the disjoint union $N\sqcup\partial V$.
+$(3)$ If $\partial V$ is non-empty, then $P\natural V$ is a quasi-invertible ((resp.
+strongly quasi-invertible)) cobordism from $M$ to $N\#\partial V$, where the boundary sum is taken for the boundary component $N$ of $P$.
+$(4)$ Every $4$-manifold obtained by attaching $1$-handles to $P$ along a connected component of $N$ is a quasi-invertible ((resp.
+strongly quasi-invertible)) cobordism from $M$ to the new boundary component.
+$(5)$ Every $4$-manifold obtained by attaching homotopically canceling pairs of $1$- and $2$-handles to $P$ along $N$ is a quasi-invertible ((resp.
+strongly quasi-invertible)) cobordism from $M$ to the new boundary component.
+$(6)$ Every $4$-manifold obtained by attaching $2$-handles to $P$ along a strongly slice link $\ell$ in $N$ with 0-framings is a quasi-invertible ((resp.
+strongly quasi-invertible)) cobordism from $M$ to the 0-surgery on $\ell$ in $N$.
 Furthermore, the quasi-invertible cobordisms given by these claims are invertible in the case where $P,L$ are respectively an invertible cobordism and an empty set.
 Proof.
-The claim (1) is straightforward, and the claims (2) and (3) are straightforward from Lemma 6.4 and the claim (1).
+For the claim (1), attach the quasi-inverse cobordism of $P$ to the exterior of the submanifold.
+The resulting manifold gives a quasi-inverse cobordism of the submanifold, implying the claim (1).
+The claims (2) and (3) are straightforward from Lemma 6.4 and the claim (1).
 For (4), by further attaching a canceling 2-handle to each 1-handle, we obtain a 4-manifold diffeomorphic to $P$.
 The claim (4) thus follows from (1).
-For (5), by further attaching the upside down cobordism of the cobordism consisting of the handle pair, we obtain a 4-manifold diffeomorphic to $P$.
+For (5), by further attaching the upside down cobordism of the cobordism consisting of the handle pairs, we obtain a 4-manifold diffeomorphic to $P$ due to Lemma 3.4.
 The claim (5) thus follows from (1).
 The claim (6) is easily checked as follows.
-The link $\ell$ in $0\times N$ bounds a disjoint union of smoothly embedded disks in $I\times N$, and the tubular neighborhood of these disks are the 2-handles attached to $P$.
-Thus the 4-manifold in (6) can be embedded into $P\cup_{\mathrm{id}_{N}}(I\times N)$, and thus the claim (6) follows from (1).
+The link $\ell$ in $0\times N$ bounds a disjoint union of smoothly embedded disks in $I\times N$, and the tubular neighborhood of these disks give the 2-handles attached to $P$.
+Hence, the 4-manifold in (6) can be embedded into $P\cup_{\mathrm{id}_{N}}(I\times N)$, and thus the claim (6) follows from (1).
+The last claim easily follows from these constructions.
 ∎
-See also Section 8 for examples of codimension-00 submanifolds of $S^{4}(L)$.
 6.2. Algebraic properties
-We study algebraic properties of quasi-invertible cobordisms. We begin with quasi-product cobordisms.
+We begin with quasi-product cobordisms.
 Lemma 6.8.
@@
 Proof.
-Let $\widehat{R}$ be the 4-manifold obtained from $R$ by performing surgeries along finitely many embedded 1-handlebodies $\natural_{k_{i}}S^{1}\times D^{3}$’s so that $\widehat{R}$ is diffeomorphic to $I\times M$.
+Let $\widehat{R}$ be the 4-manifold obtained from $R$ by performing surgeries along finitely many embedded 1-handlebodies $\natural_{k_{i}}S^{1}\times D^{3}$ so that $\widehat{R}$ is diffeomorphic to $I\times M$.
 We remove the interior of these 1-handlebodies from $R$ and denote the resulting 4-manifold by $R_{c}$.
+Note that $R=\widehat{R}=R_{c}$ in the case where $R$ is a product cobordism.
 Let ${j}:M\hookrightarrow{R}$, $j_{c}:M\hookrightarrow R_{c}$, $h:R_{c}\hookrightarrow R$, $\widehat{h}:R_{c}\hookrightarrow\widehat{R}$ and $\widehat{j}:M\hookrightarrow\widehat{R}$ be the inclusions.
-By the excision theorem and the Poincaré duality, we see that $H_{3}(\widehat{R},R_{c})\cong\oplus H_{3}(\natural_{k_{i}}(S^{2}\times D^{2}),k_{i}(S^{1}\times S^{2}))\cong 0$.
-The homology exact sequence for the pair $(\widehat{R},R_{c})$ thus shows that $\widehat{h}_{*}:H_{2}(R_{c})\to H_{2}(\widehat{R})$ is injective.
-Since the inclusion $\widehat{j}=\widehat{h}\circ j_{c}:M\hookrightarrow\widehat{R}$ is a homotopy equivalence, the injection $\widehat{h}_{*}$ and also ${j_{c}}_{*}:H_{2}(M)\to H_{2}(R_{c})$ are isomorphisms.
+[…]
 By the excision theorem and the Poincaré duality, we see that $H_{2}(R,R_{c})\cong\oplus H_{2}(\natural_{k_{i}}(S^{1}\times D^{3}),k_{i}(S^{1}\times S^{2}))\cong 0$.
@@
 Corollary 6.9.
-If $M$ is a disjoint union of rational homology $3$-spheres, then any quasi-product ((((resp.
-quasi-invertible)))) cobordism from $M$ to $N$ is strongly quasi-product ((((resp.
-strongly quasi-invertible)))).
+If $M$ is a disjoint union of rational homology $3$-spheres, then any quasi-product ((resp. quasi-invertible)) cobordism from $M$ to $N$ is strongly quasi-product ((resp. strongly quasi-invertible)).
 Proof.
@@
 In the case where $P$ is strongly quasi-invertible, we assume $R$ is strongly quasi-product.
-Furthermore, let $A^{M}_{P}$ and $B_{P}$ respectively denote the image of the homomorphism $H_{2}(M)\to H_{2}(P)$ and the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$, where these are the homomorphisms induced by the inclusions.
+Let $I^{M}_{P}$ and $K^{R}_{P}$ respectively denote the image of the homomorphism $H_{2}(M)\to H_{2}(P)$ and the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$, where these are the homomorphisms induced by the inclusions.
 Then the following lemma holds.
 Lemma 6.10.
-$(1)$ The homomorphism $H_{1}(M)\to H_{1}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=A^{M}_{P}+B_{P}$ holds.
-$(2)$ If $P$ is strongly quasi-invertible, then the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=A^{M}_{P}\oplus B_{P}$ holds.
+$(1)$ The homomorphism $H_{1}(M)\to H_{1}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=I^{M}_{P}+K^{R}_{P}$ holds.
+$(2)$ If $P$ is strongly quasi-invertible, then the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=I^{M}_{P}\oplus K^{R}_{P}$ holds.
 Proof.
 Since the homomorphism $H_{*}(M)\to H_{*}(R)$ is the composition of homomorphisms $H_{*}(M)\to H_{*}(P)$ and $H_{*}(P)\to H_{*}(R)$, the claims follow from Lemma 6.8. ∎
-Finally, we discuss algebraic properties of an attachment of a quasi-invertible cobordisms.
+Finally, we discuss algebraic effects of an attachment of a quasi-invertible cobordisms.
 Let $X$ be a 4-manifold having a boundary component orientedly diffeomorphic to $M$, and let $X^{\prime}$ denote a 4-manifold $X\cup_{M}P$.
 We attach $Q$ to $X^{\prime}$ so that the submanifold $R=P\cup_{N}Q$ is a quasi-product cobordism.
-Let $X^{\prime\prime}$ denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$.
-Let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions.
+In the case where $P$ is strongly quasi-invertible (resp.
+invertible), we assume $R$ is strongly quasi-product (resp.
+product).
+Let $X^{\prime\prime}$ denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$, and let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions.
 Then the following properties hold.
@@
 $(1)$ The homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is surjective.
-$(2)$ $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(B_{P})$.
-$(3)$ The restriction of the intersection form of $X^{\prime}$ to ${\iota_{P}}_{*}(B_{P})$ is the zero-map.
+$(2)$ $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(K^{R}_{P})$.
+$(3)$ The restriction of the intersection form of $X^{\prime}$ to ${\iota_{P}}_{*}(K^{R}_{P})$ is the zero-map.
 $(4)$ $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is surjective if and only if $P$ is $H_{2}$-surjective.
@@
 The excision theorem shows $H_{2}(X^{\prime\prime},X)\cong H_{2}(R,M)=0$.
-The homology exact sequence for the pair $(X^{\prime\prime},X)$ thus shows that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is surjective.
-$(2)$ The claim (1) and the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$ implies $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+B_{X^{\prime}}$, where $B_{X^{\prime}}$ denotes the kernel of the homomorphism $H_{2}(X^{\prime})\to H_{2}(X^{\prime\prime})$ induced by the inclusion.
-We note ${\iota_{P}}_{*}(B_{P})\subset B_{X^{\prime}}$.
-Using the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}P$ and Lemma 6.10, we can check $\iota_{*}(H_{2}(X))+B_{X^{\prime}}=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(B_{P})$, showing the claim.
+The homology exact sequence for the pair $(X^{\prime\prime},X)$ thus shows the claim.
+$(2)$ The claim (1) and the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$ implies $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+K^{X^{\prime\prime}}_{X^{\prime}}$, where $K^{X^{\prime\prime}}_{X^{\prime}}$ denotes the kernel of the homomorphism $H_{2}(X^{\prime})\to H_{2}(X^{\prime\prime})$ induced by the inclusion.
+We note ${\iota_{P}}_{*}(K^{R}_{P})\subset K^{X^{\prime\prime}}_{X^{\prime}}$.
+Using the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}P$ and Lemma 6.10, we can check $\iota_{*}(H_{2}(X))+K^{X^{\prime\prime}}_{X^{\prime}}=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(K^{R}_{P})$, showing the claim.
 $(3)$ By Lemma 6.8, the intersection form of $R$ is the zero-map, and thus the claim follows.
-$(4)$ By Lemma 6.10 and the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}P$, the homomorphism $\iota_{*}:H_{1}(X)\to H_{1}(X^{\prime})$ is injective.
-Thus, due to the homology exact sequence for the pair $(X^{\prime},X)$, the homomorphism $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is surjective if and only if $H_{2}(X^{\prime},X)=0$.
+$(4)$ Lemma 6.10 and the Mayer–Vietoris exact sequence for $X^{\prime}=X\cup_{M}P$ imply that $\iota_{*}:H_{1}(X)\to H_{1}(X^{\prime})$ is injective.
+It thus follows from the homology exact sequence for the pair $(X^{\prime},X)$ that $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is surjective if and only if $H_{2}(X^{\prime},X)=0$.
 Similarly, by Lemma 6.10, $P$ is $H_{2}$-surjective if and only if $H_{2}(P,M)=0$.
@@
 ∎
-The corollary below is straightforward.
+The useful corollary below is straightforward from the above (4).
 Corollary 6.12.
 Suppose $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is injective. Then, $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is an isomorphism if and only if $P$ is $H_{2}$-surjective.
-We also have the following decomposition.
-Lemma 6.13.
-If the intersection form of $X$ is non-degenerate ((((on $H_{2}(X)/\mathrm{Tor}$)))), then $H_{2}(X^{\prime})/\mathrm{Tor}=\iota_{*}(H_{2}(X))/\mathrm{Tor}\,\oplus\,{\iota_{P}}_{*}(B_{P})/\mathrm{Tor}$.
-Proof.
-Since every non-torsion class of $H_{2}(X)$ algebraically intersects with some class of $H_{2}(X)$ under the assumption, the claim easily follows from Lemma 6.11. ∎
 In the case where the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective, nice properties hold.
@@
 Suppose the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is injective.
-Then, $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is injective, and the decomposition $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))\oplus{\iota_{P}}_{*}(B_{P})$ holds.
+Then, $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is injective, and the decomposition $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))\oplus{\iota_{P}}_{*}(K^{R}_{P})$ holds.
 Proof.
-The injectivity of $\iota_{*}$ is clear from the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$.
-Since the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective, we easily see $\iota_{*}(H_{2}(X))\cap{\iota_{P}}_{*}(B_{P})=0$.
-Lemma 6.11 thus shows $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))\oplus{\iota_{P}}_{*}(B_{P})$.
-∎
+The injectivity is clear from the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$, and the decomposition easily follows from Lemma 6.11. ∎
 We give sufficient conditions that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective.
@@
     The intersection form of $X$ is non-degenerate, and $H_{2}(X)$ has no torsion.
-Then the homomorphisms $H_{2}(X)\to H_{2}(X^{\prime})$ and $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusions are injective.
+Then the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is injective.
 Proof.
-In the case where (i) or (ii) holds, by the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}R$ and Lemma 6.8, we see that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective, and thus the homomorphism $H_{2}(X)\to H_{2}(X^{\prime})$ is also injective.
-In the case where (iii) holds, we easily see that the condition (ii) also holds, and hence the claim holds.
+In the case where (i) or (ii) holds, the Mayer–Vietoris exact sequence for $X^{\prime\prime}=X\cup_{M}R$ implies that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective.
+In the case where (iii) holds, every class of $H_{2}(X)$ algebraically intersects with some class of $H_{2}(X)$, implying that the condition (ii) also holds.
+Hence, the claim follows.
 ∎
-The property below is useful for studying genus functions.
+The lemma below is useful for applying the results in Section 7.
 Lemma 6.15.
@@
     $P$ is strongly quasi-invertible, and $N$ is a disjoint union of rational homology $3$-spheres.
-Then, $H_{2}(P)$ and ${\iota_{P}}_{*}(B_{P})$ are torsion-free.
+Then, $H_{2}(P)$ and ${\iota_{P}}_{*}(K^{R}_{P})$ are torsion-free.
 Proof.
-By the Poincaré duality and the universal coefficient theorem, we see that $H_{2}(N)$ and $H_{2}(M)$ are torsion-free. Since $P$ is strongly quasi-inverible, $H_{2}(R)$ is also torsion-free.
+The Poincaré duality and the universal coefficient theorem imply that $H_{2}(N)$ and $H_{2}(M)$ are torsion-free. Since $P$ is strongly quasi-inverible, we see that $H_{2}(R)$ is also torsion-free.
 We consider the Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$, and we will show that the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ in the sequence is injective.
@@
 So, we prove the injectivity in the case where (ii) holds.
-In this case, we may assume $R$ is diffeomorphic to $I\times M$, and hence the inclusion $M\hookrightarrow R$ is a homotopy equivalence.
+In this case, $R$ is diffeomorphic to $I\times M$, and hence the inclusion $M\hookrightarrow R$ is a homotopy equivalence.
 Due to the sequence $M\subset P\subset R$, this fact implies that the homomorphism $H_{3}(P)\to H_{3}(R)$ induced by the inclusion is surjective.
@@
 Now we are ready to prove the claim.
-Since $H_{2}(N)$ and $H_{2}(R)$ are torsion-free, and the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ is injective, the Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$ implies that $H_{2}(P)$ is torsion-free, and hence $B_{p}$ is also torsion-free.
-By Lemma 6.10, we have $H_{2}(P)=A^{M}_{P}\oplus B_{P}$.
-The Mayer–Vietoris exact sequence for $X^{\prime}=X\cup_{M}P$ thus shows that the restriction of ${\iota_{P}}_{*}:H_{2}(P)\to H_{2}(X^{\prime})$ to $B_{P}$ is an isomorphism.
-Therefore, ${\iota_{P}}_{*}(B_{P})$ is torsion-free.
+Since $H_{2}(N)$ and $H_{2}(R)$ are torsion-free, and the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ is injective, the Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$ implies that $H_{2}(P)$ is torsion-free, and hence $K^{R}_{P}$ is also torsion-free.
+By Lemma 6.10, we have $H_{2}(P)=I^{M}_{P}\oplus K^{R}_{P}$.
+It thus follows from the Mayer–Vietoris exact sequence for $X^{\prime}=X\cup_{M}P$ that the restriction ${\iota_{P}}_{*}|_{K^{R}_{P}}:K^{R}_{P}\to{\iota_{P}}_{*}(K^{R}_{P})$ is an isomorphism.
+Therefore, ${\iota_{P}}_{*}(K^{R}_{P})$ is torsion-free.
 ∎
+In the case where the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ is injective, we still have nice properties.
+Proposition 6.16.
+Suppose the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ induced by the inclusion is injective.
+Then, $\iota_{*}:H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime})/\mathrm{Tor}$ is injective, and the decomposition $H_{2}(X^{\prime})/\mathrm{Tor}=\iota_{*}(H_{2}(X))/\mathrm{Tor}\oplus{\iota_{P}}_{*}(K^{R}_{P})/\mathrm{Tor}$ holds.
+Proof.
+The proof is similar to that of Proposition 6.13. ∎
+We give sufficient conditions that the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ is injective.
+Proposition 6.17.
+Suppose at least one of the following conditions (i)(iii) holds.
+  1. (i)
+    $P$ is strongly quasi-invertible.
+  2. (ii)
+    The homomorphism $H_{2}(M)\to H_{2}(X)/\mathrm{Tor}$ induced by the inclusion is the zero-map.
+  3. (iii)
+    The intersection form of $X$ is non-degenerate.
+Then the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ induced by the inclusion is injective.
+Proof.
+The proof is similar to that of Proposition 6.14. ∎
 7. Genus functions and quasi-invertible cobordisms
-In this section, we study behavior of genus functions of 4-manifolds with quasi-invertible cobordisms attached.
-We also show stabilities of algebraic inequivalences of genus functions under attachments of quasi-invertible cobordisms.
+In this section, we study behavior of genus functions of 4-manifolds under attachment of quasi-invertible cobordisms.
+We also show stabilities of algebraic inequivalences of genus functions under the attachments.
 Throughout this section, we allow manifolds, submanifolds, and their boundaries to be disconnected unless otherwise stated, but we still assume surfaces are connected.
 7.1. Behavior of genus functions
-We first study behavior of genus functions under attachments of quasi-invertible cobordisms.
-Throughout this section, we use the following setting.
+Throughout this subsection, we use the following setting.
 Let $M,N$ be closed 3-manifolds, and let $P$ be a quasi-invertible cobordism from $M$ to $N$.
@@
 Proof.
-(1) We first consider the case where the attached handle is a 3-handle $D^{3}\times D^{1}$.
+(1) We prove the claim in the case where the attached handle is a 3-handle $D^{3}\times D^{1}$.
+In the case of a 4-handle, the claim follows by a much simpler argument.
 Let $j:Z\hookrightarrow Z^{\prime}$ denote the inclusion.
-We note that the kernel of $j_{*}:H_{2}(Z)\to H_{2}(Z^{\prime})$ is the image of the homomorphism $H_{2}(S^{2}\times D^{1})\to H_{2}(Z)$ induced by the inclusion as seen from the Mayer–Vietoris exact sequence, where $S^{2}\times D^{1}$ denotes the attaching region of the 3-hanlde.
+We note that the kernel of $j_{*}:H_{2}(Z)\to H_{2}(Z^{\prime})$ is the image of the homomorphism $H_{2}(S^{2}\times D^{1})\to H_{2}(Z)$ induced by the inclusion as seen from the Mayer–Vietoris exact sequence, where $S^{2}\times D^{1}$ denotes the attaching region of the 3-handle.
 Let $\alpha$ be a class of $H_{2}(Z)$ and define the class $\alpha^{\prime}$ of $H_{2}(Z^{\prime})$ by $\alpha^{\prime}={j}_{*}(\alpha)$.
@@
 Therefore, we have $g=g^{\prime}$, showing that $j$ preserves the genus functions.
-In the case where the attached handle is a 4-handle, the claim follows by a much simpler argument.
 (2) Since $Z^{\prime}$ is obtained from $Z$ by attaching 3-and 4-handles, the claim follows from (1). ∎
-Now we prove Theorem 7.1.
+In the rest of this subsection, we use the following setting to prove Theorem 7.1 and its corollaries.
+We attach a quasi-inverse cobordism $Q$ of $P$ to $X^{\prime}$ so that the submanifold $R=P\cup_{N}Q$ is a quasi-product cobordism, and we denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$ by $X^{\prime\prime}$.
 Proof of Theorem 7.1.
-We attach a quasi-inverse cobordism $Q$ of $P$ to $X^{\prime}$ so that the submanifold $R=P\cup_{N}Q$ is a quasi-product cobordism.
-Let $X^{\prime\prime}$ denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$.
-The cobordism $R$ becomes diffeomorphic to $I\times M$ after performing surgeries along finitely many embedded 4-dimensional 1-handlebodies.
+The cobordism $R$ becomes diffeomorphic to $I\times M$ possibly after performing surgeries along finitely many embedded 4-dimensional 1-handlebodies.
 Thus, the 4-manifold $\widehat{X}^{\prime\prime}$ obtained from $X^{\prime\prime}$ by these surgeries is diffeomorphic to $X\cup_{\mathrm{id}_{M}}(I\times M)$.
@@
 We remove the interior of the aforementioned 1-handlebodies in $R$ from $X^{\prime\prime}$ and denote the resulting 4-manifold by $X^{\prime\prime}_{c}$.
-Let $j:X\hookrightarrow X^{\prime\prime}$, $j^{\prime}:X^{\prime}\hookrightarrow X^{\prime\prime}$, $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$, $h:X^{\prime\prime}_{c}\hookrightarrow X^{\prime\prime}$ and $\widehat{h}:X^{\prime\prime}_{c}\hookrightarrow\widehat{X}^{\prime\prime}$ be the inclusions.
-Since $j_{c}$ and $\widehat{h}$ are inclusions, we see that the inequalities $g_{X}(\alpha)\geq g_{X^{\prime\prime}_{c}}({j_{c}}_{*}(\alpha))$ and $g_{X^{\prime\prime}_{c}}(\beta)\geq g_{\widehat{X}^{\prime\prime}}(\widehat{h}_{*}(\beta))$ hold for any $\alpha\in H_{2}(X)$ and $\beta\in H_{2}(X^{\prime\prime}_{c})$.
-As noted before, $\widehat{j}=\widehat{h}\circ j_{c}:X\hookrightarrow\widehat{X}^{\prime\prime}$ preserves the genus functions.
-The above inequalities thus imply that $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$ also preserves the genus functions.
+Note that $X^{\prime\prime}=\widehat{X}^{\prime\prime}=X^{\prime\prime}_{c}$ in the case where $R$ is a product cobordism.
+Let $\hookrightarrow X^{\prime\prime}, j:XX′′j^{\prime}$$^{\prime}\hookrightarrow X^{\prime\prime}$, $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$, $h:X^{\prime\prime}_{c}\hookrightarrow X^{\prime\prime}$ and $\widehat{h}:X^{\prime\prime}_{c}\hookrightarrow\widehat{X}^{\prime\prime}$ be the inclusions.
+Since $j_{c}$ and $\widehat{h}$ are inclusions, and $\widehat{j}=\widehat{h}\circ j_{c}$ preserves the genus functions, every $\alpha\in H_{2}(X)$ satisfies
+$g_{X}(\alpha)\geq g_{X^{\prime\prime}_{c}}({j_{c}}_{*}(\alpha))\geq g_{\widehat{X}^{\prime\prime}}(\widehat{h}_{*}({j_{c}}_{*}(\alpha)))=g_{X}(\alpha),$
+showing that $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$ also preserves the genus functions.
 By Lemma 7.2, the inclusion $h:X^{\prime\prime}_{c}\hookrightarrow X^{\prime\prime}$ preserves the genus functions, and thus $j=h\circ j_{c}:X\hookrightarrow X^{\prime\prime}$ also preserves the genus functions.
-This fact implies that $\iota:X\hookrightarrow X^{\prime}$ preserves the genus functions due to the relation $j=j^{\prime}\circ\iota$.
-∎
+Due to this fact and the relation $j=j^{\prime}\circ\iota$, we see that every $\alpha\in H_{2}(X)$ satisfies
+$g_{X}(\alpha)\geq g_{X^{\prime}}({\iota}_{*}(\alpha))\geq g_{{X}^{\prime\prime}}(j^{\prime}_{*}({\iota}_{*}(\alpha)))=g_{X}(\alpha),$
+showing that $\iota:X\hookrightarrow X^{\prime}$ preserves the genus functions. ∎
 Remark 7.3.
 (1) By Theorem 7.1 and Lemma 6.7, the genus functions are preserved under attachment of a 1-handle whose attaching region belongs to a connected boundary component and also under attachment of 2-handles along a strongly slice link with 0-framings.
-(2) To prove a result similar to Theorem 7.1 for other invariants, we note that the following properties (i)–(iii) of genus functions are used in the above proof.
-(i) Every diffeomorphism preserves the genus functions.
-(ii) For any orientation preserving embedding, the homomorphism induced by the embedding does not strictly increase the values of their genus functions.
-(iii) Lemma 7.2.(1).
-Even when we do not use Lemma 7.2.(1), we can still prove Theorem 7.1 in the case where $P$ is an invertible cobordism as seen from the above proof.
+Applying this result, we can show non-sliceness for various knots in $S^{3}$.
+(2) To prove a result similar to Theorem 7.1 for other invariants in Section 11, we note that the properties of genus functions used in the above proof are the following (i)–(iii).
+(i) Every (orientation-preserving) diffeomorphism preserves the genus functions.
+(ii) For any (orientation-preserving) embedding, the homomorphism induced by the embedding does not increase the values of genus functions.
+(iii) The property stated in Lemma 7.2.(1).
+We note that, in the case where $P$ is an invertible cobordism, the above proof of Theorem 7.1 does not require Lemma 7.2.(1).
+(3) There are many examples of cobordisms that do not have the property of Theorem 7.1.
+For instance, many cobordisms do not have this property after taking a connected sum with $S^{2}\times S^{2}$.
+We can also construct examples of such homology cobordisms as follows.
+Take a (contractible) cork $(C,f)$ admitting a 4-manifold $Z$ such that its cork twist along $(C,f)$ changes the algebraic equivalent class of the genus function of $Z$.
+Note that there are various examples of such corks (e.g.
+[6], [9], [10], [37]).
+Then, $\overline{C}-\textnormal{int}\,D^{4}$ is a homology cobordism from $\partial C$ to $S^{3}$, and the attachment of this cobordism to $\overline{Z-\textnormal{int}\,C}$ using either $f$ or $\mathrm{id}_{\partial C}$ does not preserve the genus function of $\overline{Z-\textnormal{int}\,C}$ due to the assumption on $Z$.
 The corollary below is straightforward.
 Corollary 7.4.
-If ${\iota}_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is an isomorphism, then the genus function of ${X^{\prime}}$ is algebraically equivalent to that of ${X}$.
-The corollary below provides sufficient conditions that $g_{X^{\prime}}$ is algebraically equivalent to $g_{X}$.
+If ${\iota}_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is an isomorphism, then the genus function of $X^{\prime}$} is algebraically equivalent to that of X{X}.
+The corollaries below provide sufficient conditions that $g_{X^{\prime}}$ is algebraically equivalent to $g_{X}$.
+Corollary 7.5.
+If $P$ is an invertible homology cobordism from $M$ to $N$, then the genus function of $X^{\prime}$} is algebraically equivalent to that of X{X}.
+Proof.
+This is straightforward from Corollary 7.4. ∎
 Corollary 7.6.
@@
     The intersection form of $X$ is non-degenerate, and $H_{2}(X)$ has no torsion.
-Then the genus function of ${X^{\prime}}$ is algebraically equivalent to that of ${X}$.
+Then the genus function of $X^{\prime}$} is algebraically equivalent to that of X{X}.
 Proof.
-This is straightforward from Corollary 7.4, Lemma 6.11, and Proposition 6.15. ∎
+This is straightforward from Corollary 7.4, Lemma 6.11, and Propositions 6.14 and 6.13. ∎
 In the case where $P$ is $H_{2}$-surjective, the genus function $g_{X^{\prime}}$ is completely determined by $g_{X}$ and $\iota_{*}$ due to Theorem 7.1 and Lemma 6.11.
 In the case where $P$ is not $H_{2}$-surjective, we give estimates on the values of $g_{X^{\prime}}$ for all classes.
-As in Subsection 6.2, let $B_{P}$ be the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$ induced by the inclusion, where $R$ is the quasi-product cobordism defined in the proof of Theorem 7.1.
-Also, let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions.
-Due to Lemma 6.11, we have $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(B_{P})$.
+As in Subsection 6.2, let $K^{R}_{P}$ be the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$ induced by the inclusion, and let $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusion.
+Due to Lemma 6.11, we have $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(K^{R}_{P})$.
 We prove the following estimates for $g_{X^{\prime}}$.
-Proposition 7.6.
-The following inequality holds for any $\alpha\in H_{2}(X)$ and any $\beta\in B_{P}$.
+Theorem 7.7.
+Every classes $\alpha\in H_{2}(X)$ and $\beta\in K^{R}_{P}$ satisfy
 $g_{X}(\alpha)\leq g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\leq g_{X}(\alpha)+g_{P}(\beta).$
@@
 The right inequality is obvious, so we prove the left inequality.
-Let $X^{\prime\prime}$ be as in the proof of Theorem 7.1, and let $j:X\hookrightarrow X^{\prime\prime}$ and $j^{\prime}:X^{\prime}\hookrightarrow X^{\prime\prime}$ be the inclusions.
-By the definition of $B_{P}$, we see $j^{\prime}_{*}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))=j_{*}(\alpha)$, showing $g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\geq g_{X^{\prime\prime}}(j_{*}(\alpha))$.
+Let $j:X\hookrightarrow X^{\prime\prime}$ and $j^{\prime}:X^{\prime}\hookrightarrow X^{\prime\prime}$ be the inclusions.
+By the definition of $K^{R}_{P}$, we see $j^{\prime}_{*}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))=j_{*}(\alpha)$, showing $g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\geq g_{X^{\prime\prime}}(j_{*}(\alpha))$.
 Since $j_{*}$ preserves the genus functions due to Theorem 7.1, we obtain $g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\geq g_{X}(\alpha)$.
 ∎
-There are many examples of quasi-invertible cobordisms $P$ such that $g_{P}$ is the zero-map, and for such examples, $g_{X^{\prime}}$ is completely determined even when $P$ is not $H_{2}$-surjective due to Proposition 7.6.
-Similarly to ordinary genus functions, torsion-free genus functions are preserved by the inclusions as easily seen from Proposition 7.6 and Lemma 6.11.
+As seen from Lemma 6.7, there are many examples of quasi-invertible cobordisms $P$ such that $g_{P}$ is the zero-map, and for such examples, Theorem 7.7 shows that $g_{X^{\prime}}$ is completely determined by $g_{X}$ even when $P$ is not $H_{2}$-surjective.
+Under an additional assumption, the inclusion furthermore preserves the torsion-free genus functions.
 Corollary 7.8.
-The inclusion $\iota:X\hookrightarrow X^{\prime}$ preserves the torsion-free genus functions.
+Suppose the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ induced by the inclusion is injective.
+Then, the inclusion $\iota:X\hookrightarrow X^{\prime}$ preserves the torsion-free genus functions.
+Proof.
+Let $A$ be a class of $H_{2}(X)/\mathrm{Tor}$, and let $\alpha$ be a class of $H_{2}(X)$ that represents $A$ and satisfies $g_{X}(\alpha)=g^{*}_{X}(A)$. By Theorem 7.1, we see
+$g^{*}_{X}(A)=g_{X}(\alpha)=g_{X^{\prime}}({\iota}_{*}(\alpha))\geq g^{*}_{X^{\prime}}({\iota}_{*}(A)).$
+Let $\alpha^{\prime}$ be a class of $H_{2}(X^{\prime})$ that represents ${\iota}_{*}(A)\in H_{2}(X^{\prime})/\mathrm{Tor}$ and satisfies $g_{X^{\prime}}(\alpha^{\prime})=g^{*}_{X^{\prime}}({\iota}_{*}(A))$.
+Proposition 6.16 implies that there exist torsion classes $\tau\in H_{2}(X)$ and $\beta\in K^{R}_{P}$ satisfying $\alpha^{\prime}={\iota}_{*}(\alpha)+{\iota}_{*}(\tau)+{\iota_{P}}_{*}(\beta)$.
+Theorem 7.7 and the above inequality thus imply
+$g^{*}_{X^{\prime}}({\iota}_{*}(A))=g_{X^{\prime}}(\alpha^{\prime})\geq g_{X}(\alpha+\tau)\geq g^{*}_{X}(A)\geq g^{*}_{X^{\prime}}({\iota}_{*}(A)),$
+showing $g^{*}_{X^{\prime}}({\iota}_{*}(A))=g^{*}_{X}(A)$. Therefore, the inclusion $\iota:X\hookrightarrow X^{\prime}$ preserves the torsion-free genus functions. ∎
+We remark that the assumption of the above corollary holds if $P$ is strongly quasi-invertible. For other sufficient conditions, see Proposition 6.17.
 7.2. Stabilities for algebraic inequivalences of genus functions
-We next discuss stabilities for algebraic inequivalences of genus functions under attachments of quasi-invertible cobordisms.
-In the rest of this section, we use the following setting.
+Throughout this subsection, we use the following setting.
 For $i=1,2$, let $M_{i},N_{i}$ be closed 3-manifolds, let $P_{i}$ be a quasi-invertible cobordism from $M_{i}$ to $N_{i}$, and let $Q_{i}$ be its quasi-inverse cobordism.
-Also, let $X_{i}$ be a connected 4-manifolds having a boundary component diffeomorphic to $M_{i}$, and we denote a glued 4-manifold $X_{i}\cup_{M_{i}}P_{i}$ by $X^{\prime}_{i}$.
-We attach $Q_{i}$ to $X_{i}^{\prime}$ so that the submanifold $R_{i}=P_{i}\cup_{N_{i}}Q_{i}$ is a quasi-product cobordism.
-Let $X^{\prime\prime}$ denote the resulting 4-manifold $X_{i}\cup_{M_{i}}P_{i}\cup_{N_{i}}Q_{i}$.
-Let $\iota:X_{i}\hookrightarrow X^{\prime}_{i}$ and $j:X_{i}\hookrightarrow X^{\prime\prime}_{i}$ be the inclusions.
-Furthermore, we define the subgroup ${\iota_{P_{i}}}_{*}(B_{P_{i}})$ of $H_{2}(X_{i}^{\prime})$ as in Subsection 6.2.
-The theorem below shows a stability for algebraic inequivalence of genus functions.
+Also, let $X_{i}$ be a connected 4-manifold having a boundary component orientedly diffeomorphic to $M_{i}$, and we denote a glued 4-manifold $X_{i}\cup_{M_{i}}P_{i}$ by $X^{\prime}_{i}$.
+We attach $Q_{i}$ to $X_{i}^{\prime}$ so that the submanifold $R_{i}=P_{i}\cup_{N_{i}}Q_{i}$ is a quasi-product cobordism, and we denote the resulting 4-manifold $X_{i}\cup_{M_{i}}P_{i}\cup_{N_{i}}Q_{i}$ by $X^{\prime\prime}$.
+Let $\iota_{i}:X_{i}\hookrightarrow X^{\prime}_{i}$ and $j_{i}:X_{i}\hookrightarrow X^{\prime\prime}_{i}$ be the inclusions.
+Furthermore, we define the subgroup ${{\iota_{i}}_{P_{i}}}_{*}(K^{R_{i}}_{P_{i}})$ of $H_{2}(X_{i}^{\prime})$ as in Subsection 6.2.
+We will show that algebraic inequivalences of (torsion-free) genus functions are stable under attaching a quasi-invertible cobordism.
+Our results thus show that exotic 4-manifolds with algebraically inequivalent genus functions always yield many more (orientedly) exotic 4-manifolds by attaching quasi-invertible cobordisms.
+We first prove the following theorem, which shows stabilities by taking the contrapositions.
 Theorem 7.9.
 $(1)$ Suppose each ${\iota_{i}}_{*}:H_{2}(X_{i})\to H_{2}(X^{\prime}_{i})$ is an isomorphism.
-Then the genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
-$(2)$ Suppose each ${\iota_{i}}_{*}:H_{2}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime}_{i})/\mathrm{Tor}$ is an isomorphism.
-Then the torsion-free genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
+Then, the genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
+$(2)$ Suppose that each ${\iota_{i}}_{*}:H_{2}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime}_{i})/\mathrm{Tor}$ is an isomorphism and that each ${j_{i}}_{*}:H_{2}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime\prime}_{i})/\mathrm{Tor}$ is injective.
+Then, the torsion-free genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
 Proof.
@@
 ∎
-For general quasi-invertible cobordisms, the stability of algebraic inequivalences still holds under additional assumptions.
-See Propositions 6.14 and 6.15 and Lemma 6.16 for sufficient conditions that the assumption below holds.
+We remark that the assumptions of this theorem hold in the case where each $P_{i}$ is strongly quasi-invertible and $H_{2}$-surjective.
+See Propositions 6.14 and 6.17 and Lemma 6.11 for other sufficient conditions.
+For general quasi-invertible cobordisms, we give a stability of algebraic inequivalences under additional assumptions.
+See Propositions 6.13 and 6.14 and Lemma 6.15 for sufficient conditions that the assumption below holds.
 Theorem 7.10.
 Suppose the intersection form of each $X_{i}$ is non-degenerate.
-Suppose further that the groups $H_{2}(X_{1})$ and $H_{2}(X_{2})$ or the groups ${\iota_{P_{1}}}_{*}(B_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(B_{P_{2}})$ are torsion-free and that each ${j_{i}}_{*}:H_{2}(X_{i})\to H_{2}(X^{\prime\prime}_{i})$ is injective.
+Suppose further that the groups $H_{2}(X_{1})$ and $H_{2}(X_{2})$ or the groups ${\iota_{P_{1}}}_{*}(K^{R_{1}}_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(K^{R_{2}}_{P_{2}})$ are torsion-free and that each ${j_{i}}_{*}:H_{2}(X_{i})\to H_{2}(X^{\prime\prime}_{i})$ is injective.
 If the genus functions of $X^{\prime}_{1}$} and $X^{\prime}_{2}$ are algebraically equivalent, then those of $X_{1}$} and $X_{2}$ are algebraically equivalent.
-If furthermore $H_{2}(X_{1})$ and $H_{2}(X_{2})$ are torsion-free, then the restrictions of the genus functions of ${X^{\prime}_{1}}$ and $X^{\prime}_{2}$ respectively to ${\iota_{P_{1}}}_{*}(B_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(B_{P_{2}})$ are also algebraically equivalent.
+If furthermore $H_{2}(X_{1})$ and $H_{2}(X_{2})$ are torsion-free, then the restrictions of the genus functions of $X^{\prime}_{1}$} and $X^{\prime}_{2}$ respectively to ${\iota_{P_{1}}}_{*}(K^{R_{1}}_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(K^{R_{2}}_{P_{2}})$ are also algebraically equivalent.
 Proof.
-Due to Propositions 5.3 and 5.2, the claim follows from Lemma 6.11, Propositions 6.14 and 7.6 and, Theorem 7.1. ∎
-For torsion-free genus functions, a similar stability holds under a much simpler assumption.
+Due to Propositions 5.3 and 5.2, the claims follow from Lemma 6.11, Proposition 6.13 and Theorems 7.1 and $7.$ ∎
+For torsion-free genus functions, we give a similar stability under a much simpler assumption.
 Theorem 7.11.
 Suppose the intersection form of each $X_{i}$ is non-degenerate.
-If the torsion-free genus functions of ${X^{\prime}_{1}}$ and $X^{\prime}_{2}$ are algebraically equivalent, then those of ${X_{1}}$ and $X_{2}$ are algebraically equivalent.
+If the torsion-free genus functions of $X^{\prime}_{1}$} and $X^{\prime}_{2}$ are algebraically equivalent, then those of $X_{1}$} and $X_{2}$ are algebraically equivalent, and those of $X^{\prime}_{1}$} and $X^{\prime}_{2}$} restricted respectively to ${\iota_{P_{1}}}_{*}(K^{R_{1}}_{P_{1}})/\mathrm{Tor}$ and ${\iota_{P_{2}}}_{*}(K^{R_{2}}_{P_{2}})/\mathrm{Tor}$ are algebraically equivalent.
 Proof.
-Since the intersection form of each $X_{i}$ is non-degenerate, we see that $\iota_{*}:H_{"}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime}_{i})/\mathrm{Tor}$ is injective.
-Due to Proposition 5.4, the claim follows from Lemmas 6.11 and 6.13, Proposition 7.6 and Corollary 7.7.
+Due to Proposition 6.16, Theorem 7.7 and Corollary 7.8, we easily see that every $\alpha\in{\iota_{i}}_{*}(H_{2}(X_{i}))/\mathrm{Tor}$ and $\beta\in{\iota_{P_{1}}}_{*}(K^{R_{1}}_{P_{1}})/\mathrm{Tor}$ satisfy $g^{*}_{X^{\prime}_{i}}(\alpha)\leq g^{*}_{X^{\prime}_{i}}(\alpha+\beta)$ for each $i$ by an argument similar to the proof of Corollary 7.8.
+The claim thus follows from Propositions 5.3, 5.2 and 6.16 and Corollary 7.8.
 ∎
+We remark that Theorems 7.10 and 7.11 do not hold if the non-degeneracy assumption is removed as seen from Remark 8.11.
 8. Genus functions, connected sums and boundary sums
@@
 8.1. Submanifolds of $S^{4}(L)$
-Before discussing genus functions, we show that various 4-manifolds are realized as codimension-00 submanifolds of the surgered 4-manifold $S^{4}(L)$ (see Definition 6.3) for some $S^{2}$-link $L$ in $S^{4}$.
+Before discussing genus functions, we show that various 4-manifolds are realized as codimension-0 submanifolds of the surgered 4-manifold $S^{4}(L)$ (see Definition 6.3) for some $S^{2}$-link $L$ in $S^{4}$.
+Note that $S^{4}(L)=S^{4}$ in the case where $L$ is an empty set.
 Lemma 8.1.
-Let $L,L^{\prime}$ be $S^{2}$-links in $S^{4}$, and let $V,V^{\prime}$ be connected codimension-00 submanifolds of $S^{4}(L),S^{4}(L^{\prime})$ with non-empty boundaries, respectively.
+Let $L,L^{\prime}$ be $S^{2}$-links in $S^{4}$, and let $V,V^{\prime}$ be connected codimension-0 submanifolds of $S^{4}(L),S^{4}(L^{\prime})$ with non-empty boundaries, respectively.
 Then, the following hold.
-Note that $S^{4}(L)=S^{4}$ in the case where $L$ is an empty set.
-$(1)$ The boundary sum $V\natural V^{\prime}$ admits an embedding into $S^{4}(L\sqcup L^{\prime})$, where $L\sqcup L^{\prime}$ denotes the disjoint union of $L$ and $L^{\prime}$ in $S^{4}\#S^{4}\cong S^{4}$.
-$(2)$ Attach 1-handles to $V$. Then the resulting $4$-manifold admits an embedding into $S^{4}(L)$.
-$(3)$ Attach an algebraically canceling pair of $1$- and $2$-handles to $V$. Then the resulting $4$-manifold admits an embedding into $S^{4}(L)$.
-$(4)$ Attach $2$-handles to $V$ along a strongly slice link in $\partial V$ with 00-framings. Then the resulting $4$-manifold admits an embedding into $S^{4}(L)$.
+$(1)$ Any boundary sum $V\natural V^{\prime}$ admits an embedding into $S^{4}(L\sqcup L^{\prime})$, where $L\sqcup L^{\prime}$ denotes the split union of $L$ and $L^{\prime}$ in $S^{4}\#S^{4}\cong S^{4}$.
+$(2)$ If a $4$-manifold is obtained from $V$ by attaching $1$-handles along a connected component of $\partial V$, then the $4$-manifold admits an embedding into $S^{4}(L)$.
+$(3)$ If a $4$-manifold is obtained from $V$ by attaching homotopically canceling pairs of $1$- and $2$-handles, then the $4$-manifold admits an embedding into $S^{4}(L)$.
+$(4)$ If a $4$-manifold is obtained from $V$ by attaching $2$-handles along a strongly slice link in $\partial V$ with 0-framings, then the $4$-manifold admits an embedding into $S^{4}(L)$.
 $(5)$ Let $\ell$ be a link in $\partial(\natural_{n}S^{2}\times D^{2})$ that bounds a disjoint union of smoothly and properly embedded disks in $\natural_{n}S^{2}\times D^{2}$.
-Attach $2$-handles to $\natural_{n}S^{1}\times D^{3}$ along a link $\ell$ with the framings induced from the disks, where $\ell$ is regarded as a link in $\partial(\natural_{n}S^{1}\times D^{3})$ using an arbitrary diffeomorphism $\partial(\natural_{n}S^{2}\times D^{2})\to\partial(\natural_{n}S^{1}\times D^{3})$.
-Then the resulting $4$-manifold admits an embedding into $S^{4}$.
+Suppose a $4$-manifold is obtained from $\natural_{n}S^{1}\times D^{3}$ by attaching $2$-handles along a link $\ell$ with the framings induced from the disks, where $\ell$ is regarded as a link in $\partial(\natural_{n}S^{1}\times D^{3})$ using an arbitrary diffeomorphism $\partial(\natural_{n}S^{2}\times D^{2})\to\partial(\natural_{n}S^{1}\times D^{3})$.
+Then the $4$-manifold admits an embedding into $S^{4}$.
 Proof.
@@
 ∎
-Using the above lemma, we can generate numerous examples. Here we provide simple examples.
+Using the above lemma, we can generate numerous examples. Here we provide simple examples of codimension-0 submanifolds of $S^{4}$.
 Example 8.2.
 Let $\Sigma_{g}$ denote a closed surface of genus $g$.
-Then $\Sigma_{g}\times D^{2}$ $(g\geq 0)$, $\natural_{n}S^{1}\times D^{3}$ $(n\geq 0)$, Mazur-type contractible 4-manifolds, their boundary sums, and their connected sums admit embeddings into $S^{4}$ and hence realized as codimension-00 submanifolds of $S^{4}$.
-Furthermore, these 4-manifolds with 2-handles attached along strongly slice links with 0-framings are realized as codimension-00 submanifolds of $S^{4}$.
+Then $\Sigma_{g}\times D^{2}$ $(g\geq 0)$, $\natural_{n}S^{1}\times D^{3}$ $(n\geq 0)$, Mazur-type contractible 4-manifolds, their boundary sums, and their connected sums admit embeddings into $S^{4}$ and hence are realized as codimension-0 submanifolds of $S^{4}$.
+Furthermore, these 4-manifolds with 2-handles attached along strongly slice links with 0-framings are realized as codimension-0 submanifolds of $S^{4}$.
 We here give sufficient conditions that the second homology group of a submanifold of $S^{4}(L)$ is torsion-free.
@@
 $(1)$ $H_{2}(S^{4}(L))=0$.
-$(2)$ For any codimension-00 submanifold of $S^{4}$, its second homology group is torsion-free.
-$(3)$ If the boundary of a codimension-00 submanifold $Z$ of $S^{4}(L)$ is a disjoint union of rational homology $3$-spheres, then $H_{2}(Z)$ is torsion-free.
+$(2)$ For any codimension-0 submanifold $Z$ of $S^{4}$, $H_{2}(Z)$ is torsion-free.
+$(3)$ If the boundary of a codimension-0 submanifold $Z$ of $S^{4}(L)$ is a disjoint union of rational homology $3$-spheres, then $H_{2}(Z)$ is torsion-free.
 Proof.
@@
 So, we assume $\partial Z$ is non-empty.
-Since $Z-\mathrm{int}\,D^{4}$ is an invertible cobordism from $S^{3}$ to $\partial Z$, and $H_{2}(Z-\mathrm{int}\,D^{4})$ is isomorphic to $H_{2}(Z)$, the claims (2) and (3) follow from Lemma 6.16.
+Then, by Lemma 6.7, $Z-\mathrm{int}\,D^{4}$ is a quasi-invertible cobordism from $S^{3}$ to $\partial Z$, and this cobordism is invertible in the case where $Z$ is a submanifold of $S^{4}$.
+The claims (2) and (3) thus follow from Lemma 6.15, since $H_{2}(Z-\mathrm{int}\,D^{4})$ is isomorphic to $H_{2}(Z)$.
 ∎
@@
 Now we discuss genus functions.
-Throughout this subsection, let $X$ be a ((((possibly closed)))) connected 4-manifold, let $L$ be an $S^{2}$-link in $S^{4}$, and let $Z$ be a connected summand or a codimension-00 connected submanifold of the surgered 4-manifold $S^{4}(L)$.
+Throughout this subsection, let $X$ be a ((possibly closed)) connected 4-manifold, let $L$ be an $S^{2}$-link in $S^{4}$, and let $Z$ be a connected summand or a codimension-0 connected submanifold of the surgered 4-manifold $S^{4}(L)$.
 We show that genus functions are preserved under connected sums and boundary sums with $Z$.
 Theorem 8.4.
-The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the genus functions of $X$ and $X\#Z$.
+The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the genus functions of $X$ and X#ZX\#Z.
 Furthermore, in the case where $\partial X$ and $\partial Z$ are non-empty, the inclusion $X\hookrightarrow X\natural Z$ also preserves the genus functions.
-For the first statement, we identify $H_{2}(X\#Z)$ with $H_{2}(X)\oplus H_{2}(Z)$ by using the isomorphisms $H_{2}(X-\mathrm{int}\,D^{4})\to H_{2}(X)$ and $H_{2}(Z-\mathrm{int}\,D^{4})\to H_{2}(Z)$ induced by the inclusions.
+Note that, due to the definition $X\#Z=(X-\mathrm{int}\,D^{4})\cup_{\partial D^{4}}(Z-\mathrm{int}\,D^{4})$, we identify $H_{2}(X\#Z)$ with $H_{2}(X)\oplus H_{2}(Z)$ by using the isomorphisms $H_{2}(X-\mathrm{int}\,D^{4})\to H_{2}(X)$ and $H_{2}(Z-\mathrm{int}\,D^{4})\to H_{2}(Z)$ induced by the inclusions.
 Proof of Theorem 8.4.
-Since $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is diffeomorphic to $Z\#(I\times S^{3})$, Lemma 6.7 shows that $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is a (strongly) quasi-invertible cobordism from $S^{3}$ to the disjoint union $S^{3}\sqcup\partial Z$.
+Since $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is diffeomorphic to $Z\#(I\times S^{3})$, Lemma 6.7 shows that $Z-\mathrm{int}\,(D^{4}\sqcup D^{4})$ is a strongly quasi-invertible cobordism from $S^{3}$ to the disjoint union $S^{3}\sqcup\partial Z$.
 The 4-manifold $X\#(Z-\mathrm{int}\,D^{4})$ is constructed from $X-\mathrm{int}\,D^{4}$ by attaching this cobordism, so the inclusion $X-\mathrm{int}\,D^{4}\hookrightarrow X\#(Z-\mathrm{int}\,D^{4})$ preserves the genus functions due to Theorem 7.1.
-We note that the inclusions $X-\mathrm{int}\,D^{4}\hookrightarrow X$ and $Z-\mathrm{int}\,D^{4}\hookrightarrow Z$ preserve the genus functions by Lemma 7.2.
+We note that the inclusions $X-\mathrm{int}\,D^{4}\hookrightarrow X$ and $X\#(Z-\mathrm{int}\,D^{4})\hookrightarrow X\#Z$ preserve the genus functions by Lemma 7.2.
 Therefore, the inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)$ preserves the genus functions.
 Now we assume $\partial X$ and $\partial Z$ are non-empty.
-By Lemma 6.7, any boundary sum $(I\times\partial X)\natural Z$ is a (strongly) quasi-invertible cobordism, where the sum is taken along $1\times\partial X$.
+By Lemma 6.7, any boundary sum $(I\times\partial X)\natural Z$ is a strongly quasi-invertible cobordism, where the sum is taken along $1\times\partial X$.
 Since $X\natural Z$ is constructed from $X$ by attaching this cobordism, the inclusion $X\hookrightarrow X\natural Z$ preserves the genus functions due to Theorem 7.1.
@@
 In the case where $H_{2}(Z)\neq 0$, we give the following estimates.
-Proposition 8.5.
+Theorem 8.5.
 Every $\alpha\in H_{2}(X)$ and $\beta\in H_{2}(Z)$ satisfy
@@
 Proof.
-Due to the constructions of $X\#Z$ and $X\natural Z$ in the proof of Theorem 8.4, the claim follows from Proposition 7.6. ∎
-Due to the above proposition, torsion-free genus functions are preserved under connected sums and boundary sums with $Z$.
+Due to the constructions of X#ZX\#Z and $X\natural Z$ in the proof of Theorem 8.4, the claim follows from Theorem 7.7. ∎
+Due to the constructions of X#ZX\#Z and $X\natural Z$ in the proof of Theorem 8.4, Corollary 7.8 shows that the torsion-free genus functions are preserved under connected sums and boundary sums with $Z$.
 Corollary 8.6.
-The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the torsion-free genus functions of $X$ and $X\#Z$.
+The inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the torsion-free genus functions of $X$ and X#ZX\#Z.
 Furthermore, in the case where both $\partial X$ and $\partial Z$ are non-empty, the inclusion $X\hookrightarrow X\natural Z$ also preserves the torsion-free genus functions.
-Remark 8.7.
-The inclusions $H_{2}(Z)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ and $Z\hookrightarrow X\natural Z$ do not necessarily preserve the genus functions.
+We remark that the inclusions $H_{2}(Z)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ and $Z\hookrightarrow X\natural Z$ do not necessarily preserve the genus functions.
 For example, in the case where $X$ is $S^{2}\times S^{2}$ or $S^{2}\times S^{2}-\mathrm{int}\,D^{4}$, we have many counterexamples of $Z$ as easily seen from results of [10], [11].
@@
 8.3. Stabilities for algebraic inequivalences of genus functions
-Now we discuss stabilities for algebraic inequivalences of genus functions.
-In the rest of this section, for each $i=1,2$, let $X_{i}$ be a ((((possibly closed)))) connected 4-manifold, let $L_{i}$ be an $S^{2}$-link in $S^{4}$, and let $Z_{i}$ be a connected summand or a codimension-00 connected submanifold of $S^{4}(L_{i})$.
-We show that algebraic inequivalences of genus functions are stable under connected sums and boundary sums in the case where $H_{2}(Z_{i})=0$. The theorem below is a restatement of Theorem 1.9.
+Throughout this subsection, for each $i=1,2$, let $X_{i}$ be a ((possibly closed)) connected 4-manifold, let $L_{i}$ be an $S^{2}$-link in $S^{4}$, and let $Z_{i}$ be a connected summand or a codimension-0 connected submanifold of $S^{4}(L_{i})$.
+We will show that algebraic inequivalences of genus functions are stable under connected sums and boundary sums.
+Our results thus show that exotic 4-manifolds with algebraically inequivalent genus functions always yield many more (orientedly) exotic 4-manifolds by taking connected sums and/or boundary sums.
+We first show stabilities of algebraic inequivalences in the case where $H_{2}(Z_{i})=0$.
+Specifically, we prove the following theorem, which is a restatement of Theorem 1.9 and shows the stabilities by taking the contrapositions.
 Theorem 8.7.
@@
 Proof.
-Under the assumptions, the genus functions of each $X_{i}\#Z_{i}$ and $X_{i}\natural Z_{i}$ are algebraically equivalent to those of $X_{i}$ due to Theorem 8.4. The claims thus immediately follow. ∎
-The corresponding results hold for torsion-free genus functions as well.
+Under the assumptions, Theorem 8.4 shows that the genus functions of each $X_{i}\#Z_{i}$ and $X_{i}\natural Z_{i}$ are algebraically equivalent to that of $X_{i}$, showing the claims. ∎
+We also prove stabilities for torsion-free genus functions.
 Theorem 8.8.
@@
 Proof.
-Due to Corollary 8.6, the claims follow from the same argument as in the proof of Theorem 8.8. ∎
-In the case where $b_{2}(Z_{i})\neq 0$, similar stabilities still hold under additional assumptions.
-The theorem below is a refinement of Theorem 1.10.
+The claims follow from Corollary 8.6 by the same argument as in the proof of Theorem 8.7. ∎
+In the case where $b_{2}(Z_{i})\neq 0$, we prove similar stabilities of algebraic inequivalences under additional assumptions.
+Specifically, we prove the following theorem which is a refinement of Theorem 1.10.
 See Lemma 8.3 for sufficient conditions that each $H_{2}(Z_{i})$ is torsion-free.
 Theorem 8.9.
-Assume that the intersection form of each $X_{i}$ is non-degenerate and that $H_{2}(X_{i})$’s or $H_{2}(Z_{i})$’s are torsion-free. Then the following hold.
+Assume that the intersection form of each $X_{i}$ is non-degenerate and that the groups $H_{2}(X_{1})$ and $H_{2}(X_{2})$ or the groups $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are torsion-free.
+Then the following hold.
 $(1)$ Suppose that the genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent.
 Then, those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
-If furthermore $H_{2}(X_{i})$’s are torsion-free, then those of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ restricted to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
+If furthermore $H_{2}(X_{1})$ and $H_{2}(X_{2})$ are torsion-free, then the genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ restricted respectively to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
 $(2)$ Suppose that each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty and that the genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent.
 Then, those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
-If furthermore $H_{2}(X_{i})$’s are torsion-free, then those of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ restricted to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
+If furthermore $H_{2}(X_{1})$ and $H_{2}(X_{2})$ are torsion-free, then the genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ restricted respectively to $H_{2}(Z_{1})$ and $H_{2}(Z_{2})$ are algebraically equivalent.
 Proof.
-Since $H_{2}(S^{4}(L_{i}))=0$ for each $i$, the intersection form of each $Z_{i}$ is the zero-map. Due to Propositions 5.3 and 5.2, the claims thus follow from Theorem 8.4 and Proposition 8.5. ∎
-For torsion-free genus functions, the stability holds without the torsion-free assumption.
+Since $H_{2}(S^{4}(L_{i}))=0$ for each $i$, the intersection form of each $Z_{i}$ is the zero-map. Due to Propositions 5.3 and 5.2, the claims thus follow from Theorems 8.4 and 8.5. ∎
+For torsion-free genus functions, we can prove stabilities without the torsion-free assumption.
 Theorem 8.10.
 Assume the intersection form of each $X_{i}$ is non-degenerate. Then the following hold.
-$(1)$ If the torsion-free genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
-$(2)$ If each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty, and the torsion-free genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
+$(1)$ If the torsion-free genus functions of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent, and those of $X_{1}\#Z_{1}$ and $X_{2}\#Z_{2}$ restricted respectively to $H_{2}(Z_{1})/\mathrm{Tor}$ and $H_{2}(Z_{2})/\mathrm{Tor}$ are algebraically equivalent.
+$(2)$ If each $\partial X_{i}$ and $\partial Z_{i}$ are non-empty, and the torsion-free genus functions of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ are algebraically equivalent, then those of $X_{1}$ and $X_{2}$ are algebraically equivalent, and those of $X_{1}\natural Z_{1}$ and $X_{2}\natural Z_{2}$ restricted respectively to $H_{2}(Z_{1})/\mathrm{Tor}$ and $H_{2}(Z_{2})/\mathrm{Tor}$ are algebraically equivalent.
 Proof.
-Since the intersection form of each $Z_{i}$ is the zero-map, the claims follow from Propositions 5.4 and 8.5 and Corollary 8.6. ∎
+Since the intersection form of each $Z_{i}$ is the zero-map, the claims easily follow from Propositions 5.3 and 5.2, Theorem 8.5 and Corollary 8.6. ∎
 Remark 8.11.
-Theorems 8.10 and 8.11 do not hold in the case where the intersection forms of $X_{1},X_{2}$ are degenerate by the following trivial reason.
-Let $X_{1},X_{2}$ be an exotic pair of 4-manifolds with algebraically inequivalent genus functions, and assume they admits embeddings into $S^{4}$.
-Such examples were produced by Akbulut and the author ([10], [11]), see also Theorem 3.3.
+Theorems 8.9 and 8.10 do not hold in the case where the intersection forms of $X_{1},X_{2}$ are degenerate by the following simple reason.
+Let $X_{1},X_{2}$ be an exotic pair of 4-manifolds with algebraically inequivalent genus functions, and assume they admit embeddings into $S^{4}$.
+Note that there are various examples of such pairs ([10], [11]), see also Theorem 3.7.
 Set $Z_{1}=X_{2}$ and $Z_{2}=X_{1}$.
-In this case $X_{1},X_{2},Z_{1},Z_{2}$ satisfy the assumption of the above theorem except the non-degeneracy assumption, but clearly the conclusions do not hold.
-As an application of Theorem 8.8, we show that algebraic equivalent classes of genus functions do not determine homeomorphism types even within 4-manifolds having isomorphic cohomology rings and arbitrarily large $b_{2}$.
-Proposition 8.13.
-For any closed connected $4$-manifold $X$, there exist infinitely many pairwise non-homeomorphic closed $4$-manifolds with pairwise algebraically equivalent genus functions such that their cohomology rings are isomorphic to those of $X\#S^{1}\times S^{3}$.
+In this case, $X_{1},X_{2},Z_{1},Z_{2}$ satisfy the assumptions of Theorems 8.9 and 8.10 except the non-degeneracy assumption, but clearly the conclusions do not hold.
+As a simple application, we show that algebraic equivalent classes of genus functions do not determine homeomorphism types even within closed 4-manifolds having isomorphic cohomology rings and arbitrarily large $b_{2}$.
+Theorem 8.12.
+For any closed connected $4$-manifold $X$, there exist infinitely many pairwise non-homeomorphic closed $4$-manifolds with pairwise algebraically equivalent genus functions such that their cohomology rings are isomorphic to that of $X\#S^{1}\times S^{3}$.
 Proof.
@@
 Then $\pi_{1}(S^{4}-K)$ is isomorphic to $\pi_{1}(S^{3}-k)$ (see e.g.
-[18, Section 2.1]), and thus we obtain a desired infinite family of $S^{2}$-knots by varying a knot $k$.
+[19, Section 2.1]), and thus we obtain a desired infinite family of $S^{2}$-knots by varying a knot $k$.
 By the Seifert van-Kampen theorem, we see that each $\pi_{1}(S^{4}(K_{i}))$ is isomorphic to $\pi_{1}(S^{4}-K_{i})$.
@@
 Furthermore, we can easily show that the cohomology ring of $S^{4}(K_{i})$ is isomorphic to that of $S^{1}\times S^{3}$ (cf.
-[17, Section 11]).
-Now let $X$ be a closed 4-manifold, and let $Y_{i}$ be the connected sum $X\#S^{4}(K_{i})$.
-Note that $\pi_{1}(Y_{i})$ is isomorphic to the free product $\pi_{1}(X)*\pi_{1}(S^{4}(K_{i}))$.
-It thus follows from the Grushko decomposition theorem (see e.g.
-[60]) that $\pi_{1}(Y_{i})$’s are pairwise non-isomorphic, showing that $Y_{i}$’s are pairwise non-homeomorphic.
-On the other hand, it is clear that their homology groups, cohomology rings and intersection forms are isomorphic to those of $X\#S^{1}\times S^{3}$.
+[18, Section 11]).
+Note that every $S^{2}$-knot in $S^{4}$ admits a Seifert hypersurface (see e.g.
+[19, Section 2.5]).
+Now let $X$ be a closed connected 4-manifold, and let $Y_{i}$ be the connected sum $X\#S^{4}(K_{i})$.
+Since each $\pi_{1}(Y_{i})$ is isomorphic to the free product $\pi_{1}(X)*\pi_{1}(S^{4}(K_{i}))$, it follows from the Grushko decomposition theorem (see e.g.
+[62]) that the fundamental groups $\pi_{1}(Y_{i})$ are pairwise non-isomorphic, showing that the 4-manifolds $Y_{i}$ are pairwise non-homeomorphic.
+On the other hand, it is clear that their cohomology rings are isomorphic to that of $X\#S^{1}\times S^{3}$.
 Furthermore, by Theorem 8.4, their genus functions are algebraically equivalent to that of $X$ and hence to that of $X\#S^{1}\times S^{3}$.
@@
 In this section, we produce various families of exotic 4-manifolds with respect to genus functions. Note that we follow Convention 1.1.
-We first show that any $4$-dimensional $2$-handlebody can be modified into infinite order generalized corks. The theorem below is a refinement of Theorem 4.1.
+We first show that any $4$-dimensional 22-handlebody can be modified into infinite order generalized corks. Specifically, we prove the following theorem which is a refinement of Theorem 4.1.
 Theorem 9.1.
-For any $4$-dimensional $2$-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ topologically resembles $X$, admits an embedding into $X$, and is homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
-Furthermore, in the case where $H_{2}(X)$ is a non-trivial group, there exist arbitrarily many generalized corks $(X_{i},f_{i})$’s of infinite order such that $X_{i}$’s have the properties of $X^{\prime}$ and are pairwise exotic, and that the genus functions of $X_{i}$’s are pairwise algebraically inequivalent for any orientations.
+For any $4$-dimensional 22-handlebody $X$, there exists a generalized cork $(X^{\prime},f)$ of infinite order such that $X^{\prime}$ is HIHC-equivalent to $X$, admits an embedding into $X$, and is homeomorphic to a $4$-manifold that deformation retracts onto an embedded copy of $X$.
+Moreover, in the case where $H_{2}(X)\neq 0$, there exist arbitrarily many generalized corks $(X_{i},f_{i})$ of infinite order with the properties of $(X^{\prime},f)$ such that the $4$-manifolds $X_{i}$ are pairwise exotic.
+Furthermore, their genus functions are pairwise algebraically inequivalent for any orientations.
 Proof.
 Let $X$ be a 4-dimensional 2-handlebody.
-By Proposition 3.3, we have a 4-manifold $Y$ admitting a Stein structure such that $Y$ topologically resembles $X$, admits an embedding into $X$, and is homeomorphic to a 4-manifold admitting a deformation retraction onto $X$.
+By Proposition 3.6, we have a 4-manifold $Y$ admitting a Stein structure such that $Y$ is HIHC-equivalent to $X$, admits an embedding into $X$, and is homeomorphic to a 4-manifold that deformation retracts onto an embedded copy of $X$.
 Therefore, by Proposition 4.2, we obtain an infinite order generalized cork $(X^{\prime},f)$ having the required properties.
-Now we assume $H_{2}(X)$ is non-trivial.
-Then, by Theorem 3.4, we obtain arbitrarily many pairwise exotic 4-manifolds $Y_{i}$’s such that each $Y_{i}$ have the above properties of $Y$ and that the genus functions of $Y_{i}$’s are pairwise algebraically inequivalent for any orientations.
-Therefore, by the construction in the proof of Proposition 4.2, we obtain arbitrarily many generalized corks $(X_{i},f_{i})$’s of infinite order such that each $X_{i}$ have the above properties of $X^{\prime}$ and that $X_{i}$’s are pairwise homeomorphic.
-Since each $X_{i}$ is obtained by attaching an invertible homology cobordism to $Y_{i}$, the genus function of each $X_{i}$ is algebraically equivalent to that of $Y_{i}$ due to Corollary 7.5, and thus $X_{i}$’s are pairwise exotic.
+Now we assume $H_{2}(X)\neq 0$.
+Then, by Theorem 3.7, we obtain arbitrarily many pairwise exotic 4-manifolds $Y_{i}$ such that each $Y_{i}$ have the above properties of $Y$ and that the genus functions of $Y_{i}$ are pairwise algebraically inequivalent for any orientations.
+Therefore, by the construction in the proof of Proposition 4.2, we obtain arbitrarily many generalized corks $(X_{i},f_{i})$ of infinite order such that the 4-manifolds $X_{i}$ have the above properties of $X^{\prime}$ and are pairwise homeomorphic.
+Since each $X_{i}$ is obtained by attaching an invertible homology cobordism to $Y_{i}$, the genus function of each $X_{i}$ is algebraically equivalent to that of $Y_{i}$ due to Corollary 7.5, and thus the 4-manifolds $X_{i}$ are pairwise exotic.
 This completes the proof.
@@
 Proposition 9.3.
-Let $\{(C,f)\}_{f\in A}$ be a family of pairwise inequivalent generalized corks with $\partial C=M$, and assume that each $f\in A$ extends to a self-homeomorphism of $C$ that preserves the genus function.
-Then there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M}\}}$ of pairwise exotic $4$-manifolds topologically resembling $C$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $C$.
-Furthermore, they admit embeddings into $C$ and deformation retractions onto $C$.
+Let $\{(C,f)\}_{f\in A}$ be a family of pairwise inequivalent generalized corks, and assume each $f\in A$ extends to a self-homeomorphism of $C$ that preserves the genus function.
+Then, there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C}\}}$ of pairwise exotic $4$-manifolds HIHC-equivalent to $C$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $C$.
+Furthermore, they admit deformation retractions onto an embedded copy of $C$ and embeddings into $C$.
 Proof.
-By Corollary 3.10, we have a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M}\}}$ of pairwise exotic 4-manifolds having all the desired properties except those on genus functions.
-We note that $C_{\mathrm{id}_{M}}$ is obtained from $C$ by attaching an invertible cobordism $\mathbb{P}$, and each $C_{f}$ is obtained from $C_{\mathrm{id}_{M}}$ by removing $C$ and gluing it back using $f$.
+For each $f\in A\sqcup\{\mathrm{id}_{\partial C}\}$, let $C_{f}$ denote the 4-manifold $C\cup_{f}\mathbb{P}$ given in the proof of Corollary 3.13.
+By Corollary 3.13 and Lemma 3.11, the family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C}\}}$ has all the desired properties possibly except those on genus functions.
 By Corollary 7.5, the genus function of each $C_{f}$ is algebraically equivalent to that of $C$, and this equivalence is induced by the inclusion $C\hookrightarrow C_{f}$.
-Since each $f\in A$ preserves the orientation and extends to a self-homeomorphism of $C$ that preserves the genus function, this self-homeomorphism of $C$ extends to an orientation-preserving homeomorphism $C_{\mathrm{id}_{M}}\to C_{f}$ that preserves the genus functions.
+Since each $f\in A$ preserves the orientation and extends to a self-homeomorphism of $C$ that preserves the genus function, this self-homeomorphism of $C$ extends to an orientation-preserving homeomorphism $C_{\mathrm{id}_{\partial C}}\to C_{f}$ that preserves the genus functions.
+Therefore, the family has the desired properties on genus functions as well.
 ∎
-Now we will prove results on exotic 4-manifolds given in Section 1. The theorem below is a refinement of Theorem 1.3.
+Now we are ready to produce various infinite families of exotic 4-manifolds with pairwise equivalent genus functions. We here prove the following theorem which is a refinement of Theorem 1.3.
 Theorem 9.4.
-For any $4$-dimensional $2$-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise orientedly equivalent.
-Furthermore, they admit embeddings into $X$ and are homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
+For any $4$-dimensional 22-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ such that their genus functions are pairwise orientedly equivalent.
+Furthermore, these $4$-manifolds admit embeddings into $X$ and are homeomorphic to a $4$-manifold that deformation retracts onto an embedded copy of $X$.
 Proof.
@@
 Theorem 9.5.
-Suppose that a $4$-manifold $X$ admits an embedding into a symplectic $4$-manifold with weakly convex boundary.
-Then, there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$.
-Furthermore, they admit embeddings into $X$ and are homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
+For any $4$-manifold $X$ admitting an embedding into a symplectic $4$-manifold with weakly convex boundary, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$.
+Furthermore, these $4$-manifolds admit deformation retractions onto embedded copies of $X$ and embeddings into $X$.
 Proof.
 This is straightforward from Propositions 4.2 and 9.3. ∎
-The corollary below is a slight refinement of Corollary 1.6.
+The above theorem implies the following corollary, which is a slight refinement of Corollary 1.6.
 Corollary 9.6.
-$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((((with boundary)))) such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
-$(2)$ If $X$ is the total space of a $D^{2}$-bundle over a closed surface, then there exist infinitely many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$.
+$(1)$ For each positive integer $n$, there exist infinitely many pairwise exotic simply connected $4$-manifolds ((with boundary)) such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $n\mathbb{C}P^{2}$.
+$(2)$ For any $D^{2}$-bundle over a closed surface, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to the total space $X$ of the bundle such that their genus functions are pairwise orientedly equivalent and are algebraically equivalent to that of $X$.
 Proof.
-(1) Since $D^{4}$ is a Stein manifold, its blow-ups $n\overline{\mathbb{C}P^{2}}-\textnormal{int}D^{4}$ is a symplectic $4$-manifold with weakly convex boundary.
+(1) Since $D^{4}$ admits a Stein structure, its blow-ups $n\overline{\mathbb{C}P^{2}}-\textnormal{int}D^{4}$ is a symplectic $4$-manifold with weakly convex boundary.
 Also, the genus function of $n\overline{\mathbb{C}P^{2}}-\textnormal{int}D^{4}$ is equivalent to that of $n\overline{\mathbb{C}P^{2}}$ due to Lemma 7.2.
 The claim (1) thus follows from Theorem 9.5.
-(2) Let $X$ denote the total space of $D^{2}$-bundle over a closed surface of genus $g$ with Euler number $n$.
-Then $X$ admits a Stein structure for $n\leq 2g-2$ (see [38, Exercises 11.2.5.(a)]).
+(2) Let $X$ denote the total space of the $D^{2}$-bundle over a closed surface of genus $g$ with Euler number $n$.
+Then $X$ admits a Stein structure for $n\leq 2g-2$ (see [39, Exercises 11.2.5.(a)]).
 Hence the claim (2) follows from Theorem 9.5 in the case where $n\leq 2g-2$.
-In the case where $n\geq 2g-1$ and $g\geq 1$, the claim (2) follows from the case where $n\leq 2g-2$ by reversing the orientation.
-The same argument also shows the claim (2) in the case where $n\geq 2$ and $g=0$.
+In the case where $n\geq 2-2g$, the claim (2) follows from the case where $n\leq 2g-2$ by reversing the orientation.
 The remaining cases are $(g,n)=(0,0),(0,\pm 1)$.
@@
 Theorem 9.7.
-There exist infinitely many pairwise exotic simply connected $4$-manifolds such that their genus functions are pairwise algebraically inequivalent for any orientations and that each of their genus functions is orientedly equivalent to those of another infinitely many pairwise exotic $4$-manifolds.
+There exist infinitely many pairwise exotic simply connected $4$-manifolds whose genus functions are pairwise algebraically inequivalent for any orientations such that, for each $4$-manifold $X$ among them, there exist infinitely many pairwise exotic $4$-manifolds homeomorphic to $X$ whose genus functions are all orientedly equivalent to that of $X$.
 Proof.
@@
 Such examples were obtained by Akbulut and the author in [12] and the author in [70].
-By Proposition 4.2, we obtain an infinite family $\{(Y_{i},f_{i})\}_{i\in\mathbb{Z}}$ of infinite order generalized corks such that each $Y_{i}$ is obtained from $X_{i}$ by attaching the same invertible homology cobordism, and that $Y_{i}$’s are pairwise homeomorphic.
+By the construction in the proof of Proposition 4.2, we obtain an infinite family $\{(Y_{i},f_{i})\}_{i\in\mathbb{Z}}$ of infinite order generalized corks such that each $Y_{i}$ is obtained from $X_{i}$ by attaching the same invertible homology cobordism, and that the 4-manifolds $Y_{i}$ are pairwise homeomorphic.
 By Corollary 7.5, the genus function of each $Y_{i}$ is algebraically equivalent to that of $X_{i}$.
 Due to the proof of Proposition 9.3, by attaching another fixed invertible homology cobordism to each $Y_{i}$ and twisting $Y_{i}$ in the resulting 4-manifold $X^{\prime}_{i,0}$ using each $f_{i}^{j}$, we obtain an infinite family $\{X^{\prime}_{i,j}\}_{(i,j)\in\mathbb{Z}\times\mathbb{Z}}$ of pairwise exotic simply connected 4-manifolds.
-By Proposition 9.3, for each fixed $i$, the genus functions of $X^{\prime}_{i,j}$’s are pairwise equivalent, and furthermore for each fixed $j$, the genus functions of $X^{\prime}_{i,j}$’s are pairwise algebraically inequivalent for any orientations.
+By Proposition 9.3, for each fixed $i$, the genus functions of the 4-manifolds $X^{\prime}_{i,j}$ are pairwise equivalent, and furthermore for each fixed $j$, the genus functions of $X^{\prime}_{i,j}$ are pairwise algebraically inequivalent for any orientations.
 ∎
@@
 For example, any finitely presented group is realized as the fundamental group of infinitely many exotic 4-manifolds satisfying the properties in the above theorem.
-Here we construct similar examples for much more homeomorphism types, by relaxing the condition “infinitely many” to “arbitrarily many”. The theorem below is a refinement of Theorem 1.8.
+Here we construct similar examples for much more homeomorphism types, by relaxing the condition “infinitely many” to “arbitrarily many”.
+Specifically, we prove the following theorem which is a refinement of Theorem 1.8.
 Theorem 9.8.
-For any $4$-dimensional $2$-handlebody $X$ with non-trivial second homology group, there exist arbitrarily many pairwise exotic $4$-manifolds topologically resembling $X$ such that their genus functions are pairwise algebraically inequivalent for any orientations and that each of their genus functions is orientedly equivalent to those of infinitely many pairwise exotic $4$-manifolds.
-Furthermore, all of them admit embeddings into $X$ and are homeomorphic to a $4$-manifold admitting a deformation retraction onto $X$.
+For any $4$-dimensional 22-handlebody $X$ with $H_{2}(X)\neq 0$, there exist arbitrarily many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ whose genus functions are pairwise algebraically inequivalent for any orientations, such that each $4$-manifold $X^{\prime}$ among them satisfies the following.
+  • There exist infinitely many pairwise exotic $4$-manifolds homeomorphic to $X^{\prime}$ whose genus functions are all orientedly equivalent to that of $X^{\prime}$.
+  • $X^{\prime}$ and the above infinitely many $4$-manifolds admit embeddings into $X$ and are homeomorphic to a $4$-manifold that deformation retracts onto an embedded copy of $X$.
 Proof.
 Using infinite order generalized corks of Theorem 9.1 as a building block, we can prove the claim similarly to the proof of Theorem 9.7. ∎
-As an application of our results on genus functions, we give exotic 4-manifolds which are stably exotic under boundary sums with arbitrary Stein manifolds having non-degenerate intersection forms.
-The theorem below is a restatement of Theorem 1.11.
+As an application of our results and Theorem 3.7, we give exotic 4-manifolds which are stably exotic under boundary sums with arbitrary Stein manifolds having non-degenerate intersection forms.
+Specifically, we prove the following theorem which is a refinement of Theorem 1.11.
 Theorem 9.9.
-For any positive integer $n$, there exist pairwise exotic simply connected $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ admitting Stein structures such that, for any pairwise orientedly homeomorphic $4$-manifolds $Y_{1},Y_{2},\dots,Y_{n}$ having non-degenerate intersection forms and admitting Stein structures, the boundary sums $X_{1}\natural Y_{1}$, $X_{2}\natural Y_{2}$, $\dots$, $X_{n}\natural Y_{n}$ are pairwise exotic.
+Let $X$ be a $4$-dimensional 22-handlebody with $H_{2}(X)\neq 0$ whose intersection form is represented by a zero-matrix.
+Then for any positive integer $n$, there exist pairwise exotic $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ HIHC-equivalent to $X$ and admitting Stein structures such that, for any pairwise orientedly homeomorphic $4$-manifolds $Y_{1},Y_{2},\dots,Y_{n}$ having non-degenerate intersection forms and admitting Stein structures, the boundary sums $X_{1}\natural Y_{1}$, $X_{2}\natural Y_{2}$, $\dots$, $X_{n}\natural Y_{n}$ are pairwise exotic.
 Proof.
-Let $X$ be a 4-dimensional 2-handlebody with non-trivial second homology group such that its intersection form is represented by a zero-matrix, and let $Y_{1},Y_{2},\dots,Y_{n}$ be 4-manifolds having non-degenerate intersection forms and admitting Stein structures.
-By Theorem 3.4, we have pairwise exotic $4$-manifolds $X_{1},X_{2},\dots,X_{n}$ admitting Stein structures such that they topologically resemble $X$ and that their genus functions are pairwise algebraically inequivalent for any orientations.
-Since each $X_{i}$ and $Y_{i}$ admit Stein structures, we can apply the argument on estimating genus functions in the proof of Theorem 3.4, and thus we see that the restrictions of the genus functions of $X_{i}\natural Y_{i}$’s to $H_{2}(X_{i})$’s remain pairwise algebraically inequivalent.
-Note that the second homology group of any 2-handlebody is torsion-free.
-It thus follows from Proposition 5.2 that the genus functions of $X_{i}\natural Y_{i}$’s are pairwise inequivalent.
+By Theorem 3.7, we have pairwise exotic $4$-manifolds $X_{1},\dots,X_{n}$ HIHC-equivalent to $X$ and admitting Stein structures such that their genus functions are pairwise algebraically inequivalent for any orientations.
+Let $Y_{1},\dots,Y_{n}$ be pairwise orientedly homeomorphic 4-manifolds having non-degenerate intersection forms and admitting Stein structures.
+Since each $X_{i}$ and $Y_{i}$ admit Stein structures, we can apply the argument on the genus functions of the 4-manifolds $X_{i}$ in the proof of Theorem 3.7 to the restrictions of those of the 4-manifolds $X_{i}\natural Y_{i}$ to $H_{2}(X_{i})$ in the exactly same way.
+We thus see that the restrictions of the genus functions of $X_{i}\natural Y_{i}$ to $H_{2}(X_{i})$ remain pairwise algebraically inequivalent for any orientations.
+Since the second homology group of any 2-handlebody is torsion-free, it follows from Proposition 5.2 that the genus functions of $X_{i}\natural Y_{i}$ are pairwise algebraically inequivalent for any orientations.
+Therefore, the 4-manifolds $X_{i}\natural Y_{i}$ are pairwise exotic.
 ∎
-As seen from the above proof, we can realize various homeomorphism types as those of $X_{i}$’s.
+There are rich examples of 4-manifolds satisfying the above assumption of $X$, hence we can realize various homeomorphism types including non-simply connected ones as those of the above $X_{1},\dots,X_{n}$.
 10. Exotically knotted submanifolds
-As an application of our exotic 4-manifolds in the last section, we modify a submanifold of a 4-manifold with codimension less than two into infinitely many pairwise exotically knotted submanifolds.
-For a codimension-00 submanifold $X$ of a 4-manifold $Z$, the closure of $Z-X$ will be called the exterior of $X$ and denoted by $E(X)$.
-To state our results shortly, we use the following terminologies.
+As an application of our exotic 4-manifolds in the last section, we modify a submanifold of a 4-manifold with codimension at most one into infinitely many pairwise exotically knotted submanifolds.
+Note that we follow Convention 1.1.
+For a codimension-0 submanifold $X$ of a 4-manifold $Z$, the closure of ZXZ-X is called the exterior of $X$ and denoted by $E(X)$. To state our results shortly, we use the following terminologies.
 Definition 10.1.
 Let $X_{1},X_{2}$ be (possibly non-orientable) submanifolds of a (possibly non-orientable) 4-manifold $Z$.
-$(1)$ The pair $X_{1},X_{2}$ will be called exotically knotted if they have the following properties.
+$(1)$ The pair $X_{1},X_{2}$ is called exotically knotted if they have the following properties.
   • (i)
@@
     $X_{1}$ is not smoothly isotopic to $X_{2}$. (We do not assume a smooth isotopy fixes $\partial Z$ pointwise.)
-$(2)$ The pair $X_{1},X_{2}$ will be called exotically embedded if they have the following properties.
+$(2)$ The pair $X_{1},X_{2}$ is called exotically embedded if they have the following properties.
   • (i)
@@
     There exists no self-diffeomorphism of $Z$ that maps $X_{1}$ to $X_{2}$.
-We produce exotically knotted submanifolds of codimension-0. The theorem below is a refinement of Theorem 1.12.
+In the above definition, we require the conditions on orientations and the boundary so that our examples have stronger properties, but our results are still new without these extra conditions.
+Also, our results are new as exotically knotted submanifolds and also as exotically embedded submanifolds.
+We first produce exotically knotted submanifolds of codimension-0. Specifically, we prove the following theorem which is a refinement of Theorem 1.12.
 Theorem 10.2.
-Let $X$ be a codimension-00 submanifold of a $4$-manifold $Z$.
-These manifolds are possibly non-orientable and may have ((((possibly disconnected)))) boundaries.
-Suppose the exterior of $X$ is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary.
-Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-00 submanifolds topologically resembling $X$ such that each of these $4$-manifolds contains $X$ as a submanifold and deformation retracts to $X$.
+Let $X$ be a codimension-0 submanifold of a $4$-manifold $Z$.
+These manifolds may be non-orientable and may have ((possibly disconnected)) boundaries.
+Suppose the exterior of $X$ either is a ((orientable)) 22-handlebody or admits an embedding into a symplectic $4$-manifold with weakly convex boundary, and assume the exterior has connected boundary.
+Then, $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-0 submanifolds HIHC-equivalent to $X$.
+Furthermore, each of these submanifolds contains $X$ as a submanifold and admits a deformation retraction onto $X$.
 To prove this theorem, we give a method for producing exotically knotted submanifolds by using generalized corks obtained from contractible corks.
-We use the following setting only for the lemma and the proposition below.
-Let $\{(C_{0},f)\}_{f\in A}$ be a family of contractible corks with connected boundary having a fixed embedded 3-ball $B_{0}$ in $\partial C_{0}$ such that every $f\in A$ is identity on a neighborhood of the 3-ball $B_{0}$.
-Let $Y$ be a 4-manifold with $\partial Y=M$.
+We use the following setting only for the forthcoming Lemma 10.3 and Proposition 10.4.
+Let $\{(C_{0},f)\}_{f\in A}$ be a family of pairwise inequivalent contractible corks having a fixed embedded 3-ball $B_{0}$ in $\partial C_{0}$ such that every $f\in A$ is identity on a neighborhood of $B_{0}$.
+Let $Y$ be a 4-manifold with non-empty boundary $M$.
 Take the boundary sum $(I\times M)\natural C_{0}$ for the boundary component $1\times M$ using the 3-ball $B_{0}$ of $\partial C_{0}$.
@@
 We set $M^{\prime}=M\#\partial C_{0}$.
-Abusing the notation, let $f$ denote the self-diffeomorphism of $M^{\prime}$ that extends the self-diffeomorphism $f$ of $\partial C_{0}$ as the identity on the rest of $M^{\prime}$.
-We denote the glued 4-manifold $Y\cup_{\mathrm{id}_{M}}P$ by $C$, and assume $\{(C,f)\}_{f\in A}$ is a family of pairwise inequivalent generalized corks.
-We furthermore assume that, for each $f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}$, the twisted double $C_{0}\cup_{f}\overline{C_{0}}$ of $C_{0}$ is diffeomorphic to $S^{4}$.
-We identify the double of $I\times M$ with $[0,2]\times M$.
-Then, for each $f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}$, the twisted double $P\cup_{f}\overline{P}$ can be regarded as the connected sum of $[0,2]\times M$ and the twisted double $C_{0}\cup_{f}\overline{C_{0}}$.
-Since $f$ is identity on a neighborhood of the 3-ball $B_{0}$ in $\partial C_{0}$, it follows that the double $P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ has an embedded 4-ball $D_{P}$ such that every twisted double $P\cup_{f}\overline{P}$ is obtained from $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{P}$ by attaching a “twisted partial double” of $C_{0}$, which is obtained from $C_{0}$ by attaching $\overline{C_{0}}$ using the restriction of $f$ to the exterior of $B_{0}$.
-We here observe the following lemma.
+For each $f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}$, let $\widetilde{f}$ denote the self-diffeomorphism of $M^{\prime}$ that extends the self-diffeomorphism $f$ of $\partial C_{0}$ as the identity on the rest of $M^{\prime}$.
+We denote the glued 4-manifold $Y\cup_{\mathrm{id}_{M}}P$ by $C$, and assume that $\{(C,\widetilde{f})\}_{f\in{A}}$ is a family of pairwise inequivalent generalized corks.
+We furthermore assume that, for each $f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}$, the twisted double $C_{0}\cup_{{f}}\overline{C_{0}}$ of $C_{0}$ is diffeomorphic to $S^{4}$.
+Let $D_{C_{0}}$ be the “partial double” $C_{0}\cup_{\mathrm{id}_{\partial C_{0}-\mathrm{int}\,B_{0}}}\overline{C_{0}}$ of $C_{0}$, which is obtained from $C_{0}$ by attaching $\overline{C_{0}}$ using the identity map on the exterior of $B_{0}$.
+Clearly, $D_{C_{0}}$ is a submanifold of the double $P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$.
+Due to the assumption on the double of $C_{0}$, the disk theorem implies that $D_{C_{0}}$ is diffeomorphic to the 4-ball.
+Since each $f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}$ is identity on a neighborhood of the 3-ball $B_{0}$ in $\partial C_{0}$, it follows that the twisted double $P\cup_{\widetilde{f}}\overline{P}$ is obtained from $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{C_{0}}$ by attaching the “twisted partial double” $C_{0}\cup_{f|_{\partial C_{0}-\mathrm{int}\,B_{0}}}\overline{C_{0}}$ of $C_{0}$, which is obtained from $C_{0}$ by attaching $\overline{C_{0}}$ using the restriction of $f$ to the exterior of $B_{0}$.
+We note that each twisted double $P\cup_{\widetilde{f}}\overline{P}$ is diffeomorphic to the connected sum of $I\times M$ and the twisted double $C_{0}\cup_{f}\overline{C_{0}}$.
 Lemma 10.3.
 $(1)$ $P$ admits a deformation retraction onto the boundary component $M$.
-$(2)$ For each $f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}$, there exists a diffeomorphism $P\cup_{f}\overline{P}\to I\times M$ such that it fixes the boundary $M\sqcup\overline{M}$ pointwise and that its restriction to the submanifold $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{P}$ is independent of $f$.
+$(2)$ For each $f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}$, there exists a diffeomorphism $P\cup_{\widetilde{f}}\overline{P}\to I\times M$ such that it fixes the boundary $M\sqcup\overline{M}$ pointwise and that its restriction to the submanifold $(P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P})-\textnormal{int}\,D_{C_{0}}$ is independent of $f$.
 Proof.
@@
 Since each twisted double $C_{0}\cup_{f}\overline{C_{0}}$ is diffeomorphic to $S^{4}$, the disk theorem shows that the twisted partial double of $C_{0}$ is diffeomorphic to $D^{4}$.
-The claim (2) thus follows from the decomposition of the twisted double of $P$ in the paragraph preceding this lemma.
+Thus, every diffeomorphism between the boundary of the twisted partial double and $\partial D^{4}$ extends to a diffeomorphism between their interiors.
+Therefore, the claim (2) follows from the aforementioned decomposition of the twisted double of $P$.
 ∎
+The following proposition gives exotically knotted codimension-0 submanifolds.
 Proposition 10.4.
-Let $X$ be a codimension-00 submanifold of a $4$-manifold $Z$.
-These manifolds are possibly non-orientable and may have ((((possibly disconnected)))) boundaries.
-If the exterior $E(X)$ is diffeomorphic to $Y$, then $Z$ admits a family $\{X_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M^{\prime}}\}}$ of pairwise exotically knotted codimension-00 submanifolds topologically resembling $X$ such that each $X_{f}$ contains $X$ as a submanifold and admits a deformation retraction onto $X$.
+Let $X$ be a codimension-0 submanifold of a $4$-manifold $Z$.
+These manifolds may be non-orientable and may have ((possibly disconnected)) boundaries.
+If the exterior $E(X)$ is diffeomorphic to the $4$-manifold $Y$, then $Z$ admits a family $\{X_{\widetilde{f}}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C_{0}}\}}$ of pairwise exotically knotted and exotically embedded codimension-0 submanifolds HIHC-equivalent to $X$.
+Furthermore, each $X_{\widetilde{f}}$ contains $X$ as a submanifold and admits a deformation retraction onto $X$.
 Proof.
@@
 Since the double $\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}$ admits a diffeomorphism to $I\times M^{\prime}$ that fixes the boundary $\overline{M^{\prime}}\sqcup M^{\prime}$ pointwise (see Remark 3.12), the submanifold $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ admits a diffeomorphism to the double $P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ that fixes the boundary $\overline{M}\sqcup M$ pointwise.
-Let $\widehat{D}_{P}$ denote the preimage of $D_{P}$ under this diffeomorphism.
+Let $\widehat{D}_{C_{0}}$ denote the preimage of $D_{C_{0}}$ under this diffeomorphism.
 By Lemma 10.3, the 4-manifold $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ is diffeomorphic to $Y$ and hence to the exterior $E(X)$, and this diffeomorphism $\varphi_{\mathrm{id}_{M^{\prime}}}:\widehat{Y}_{\mathrm{id}_{M^{\prime}}}\to E(X)$ fixes the boundary $M$ pointwise.
-We set $D_{P}^{0}=\varphi_{\mathrm{id}_{M^{\prime}}}(\widehat{D}_{P})$.
-For each $f\in A$, we apply generalized cork twists to $S_{\mathrm{id}_{M^{\prime}}}$ and $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ along $(C,f)$, and denote the resulting 4-manifolds $C\cup_{f}\mathbb{P}$ and $C\cup_{f}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ by $S_{f}$ and $\widehat{Y}_{f}$, respectively.
-By the diffeomorphism $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}\to P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ in the last paragraph, each cobordism $P\cup_{f}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ is obtained from the exterior of $\widehat{D}_{P}$ in $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ by attaching a homotopy 4-ball diffeomorphic to the partial twisted double of $C_{0}$.
-Thus, by Lemma 10.3, we have a diffeomorphism $P\cup_{f}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}\to P\cup_{{\mathrm{id}_{M^{\prime}}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ that fixes the exterior of $\widehat{D}_{P}$ and hence the boundary $\overline{M}\sqcup M$ pointwise.
-Extending this diffeomorphism trivially, we obtain a diffeomorphism $\psi_{f}:\widehat{Y}_{f}\to\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that is identity on $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{P}$.
-Let $\varphi_{f}:\widehat{Y}_{f}\to E(X)$ denote the diffeomorphism defined by $\varphi_{f}=\varphi_{\mathrm{id}_{M^{\prime}}}\circ\psi_{f}$.
-For each $f\in A\sqcup\{{\mathrm{id}_{M^{\prime}}}\}$, we denote the submanifolds $\varphi_{f}(S_{f})$, $\varphi_{f}(\overline{\mathbb{P}})$ and $\varphi_{f}(\overline{P})$ of $E(X)$ respectively by $S_{f}^{0}$, ¯$\overline{\mathbb{P}}^{0}_{f}$ and $\overline{P}^{0}_{f}$.
-Let $X_{f}$ be the submanifold ¯$\overline{\mathbb{P}}^{0}_{f}\cup\overline{P}^{0}_{f}\cup X$ of $Z$.
-Then, $X_{f}$’s are pairwise diffeomorphic since each $\varphi_{f}$ fixes the boundary $M$ pointwise.
-Note that they are pairwise orientedly diffeomorphic in the case where $Z$ is orientable.
-Clearly, the submanifold $\overline{\mathbb{P}^{0}_{f}}\cup\overline{P}^{0}_{f}$ of $Z$ is an invertible homology cobordism from $\overline{M}$ to $\overline{N}$, and each $X_{f}$ is obtained from $X$ by attaching this cobordism along $\overline{M}$.
-We thus see that each $X_{f}$ deformation retracts to $X$ and topologically resembles $X$ by Lemmas 3.8 and 10.3.
-We will show that the pairs $X_{f}$’s are pairwise exotically knotted and exotically embedded.
-We note that $Z=S_{f}^{0}\cup X_{f}$ for each $f\in A\sqcup\{{\mathrm{id}_{M^{\prime}}}\}$.
-By Theorem 3.6, the exteriors of $X_{f}$’s are pairwise exotic, and thus the pairs $(Z,X_{f})$’s are pairwise non-diffeomorphic.
-Consequently, $X_{f}$’s are pairwise smoothly non-isotopic.
-In the rest, we will prove that $X_{f}$’s are pairwise topologically ambient isotopic by showing that each $X_{f}$ is topologically ambient isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$.
-Due to the assumption, each $f^{-1}\in A$ extends to a self-homeomorphism of $C$ that is identity on the exterior of the contractible 4-manifold $C_{0}$.
-Extending this homeomorphism trivially, we get a homeomorphism $\widehat{Y}_{f}\to\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that maps $S_{f}$ to $S_{\mathrm{id}_{M^{\prime}}}$.
-Note that this homeomorphism is identity on the exterior $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{P}$.
-Composing this homeomorphism with the diffeomorphism $\psi_{f}$, we obtain a self-homeomorphism of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that fixes the $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{P}$ and hence the boundary $M$ pointwise.
-Therefore, by composing this self-homeomorphism of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ with the diffeomorphisms $\varphi_{\mathrm{id}_{M^{\prime}}}$ and $\varphi_{\mathrm{id}_{M^{\prime}}}^{-1}$, we obtain a self-homeomorphism of $E(X)$ that maps $S_{f}^{0}$ to $S_{{\mathrm{id}_{M^{\prime}}}}^{0}$ and fixes $E(X)-\textnormal{int}\,D_{P}^{0}$ and the boundary $M$ pointwise.
-Extending this self-homeomorphism trivially, we obtain a self-homeomorphism ${\Phi}_{f}$ of $Z$ that maps $X_{f}$ to $X_{{\mathrm{id}_{M^{\prime}}}}$ and fixes the exterior of $D_{P}^{0}$ pointwise.
-Note that ${\Phi}_{f}$ fixes $\partial Z$ pointwise in the case where $\partial Z$ is non-empty.
-The restriction of ${\Phi}_{f}$ to $D_{P}^{0}$ fixes the boundary pointwise.
-Since $D_{P}^{0}$ is diffeomorphic to $D^{4}$, it thus follows from a result of Orson and Powell [55, Corollary D] that the restriction to $D_{P}^{0}$ is topologically isotopic to the identity map on $D_{P}^{0}$ fixing the boundary pointwise.
-Thus, ${\Phi}_{f}$ is topologically isotopic to the identity map on $Z$.
+We set $D_{C_{0}}^{0}=\varphi_{\mathrm{id}_{M^{\prime}}}(\widehat{D}_{C_{0}})$.
+[…]
+[…]
+[…]
+In the rest, we will prove that the submanifolds $X_{\widetilde{f}}$ are pairwise topologically ambient isotopic in $Z$ by showing that each $X_{\widetilde{f}}$ is topologically ambient isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$.
+Due to the assumption, each $\widetilde{f}$} extends to a self-homeomorphism of $C$ that is identity on the exterior of the contractible 4-manifold $C_{0}$.
+Extending this homeomorphism as the identity on the rest of $\widehat{Y}_{\widetilde{f}}$, we get a homeomorphism $\widehat{Y}_{\widetilde{f}}\to\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that maps $S_{\widetilde{f}}$ to $S_{\mathrm{id}_{M^{\prime}}}$.
+Note that this homeomorphism is identity on the exterior $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{C_{0}}$ of $\widehat{D}_{C_{0}}$.
+Composing this homeomorphism with the diffeomorphisms $\varphi_{\widetilde{f}}^{-1}$ and $\varphi_{\mathrm{id}_{M^{\prime}}}$, we obtain a self-homeomorphism of $E(X)$ that maps $S^{0}_{\widetilde{f}}$ to $S^{0}_{{\mathrm{id}_{M^{\prime}}}}$ and fixes $E(X)-\textnormal{int}\,D_{C_{0}}^{0}$ and hence the boundary $M$ of $E(X)$ pointwise.
+Extending this self-homeomorphism as the identity on the rest of $Z$, we obtain a self-homeomorphism ${\Phi}_{\widetilde{f}}$ of $Z$ that maps $X_{\widetilde{f}}$ to $X_{{\mathrm{id}_{M^{\prime}}}}$ and fixes the exterior of $D_{C_{0}}^{0}$ pointwise.
+Note that ${\Phi}_{\widetilde{f}}$ fixes $\partial Z$ pointwise in the case where $\partial Z$ is non-empty.
+Clearly, the restriction of ${\Phi}_{\widetilde{f}}$ to $D_{C_{0}}^{0}$ fixes $\partial D_{C_{0}}^{0}$ pointwise.
+Since $D_{C_{0}}^{0}$ is diffeomorphic to $D^{4}$, the Alexander trick [13] shows that the restriction to $D_{C_{0}}^{0}$ is topologically isotopic to the identity map on $D_{C_{0}}^{0}$ fixing $\partial D_{C_{0}}^{0}$ pointwise.
+Thus, ${\Phi}_{\widetilde{f}}$ is topologically isotopic to the identity map on $Z$.
 Note that this isotopy fixes $\partial Z$ pointwise in the case where $\partial Z$ is non-empty.
-Therefore, each submanifold $X_{f}$ of $Z$ is topologically isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$.
+Therefore, each submanifold $X_{\widetilde{f}}$ of $Z$ is topologically ambient isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$.
 This completes the proof.
@@
 Let $Y$ be a 4-manifold diffeomorphic to the exterior $E(X)$.
-We first consider the case where $Y$ is diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary.
+We first consider the case where $Y$ admits an embedding into a symplectic $4$-manifold with weakly convex boundary.
 Let $(C_{0},\tau)$ be an infinite order contractible cork of Gompf (see Subsection 3.5).
 By results of Akbulut [5], Tange [66] and Naylor and Schwartz [57], the twisted double $C_{0}\cup_{\tau^{i}}\overline{C_{0}}$ is diffeomorphic to $S^{4}$ for each integer $i$.
-Therefore, we can apply Proposition 10.4 using a generalized cork obtained from $Y$ by Proposition 4.2.
+Therefore, we can apply Proposition 10.4 using a generalized cork obtained from $Y$ in the proof of Proposition 4.2.
 The claim thus follows in this case.
-In the case where $Y$ is a 2-handlebody, we first modify $Y$ into a Stein handlebody $Y^{\prime}$ by using Proposition 3.3.
-Note that $Y$ is obtained from $Y^{\prime}$ by attaching the upside down cobordism of an invertible cobordism.
-Furthermore, $\partial Y$ is a deformation retract of this invertible homology cobordism.
-We attach this invertible cobordism to $X$ and denote the resulting 4-manifold by $X^{\prime}$.
-We can regard $X^{\prime}$ as a submanifold of $Z$ containing $X$ as a submanifold.
-Clearly, $X$ is a deformation retract of $X^{\prime}$.
-Therefore, we can prove the claim by applying the argument of the first case to the submanifold $X^{\prime}$.
+We next consider the case where $Y$ is a 2-handlebody.
+In this case, we modify $Y$ into a Stein handlebody $Y^{\prime}$ by using Proposition 3.6.
+Note that $Y$ is obtained from $Y^{\prime}$ by attaching the upside down cobordism of an invertible cobordism $P$.
+Furthermore, $P$ admits a deformation retraction onto the negative boundary $\partial Y$.
+Since $Y$ is diffeomorphic to $E(X)$, we may regard $P$ as a submanifold of $E(X)$ such that the negative boundary of $P$ is $\partial E(X)$.
+Then, the union $X^{\prime}=X\cup P$ is a submanifold of $Z$.
+Clearly, $X^{\prime}$ contains $X$ as a submanifold and admits a deformation retraction onto $X$.
+Since the exterior of $X^{\prime}$ in $Z$ is diffeomorphic to the Stein manifold $Y^{\prime}$, we can prove the desired claim by applying that of the first case to $X^{\prime}$.
 ∎
 Remark 10.5.
-As shown in the proof, each $X_{i}$ in Theorem 10.2 is obtained from $X$ by attaching an invertible homology cobordism, and thus the genus function of $X_{i}$ is algebraically equivalent to that of $X$.
+For Theorem 10.2 and Proposition 10.4, the exotically knotted submanifolds of $Z$ are obtained from $X$ by attaching invertible homology cobordisms embedded in $Z$, and hence they admit embeddings into $X$.
+It thus follows from Corollary 7.5 that the genus functions of these exotically knotted submanifolds are algebraically equivalent to that of $X$.
 We next discuss codimension-1 submanifolds. We use the following terminology.
 Definition 10.6.
-Let $M_{1},M_{2}$ be codimension-1 submanifolds of a (possibly non-orientable) 4-manifold $Z$.
-The manifold $Z$ is possibly non-orientable and may have ((((possibly disconnected)))) boundary.
-The submanifold $M_{1}$ will be called homology cobordant to $M_{2}$ through $Z$ if $Z$ admits a codimension-00 submanifold $H$ with $\partial H=\overline{M_{1}}\sqcup M_{2}$ such that $H$ is a homology cobordism from $M_{1}$ to $M_{2}$.
+Let $Z$ be a (possibly non-orientable) 4-manifold which may have (possibly disconnected) boundary, and let $M_{1},M_{2}$ be codimension-1 submanifolds of $Z$.
+The submanifold $M_{1}$ is called homology cobordant to $M_{2}$ through $Z$ if $Z$ admits a codimension-0 (oriented) submanifold $P$ with $\partial P=\overline{M_{1}}\sqcup M_{2}$ such that $P$ is a homology cobordism from $M_{1}$ to $M_{2}$.
 As an application of Theorem 10.2, we modify separating codimension-$1$ submanifolds into infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds.
-The theorem below is a refinement of Theorem 1.14.
+Specifically, we prove the following theorem which is a refinement of Theorem 1.14.
 Theorem 10.7.
-Let $Z$ be a ((((possibly non-orientable)))) $4$-manifold which may have ((((possibly disconnected)))) boundary, and let $M$ be a closed $3$-manifold embedded in the interior of $Z$ as a separating submanifold.
-Suppose the exterior of the tubular neighborhood of $M$ has a connected component which is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary.
-Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$ through $Z$.
+Let $M$ be a closed $3$-manifold embedded in the interior of a ((possibly non-orientable)) 44-manifold $Z$ which may have ((possibly disconnected)) boundary.
+Suppose the exterior of the tubular neighborhood of $M$ has a connected component with connected boundary which either is a $4$-dimensional ((orientable)) 22-handlebody or admits an embedding into a symplectic $4$-manifold with weakly convex boundary.
+Then, $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$ through $Z$.
 Proof.
-Due to the assumption, we can apply Theorem 10.2 to a connected component $X$ of the exterior of the tubular neighborhood for $M$.
-Let $X_{i}$’s be the resulting infinitely many pairwise exotically knotted and exotically embedded codimension-0 submanifolds of $Z$, and let each $M_{i}$ be the boundary $\partial E(X_{i})$ of the exterior $E(X_{i})$.
+Due to the assumption, we can apply Theorem 10.2 to a component $X$ of the exterior of the tubular neighborhood for $M$.
+Note that $\partial E(X)$ is diffeomorphic to $M$.
+Let $X_{i}$ $(i\in\mathbb{Z})$ be the resulting infinitely many pairwise exotically knotted and exotically embedded codimension-0 (possibly non-orientable) submanifolds of $Z$, and let each $M_{i}$ be the boundary $\partial E(X_{i})$ of the exterior $E(X_{i})$.
+We fix an orientation of the exterior $E(X)$ and equip each $E(X_{i})\subset E(X)$ with the orientation induced from $E(X)$.
+In the case where the orientaion of $M$ is that of $\partial E(X)$ (resp.
+$\partial\overline{E(X)}$), we equip each $M_{i}$ with the orientaion of $\partial E(X_{i})$ (resp.
+$\partial\overline{E(X_{i})}$).
+Clearly, the 3-manifolds $M_{i}$ are pairwise orientedly diffeomorphic.
 By the proof of Theorem 10.2, each $M_{i}$ is homology cobordant to $M$ through $Z$.
-Since $X_{i}$’s are pairwise topologically ambient isotopic, their boundary components $M_{i}$’s are also pairwise topologically ambient isotopic.
-By the proof of Theorem 10.2, the exteriors $E(X_{i})$’s are pairwise non-diffeomorphic, and $X_{i}$’s are pairwise diffeomorphic.
+Since the submanifolds $X_{i}$ are pairwise topologically ambient isotopic, their boundary components $M_{i}$ are also pairwise topologically ambient isotopic.
+We will show that the pairs $(Z,M_{i})$ are pairwise non-diffeomorphic.
+By the proof of Theorem 10.2, the exteriors $E(X_{i})$ are pairwise exotic, and the 4-manifolds $X_{i}$ are pairwise diffeomorphic.
 Note that the exterior of the tubular neighborhood of each $M_{i}$ is the disconnected 4-manifold $E(X_{i})\sqcup X_{i}$.
-If a pair $(Z,M_{i})$ is diffeomorphic to some $(Z,M_{j})$, then $E(X_{i})\sqcup X_{i}$ is diffeomorphic to $E(X_{j})\sqcup X_{j}$ due to the uniqueness of the tubular neighborhood.
-However, in the case where $E(X_{i})$ is not homeomorphic to $X_{i}$, it follows that all $(E(X_{i})\sqcup X_{i})$’s are pairwise non-diffeomorphic, and thus the pairs $(Z,M_{i})$’s are pairwise non-diffeomorphic.
-Therefore, $M_{i}$’s are pairwise exotically embedded and exotically knotted.
-In the case where $E(X_{i})$ is homeomorphic to $X_{i}$, at least infinitely many of $(E(X_{i})\sqcup X_{i})$’s are pairwise non-diffeomorphic by the pigeonhole principle.
-Therefore, at least infinitely many of $M_{i}$’s are pairwise exotically embedded and exotically knotted.
+Now suppose to the contrary that a pair $(Z,M_{i})$ is diffeomorphic to some $(Z,M_{j})$ with $i\neq j$.
+Then, $E(X_{i})\sqcup X_{i}$ is diffeomorphic to $E(X_{j})\sqcup X_{j}$ due to the uniqueness of the tubular neighborhood.
+Since $E(X_{i})$ is not diffeomorphic to $E(X_{j})$, we see that $E(X_{i})$ is diffeomorphic to $X_{j}$ and hence to $X_{i}$ and that $X_{i}$ is diffeomorphic to $E(X_{j})$.
+Consequently, $E(X_{i})$ is diffeomorphic to $E(X_{j})$, giving a contradiction.
+Thus, the pairs $(Z,M_{i})$ are pairwise non-diffeomorphic.
+Therefore, the 3-manifolds $M_{i}$ are pairwise exotically embedded and exotically knotted in $Z$.
 ∎
-The corollaries below are restatements of Corollaries 1.15 and 1.16.
+We obtain the following corollary which is a restatement of Corollary 1.15.
 Corollary 10.8.
-Suppose a closed $3$-manifold admits an embedding into $S^{4}$.
-Then every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds homology cobordant to the $3$-manifold.
+For any closed $3$-manifold $M$ embedded in $S^{4}$, every ((possibly non-orientable)) 44-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$.
 Proof.
@@
 Since $M$ is separating, the boundary $\partial Y$ is diffeomorphic to $M$.
-Also, $Y$ admits an embedding into $D^{4}$ and hence into any symplectic 4-manifold with weakly convex boundary.
-Now let $Z$ be an arbitrary 4-manifold.
-Then $Y$ admits an embedding into the interior of $Z$.
-Let $M$ denote the boundary of $Y$.
-Then we can apply Theorem 10.7 to this $M$, and thus the claim follows.
+Also, $Y$ admits an embedding into $D^{4}$ and hence into a symplectic 4-manifold with weakly convex boundary.
+Now let $Z$ be an arbitrary (possibly non-orientable) 4-manifold.
+Then $Y$ admits an embedding into the interior of $Z$, and thus we can apply Theorem 10.7 to the submanifold $\partial Y$ of $Z$.
+Therefore, the claim follows.
 ∎
+The above corollary implies the following corollary which is a refinement of Corollary 1.16.
 Corollary 10.9.
-There exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold of them, every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to the $3$-manifold.
+For each integer $n\geq 0$, there exists a closed $3$-manifold $M$ with $b_{1}=n$ such that every ((possibly non-orientable)) 44-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$.
 Proof.
-There are infinitely many closed 3-manifolds admitting embeddings into $S^{4}$ such that their homology cobordism classes are pairwise distinct.
-For example, $n(S^{1}\times S^{2})$ admits an embedding into $S^{4}$ for each $n\geq 0$.
+As is well-known, for each integer $n\geq 0$, there are various closed 3-manifolds with $b_{1}=n$ admitting embeddings into $S^{4}$.
+For example, each $n(S^{1}\times S^{2})$ is such a 3-manifold.
 The claim thus follows from Corollary 10.8.
 ∎
-The corollary below is a refinement of Corollary 1.18.
-Note that every closed 3-manifold embedded in a simply connected closed 4-manifold is separating as easily seen from the Mayer–Vietoris exact sequence.
+For a given closed 3-manifold, here we consider the question of what is the minimum second Betti number of a closed 4-manifold which contains a pair of exotically knotted (or exotically embedded) separating submanifolds diffeomorphic to the 3-manifold.
+Since every closed 3-manifold embedded in a simply connected closed 4-manifold is separating as follows from the Mayer–Vietoris exact sequence, this question is an extension of the question mentioned in the paragraph preceding Corollary 1.17, where the ambient 4-manifold is assumed to be simply connected.
+As an application of Theorem 10.7, we show that every non-negative integer is realized as such a minimum number, giving a refinement of Corollary 1.17.
 Corollary 10.10.
-For each non-negative even integer $n$, there exists a closed $3$-manifold $M$ and a simply connected closed $4$-manifold $Z$ with $b_{2}=n$ such that $Z$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$ and that $M$ admits no embedding into any closed $4$-manifold with $b_{2}<n$ as a separating submanifold.
+For each even integer $n\geq 0$ and each integer $m$ with $m\geq n/2$, there exists a closed $3$-manifold $M$ with $b_{1}=m$ such that a ((simply connected)) closed $4$-manifold with $b_{2}=n$ contains infinitely many pairwise exotically knotted and exotically embedded separating submanifolds diffeomorphic to $M$, but no closed $4$-manifold with $b_{2}<n$ contains a separating submanifold diffeomorphic to $M$.
 Proof.
-Let $n$ be a non-negative even integer, and set $r=2/n$.
-In the $n=0$ case, the claim immediately follows from Corollary 10.9, so we prove the $n>0$ case.
+We set $r=n/2\geq 0$ and $k=m-r\geq 0$.
+In the $n=0$ case, the claim follows from Corollary 10.9, so we prove the $n>0$ case.
 Let $K$ be an invertible knot in $S^{3}$ with $|\sigma(K)|\geq 2n-1$, where $\sigma(K)$ denotes the signature of $K$.
-Let $Y_{1}$ be the 4-manifold obtained from $D^{4}$ by attaching a 2-handle along $K$ with 00-framing, and let $M_{r}$ be the boundary of the $r$-fold boundary sum $Y_{r}$ of $Y_{1}$.
-Note that $M_{r}$ is the $r$-fold connected sum of $M_{1}$.
-This 3-manifold $M_{r}$ admits an embedding into the closed 4-manifold $r(S^{2}\times S^{2})$ by taking the double of $Y_{r}$.
-By Theorem 10.7, we obtain a closed 3-manifold $M^{\prime}_{r}$ homology cobordant to $M_{r}$ such that $r(S^{2}\times S^{2})$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M^{\prime}_{r}$.
-Note that $r(S^{2}\times S^{2})$ is simply connected and satisfies $b_{2}=n$.
-We will show that $M^{\prime}_{r}$ does not admit any embedding into a closed 4-manifold with $b_{2}<n$ as a separating submanifold.
-Since $M^{\prime}_{r}$ is homology cobordant to $M_{r}$, we have a homology cobordism from $M^{\prime}_{r}$ to $M^{\prime}_{r}$ that contains $M_{r}$ in the interior as a separating submanifold.
-Now suppose to the contrary that $M^{\prime}_{r}$ is diffeomorphic to a separating submanifold of a closed 4-manifold with $b_{2}<n$.
-We remove the tubular neighborhood of $M^{\prime}_{r}$ from the closed 4-manifold and glue the homology cobordism from $M^{\prime}_{r}$ to $M^{\prime}_{r}$ along the resulting boundary.
-As seen from the Mayer–Vietoris exact sequence, the resulting closed 4-manifold satisfies $b_{2}<n$.
-Clearly, $M_{r}$ is embedded in the resulting closed 4-manifold as a separating submanifold.
-This contradicts a result of Kawauchi that $M_{r}$ admits no such an embedding (see [43, the proof of Theorem 2.5]).
+For example, many torus knots satisfy this condition.
+Let $Y$ be the 4-manifold obtained from $D^{4}$ by attaching a 2-handle along $K$ with 0-framing, and let $Y_{r,k}$ be the boundary sum of $r$ copies of $Y$ and $k$ copies of $S^{2}\times D^{2}$.
+We denote by $M_{r,k}$ the boundary $\partial Y_{r,k}$.
+We can embed each 4-manifold $Y_{r,k}$ and hence $M_{r,k}$ into the closed 4-manifold $r(S^{2}\times S^{2})$ by taking the double of $Y_{r,0}$, which contains $Y_{r,k}$ as a submanifold.
+By Theorem 10.7, we obtain a closed 3-manifold $M^{\prime}_{r,k}$ homology cobordant to $M_{r,k}$ such that $r(S^{2}\times S^{2})$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M^{\prime}_{r,k}$.
+As easily seen, we have $b_{1}(M^{\prime}_{r,k})=b_{1}(M_{r,k})=r+k=m$.
+Note that $r(S^{2}\times S^{2})$ is simply connected and satisfies $b_{2}=2r=n$.
+Since $M^{\prime}_{r,k}$ is homology cobordant to $M_{r,k}$, we have a homology cobordism $P_{r,k}$ from $M^{\prime}_{r,k}$ to $M_{r,k}=M_{r,0}\#k(S^{2}\times S^{1})$.
+By attaching 3-handles to $P_{r,k}$, we obtain a cobordism from $M^{\prime}_{r,k}$ to $M_{r,0}$.
+By further attaching 1-handles to this cobordism and then attaching the upside down cobordism of $P_{r,k}$, we obtain a cobordism $P^{\prime}_{r,k}$ from $M^{\prime}_{r,k}$ to itself whose interior contains $M_{r,0}$ as a separating submanifold.
+Now we show that each $M^{\prime}_{r,k}$ admits no embedding into a closed 4-manifold with $b_{2}<n$ as a separating submanifold.
+Suppose to the contrary that $M^{\prime}_{r,k}$ is diffeomorphic to a separating submanifold of a closed 4-manifold with $b_{2}<n$.
+We remove the tubular neighborhood of the submanifold from the closed 4-manifold and glue the cobordism $P^{\prime}_{r,k}$ along the resulting boundary.
+The resulting closed 4-manifold still satisfies $b_{2}<n$ as follows from the Mayer–Vietoris exact sequence, and furthermore it contains $M_{r,0}$ as a separating submanifold.
+This contradicts a result of Kawauchi that $M_{r,0}$ admits no such embedding ([44, the proof of Theorem 2.5]).
 Therefore, the desired claim follows.
 ∎
+As shown in the above proof, the simply connected closed $4$-manifold with $b_{2}=n$ in Corollary 10.10 can be taken as $\frac{n}{2}(S^{2}\times S^{2})$ for each $n$.
 11. Diffeomorphism invariants of genus function type
 In this section, we introduce a notion of genus function type for diffeomorphism invariants of 4-manifolds and give examples of such invariants which can detect exotic 4-manifolds.
-Furthermore, we show that an invariant of genus function type yields lower bounds for values of genus functions and shares many properties of genus functions.
+Furthermore, we show that any invariant of genus function type yields lower bounds for values of genus functions and shares many properties of genus functions.
 Throughout this section, we allow boundaries of $4$-manifolds to be disconnected.
 11.1. Definitions and examples
-To state the notion, we consider the category of 4-manifolds whose morphisms are orientation preserving diffeomorphisms.
+To state the notion, we consider the category of 4-manifolds whose morphisms are orientation-preserving diffeomorphisms.
 Note that manifolds are oriented due to our convention.
@@
 Let $R$ be an ordered set.
-A functor $G$ from the category of 4-manifolds to the category of $R$-valued maps will be called an oriented diffeomorphism invariant of genus function type if the following conditions hold.
+A functor $G$ from the category of 4-manifolds to the category of $R$-valued maps is called an oriented diffeomorphism invariant of genus function type if the following conditions hold.
   • (i)
-    $G$ assigns to each $4$-manifold $X$ a map $G_{X}:H_{2}(X)\to R$, which will be called the $G$ invariant of $X$.
+    $G$ assigns to each $4$-manifold $X$ a map $G_{X}:H_{2}(X)\to R$, which we call the $G$ invariant of $X$.
   • (ii)
     Let $X,Y$ be arbitrary 4-manifolds.
-For any orientation preserving embedding $\varphi:X\to Y$ and any class $\alpha\in H_{2}(X)$, the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ satisfies $G_{X}(\alpha)\geq G_{Y}(\varphi_{*}(\alpha))$.
-If furthermore $G$ satisfies the condition (ii) for any embedding $\varphi$ that does not necessarily preserves the orientations, then $G$ will be called a ((((unoriented)))) diffeomorphism invariant of genus function type.
+For any orientation-preserving embedding $\varphi:X\to Y$ and any class $\alpha\in H_{2}(X)$, the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ satisfies $G_{X}(\alpha)\geq G_{Y}(\varphi_{*}(\alpha))$.
+If furthermore $G$ satisfies the condition (ii) for any embedding $\varphi$ that does not necessarily preserves the orientations, then $G$ is called a ((unoriented)) diffeomorphism invariant of genus function type.
 Remark 11.2.
@@
 Definition 11.3.
-Let $G$ be a (possibly oriented) diffeomorphism invariant of genus function type, and let $X,Y$ be $4$-manifolds.
-(1) We will say that a continuous map $\varphi:X\to Y$ preserves $G$ invariants if the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ preserves the values of the $G$ invariants.
-(2) The $G$ invariants of $X$ and $Y$ will be called equivalent if there exists a homeomorphism $\varphi:X\to Y$ that preserves the $G$ invariants.
-If $\varphi$ furthermore preserves the orientations, then the $G$ invariants of $X$ and $Y$ will be called orientedly equivalent.
-(3) The $G$ invariants of $X$ and $Y$ will be called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the intersection forms and $G$ invariants.
+Let $G$ be an oriented diffeomorphism invariant of genus function type, and let $X,Y$ be $4$-manifolds.
+(1) We will say that a continuous map $\varphi:X\to Y$ preserves the $G$ invariants if the induced homomorphism $\varphi_{*}:H_{2}(X)\to H_{2}(Y)$ preserves the values of the $G$ invariants.
+(2) The $G$ invariants of $X$ and $Y$ are called equivalent if there exists a homeomorphism $\varphi:X\to Y$ that preserves the $G$ invariants.
+If $\varphi$ furthermore preserves the orientations, then the $G$ invariants of $X$ and $Y$ are called orientedly equivalent.
+(3) The $G$ invariants of $X$ and $Y$ are called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the intersection forms and $G$ invariants.
 Due to Remark 11.2, $G$ invariants of orientedly diffeomorphic 4-manifolds are orientedly equivalent.
 Also, if $G$ invariants of two 4-manifolds are orientedly equivalent, then they are algebraically equivalent.
-Furthermore, if $G$ is an unoriented diffeomorphism invariant of genus function type, then $G$ invariants of unorientedly diffeomorphic 4-manifolds are equivalent.
-We also introduce the following terminologies, extracting properties of genus functions.
+Furthermore, if $G$ is an unoriented diffeomorphism invariant of genus function type, then $G$ invariants of (unorientedly) diffeomorphic 4-manifolds are equivalent.
+We also define “torsion-free $G$ invariants” and related terminologies in the same way as for the torsion-free genus functions.
+We introduce the following terminologies extracting properties of genus functions.
 Definition 11.4.
 Let $G$ be an oriented diffeomorphism invariant of genus function type.
-(1) Let $R$ be an additive semigroup and suppose that the $G$ invariant of each 4-manifold is an $R$-valued map.
-We will say that $G$ is subadditive with respect to gluing if $G$ satisfies the following property for any 4-manifold $X$ with non-empty (possibly disconnected) boundary $M$ and for any cobordism $P$ from $M$ to a (possibly disconnected) closed 3-manifold.
-Set $X^{\prime}=X\cup_{M}P$, and let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions.
-Then the inequality $G_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\leq G_{X}(\alpha)+G_{P}(\beta)$ holds for any $\alpha\in H_{2}(X)$ and any $\beta\in H_{2}(P)$.
-(2) We will say that $G$ is preserved under attachment of a $3$-handle ((((resp.
-a $4$-handle)))) if the following property holds.
-For any 4-manifold $X^{\prime}$ obtained from a 4-manifold $X$ by attaching a 3-handle (resp.
+(1) Let $R$ be an additive magma (i.e., a set equipped with an additve binary operation ++) and suppose that the $G$ invariant of each 4-manifold is an $R$-valued map.
+We will say that $G$ is subadditive with respect to gluing if $G$ has the following property.
+Let $X^{\prime}$ be a 4-manifold obtained from a 4-manifold $X$ by attaching a 4-dimensional cobordism $P$, and let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions.
+Then every $\alpha\in H_{2}(X)$ and $\beta\in H_{2}(P)$ satisfy $G_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\leq G_{X}(\alpha)+G_{P}(\beta)$.
+(2) We will say that $G$ is preserved under attachment of a $3$-handle ((resp.
+a 44-handle)) if for any 4-manifold $X^{\prime}$ obtained from a 4-manifold $X$ by attaching a 3-handle (resp.
 a 4-handle), the inclusion $X\hookrightarrow X^{\prime}$ preserves the $G$ invariants.
-Genus functions clearly have the above subadditivity and are preserved by 3- and 4-handles due to Lemma 7.2.
-We remark that all the properties of genus functions used for the proofs of the results in Sections 7 and 8 are only those in Definitions 11.1 and 11.4.
-We also define “torsion-free $G$ invariants” and related terminologies in the same way as the torsion-free genus functions.
-The genus functions have a property that $g_{X}(0)=g_{D^{4}}(0)$ for any 4-manifold $X$.
-This property gives the following simple obstructions for having the properties on 3- and 4-handles.
-As we will see in Example 11.6, these obstructions are effective.
-Lemma 11.5.
-$(1)$ If $G$ is preserved under attachment of a $4$-handle, then $G_{D^{4}}(0)=G_{S^{4}}(0)$.
-$(2)$ If $G$ is preserved under attachment of a $3$-handle, then every $\alpha\in H_{2}(S^{2}\times D^{2})$ satisfies $G_{S^{2}\times D^{2}}(\alpha)=G_{D^{4}}(0)=G_{S^{1}\times S^{3}-\textnormal{int}D^{4}}(0)$.
-Proof.
-The claim (1) and the equality $G_{S^{2}\times D^{2}}(\alpha)=G_{D^{4}}(0)$ are straightforward.
-Since attachment of a 1-handle preserves $G$ invariants due to Remark 11.11, the equality $G_{D^{4}}(0)=G_{S^{1}\times S^{3}-\textnormal{int}D^{4}}(0)$ follows.
-∎
-Other than genus functions, we give the following examples of genus function type which can detect exotic 4-manifolds.
+(3) Let $P$ be a 4-dimensional cobordism.
+We will say that $G$ is preserved under attachment of $P$ if for any 4-manifold $X^{\prime}$ obtained from a 4-manifold $X$ by attaching the cobordism $P$, the inclusion $X\hookrightarrow X^{\prime}$ preserves the $G$ invariants.
+(4) Let $Z$ be a $4$-manifold.
+We will say that $G$ is preserved under connected sum with $Z$ if for any 4-manifold $X$, the inclusion $H_{2}(X)\hookrightarrow H_{2}(X\#Z)=H_{2}(X)\oplus H_{2}(Z)$ preserves the $G$ invariants of $X$ and X#ZX\#Z.
+In the case where $\partial Z$ is non-empty, we will say that $G$ is preserved under boundary sum with $Z$ if for any 4-manifold $X$ with non-empty boundary and for any choice of a boundary sum $X\natural Z$, the inclusion $X\hookrightarrow X\natural Z$ preserves the $G$ invariants.
+The genus function clearly has the above subadditivity and is preserved under attachment of a 3- and a 4-handle due to Lemma 7.2.
+All the properties of the genus function used in the proofs of the results in Sections 7 and 8 are only those in Definitions 11.1 and 11.4.
+It might be worth to remark that the genus function furthermore has the property that every 4-manifold $X$ satisfies $g_{X}(0)=g_{D^{4}}(0)$.
+We give other examples of oriented diffeomorphism invariants of genus function type which can detect exotic 4-manifolds.
 Example 11.5.
-(1) For a 4-manifold $X$, we define the map $G_{X}:H_{2}(X)\to\mathbb{Z}\cup\{-\infty\}$ as follows.
-Suppose $X$ admits an orientation preserving embedding $\varphi:X\to Z$ into a closed 4-manifold $Z$ with $b_{2}^{+}>1$ having a non-vanishing Seiberg–Witten invariant.
-We furthermore assume $Z$ is of simple type in the sense of [56].
-Let $K\in H^{2}(Z)$ be a Seiberg-Witten basic class of $Z$.
-Then, for $\alpha\in H_{2}(X)$, we define $G_{X}(\alpha)$ as the maximal value of the integer $|\langle K,\varphi_{*}(\alpha)\rangle|$ for all $Z$, $\varphi$ and $K$.
-Note that $G_{X}(\alpha)<\infty$ for any $\alpha$ due to the adjunction inequality ([46], [56]).
-In the case where $X$ admits no embedding into such a closed 4-manifold $Z$, we set $G_{X}(\alpha)=-\infty$ for any $\alpha\in H_{2}(X)$.
-It is easy to see that $G_{X}$’s form an oriented diffeomorphism invariant $G$ of genus function type.
-This invariant can detect various families of orientedly exotic 4-manifolds, as implicitly shown in [11], [12], [67], [68] and [70].
-We can define several variations of the above invariant $G$ by using the Seiberg–Witten invariant and the Bauer–Furuta invariant.
+(1) For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, let $E^{SW}_{X}(\alpha)$ be the maximum value of an integer $|\langle K,\varphi_{*}(\alpha)\rangle|$, where $\varphi:X\to Z$ is an orientation-preserving embedding of $X$ into a closed 4-manifold $Z$ with $b_{2}^{+}>1$ having non-vanishing Seiberg–Witten invariant, and $K$ is a Seiberg–Witten basic class of $Z$.
+We furthermore assume $Z$ is of Seiberg–Witten simple type in the sense of [58] so that the adjunction inequalities ([48], [58]) guarantee $E^{SW}_{X}(\alpha)<\infty$.
+In the case where $X$ admits no embedding into such a closed 4-manifold $Z$, we set $E^{SW}_{X}(\alpha)=$-$\infty$.
+Then, the resulting map $E^{SW}_{X}:H_{2}(X)\to\mathbb{Z}\cup\{-\infty\}$ gives an oriented diffeomorphism invariant of genus function type by varying $X$.
+Algebraic equivalent classes of this invariant ${E}^{SW}$ can detect various families of orientedly exotic 4-manifolds, as implicitly shown in [10], [11], [12], [69], [70] and [72].
+This invariant ${E}^{SW}$ clearly has the subadditivity with respect to gluing, but it is not preserved under attachment of a 4-handle, since $E^{SW}_{D^{4}}(0)=0$ and $E^{SW}_{S^{4}}(0)=$-$\infty$.
+We can define several variations of this invariant by using the Seiberg–Witten invariant and the Bauer–Furuta invariant.
 These invariants will be further discussed in [74].
-The above $G$ clearly has the subadditivity with respect to gluing, but it is preserved under an attachment of neither 3- nor 4-handles.
-The claim for a 4-handle follows from the equalities $G_{D^{4}}(0)=0$ and $G_{S^{4}}(0)=-\infty$ due to Lemma 11.5.
-Similarly, the claim for a 3-handle follows from the equalities $G_{D^{4}}(0)=0$ and $G_{S^{1}\times S^{3}-\textnormal{int}D^{4}}(0)=-\infty$ due to Lemma 11.5.
-Note that for any closed 4-manifold $Z$ with $b_{2}^{+}>1$, the Seiberg–Witten invariant of $Z\#(S^{1}\times S^{3})$ vanishes, since the parities of $b_{2}^{+}-b_{1}$ of these two 4-manifolds are distinct.
 (2) Recently, Ren and Willis [60] introduced an interesting new invariant of 4-manifolds called the lasagna $s$-invariant by using skein lasagna modules.
 The lasagna $s$-invariant (with empty boundary link) of a 4-manifold $X$ is a map $s_{X}:H_{2}(X)\to\mathbb{Z}\cup\{-\infty\}$.
-According to their result [58, Proposition 1.11], this invariant is an oriented diffeomorphism invariant of genus function type.
+According to their result [60, Proposition 1.11], this invariant is an oriented diffeomorphism invariant of genus function type.
 Furthermore, they showed that this invariant can detect many orientedly exotic pairs of 4-manifolds.
+(3) In the rest of this section, an immersed 2-sphere means the one whose singular points are only finitely many transverse double points.
+For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, let $g^{\mathrm{dp}}_{X}(\alpha)$ be the minimum number of double points of an immersed 2-sphere in $X$ representing $\alpha$.
+In the case where $\alpha$ is not represented by any immersed 2-sphere, we set $g^{\mathrm{dp}}_{X}(\alpha)=\infty$.
+Then, the resulting map $g^{\mathrm{dp}}_{X}:H_{2}(X)\to\mathbb{Z}\cup\{\infty\}$ gives an unoriented diffeomorphism invariant of genus function type by varying $X$.
+Similarly to the genus function, determining the values of $g^{\mathrm{dp}}_{X}$ has been an interesting problem, and it is in general quite hard (see e.g [29], [61]).
+As is well-known, in the case where $X$ is simply connected, every $\alpha\in H_{2}(X)$ satisfies $g^{\mathrm{dp}}_{X}(\alpha)<\infty$ (see [39, Remark 1.2.4.]).
+By contrast, in the case where $\pi_{2}(X)=0$, every non-zero class $\alpha\in H_{2}(X)$ satisfies $g^{\mathrm{dp}}_{X}(\alpha)=\infty$.
+In either case, every $\alpha\in H_{2}(X)$ satisfies $g^{\mathrm{dp}}_{X}(\alpha)\geq g_{X}(\alpha)$, since any double point can be eliminated by increasing the genus (see e.g.
+[39, p.38]).
+Using this fact, we easily see that algebraic equivalent classes of $g^{\mathrm{dp}}$ can detect many of infinite families of exotic 4-manifolds in [12], [69], [72].
+The invariant $g^{\mathrm{dp}}$ clearly has the subadditivity with respect to gluing, and we can show that this invariant is preserved under attachment of a 3- and a 4-handle, similarly to the proof of Lemma 7.2.
 (4) Let $G$ be an oriented diffeomorphism invariant of genus function type.
 We can modify $G$ into an unoriented invariant, for example, as follows.
-For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, we set $G^{\prime}_{X}(\alpha)=\max\{G_{X}(\alpha),G_{\overline{X}}(\alpha)\}$.
-The resulting $G^{\prime}$ is an unoriented diffeomorphism invariant of genus function type.
-For example, in the case of the invariant defined in (1), the unoriented version can distinguish various families of unorientedly exotic 4-manifolds as implicitly shown in [11] and [68].
+For a 4-manifold $X$, we set $\overline{G}_{X}=G_{\overline{X}}$.
+Note that the resulting orientation reversal invariant $\overline{G}$ is also of genus function type.
+For a class $\alpha\in H_{2}(X)$, we set $|G|_{X}(\alpha)=\max\{G_{X}(\alpha),G_{\overline{X}}(\alpha)\}$.
+The resulting $|G|$ is an unoriented diffeomorphism invariant of genus function type.
+For example, in the case of the invariant $E^{SW}$ defined in (1), the unoriented version $|E^{SW}|$ can distinguish various families of unorientedly exotic 4-manifolds as implicitly shown in [11] and [70].
 11.2. Properties
@@
 For integers $g,n$ with $g\geq 0$, let $S(g,n)$ be the total space of the $D^{2}$-bundle over the closed surface of genus $g$ with Euler number $n$, and let $\gamma(g,n)$ denote a generator of $H_{2}(S(g,n))(\cong\mathbb{Z})$.
-For $(r,n)\in R\times\mathbb{Z}$, we define an element $A_{G}(r,n)$ of ${}\mathbb{Z}\cup\{\infty\}$ as follows.
-In the case where $G_{S(g,n)}(\gamma(g,n))<r<G_{S(g+1,n)}(\gamma(g+1,n))$ for some integer $g\geq 0$, we set $A_{G}(r,n)=g+1$.
-In the case where $r=G_{S(g,n)}(\gamma(g,n))$ for some integer $g\geq 0$, we set
-$A_{G}(r,n)=\min\{g\in\mathbb{Z}_{\geq 0}\mid r=G_{S(g,n)}(\gamma(g,n))\}.$
-In the case where $r<G_{S(g,n)}(\gamma(g,n))$ (resp.
-$r>G_{S(g,n)}(\gamma(g,n))$) for any integer $g\geq 0$, we set $A_{G}(r,n)=0$ (resp.
-$A_{G}(r,n)=\infty$).
-The resulting function $A_{G}:R\times\mathbb{Z}\to\mathbb{Z}\cup\{\infty\}$ is well-defined due to the following lemma.
+To define the function $A_{G}$, we observe the following lemma.
 Lemma 11.6.
@@
 Proof.
-As easily seen from a handlebody diagram, the 4-manifold $S(g,n)$ admits an orientation-preserving self-diffeomorphism whose induced isomorphism maps $\gamma(g,n)$ to $-\gamma(g,n)$.
+As easily seen from a handlebody diagram, each $S(g,n)$ admits an orientation-preserving self-diffeomorphism whose induced isomorphism maps $\gamma(g,n)$ to $\gamma(g,n)$.
 The claim (1) thus follows.
-Due to the assumption $g_{2}\geq g_{1}$, we easily see that $\gamma(g_{1},n)$ is represented by a closed surface of genus $g_{2}$ with self-intersection number $n$, which is obtained from the base surface of $S(g_{1},n)$ by taking a connected sum with a locally embedded surface.
-We thus have an embedding of $S(g_{2},n)$ into $S(g_{1},n)$ whose induced homomorphism maps $\gamma(g_{2},n)$ to $\gamma(g_{1},n)$ or to $-\gamma(g_{1},n)$.
+Due to the assumption $g_{2}\geq g_{1}$, we easily see that $\gamma(g_{1},n)$ is represented by a closed surface of genus $g_{2}$ with self-intersection number $n$, which is obtained from that of genus $g_{1}$ by taking a connected sum with a locally embedded surface.
+We thus have an embedding of $S(g_{2},n)$ into $S(g_{1},n)$ whose induced homomorphism maps $\gamma(g_{2},n)$ to $\gamma(g_{1},n)$ or to $\gamma(g_{1},n)$.
 The claim (2) thus follows from the definition of genus function type and the claim (1).
 ∎
+For $(r,n)\in R\times\mathbb{Z}$, we define an element $A_{G}(r,n)$ of $\mathbb{Z}\cup\{\infty\}$ as follows. In the case where $r=G_{S(g,n)}(\gamma(g,n))$ for some integer $g\geq 0$, we set
+$A_{G}(r,n)=\min\{g\in\mathbb{Z}_{\geq 0}\mid r=G_{S(g,n)}(\gamma(g,n))\}.$
+In the case where $G_{S(g,n)}(\gamma(g,n))<r<G_{S(g+1,n)}(\gamma(g+1,n))$ for some integer $g\geq 0$, we set $A_{G}(r,n)=g+1$.
+In the case where $r<G_{S(g,n)}(\gamma(g,n))$ (resp.
+$r>G_{S(g,n)}(\gamma(g,n))$) for any integer $g\geq 0$, we set $A_{G}(r,n)=0$ (resp.
+$A_{G}(r,n)=\infty$).
+The resulting function $A_{G}:R\times\mathbb{Z}\to\mathbb{Z}\cup\{\infty\}$ is well-defined due to Lemma 11.6.
 We prove that the function $A_{G}$ gives lower bounds for values of genus functions.
@@
 In the above claim, $\alpha\cdot\alpha$ denotes the self-intersection number of $\alpha$.
-We remark that this inequality can be sharp, depending on the invariant $G$.
-For example, for the invariant in Example 11.6.(1) defined by using the Seiberg–Witten invariant, there are many examples of 4-manifolds and their homology classes for which the above bound gives sharp estimates, realizing arbitrary large values of genus functions.
-This claim can be easily checked using the adjunction inequality for the Seiberg–Witten invariant.
 Proof of Proposition 11.7.
 We set $g_{0}=g_{X}(\alpha)$.
-Then $\alpha$ is represented by a closed surface of genus $g_{0}$ with the self-intersection number $n$.
-We thus have an embedding of $S(g_{0},n)$ into $X$ whose induced homomorphism maps $\gamma(g_{0},n)$ to $\alpha$, showing that $G_{S(g_{0},n)}(\gamma(g_{0},n))\geq G_{X}(\alpha)$.
+Then $\alpha$ is represented by a closed surface of genus $g_{0}$.
+Let $n$ be its self-intersection number.
+We have an embedding of $S(g_{0},n)$ into $X$ whose induced homomorphism maps $\gamma(g_{0},n)$ to $\alpha$, showing that $G_{S(g_{0},n)}(\gamma(g_{0},n))\geq G_{X}(\alpha)$.
 This inequality and the definition of $A_{G}(G_{X}(\alpha),\alpha\cdot\alpha)$ shows $g_{0}\geq A_{G}(G_{X}(\alpha),\alpha\cdot\alpha)$.
@@
 ∎
-We note that for each fixed $n$, the function $A_{G}(\;,n):R\to\mathbb{Z}$ is non-decreasing.
-We observe that this function is the best possible one among similar functions derived from $G$ in the following sense.
-Proposition 11.9.
-Let $n$ be an integer.
-Suppose a non-increasing function $A:R\to\mathbb{Z}\cup\{\infty\}$ satisfies $g_{X}(\alpha)\geq A(G_{X}(\alpha))$ for any $4$-manifold $X$ and any $\alpha\in H_{2}(X)$ with $\alpha\cdot\alpha=n$.
-Then the inequality $A_{G}(G_{X}(\alpha),n)\geq A(G_{X}(\alpha))$ holds for any $4$-manifold $X$ and any $\alpha\in H_{2}(X)$ with $\alpha\cdot\alpha=n$.
-Proof.
-Suppose to the contrary that the inequality $A_{G}(G_{X}(\alpha),n)<A(G_{X}(\alpha))$ holds for some 4-manifold $X$ and some $\alpha\in H_{2}(X)$ with $\alpha\cdot\alpha=n$.
-We set $g_{A}=A(G_{X}(\alpha))$.
-Then, we have $g_{A}>A_{G}(G_{X}(\alpha),n)$, which implies $G_{S(g_{A}-1,n)}(\gamma(g_{A}-1,n))\geq G_{X}(\alpha)$.
-By the assumptions on $A$, this inequality implies $g_{A}-1\geq A(G_{S(g_{A}-1,n)}(\gamma(g_{A}-1,n)))\geq g_{A}$, giving a contradiction.
-Therefore, the claim follows.
-∎
+The inequality in Proposition 11.7 can be sharp, depending on the invariant $G$.
+For example, in the case of the genus function, the equality always holds.
+For the invariant $E^{SW}$ defined in Example 11.5.(1) by using the Seiberg–Witten invariant, we can easily determine all the values of the function $A_{E^{SW}}$ using the adjunction inequality.
+Furthermore, we easily see that there are many examples of 4-manifolds and their homology classes for which the above bound is sharp.
 We next show that $G$ is preserved under attachment of a (quasi-)invertible cobordism.
 Theorem 11.8.
-Let $M,N$ be ((((possibly disconnected)))) closed $3$-manifolds, let $P$ be a cobordism from $M$ to $N$, and let $X$ be a $4$-manifold with $\partial X=M$. Then the following hold.
-$(1)$ If $P$ is invertible, then the inclusion $X\hookrightarrow X\cup_{M}P$ preserves the $G$ invariants.
-$(2)$ Suppose that $G$ is preserved by attachments of $3$- and $4$-handles. If $P$ is quasi-invertible, then the inclusion $X\hookrightarrow X\cup_{M}P$ preserves the $G$ invariants.
+$(1)$ $G$ is preserved under attachment of any invertible cobordism and also under boundary sum with any codimension-0 submanifold of $S^{4}$ having non-empty boundary.
+$(2)$ If $G$ is preserved under attachment of a $4$-handle, then $G$ is preserved under connected sum with any codimension-0 submanifold of $S^{4}$.
+$(3)$ Suppose $G$ is preserved under attachment of a $3$- and a $4$-handle.
+Then, $G$ is preserved under attachment of any quasi-invertible cobordism.
+Furthermore, for any $S^{2}$-link $L$, the invariant $G$ is preserved under boundary sum with any codimension-0 submanifold of $S^{4}(L)$ having non-empty boundary and also under connected sum with any codimension-0 submanifold of $S^{4}(L)$.
 Proof.
-We can prove these claims similarly to Theorem 7.1. See also Remark 7.3. ∎
+We can prove these claims similarly to the proofs of Theorems 7.1 and 8.4. See also Remark 7.3. ∎
 Remark 11.9.
@@
 (2) The conclusion of Theorem 11.8.(3) still holds even if we replace the assumption on 3- and 4-handles with the weaker assumption that $G$ is preserved under attachment of $\natural_{n}(S^{1}\times D^{3})$ along $n(S^{1}\times S^{2})$ for any $n\geq 0$.
-We show that the stabilities of algebraic inequivalences also hold for $G$ invariants. More precisely, we prove the following theorem for attachments of (quasi-)invertible cobordisms.
+Using oriented diffeomorphism invariants of genus function type, we can characterize $S^{4}$ and its submanifolds as follows.
+Proposition 11.10.
+$(1)$ A $4$-manifold is diffeomorphic to $S^{4}$ if and only if any oriented diffeomorphism invariant of genus function type is preserved under connected sum with the $4$-manifold.
+$(2)$ A $4$-manifold with non-empty boundary admits an embedding into $S^{4}$ if and only if any oriented diffeomorphism invariant of genus function type is preserved under boundary sum with the $4$-manifold.
+Proof.
+The “only if” part of the claim (1) is straightforward, and that of (2) is Theorem 11.8. In the rest, we prove the “if” part of each claim.
+(1) For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, we set $D^{S^{4}}_{X}(\alpha)=0$ (resp.
+$D^{S^{4}}_{X}(\alpha)=\infty$) in the case where $X$ is diffeomorphic (resp.
+not diffeomorphic) to $S^{4}$.
+Then, the resulting map $D^{S^{4}}_{X}:H_{2}(X)\to\{0,\infty\}$ gives an oriented diffeomorphism invariant of genus function type.
+Now let $Z$ be a 4-manifold such that any oriented diffeomorphism invariant of genus function type is preserved under connected sum with $Z$.
+Then, $Z$ satisfies $D^{S^{4}}_{S^{4}\#Z}(0)=D^{S^{4}}_{S^{4}}(0)=0$, and hence $Z$ is diffeomorphic to $S^{4}$.
+Therefore, the claim (1) follows.
+(2) For a 4-manifold $X$ and a class $\alpha\in H_{2}(X)$, we set $E^{S^{4}}_{X}(\alpha)=0$ (resp.
+$E^{S^{4}}_{X}(\alpha)=$-$\infty$) in the case where $X$ admits an embedding (resp.
+no embedding) into $S^{4}$ that preserves the orientations.
+Then, the resulting map $E^{S^{4}}_{X}:H_{2}(X)\to\{-\infty,0\}$ gives an oriented diffeomorphism invariant of genus function type.
+Now let $Z$ be a 4-manifold with non-empty boundary such that any oriented diffeomorphism invariant of genus function type is preserved under boundary sum with $Z$.
+Then, $Z$ satisfies $E^{S^{4}}_{D^{4}\natural Z}(0)=E^{S^{4}}_{D^{4}}(0)=0$, and hence $Z$ admits an embedding into $S^{4}$.
+Therefore, the claim (2) follows.
+∎
+Remark 11.11.
+(1) Similarly to the above proof, we see that for any closed 4-manifold, its diffeomorphism type and oriented diffeomorphism type are determined by oriented diffeomorphism invariants of genus function type similar to $D^{S^{4}}$.
+(2) For a closed (possibly disconnected) 3-manifold $M$, we define an oriented diffeomorphism invariant $E^{I\times M}$ in the same way as the above $E^{S^{4}}$.
+Using this invariant similarly to the above proof, we can prove that a cobordism with negative boundary $M$ admits an embedding into $I\times M$ that preserves the orientations if any oriented diffeomorphism invariant of genus function type is preserved under attachment of the cobordism.
+We show that the stabilities of algebraic inequivalences also hold for $G$ invariants and their torsion-free versions.
+More precisely, we prove the following theorem for attachments of (quasi-)invertible cobordisms.
 Theorem 11.12.
-$(1)$ Theorems 7.1, 7.8, 7.9 and 7.10 and Corollaries 7.4, 7.5 and 7.7 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $P$ is invertible.
-$(2)$ Suppose that $G$ is preserved under attachments of $3$- and $4$-handles.
-Then, Theorems 7.1, 7.8, 7.9 and 7.10 and Corollaries 7.4, 7.5 and 7.7 hold even when “genus functions” in each statement is replaced with “$G$ invariants”.
+$(1)$ Theorems 7.9, 7.10 and 7.11 and Corollaries 7.4, 7.5 and 7.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where the cobordism $P$ in Subsection 7.1 is invertible.
+$(2)$ Suppose that $G$ is preserved under attachment of a $3$- and a $4$-handle.
+Then, Theorems 7.9, 7.10 and 7.11 and Corollaries 7.4 and 7.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”.
 Proof.
-Due to Theorem 11.10.(2), we see that the left inequality in Proposition 7.6 holds for $G$ invariants by the identical argument.
-Therefore, we can prove the claims of (2) by the identical arguments.
-Due to Theorem 11.10.(1), we can prove the claims of (1) in the same way.
+Due to Theorem 11.8, we see that the left inequality in Theorem 7.7 and Corollary 7.8 holds for $G$ invariants under the above assumptions by the identical argument.
+Therefore, we can prove the claims by the identical arguments.
 ∎
@@
 Theorem 11.13.
-$(1)$ For boundary sums, Theorems 8.4, 8.8, 8.9, 8.10 and 8.11 and Corollary 8.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets.
+$(1)$ For boundary sums, Theorems 8.7, 8.8, 8.9 and 8.10 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets.
 $(2)$ Suppose that $G$ is preserved under attachment of a $4$-handle.
-Then, Theorems 8.4, 8.8, 8.9, 8.10 and 8.11 and Corollary 8.6 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets.
-If furthermore $G$ is preserved by attachment of a $3$-handle, then these claims and Proposition 8.13 for $G$ invariants hold without any condition on $L$, $L_{1}$ and $L_{2}$.
+Then, Theorems 8.7, 8.8, 8.9 and 8.10 hold even when “genus functions” in each statement is replaced with “$G$ invariants”, in the case where $L$, $L_{1}$ and $L_{2}$ are empty sets.
+If furthermore $G$ is preserved under attachment of a $3$-handle, then these claims and Theorem 8.12 for $G$ invariants hold without any condition on $L$, $L_{1}$ and $L_{2}$.
 Proof.
-Due to Theorem 11.10, we see that the second claim of Theorem 8.4 and the left inequality in the second claim of Proposition 8.5 hold for $G$ invariants by the identical arguments.
+Due to Theorem 11.8, we see that the left inequality in the second claim of Theorem 8.5 and the second claim of Corollary 8.6 hold for $G$ invariants by the identical arguments.
 Therefore, we can prove the claim (1) by the identical arguments.
-In the case where $G$ is preserved under attachment of a $4$-handle, we can prove the claim (2) in the same way.
-We remark that the left inequality in the first claim of Proposition 8.5 also holds for $G$ invariants in this case.
+We can similarly prove the claim (2) under the above assumptions.
 ∎
-The above theorems imply that if algebraic equivalent classes of $G$ invariants can detect a family of exotic 4-manifolds, then they can detect infinitely many families of exotic 4-manifolds.
+The above theorems imply that if algebraic equivalent classes of $G$ invariants can detect a family of orientedly exotic 4-manifolds, then they can detect infinitely many families of orientedly exotic 4-manifolds.
 Corollary 11.14.
 Suppose $G$ admits a family $\{X_{i}\}_{i\in A}$ of pairwise orientedly homeomorphic $4$-manifolds whose $G$ invariants are pairwise algebraically inequivalent. Then the following hold.
-$(1)$ If the members $X_{i}$’s have non-empty boundary, then there exists a doubly indexed family $\{X_{i,j}\}_{(i,j)\in A\times\mathbb{Z}}$ of $4$-manifolds having the following properties.
+$(1)$ If the members $X_{i}$ have non-empty boundary, then there exists a doubly indexed family $\{X_{i,j}\}_{(i,j)\in A\times\mathbb{Z}}$ of $4$-manifolds having the following properties.
   • For each fixed $j\in\mathbb{Z}$, the members of $\{X_{i,j}\}_{i\in A}$ are pairwise orientedly exotic and have pairwise algebraically inequivalent $G$ invariants.
   • For each fixed $i\in A$, the members of $\{X_{i,j}\}_{j\in\mathbb{Z}}$ are pairwise not homeomorphic.
-$(2)$ If $G$ is preserved by a $4$-handle, then the conclusion of the above $(1)$ holds even in the case where $X_{i}$’s are closed $4$-manifolds.
+$(2)$ If $G$ is preserved under attachment of a $4$-handle, then the conclusion of the above $(1)$ holds even in the case where each $X_{i}$ has no boundary.
+$(3)$ If $G$ is preserved under attachment of a $3$- and a $4$-handle, and each $X_{i}$ has no boundary, then the conclusion of the above $(1)$ still holds, and furthermore the members $X_{i,j}$ can be chosen to be closed $4$-manifolds.
 Proof.
 We first consider the case (1).
-By Theorems 11.12 and 11.13, we can construct many examples of desired families, for example, by taking boundary sums of $X_{i}$’s with $S^{1}\times D^{3}$’s.
-Their homeomorphism types are detected by their first homology groups.
-In the (2) case, we can similarly construct many examples of desired families, for example, by taking connected sums of $X_{i}$’s with $S^{1}\times D^{3}$’s.
-We remark that, if $G$ is furthermore preserved by a 3-handle, then we can give desired families of closed 4-manifolds as well, by taking connected sums with $S^{4}(L)$’s.
+By Theorems 11.12 and 11.13, we can construct many examples of desired families, for example, by taking boundary sums of the 4-manifolds $X_{i}$ with codimension-0 submanifolds of $S^{4}$.
+See also Lemma 8.1.
+In the (2) case, we can similarly construct many examples of desired families by taking connected sums instead of boundary sums.
+In the (3) case, we can give desired families of closed 4-manifolds by taking connected sums with $S^{4}(L)$ for an $S^{2}$-link $L$ in $S^{4}$.
 ∎
-In contrast to the above corollary, we also show that there exist various infinite families of exotic 4-manifolds which can not be detected by their $G$ invariants.
-More precisely, we show the following theorem.
+As shown in Remark 11.11, every exotic pair of closed 4-manifolds admits an oriented diffeomorphism invariant of genus function type that can detect the pair.
+By contrast, we show that there exist various infinite families of exotic 4-manifolds which can not be detected by any such invariant.
 Theorem 11.15.
-Proposition 9.3 and Theorems 9.4 and 9.5 hold even when “genus functions” in each statement is replaced with “$G$ invariants”.
-Proof.
-Due to Theorem 11.10, the proofs are identical with the original ones. Note that the boundary of a 4-manifold is assumed to be connected as in the original claims. ∎
+For any $4$-dimensional 22-handlebody $X$, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ such that, for any oriented diffeomorphism invariant $G$ of genus function type, their $G$ invariants are pairwise equivalent.
+Theorem 11.16.
+Suppose that a $4$-manifold $X$ with connected boundary admits an embedding into a symplectic $4$-manifold with weakly convex boundary.
+Then, there exist infinitely many pairwise exotic $4$-manifolds HIHC-equivalent to $X$ such that, for any oriented diffeomorphism invariant $G$ of genus function type, their $G$ invariants are pairwise equivalent and are algebraically equivalent to that of $X$.
+Proof of Theorems 11.15 and 11.16.
+Due to Theorem 11.8, the proofs are identical with those of Theorems 9.4 and 9.5. ∎
+We close this section by making a few remarks.
 Remark 11.17.
@@
 Topological genus functions clearly form an (unoriented) diffeomorphism invariant of genus function type.
-Due to the transversality theorem for locally flat proper submanifolds (see [15, Theorem 1.5] and [57]), the proof of Lemma 7.2 shows that topological genus functions are preserved by 3- and 4-handles.
+Due to the transversality theorem for locally flat proper submanifolds (see [16, Theorem 1.5] and [59]), the proof of Lemma 7.2 shows that topological genus functions are preserved under attachment of a 3- and a 4-handle.
 Since two homeomorphic 4-manifolds have equivalent topological genus functions, they cannot be used for detecting exotic 4-manifolds.
 However, results in this section for topological genus functions might be useful for determining their values and also for detecting homeomorphism types by using stabilities of algebraic inequivalences.
-$(2)$ Due to our convention, an invariant $G$ of genus function type is defined only for compact oriented 4-manifolds, but $G$ can be defined for non-compact 4-manifolds as well at least for genus functions.
-For such an invariant $G$, it easily follows from the definition of genus function type that the $G$ invariants of a 4-manifold and its interior are algebraically equivalent.
+$(2)$ Due to our convention, we defined the notion of genus function type only for (oriented) diffeomorphism invariants of compact 4-manifolds, but in the same manner, we can define the notion more generally for invariants of 4-manifolds which are not necessarily compact.
+For example, the genus function and the invariants $E^{SW}$ and $g^{\mathrm{dp}}$ defined in Example 11.5 are of genus function type in the latter sense.
+For such an invariant $G$, we can show that the $G$ invariants of a 4-manifold and its interior are algebraically equivalent similarly to the proof of Theorem 7.1.
 Thus, for any two 4-manifolds having algebraically inequivalent $G$ invariants, their interiors are not orientedly diffeomorphic.
+On the other hand, we can also define the notion of genus function type for (oriented) diffeomorphism invariants of compact 4-manifolds having connected (possibly empty) boundary.
+For such invariants, our results in this section still hold for 4-manifolds with connected (possibly empty) boundary, as easily seen from the proofs.